A Complete Bibliography of Publications in the ORSA Journal on Computing and the INFORMS Journal on Computing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

14 October 2025 Version 1.40

Title word cross-reference

(2/3)n^{3} [762]. (s, S) [1250]. 0 [422, 1319, 1929, 401, 414, 159, 1588]. {0, 1/2} [728]. 1 [422, 1319, 466, 142, 1929, 401, 414, 159, 143, 1588]. 2 [408]. 2^{31} − 1 [1029]. 3 [1556, 609]. ^{∗} [1216]. ^{+} [1844]. ^{K} [1286]. _{t} [1286, 1090]. b [265]. β [1356]. BMAP/G/1/N [1272]. D [1920]. ℓ_{0} [1910, 1536]. ℓ_{q}(0 < q < 1) [1535]. GI/G/1 [750]. GI/G/c [622]. GI^{X}/Geom/m [485]. ∞ [1286, 53, 591, 590]. K [790, 2056, 1071, 1117, 314, 1115, 692, 1649, 1891, 1230, 1106, 816, 1327, 1036]. k ≥ 3 [692]. L [1843]. L_{0} [1778]. L_{1} [1511]. m [1327, 1898]. M(t)/M/s(t) [726]. M/D_{N}/1 [717]. M/E_{k}/1 [813]. M/G/1 [579]. M/M/c [718]. M^{X}/G^{Y} /1/K + B [678]. F_{2} [802]. N [1521, 1303]. N − 1 [1653]. O(ln n) [1328]. O(n log log n) [743]. O(n log n) [316]. p [572, 375, 1293, 974, 924, 1448, 2077, 1255, 251, 699, 1772, 1903]. ϕ [816]. r [1686]. r_{max} [1550]. T [1342, 1538].

-2 [1303]. -Adaptability [2056]. -Center [1448, 572]. -Closeness [1342].

-Colorable [1649]. -Connected [1327]. -Connectivity [314]. -Constrained [1536]. -Convolutions [1962]. -core [1230]. -Corner [466]. -cut [1036]. -Cuts [728]. -cycle [692]. -D [1556, 609]. -Depot [1891]. -Dispersion [2077]. -Distributed [1538]. -Fold [1327]. -Hub [1903, 375, 924]. -Interdiction [1686]. -Layer [408]. -Linear [802]. -Matching [265]. -Median [142, 974, 1293, 1255, 1106]. -medians [1115]. -Nearest [1071]. -Norm [1511]. -Optimal [1920]. -Percent [1772]. -Persistent [251]. -Regularized [1778]. -Robust [1356]. -Segment [143]. -Shaped [1843]. -way [1117].

/M [1090]. /Ph [1286]. /s/c [1090].

0-1 [582, 964]. 0/1/ [53].

1 [1959, 131, 89, 289]. 1-Type [946]. 11 [498]. 19 [1844, 1777].

2 [535, 158, 272]. 2-approximation [1197]. 2-OPT [2089]. 2003 [850]. 2016 [1553]. 28 [1553].

3/2 [1197]. 3/2-approximation [1197]. 3D [775].

8 [75].

9 [483]. 99a [483].

abandonment [1227]. Abandonments [1613]. absolute [1028]. Abstract [790]. AC [1845, 1143]. accelerate [1283, 631]. Accelerated [1762]. Accelerating [853, 1780, 1310]. Acceleration [24]. Acceptance [273]. Accepting [273]. Access [199, 261, 1985, 792, 1152]. Accounts [1463]. Accuracy [1709, 1440, 1280, 1006]. accurate [532]. Achieving [1189, 1201]. ACO [822]. acquire [367]. Acquiring [210]. Acquisition [1990, 429, 257]. Across [1463, 1196, 1045, 1176]. Actionable [543]. Active [1295, 2015, 1638, 1127]. Activity [451, 1586, 1067, 968]. Acuity [1842]. Acuity-Based [1842]. Adaptability [1987, 2056]. Adaptable [1441]. Adaptation [1134]. adapting [1308]. Adaptive [1475, 1446, 425, 1484, 58, 298, 1987, 1775, 1737, 1717, 1974, 1680, 910, 1808, 1935, 1323, 1338, 1317, 1243, 1262, 421, 1043]. Addendum [483]. Adding [358]. Additional [515]. Additive [430, 1362, 700, 1778]. Addressing [1534, 650]. Adiabatic [1636]. adjoint [675]. Adjustable [1919, 1526, 1882, 1736, 1703, 1247]. Adjusted [956]. Adjustment [128, 1950]. Admission [1683]. Admissions [1844]. Adoption [2002, 1642]. Advance [153, 1683]. Advanced [1892, 830, 1422, 382]. Advances [1746, 842]. Adversaries [480]. Advertisements [1369, 573]. Advertising [1761, 773, 1176]. Aerial [1823]. Affine [27, 115, 380]. Against [480, 1596]. Agenda [892]. Agent [931]. Agent-Based [931]. Agents [761]. Aggregate [1521, 1120, 1289]. Aggregated [1678, 1837]. Aggregation [245, 1511, 134, 1932, 325, 1092, 1274, 1238]. Aggregation/Disaggregation [245]. aggressive [733]. agreeably [423]. agreement [1007]. AGV [1637]. AGV-Assisted [1637]. Ahead [802, 1191]. Ahmed [1403]. AI [1864, 881]. Aid [11]. aided [675, 577]. AILS [1974]. AILS-II [1974]. AIMS [735]. Air [400, 1661, 1237]. Aircraft [790, 1834, 1130]. Airline [916, 569, 922, 818]. Airspace [400]. Alfonso [1612]. Algebraic [253, 1706, 524, 836]. Algorithm [10, 745, 1216, 222, 1415, 1922, 1967, 40, 153, 2078, 1804, 1394, 1574, 1048, 1701, 1890, 286, 1195, 1204, 400, 246, 340, 167, 1324, 1951, 69, 1889, 955, 190, 528, 1883, 135, 957, 303, 275, 245, 781, 318, 1465, 1726, 2092, 701, 313, 2027, 2026, 75, 1616, 1975, 1363, 1891, 1800, 301, 1981, 1818, 1532, 1071, 1318, 1522, 1699, 1486, 27, 1037, 248, 279, 283, 211, 99, 236, 265, 1509, 2091, 1423, 281, 234, 1577, 200, 127, 446, 285, 119, 756, 1320, 1303, 1780, 282, 1370, 1770, 23, 1447, 284, 1576, 60, 1885, 1057, 1570, 1677, 1792, 1943]. Algorithm [1328, 2081, 1594, 316, 1907, 372, 1008, 1292, 1060, 641, 557, 1115, 898, 361, 503, 1101, 1260, 363, 759, 552, 389, 976, 963, 901, 1293, 407, 1257, 532, 1305, 1026, 786, 723, 757, 1023, 417, 326, 874, 961, 795, 567, 1012, 1255, 334, 576, 1129, 1047, 852, 1034, 1214, 1049, 494, 1058, 959, 762, 793, 776, 805, 911, 1184, 1051, 402, 1014, 889, 579, 785, 592, 1224, 807, 980, 507, 1043, 1197, 988, 1238]. algorithm-based [532]. Algorithmic [718, 16, 1160]. Algorithms [205, 310, 343, 187, 209, 139, 739, 299, 148, 208, 1615, 140, 1768, 426, 1610, 787, 1730, 1359, 180, 1639, 117, 443, 1412, 1549, 207, 667, 1839, 142, 271, 1413, 909, 473, 270, 2025, 233, 945, 1675, 1628, 789, 2032, 492, 1421, 202, 365, 486, 344, 1448, 280, 2022, 266, 112, 345, 1635, 1920, 2053, 1798, 1099, 1514, 1674, 152, 913, 929, 120, 87, 1282, 1729, 184, 2052, 1550, 124, 1712, 342, 347, 86, 452, 346, 8, 210, 1464, 1796, 1395, 1602, 36, 195, 348, 1422, 1858, 1356, 1697, 1593, 962, 166, 1218, 1102, 1121, 817, 607, 475, 896, 967, 798, 838]. algorithms [1200, 1182, 1138, 1076, 442, 924, 599, 1075, 1278, 1199, 630, 808, 698, 1311, 1150, 868, 903, 367, 971, 626, 1106, 969, 840, 650, 511, 1236]. Alias [640]. Aligning [1877]. Alignment [1547, 607, 606]. All-Pairs [1438]. all different} [477]. Alliant [75]. Allocation [1755, 1878, 819, 1734, 1203, 2036, 486, 950, 1435, 1457, 1577, 1223, 1517, 260, 156, 1643, 1370, 1695, 1842, 1458, 1576, 1459, 1806, 372, 777, 1003, 716, 1249, 1248, 760, 1176, 1236]. allocations [1256]. Allowing [464]. almost [330]. ALPO [161]. alternate [1276]. Alternating [1791, 2068, 2038, 158, 1333, 1515, 1937]. Alternative [122, 1543, 2031, 720, 130, 289]. Alternatives [1898]. Ambiguity [1716, 1902, 1655, 1882]. Ambiguous [1399, 1083]. ambulance [904, 1235]. Among [1024, 1088, 1722]. Analysis [940, 1046, 1957, 1360, 67, 392, 1530, 154, 378, 117, 443, 955, 473, 149, 462, 1802, 250, 986, 2026, 98, 486, 280, 1426, 34, 1229, 1841, 2053, 2098, 1779, 2058, 241, 1429, 12, 229, 282, 640, 1464, 542, 493, 956, 292, 413, 1826, 752, 1089, 1711, 914, 718, 1272, 894, 678, 485, 335, 329, 432, 580, 600, 722, 533, 483, 1206, 804, 1031, 922, 1152]. Analytic [1618, 1676, 633]. analytic-center [633]. Analytical [799, 1279]. Analytics [1990, 1329, 1491]. Analyzing [256, 410, 944, 702, 1641]. Anchoring [220].

Ancillary [1524]. Angle [860]. Annealed [291]. Annealing [1326, 866, 96, 46, 87, 132, 407, 714]. Annotated [4]. annotation [1127]. Anomalies [1767]. Anomalous [24]. anomaly [960]. Anonymous [918, 540]. ant [460]. antipiracy [1269]. Anything [368]. AO [1216]. Aperture [1732, 1021]. APMOD [433]. Application [1769, 1476, 1394, 2037, 942, 1388, 163, 1861, 1717, 1541, 1336, 291, 229, 1219, 1606, 1947, 1932, 376, 620, 838, 380, 1055, 1023, 326, 930, 1162, 455, 367, 968, 583, 809, 555, 1236, 1265]. Applications [1573, 1461, 332, 394, 1639, 1734, 53, 1902, 1313, 849, 244, 1761, 112, 1798, 116, 1655, 1908, 960, 458, 661, 1105, 1502]. Applied [87, 31]. Applying [179, 329, 483]. Appointment [1782, 1325, 1872, 1988, 1042]. Appointment-Generated [1325]. Appreciation [469, 561, 655, 703, 767, 825, 877, 932, 982, 1032, 1094, 1155, 1208, 1263, 1312, 1364, 1416, 1493, 1607, 1812, 1912, 2028, 321, 404, 529, 610]. Approach [1409, 1369, 1640, 1475, 833, 936, 1879, 1896, 1709, 1168, 1516, 1617, 1484, 1437, 1140, 1664, 917, 1856, 35, 1401, 338, 238, 1291, 1944, 682, 501, 96, 107, 1883, 1537, 1518, 451, 1782, 1367, 1586, 1781, 1552, 893, 2077, 993, 2054, 1229, 2044, 1868, 1950, 1694, 1137, 551, 257, 312, 1916, 1872, 354, 1385, 521, 2041, 958, 1830, 188, 210, 1977, 264, 1722, 2014, 1345, 292, 1342, 1644, 1860, 1827, 568, 1811, 1763, 121, 1641, 1653, 1976, 370, 1822, 1867, 1463, 1559, 1089, 1087, 814, 725, 669, 708, 325, 1196, 322, 364, 1124, 1040, 1262, 602, 633, 1239, 375, 670, 1241]. approach [437, 532, 1205, 966, 839, 533, 1174, 1015, 1110, 796, 1001, 665, 548, 1254, 674, 324, 1180, 922, 869, 1103, 1160, 970, 578, 1010, 1226, 1265, 1141]. Approaches [1662, 1485, 1590, 1079, 278, 130, 1349, 1203, 1902, 91, 1600, 1838, 289, 1164, 1375, 1899, 1629, 1702, 1449, 1671, 1166, 1742, 2075, 1120, 1300, 778, 907, 1052, 435, 1289, 384, 800]. approximability [482]. Approximate [392, 1765, 2015, 1504, 2026, 1966, 904, 115, 881, 1320, 1976, 1227, 1191, 436, 883, 1194, 1103]. approximately [589]. Approximating [376, 1093, 1698, 313, 1453, 1060, 734, 971, 978]. Approximation [1475, 180, 686, 1143, 1664, 1487, 466, 813, 1492, 2094, 2090, 1628, 1584, 1787, 1891, 1075, 1920, 2053, 1231, 1908, 1550, 1179, 1754, 1613, 8, 66, 1370, 284, 1772, 1943, 1328, 2081, 926, 1275, 607, 1302, 1115, 622, 1097, 1122, 726, 467, 599, 1311, 1132, 626, 969, 1279, 820, 511, 465, 423, 445, 1197, 1238]. approximation-based [926]. Approximations [1716, 249, 2040, 1306, 1858, 1083, 1054, 1281, 1044, 1277, 883, 1177, 662]. Arbitrary [362, 835]. Arborescence [190]. Arc [1149, 908, 1336]. Arc-Routing [1336]. Arcflow [1846]. Architecture [179, 2045]. arcs [906]. Area [1357, 96, 2085]. Arising [6, 1873, 1885]. Armed [1621]. Arrangement [937]. Array [686]. Arrests [1611]. Arrival [1325, 1386, 875, 1130, 660]. arrival-departure [1130]. Arrivals [1396, 678, 485, 1153]. Art [331, 332, 159, 637, 191, 182, 868]. ARTAFIT [810]. Article [148, 393, 513, 168, 215, 191, 279, 881, 342]. articles [383]. Artificial [1648]. Ascend [378]. Ascending [799]. Ascent [1951, 1339, 437, 366]. Ascent-Based [1339]. Ash [1552]. Asian [1354].

Assemble [209, 1192]. assemble-to-order [1192]. Assembly [822, 1373, 2092, 504, 354, 353, 568, 2075, 1630, 390, 1058, 1014, 413]. Assembly-line [504]. Assembly-System [568]. Assessing [1277, 1041]. assessment [885, 729]. Assets [1380]. Assignment [745, 766, 176, 1534, 1412, 98, 908, 1332, 31, 1464, 23, 1564, 635, 938, 696, 552, 1118, 421, 1249, 1248, 1305, 999, 356, 1081, 436, 548, 804, 455, 1116, 677, 922, 583, 578, 1148, 1181]. Assignments [163, 690, 128]. Assisted [1841, 1637, 1837, 2095]. Associated [1877]. Association [40, 1125, 2042, 773]. Assortment [1825, 1941, 1471, 829, 1196]. Asymmetric [1852, 557]. Asymmetry [1435]. Asymptotic [255, 989, 1227]. Asymptotically [1367, 1898, 829]. Asynchronous [176, 760]. ATM [476]. Attack [1747, 993, 1881]. Attack-Based [1881]. Attacker [1601]. Attribute [399]. Auction [1852, 75, 910, 981]. Auctions [799, 2006, 1790, 550, 909, 952, 653]. Audio [685]. auditing [1157]. Augmented [1568, 278, 1792, 1582, 669, 1128, 729, 649]. Augmenting [1802]. Author [29, 470, 562, 611, 656, 704, 768, 826, 878, 933, 983]. Auto [1369]. Autocorrelated [100]. automata [1040, 594, 791]. Automated [761, 1687, 431, 892, 1251, 927, 735, 952, 660]. Automatic [1186, 338, 337, 844]. Automatically [1558]. Automating [985]. Automation [1923, 1335]. Automaton [90]. Autonomous [2062, 2095]. Autonomy [1337]. Autoregression [956]. Autoregression-Adjusted [956]. Autoregressive [368]. Autoregressive-to-Anything [368]. Availability [1620, 1924, 954, 1091]. Average [175, 2089, 1908, 510]. Averaging [2059, 1544]. Averse [1348, 1319, 1086, 1717, 1561, 1828, 1245, 2086, 1217, 1284, 1236]. Aversion [1692]. avoiding [1018]. Aware [1639, 2022, 1603]. Awareness [1709, 2002].

Baby [149]. Bachelor [273]. Backbone [52, 365, 198]. Backbones [1479]. Backdoor [1085]. Backlog [2070]. backtracking [526]. Backup [1806]. backward [1271]. Bad [327]. Balance [1562, 1132]. Balanced [843, 1204, 163, 1397, 967, 1004]. Balancing [209, 269, 822, 1614, 2092, 1985, 353, 41, 2075, 1630, 1003, 1014]. Bandit [1468]. Bandits [1409, 953, 793]. Bandwidth [1161, 643, 1076]. Banner [573]. Bar [1846]. Bargaining [1761]. Barrier [155, 1550, 1328]. Base [1556, 3, 690, 1071, 322]. Based [1216, 222, 1662, 1957, 931, 1360, 1543, 2078, 1804, 1897, 1926, 1485, 1437, 1895, 2005, 1856, 1412, 2038, 238, 1713, 1758, 942, 1291, 1357, 682, 2059, 1413, 1614, 2002, 1518, 1600, 1492, 1787, 944, 701, 43, 85, 431, 2013, 1937, 1866, 2077, 2054, 1339, 496, 863, 1673, 1989, 2001, 1558, 1788, 700, 1486, 1741, 1759, 236, 1372, 295, 1335, 1872, 1958, 1526, 101, 1691, 1542, 1636, 221, 1411, 1656, 210, 17, 264, 1722, 2014, 544, 186, 1945, 688, 1842, 23, 1660, 1870, 1604, 1881, 1858, 1057, 1697, 1725, 2010, 1323, 1463, 1630, 1582, 1317, 166, 926, 1120, 1121, 1275, 607, 928, 1124]. based [824, 1139, 1060, 960, 1133, 1280, 623, 778, 774, 1101, 1067, 459, 801, 620, 838, 1239, 375, 1011, 532, 1248, 999, 848, 794, 442, 653, 418, 722, 487, 1110, 630, 665, 674, 1502, 367, 452, 990, 1194, 626, 663, 333, 1010, 1459, 864]. Bases [1402, 212, 70, 264, 59, 382]. Basic [1979, 1819, 250]. Basis [129, 167, 457, 412, 753, 859]. Batch [1396, 351, 252, 1702, 493, 956, 678, 485]. batched [481]. Batches [1383, 479]. batching [858]. batteries [1237]. battery [1191]. Bayesian [1713, 1518, 1868, 188, 1942, 1860, 1820]. Beam [833, 822, 860, 1999]. Beam-ACO [822]. Beam-Angle [860]. Beds [1842]. Beekeeping [1522]. Behavior [228, 1482, 1868, 407]. Behaviors [1707, 1558, 1705]. Behaviour [1384]. Belief [1362]. beliefs [873]. Bellman [757]. Benchmark [535, 1336, 1834]. Benchmarking [1759, 1320]. Benders [1316, 2087, 1268, 1035, 1715, 1563, 1856, 1758, 1011, 1138, 1600, 2092, 1673, 1741, 1220, 729, 853, 1780, 922, 1630]. Bernoulli [1309]. Best [1543, 866, 1791, 1352, 1537, 1178, 344, 1822, 716]. Best-Fit [866]. Between [2036, 245, 1833, 571, 263, 1499, 1021, 383]. Bi [1413, 1164, 1372, 724, 1414]. Bi-Objective [1164, 1372, 1413, 724, 1414]. Bias [1751]. biased [1105]. Biases [541]. Bibliography [4]. Bicriteria [924, 1044]. bicriterion [804]. bid [909, 820]. Bidder [799, 2006]. bidders [981]. Bidding [1369, 952, 1191]. Bidirectional [872, 1668, 353]. Bids [1790]. Bifidelity [1793]. bifurcation [939, 919]. big [1261]. big-bucket [1261]. Bike [2097]. Bilevel [1715, 1232, 1670, 1933, 1902, 2027, 1781, 1515, 1989, 1490, 1605, 980]. BilevelJuMP.jl [1933]. Bilinear [1965, 1699, 1464, 1757]. Bimatrix [1540, 1991]. Bin [1922, 1563, 1424, 549, 1745, 1674, 1051, 1606, 1447, 1454, 801, 963, 435, 665, 913, 710]. Bin-Packing [1922, 549, 1674, 1447, 665, 710]. Binarized [895]. Binary [1662, 1957, 2004, 1113, 1527, 1048, 1877, 1664, 2023, 1600, 1794, 1421, 1592, 965, 744, 1799, 1144, 1763, 1057, 1964, 1212, 997, 1081, 1214]. Binary-Encoded [744]. Biobjective [1334, 1665, 1415, 1195, 1204, 1487, 1421, 1419, 966, 809]. Biochips [972]. bioinformatics [596]. biology [596, 586]. biology/bioinformatics [596]. Biomanufacturing [1820]. Biomedical [2091]. Bipartite [944, 788, 1723, 1422]. Bipath [2010]. Birth [946, 1712, 987]. birth-death [987]. Bisection [144, 456]. Bit [2055]. Bivariate [1525, 985, 1005, 889]. Black [1407, 1881]. Black-Box [1407, 1881]. Blending [1735]. Block [1343, 1333, 1089, 600, 1084]. blocking [465]. Board [2047, 2063, 2083, 2100, 675, 613, 615, 628, 645, 654, 659, 1209, 1222, 1431, 1452, 1472, 1494, 1528, 1554, 1580, 1608, 1813, 1650, 1688, 1719, 1752, 1783, 1831, 1849, 1862, 1875, 1893, 1913, 1930, 1954, 1968, 1983, 1995, 2016, 2029]. Body [639]. Book [72, 37, 93, 83, 63, 1077, 1169, 113]. booking [1084]. Boolean [2021, 1788, 510, 1422]. Boosted [1571]. Boosting [2072]. Bootstrap [1895, 1977]. bootstrapping [661]. Borer [1552]. Both [1558, 1283, 241]. Bottleneck [314, 285, 1177]. Bound [1665, 936, 1574, 246, 1735, 1815, 2005, 1324, 323, 1929, 2015, 1413, 451, 1666, 318, 1465, 1747, 1339, 1322, 189, 1414, 1385, 1526, 1799, 353, 5, 1398, 1576, 1570, 1594, 962, 387, 1292, 666, 1260, 1007, 1026, 999, 997, 324, 367, 402, 1014, 809, 807]. Bounded [1886, 478, 2011, 446, 1166, 1230, 698]. Bounded-Diameter [1166]. Bounding [1442, 2004, 97, 700, 1466, 1161, 1234]. Bounds [21, 146, 139, 1427, 1418, 1113, 937, 1484, 249, 1909, 861, 1668, 1354, 639, 60, 1582, 1064, 1243, 954, 1297, 377, 1179, 468, 1053]. Box [1407, 1881, 1204, 1587]. Boxed [1419]. Brachytherapy [1371]. Branch [1665, 728, 936, 1704, 1574, 386, 246, 133, 1815, 938, 547, 2005, 1324, 1456, 323, 1929, 1413, 1569, 451, 1666, 1819, 1055, 77, 318, 2021, 1679, 1363, 1800, 301, 1981, 1339, 1817, 1532, 1699, 1037, 1509, 1423, 1414, 354, 1385, 939, 1233, 1526, 1380, 1799, 353, 5, 1510, 1770, 1447, 568, 1742, 1454, 1570, 1697, 2010, 2095, 691, 1594, 962, 1102, 387, 1292, 489, 666, 1260, 1239, 1007, 963, 1293, 374, 1138, 1026, 647, 648, 1023, 1015, 1078, 1255, 803, 1214, 324, 367, 1051, 402, 1014, 809, 807, 864]. Branch-and-Bound [1665, 936, 1574, 1815, 2005, 1324, 1929, 451, 1666, 318, 1339, 1414, 1385, 1526, 1799, 353, 5, 1570, 1594, 962, 387, 1292, 666, 1260, 1007, 1026, 324, 809, 807]. Branch-and-Cut [728, 1704, 1413, 77, 1699, 1037, 1510, 1770, 568, 691, 1102, 374, 1138, 1023, 1015]. Branch-and-Price [1569, 1363, 1800, 1981, 1509, 1742, 1697, 2095, 1055, 1233, 489, 963, 1255, 1214, 1051]. Branch-and-Price-and-Cut [1532, 1447, 2010, 1239]. Branch-Cut-and-Price [1817, 1423]. Branch-Decomposition [547]. branch-decompositions [1078]. Branch-Price-and-Cut [1456, 1819, 1679]. Branching [1418, 1085, 1883, 1363, 1329, 1588, 1275, 1142, 853]. Brave [637]. Breaches [780]. Breaking [1550, 1328, 837]. Brick [1603]. Brick-and-Mortar [1603]. Bridges [365]. Bridging [571]. Broad [1460]. Broadband [1154]. broadcast [803]. broadcasting [1015]. Browser [702]. Browsing [1558]. bsnsing [1788]. Bucket [1553, 1261, 1211, 994]. Budget [695, 819, 126, 1468, 1789, 1643, 1172, 1176]. Budget-constrained [695]. Budget-Limited [1468]. buffer [678]. Building [1703, 1594, 532]. Bulk [131, 1300, 1153]. Bulk-Service [131]. Bullet [172]. Bundle [893, 1737, 1283, 437]. Business [1990, 543, 1502]. busy [718]. buy [1300]. buy-at-bulk [1300].

C [1005, 1379, 1786, 1059, 1090]. C-NORTA [1005]. C-VNS [1379]. Ca [121]. Ca/M/s/m [121]. Cable [1532]. Cable-Routing [1532]. Caching [2005, 680, 682, 702]. Calculating [508]. Calculation [1428]. Calibrating [1751]. Calibration [1518, 1923]. Call [1864]. Can [29]. Canadian [1216]. Cancellation [92, 559]. Cancer [845]. Candidate [1837]. Cannot [2078]. Cannot-Link [2078]. Capable [746]. Capacitated [305, 142, 303, 1448, 34, 1974, 1731, 236, 519, 240, 1329, 1634, 2095, 695, 1101, 938, 459, 1241, 1297, 689, 1026, 907, 1199, 699, 792, 1159, 1185]. Capacities [265, 708, 1288]. Capacity [1599, 1726, 1679, 486, 908, 1975, 26, 1370, 2086, 1826, 1619, 349, 622, 905, 1011, 417]. capacity- [1011]. Capital [1928]. Capture [1707]. Card [1881]. Cardiac [1611]. Cardinality [1791, 1953, 870, 1141]. Cardinality-Constrained [1791]. cardioverter [1135]. Care [1683, 1622]. Carl [461]. Carlo [1122, 1151, 1714, 1928, 2065, 1180, 865]. Carrier [238]. Cartesian [1276].

Cascade [1596, 473]. Case [1525, 160, 1747, 431, 2089, 34, 1514, 1332, 1872, 1545, 290, 25, 110, 240, 1588, 824, 1188, 850]. Cases [35]. Catalog [307, 1771, 754]. Causal [1794, 1708, 1132]. Cause [2044]. CD [251]. CDMA [690]. cellular [1276]. censored [1105]. Center [1676, 143, 1448, 1903, 633, 572, 874]. Centers [16, 1302]. central [1254]. Centrality [1617, 1387]. certain [560]. Certificates [2008]. Chain [2074, 42, 65, 6, 245, 975, 1321, 602, 762, 1180, 578]. Chained [536]. Chains [145, 466, 178, 1313, 946, 248, 12, 134, 914, 594, 747]. Challenges [1815, 1785, 283, 424, 1054]. Chance [1568, 1154, 1399, 1549, 1178, 2040, 1566, 1908, 1177, 1144, 1606, 1083, 1454, 1703, 1903, 1281, 925]. Chance-Constrained [1568, 1154, 1399, 1549, 1566, 1908, 1144, 1606, 1454, 1703, 1903, 1177, 1281, 925]. Change [350]. Change-of-Variable [350]. changes [1170]. Changing [448]. Characteristic [1710]. Characteristics [1440, 75, 1823]. Characterization [1949]. Characterizing [1586]. Charge [1025, 908, 907, 1055, 1159]. Charging [2036, 1583, 1892, 2066]. Charts [1846]. Checking [1906]. checkpointing [1188]. Checks [111]. Chemoradiotherapy [1330]. Chemotherapy [1939]. Choice [1707, 1421, 1892, 2054, 2001, 1486, 1123]. Choice-Based [2054, 2001]. Choices [1690]. Cholesky [1200, 193]. Chordless [1729]. Chubanov [1119]. Circles [1866]. Circuit [2021, 128, 675]. Circuits [676, 2020, 674]. Cities [1552]. Class [1662, 1632, 1313, 275, 273, 100, 1956, 1460, 1695, 1867, 829, 372, 899, 422, 617, 492, 1126, 435, 1184, 980, 925]. Class-Imbalance [1867]. classes [475]. Classic [953]. Classical [1531, 2022]. Classification [1218, 471, 1538, 1134, 1203, 71, 958, 210, 806, 749, 961, 1092, 577]. Classifiers [1957, 428, 1127]. Classifying [1098]. Clause [3]. Client [1924]. climbing [650]. Clinic [1325]. Clinical [1475, 1680]. Clique [1465, 1322, 1466, 290, 1167, 698]. Cliques [1569]. Cloaking [1359]. clock [981]. clock-and-offer [981]. Close [1484, 1124, 1260]. Close-Enough [1484, 1124, 1260]. closed [1279, 465]. Closeness [1387, 1342]. Closest [1780, 605]. closest-string [605]. Closure [1670, 1093]. closures [1261]. Cloud [1202, 1080, 1228]. Cluster [671, 259, 1639, 683, 854, 1398, 960, 383]. Cluster-Aware [1639]. cluster-based [960]. Clustering [1952, 2078, 207, 1137, 1739, 544, 2069, 1066, 364, 324, 1106]. Clusters [1412, 601, 1803, 2079]. clusterwise [1282]. CM [158, 272]. CM-2 [158, 272]. CNF [263]. Co [1802, 1524, 1727, 539]. Co-Editors [539]. Co-Movement [1727]. Co-Optimizing [1524]. Coclustering [1706]. Code [213]. Codesign [673]. coding [967]. Coefficients [1507]. Cognitive [1482]. cognitively [1193]. Coherent [2061]. Cohort [1802]. Collaboration [1291]. Collaborative [944, 2092, 1298, 1139]. Collecting [1915, 1324, 1164, 1339, 1663]. Collection [918, 1823, 593, 538]. Collimators [845]. colony [460]. Color [967, 1674, 670]. Color-coding [967]. Colorable [1649]. Coloring [2031, 744, 312, 1018, 88, 1167, 1667, 844, 991, 796, 1142, 1214]. colors [1018]. Column [1066, 1158, 1678, 278, 942, 78, 1861, 312, 1335, 1982, 1748, 1963, 1667, 411, 1101, 1067, 670, 1221, 991, 534, 839, 1274, 818, 631, 450, 502]. Column-generation [1158, 670]. Column-Splitting [278]. Combating [1269, 930]. Combination [1939, 979, 1071, 929, 1335, 508]. Combinatorial [209, 1500, 139, 1743, 1341, 298, 1563, 1991, 713, 160, 2092, 1966, 1541, 1391, 1633, 1576, 550, 824, 385, 909, 586, 1258, 653, 406]. Combine [1917]. Combined [745, 587, 1647, 1958, 1017]. Combining [1068, 935, 2072, 907, 157, 1671, 502]. Commentary [169, 343, 170, 216, 192, 217, 394, 395, 396, 218, 149, 219, 344, 280, 345, 171, 150, 151, 152, 281, 172, 346, 883, 193, 282, 194, 882, 173, 195]. Commerce [398, 624]. Commit [1483]. Commit-to-Delivery [1483]. Commitment [1361, 1474, 1567, 1810, 1982, 1548, 1303]. Commodity [1818, 1196]. Common [189, 502]. Commonsense [1229]. Communication [1381, 122, 1480, 323, 849, 730, 634, 871]. Communication-Constrained [1480]. Community [398]. Compact [1936, 1908]. Companies [307]. Comparing [1839]. Comparison [187, 252, 289, 87, 86, 1377, 1672, 726, 511]. Comparisons [2035, 245, 44]. compensation [681]. Competition [892]. Competitive [1140, 2054, 636]. Complement [201]. Complementarity [713, 1799, 1266, 847, 494]. Complementation [248]. Complete [1676, 190]. Completely [248]. Completion [1879, 311]. Completion-Time [311]. Complex [1457, 660]. Complexity [430, 458, 197, 414, 1838, 1514, 211, 1132, 370, 599, 1278, 1104, 510, 577]. Compliance [1073, 1235]. Component [863, 2053, 1801]. Component-Based [863]. Components [67, 746, 1457]. Composable [1948]. Composite [428, 294]. Composition [1788]. compressed [714]. compressed-annealing [714]. Compressor [1401, 1368]. Comput [498, 483]. Computable [1354]. Computation [305, 332, 286, 2071, 683, 520, 687, 2015, 178, 244, 128, 737, 1243, 1091, 844, 401, 987, 1109, 1270, 816, 1235]. Computational [205, 310, 728, 79, 996, 148, 1590, 1516, 1348, 443, 1791, 1147, 720, 38, 233, 462, 1838, 1785, 1794, 600, 1164, 44, 191, 729, 948, 86, 186, 639, 284, 857, 121, 2099, 475, 1272, 981, 967, 885, 664, 596, 586, 1054, 1281, 419, 964, 793, 1104, 1181]. Computationally [1716, 1407, 1837, 1299]. Computations [940, 160, 1059, 32, 750]. Compute [1918, 1936, 1484, 1696, 1911, 1802, 1490, 579, 820]. Computer [671, 1498, 44, 1751, 198, 577, 675]. Computer-aided [577, 675]. Computers [268]. Computing [438, 995, 821, 790, 779, 337, 1404, 1553, 131, 1396, 1840, 2057, 1851, 2018, 416, 740, 1740, 1153, 1598, 149, 2008, 245, 1785, 1202, 49, 89, 1515, 861, 1256, 1173, 2085, 953, 234, 948, 59, 860, 1327, 2010, 1315, 1645, 1246, 954, 1080, 377, 786, 566, 674, 1864, 1610, 1851, 1340]. Concave [1146, 2052]. concavity [885]. Concurrency [260]. Concurrent [8, 504]. condition [777]. Conditional [1768, 1122, 2080]. Conditionally [1481]. Conditioning [1714]. Conditions [1399, 732, 1985, 989, 559]. Cone [1769, 1623]. Cones [222]. Confidence [2078, 2080, 1100]. Confidence-Based [2078]. confidential [388]. Configuration [1741].

Conflict [1530, 1834, 1546, 1292, 1175]. conflict-directed [1175]. Conflict-Driven [1546]. Conflicts [1562, 963, 913, 1051]. conformal [741]. Conformation [609]. conformations [1145]. Congestion [1769, 156, 905, 636]. Conic [1773, 2038, 1612, 1064, 1283, 807]. conical [389]. Conjugate [115]. conjunctive [1183]. Connect [1999]. Connected [1168, 1660, 1327, 1328, 2081, 928, 1138, 966, 1052]. Connecting [1615]. Connection [158, 272, 41]. Connections [1499, 383]. Connectivity [314, 1347]. CONOPT [213]. Consensus [1937, 260, 589]. Consensus-Based [1937]. Consequences [1690]. Conservatism [1489]. considerations [817, 1249, 1248]. Considering [1825]. Consistency [221, 325, 1267, 1201]. Consistent [1986, 1226]. Constant [1526, 1772, 1065]. Constrained [1568, 122, 1527, 1480, 1154, 298, 1791, 1399, 1549, 199, 135, 2025, 159, 1795, 1178, 1750, 2032, 1535, 2040, 1566, 1937, 1941, 1638, 155, 120, 1810, 1536, 1908, 285, 1882, 1144, 1606, 1548, 1953, 1117, 1454, 1703, 1903, 1513, 691, 695, 1035, 1133, 623, 1065, 620, 1011, 789, 1122, 1072, 588, 1281, 1052, 1274, 1116, 1303, 1177, 832, 925]. Constraint [1846, 80, 522, 1487, 620, 126, 523, 252, 1421, 1629, 856, 1638, 1871, 295, 1274, 990, 1998, 95, 525, 477, 2010, 2075, 1499, 1212, 935, 1008, 386, 894, 1501, 589, 385, 1267, 1239, 771, 524, 991, 847, 1001, 646, 527, 776, 710, 384, 677, 869, 383]. Constraint-based [990]. constraint-programming [646]. constraint-programming-based [1239]. Constraint-propagation-based [620]. Constraints [1425, 790, 314, 1682, 2078, 1461, 1232, 1735, 1073, 61, 1798, 116, 1694, 1407, 1562, 1748, 1576, 2007, 1083, 1570, 1697, 1792, 1623, 1797, 708, 607, 1064, 886, 335, 670, 1172, 377, 534, 692, 837, 1112, 997, 1081, 527, 1201, 923, 1308, 1141]. Constructing [1114, 2090, 752]. Construction [1086]. constructive [421]. Consumption [1401, 1116]. container [901, 805, 1104]. Content [685, 1558, 1304]. Context [625, 960, 1193]. Context-independent [625]. context-sensitive [1193]. context-tree [960]. Contextual [1857]. Contingent [1868]. continued [438]. Continuous [2012, 271, 1357, 1453, 2077, 1925, 26, 1577, 1379, 1323, 914, 996, 1268, 1115, 583, 1082]. continuous-time [914]. contrast [1110]. Contrasting [507]. Control [1900, 2037, 271, 2023, 232, 1552, 223, 2082, 260, 1320, 1873, 1703, 1198, 634, 380, 1189, 367, 990, 1163, 869, 1152]. Controllable [316, 1219]. Controlled [46, 223, 919]. Controlling [2058]. Convergence [108, 230, 1841, 493, 2009, 1136, 980]. Convergent [1639, 1798, 961]. Convex [1960, 1804, 1519, 1896, 1461, 331, 1253, 1715, 1840, 1441, 2008, 2090, 1444, 788, 1571, 120, 2052, 1044, 1736, 1430, 1576, 1870, 926, 1261, 1283, 1060, 1115, 1266, 847, 1136, 971, 1277, 1308, 978, 1160, 1224]. Convexification [1769, 1378]. Convexity [885]. convexized [1036]. Convolutions [1962]. Cooperation [1950]. cooperative [760, 1256]. Coordinate [1343]. Coordinating [2036]. Coordination [1682]. Coping [1344]. copula [835, 1114]. CORAL [1012]. Core [998, 1486, 874, 961, 1230]. Core-Based [1486]. Corner [466]. Cornuéjols [457]. Coronaviruses [597].

Correcting [1616]. Correctional [1618]. Correlate [1463]. Correlated [1405, 873, 716, 1250]. Correlation [1952, 835, 1114, 816]. correlations [495, 862]. Corrigendum [1553]. Cosine [2043]. Cost [1821, 2004, 2034, 1507, 1439, 1824, 1802, 1112, 1887, 1990, 1950, 429, 1509, 200, 24, 1280, 557, 437, 664, 1256, 1190, 583]. cost-decomposition [557]. Cost-Effective [1824, 1887, 1990]. Cost-Effectiveness [1802]. Cost-sensitive [1112]. Costs [1926, 1146, 1941, 1483, 906, 1199, 793]. Counterfactual [1993]. Counterpart [1888]. Counterparts [1573]. counting [1201, 1183]. Counts [310]. Coupled [1585, 1296]. Course [1459]. Covariance [1536, 277]. Covariates [1841, 1597]. Cover [414, 1965, 478, 1943, 1582, 1064, 401]. Coverage [1899, 1916, 1772, 1273, 695, 1304]. Covering [1485, 1022, 1916, 929, 1131, 839, 665, 1274, 1108]. Covering-Based [1485]. Covering-Type [929]. COVID [1844, 1777]. COVID-19 [1844, 1777]. Coxian [786]. Coxianization [945]. CP [492]. Crane [1945]. Crashing [451]. creating [382]. Credit [2098, 427, 1881, 1151]. Credit-Risk [427]. Crew [569, 818]. Criteria [299, 1539, 482, 712]. Criterion [1195, 1204, 1699, 1419, 315, 1763]. Critical [1691, 1387, 1917, 1220]. Critically [1054]. crooked [1147]. Cross [1825, 1726, 220, 1646, 1147, 742]. cross-entropy [742]. Cross-Fertilization [220]. Cross-Selling [1825]. Cross-Validation [1646]. Crossing [408]. crossover [1058]. CSMA [251]. CSMA/CD [251]. Cumulative [2068, 1906, 783]. Cunningham [1078]. Current [4]. Curves [823]. Customer [1707, 1578]. Customer-Driven [1578]. Customers [397]. Customized [1916]. Cut [728, 1704, 133, 1456, 1413, 1819, 77, 2021, 1592, 1679, 1361, 301, 1817, 1532, 1699, 830, 1037, 1423, 354, 1956, 1780, 1510, 1770, 1345, 1447, 568, 1098, 2010, 691, 1102, 1101, 633, 1239, 374, 1138, 1023, 1015, 1117, 864, 1036, 1352]. cut-and-column [1101]. Cut-Generating [1956]. Cuts [728, 1470, 1749, 876, 558, 1111, 1904, 1567, 1382, 1604, 1261, 977, 1268, 1147, 699, 631]. Cutting [1676, 936, 1756, 866, 1815, 1355, 535, 1424, 135, 303, 14, 893, 1861, 1776, 1224, 1582, 731, 821, 709, 1246, 973, 1039, 363, 439, 620, 633, 838, 1257, 1259, 498, 789, 566, 1296, 1159]. Cutting-Plane [1582, 633, 838, 1159]. Cutting-Planes [1776]. Cutting-Stock [535, 893, 731, 1296]. cutting/packing [1257]. CVaR [1252, 1230]. Cycle [478, 1592, 1729, 823, 1582, 1495, 648, 692]. Cycle-Cover [478]. cycles [1018]. Cyclic [228, 1964].

D [1556, 609]. daily [1170]. Dantzig [1074]. Darts [1764]. Data [10, 1016, 1864, 1030, 1957, 40, 259, 1756, 425, 1730, 755, 1825, 1895, 2070, 2073, 1538, 103, 1611, 341, 448, 1399, 679, 955, 272, 1291, 1944, 682, 212, 1775, 1854, 1572, 1325, 918, 1652, 1923, 1920, 128, 1539, 1990, 2097, 358, 1511, 2076, 1000, 1777, 544, 1366, 1393, 1342, 1860, 1430, 1083, 1491, 541, 752, 1594, 1196, 322, 424, 960, 952, 1092, 1502, 1106, 1010]. Data-Based [1895]. Data-Driven [1756, 1825, 2070, 1944, 1775, 1572, 1325, 1539, 2097, 1366, 1854, 952]. Data-Gathering [1730]. Data-Mining-Based [682]. Data-mining-driven [1000]. Data-Parallel [272]. Database [2034, 163, 262, 68, 260, 188, 388, 1157]. Databases [431]. Date [189, 502]. dates [330, 677, 1307, 1219]. Dawande [457]. DC [1951]. Deadlines [1324, 677, 1219]. Deadlock [338]. deal [1280]. Death [946, 1712, 987]. decade [406]. Decentralized [2007]. Decision [1384, 268, 1113, 1213, 1527, 1877, 1765, 1746, 58, 774, 2005, 2068, 2037, 1944, 1902, 1600, 2094, 1782, 1547, 2013, 496, 1788, 1733, 1868, 949, 211, 358, 1655, 184, 521, 1526, 1597, 1382, 1257, 532, 723, 1112, 795, 567, 1214, 883, 832, 618]. Decision-Dependent [1782, 1655]. Decision-tree-based [774]. Decisions [1845, 916, 427, 1366]. Decomposable [248, 1545, 2007]. decompose [634]. Decomposing [1569, 1585]. Decomposition [1316, 1662, 1568, 67, 1527, 2087, 425, 746, 1715, 1617, 2033, 1373, 547, 1563, 1856, 1319, 1549, 1758, 916, 51, 1507, 1441, 1600, 1775, 1726, 2092, 1673, 1506, 1629, 1741, 1332, 127, 1853, 1780, 1945, 1215, 1236, 1630, 1317, 1907, 1139, 1035, 557, 1293, 1011, 858, 1241, 1138, 997, 417, 1220, 1150, 729, 710, 455, 853, 1184, 1276, 663, 922, 864]. Decomposition-Based [1662, 1139, 864]. decompositions [647, 648, 1078, 1185]. Decorous [937]. Decreasing [1805]. Decremental [1979]. Dedicated [1981]. Deep [1949, 2070, 1658, 1869, 1642]. Deepening [1668]. Defender [1601]. Defender-Attacker [1601]. defense [1237]. defibrillator [1135]. Define [1998]. Defined [1960, 1836, 918]. deflected [818]. degenerate [977, 976]. Degraeve [498]. Degree [1617]. Delay [1988, 730, 1076]. Delayed [1570]. Deletion [1310]. Delineation [1371]. Deliveries [1818]. Delivery [845, 1437, 1891, 1483, 1395, 2095, 764, 1278, 1021]. delta [824]. deluge [1262]. Demand [940, 1755, 1878, 1709, 1825, 2067, 975, 1351, 519, 2041, 464, 1460, 1806, 1578, 777, 1273, 695, 623, 617, 1042, 2002, 1105, 1116, 990]. demand-constrained [623]. demand-responsive [617]. Demands [1843, 1309, 1250]. denial [1157]. Dense [272, 98, 2079]. Denser [743]. Density [1714, 1393]. Departments [1351]. Departure [1090, 1130]. Dependence [2026, 1575, 835, 1114]. Dependent [1046, 1565, 1675, 1782, 252, 1679, 1346, 1809, 1817, 1090, 1655, 1613, 132, 777, 998, 1112]. deployment [1104]. Depot [1891]. depots [1197]. Depth [1063, 361]. depth-first [361]. Derivable [263]. Derivation [1453]. Derivative [1445, 1986]. Derivatives [286]. Derive [1957]. Descent [1343, 1951, 2024, 2009]. Described [244]. Description [685, 797]. Design [1475, 1618, 1425, 122, 779, 2087, 746, 305, 490, 754, 199, 2060, 52, 473, 2002, 1766, 68, 74, 2027, 26, 1375, 1657, 1541, 1337, 910, 1223, 519, 101, 1542, 464, 186, 36, 1479, 1374, 1458, 568, 1621, 1585, 675, 1156, 668, 941, 1035, 459, 363, 905, 716, 1075, 1281, 630, 1190, 384, 792, 922, 869, 1159, 1148]. Designed [187]. Designing [634, 219, 637, 1300, 366, 1280]. Desktop [1336]. Detailed [1333]. Detecting [1767, 1722, 1098, 1642, 1917, 1110]. Detection [2005, 338, 1376, 1658, 1350, 993, 1946, 780, 2076, 1526, 95, 1881, 960, 885, 1244].

Determinants [277]. Determination [1643, 325, 604, 653]. Determine [1676]. Determined [102]. determining [716]. Deterministic [1854, 1374]. Deterring [1946]. Developing [343, 1227]. Development [57, 11, 2045]. Diagnosing [287, 91]. diagnosis [1310]. Diagnostic [1839]. Diagonal [2098]. Diagram [1527, 2005, 1600]. Diagram-Based [2005]. Diagrams [1113, 1213, 1877, 1765, 1746, 138, 1547, 2013, 1382, 348, 1214]. Dial [942, 1008, 577]. Dial-a-Flight [942]. dial-a-ride [1008, 577]. Diameter [1803, 1166]. Difference [1754]. differences [1110]. Different [1470]. Differentiable [1940, 1936, 1696]. Differentially [1468, 1467]. differentiation [337, 844]. Diffusion [1516, 1198]. Digital [972]. Digraphs [190]. Dimension [865, 1884]. Dimensional [154, 1563, 137, 549, 1397, 45, 881, 544, 1447, 1697, 1778, 2096, 1016, 731, 709, 1246, 641, 801, 1495, 1257, 1259, 560, 789, 435, 710, 971, 1043]. Dimensions [1344, 2082, 806, 1136]. Direct [426, 245, 1732, 1977, 1128, 1284, 1021]. Directed [87, 1175, 630]. Direction [2068, 2038, 1333, 1515, 1937]. Directions [174]. Directors [613, 615, 628, 645, 654, 659]. Dirichlet [1759]. Disaggregation [245]. discipline [726]. discontinuous [1225]. discounts [636]. Discovering [773]. Discovery [684, 1675, 428, 1010, 1045, 774, 959, 1127]. Discrepancy [570, 700]. Discrepancy-Based [700]. Discrete [169, 1769, 1845, 1939, 1442, 67, 1353, 1213, 1669, 1879, 1746, 2033, 369, 2057, 1456, 338, 168, 270, 1707, 1892, 1919, 2082, 90, 1560, 1873, 640, 1508, 173, 1885, 750, 675, 835, 1114, 485, 661, 1270, 1106, 650, 1043, 1036]. Discrete-State [369]. discrete-time [485]. Discretizable [1629]. Discretization [1675]. Discriminant [102, 292, 797, 533]. discrimination [1133]. Disease [1330]. Disentangled [1946]. Disjoint [139, 243, 1757]. Disjunctive [1353, 997, 1807, 1160, 1224]. Disk [814]. Dispatching [1504]. Dispersion [2077, 1711, 1201]. Disruption [2011]. Disruptions [1821, 1635, 1653, 1044]. Disruptive [1806]. Distance [362, 2098, 1629, 1691, 1011]. Distance-Based [1691]. distance-constrained [1011]. Distinguishing [1728]. Distortion [1598]. Distributed [394, 1538, 369, 1961, 1600, 128, 1613, 156, 1467, 941, 1301, 757, 1163, 413]. Distribution [227, 1836, 1480, 2039, 238, 403, 872, 1657, 1978, 1653, 954, 1091, 508, 661, 1153, 1075, 579, 717]. Distributionally [1716, 1734, 1631, 1944, 2032, 2040, 1888, 1655, 1882, 1828, 1458, 1963, 1356, 1703, 1903]. Distributions [239, 131, 157, 835, 1114, 1005, 1306, 1170, 988]. Districting [1149, 1731]. Diverse [1918, 505, 1304]. Diversion [1946]. DiversiTree [1918]. Diversity [1521, 1120, 1289]. Dividing [1024, 1088]. Diving [1380]. DNA [587]. DNF [263]. DNN [1738]. Do [29]. Document [702]. Document-Duplication [702]. Does [323, 944, 445]. Domain [1134, 2075, 386, 1193, 1201]. Domain-Independent [2075]. Domains [501]. Dominance [2005, 1819]. Dominating [1168, 1693, 1660, 1327, 1328, 2081, 1138]. Domination [1455]. domino [740]. domino-parity [740]. Don’t [1835]. Dose [1478, 1371, 1694, 1680, 860]. Dose-Finding [1680]. Dose-Response [1478]. Dose-Volume [1371]. Dosing [1939]. Dots [1999]. Double [1037, 210]. Double-Layered [210]. downside [1184]. Downstream [1842]. Downtime [1441]. DRG [1203]. Driven [1477, 1756, 1825, 2070, 1944, 1775, 1572, 1325, 1539, 2097, 1469, 1366, 1546, 1641, 1578, 1172, 1854, 952, 1000]. Drivers [1887]. Drone [1979, 1385, 1436]. drug [959]. Dual [1427, 928, 2004, 1418, 1890, 444, 1900, 165, 57, 238, 1505, 1738, 1221, 109, 1095, 1861, 266, 1339, 27, 9, 70, 265, 200, 1671, 1801, 1763, 1060, 437, 758, 630, 366, 800, 707, 697, 631]. dual-ascent [437]. Dual-based [928]. dual-filter [800]. Dual-Optimal [70]. Dual-Sourcing [1900]. dual-threshold [800]. Duality [1243, 1989, 1545, 166]. Duality-Based [166]. Due [189, 330, 1307, 502]. Duplication [702]. Duration [118, 1830]. During [1777]. Dynamic [607, 790, 1890, 394, 787, 1556, 1678, 1401, 1505, 942, 137, 916, 1042, 271, 1551, 1172, 1492, 1764, 1675, 1504, 3, 357, 1809, 496, 1368, 22, 1680, 45, 1223, 1872, 881, 1671, 1320, 156, 1643, 1104, 1801, 1727, 662, 1723, 1345, 1869, 1976, 1265, 2075, 1820, 763, 943, 475, 1008, 349, 361, 998, 1080, 858, 1107, 1288, 1191, 904, 852, 776, 883, 445, 1041, 1036]. Dynamic-programming [662, 475]. Dynamically [1668]. Dynamics [2051]. dynasearch [503].

e-business [1502]. e-commerce [624]. E-Logistics [1657]. Earliness [318, 330, 840, 1165]. Earliness-tardiness [330, 840]. Earliness/Tardiness [1165]. Early [666, 1203, 1733]. Easy [1865, 754]. Economic [52, 1928, 1731, 909]. Edge [1906, 1465, 783, 265, 1230, 1058]. Edge-Finding [1906]. editor [596, 125, 181, 224, 300, 553, 614, 627, 658, 671, 770, 897, 947, 1996, 683, 522, 736, 751, 811, 854, 841, 867, 880, 921, 1020, 147, 203, 1, 7, 13, 20, 28, 39, 48, 56, 64, 73, 84, 94, 105, 114, 1340, 441, 447, 454, 472, 484, 491, 499, 506, 512, 530, 545, 564, 574, 584, 693, 721, 706, 1365, 1389, 1417, 1432, 1473, 1503, 1529, 1555, 1581, 1609, 1651, 1689, 1720, 1753, 1784, 1814, 1832, 1850, 1863, 1876, 1894, 1914, 1931, 1955, 1969, 1984, 2017, 2030, 2048, 2064, 2084, 1061, 1033]. Editorial [2047, 2063, 2083, 2100, 1209, 1222, 1431, 1452, 1472, 1494, 1528, 1554, 1580, 1608, 1813, 1650, 1688, 1719, 1752, 1783, 1831, 1849, 1862, 1875, 1893, 1913, 1930, 1954, 1968, 1983, 1995, 2016, 2029]. Editors [613, 615, 628, 645, 654, 659, 136, 539]. Effect [1821, 2098, 385]. Effective [1618, 1824, 1887, 1990, 1514, 1127, 644, 1227, 818]. Effectiveness [1802]. Effects [2093, 702, 1708, 1269]. efficacy [919]. Efficiency [175, 1003, 1021, 919]. Efficient [1334, 737, 750, 1415, 835, 1272, 1724, 400, 1145, 819, 1252, 1716, 242, 293, 526, 1496, 1498, 1080, 142, 2060, 190, 844, 802, 2061, 2062, 987, 244, 2019, 1616, 966, 2077, 1635, 1798, 2065, 1378, 1517, 1873, 540, 1411, 11, 2043, 1706, 1953, 1686, 1089, 1964, 1087, 372, 458, 954, 1495, 417, 1096, 970]. Efficiently [1918, 1615, 2090, 429, 1792]. Effort [179]. eigenvalues [560]. Eigenvectors [676]. ejection [747, 578]. Elasticity [975]. Elections [1404]. Electric [1892, 2066, 1338]. Electrical [671, 2035]. Electricity [1718, 1523, 2045, 1191]. Electronic [398, 681]. eliminating [906]. elimination [692]. Eliminations [863]. Ellipsoid [1315]. Ellipsoidal [1466]. Embedded [1571, 334, 576]. Embedding [728, 1538, 1636, 1641, 1463]. Embedding-Based [1463]. Emerald [1552]. Emergency [1351]. Empirical [799, 153, 175, 117, 98, 486, 908, 604, 580]. Enabled [855, 1229, 1733]. Encoded [744]. encoding [871]. End [1658, 2045]. End-to-End [1658, 2045]. Endogenous [2067]. Energy [601, 1524, 1320, 1015, 1031]. Engagements [1857]. Engineering [671, 1391]. Engines [720]. Enhanced [1424, 2038, 1130, 1086, 1550, 464, 1708, 864, 1328, 806]. Enhancement [866, 971]. Enhancements [16, 1854, 1294]. Enhancing [1929, 707]. Enough [1484, 1124, 1260]. Enriching [1391]. Enrichment [900]. Ensemble [1798, 1071, 1859, 1947, 1867]. Ensembles [2055, 752]. Enterprise [1030]. Entity [1406]. Entropy [1840, 2019, 742]. Enumeration [1276, 1508, 712]. Envelopment [955, 752]. Environment [160, 673, 566, 333]. Environments [448]. EP [234]. Epidemics [1844]. Epsilon [1487]. Epsilon-Constraint [1487]. Equal [1866, 330, 1034]. Equalities [95]. Equality [1935]. Equations [6, 115, 1379, 1264, 621]. Equilibria [1540, 1852, 1936, 1696, 1991, 1889, 1911]. equilibrium [560]. Equipment [1777]. equity [1003]. Equivalent [965, 325]. Erlang [157]. Erratum [247]. Error [189, 1511, 1354, 2096, 1200]. Essential [1408]. estimates [666]. Estimating [227, 341, 1666, 1151, 1444, 228, 823]. Estimation [1640, 362, 1392, 824, 1825, 1895, 1441, 2025, 403, 1707, 1714, 2003, 1536, 1587, 2069, 843, 1038, 742]. Estimation-based [824]. Estimator [17]. Estimators [1445, 230, 481, 737, 1128, 497]. ETAQA [727]. Euclidean [293, 874, 373, 86, 284]. Evacuation [1504]. Evaluating [1477, 1839, 1405]. Evaluation [810, 1352, 720, 262, 228, 2098, 427, 542, 688, 1881, 381, 808]. evaluations [824, 978]. Event [169, 67, 1394, 1481, 162, 168, 1687, 1658, 1318, 616, 11, 173, 742, 559, 930]. Event-Graph [616]. Events [1586, 244, 1806]. Evidence [197]. Evolution [517, 1004, 1070]. Evolutionary [863, 521, 756, 1943, 723, 734, 924, 868, 980]. Evolving [516]. Exact [1216, 1415, 829, 1300, 1922, 1394, 896, 415, 1730, 1980, 1079, 1664, 2031, 2072, 1972, 1927, 528, 1249, 991, 1076, 1628, 2000, 1465, 1547, 1592, 1679, 2089, 1164, 1375, 1941, 1818, 1920, 2053, 1779, 1522, 436, 1486, 551, 1199, 236, 698, 929, 2091, 1577, 1729, 1702, 1739, 2088, 1663, 88, 1796, 1992, 1356, 1593, 1101, 798, 1003, 375, 870, 1257, 1182, 907, 599, 1012, 1047, 1294, 840, 1053]. Exactly [1762]. exam [1262]. examination [778]. Example [44]. Examples [1046, 263]. Exchange [1557, 1681]. Exciting [2069]. Execution [559]. Exhaustive [958, 726]. Existence [541, 445]. Expanded [942, 2091]. Expanding [1628]. Expansion [1253, 1480, 1726, 1338]. Expectation [1792]. Expected [1836, 1525, 1059, 695, 888, 1235]. Expediting [29]. Expensive [1407, 1837, 1299, 753]. Experience [303, 186, 284, 857]. Experiment [2060]. Experimental [228, 279, 281, 290, 1464, 511, 407, 726, 1206, 804, 626]. Experimenting [214]. Experiments [227, 148, 977, 1615, 1847, 107, 637, 1362, 948, 134, 1147, 959]. Expert [148, 232, 429, 1280]. ExpertRNA [1760]. explanation [367]. explanation-based [367]. Explanations [1993]. exploitative [843]. Exploiting [1212, 1961, 1512, 212, 109, 588, 795, 412]. explorative [843]. Exponential [1512, 349]. exposition [442]. exposure [1013]. Expressing [253]. Expressions [1904]. Extended [1692, 1345, 1206]. Extending [524, 1624, 836]. Extension [1957, 1064]. Extensions [2039, 1410]. extra [631]. extracting [1304]. Extraction [845, 1860]. Extremal [1711]. Extreme [801, 1567, 911].

face [981]. Facetial [1592]. Facilities [1618, 1088, 784]. Facility [308, 1902, 2054, 1635, 2001, 1779, 1541, 1509, 86, 1947, 1087, 1309, 668, 928, 1026, 1288, 606, 1052, 1311, 834, 969, 1148]. facility-layout [668]. facility-location [606]. Fact [169]. Factor [1759, 1411, 919]. Factorization [1709, 2073, 1376, 1927, 1402, 202, 576]. Factorizations [59, 1200]. Failure [1994]. Failures [1880, 1230]. Fair [480]. Fairness [1985]. Family [1981]. Fare [1370]. Fare-Locking [1370]. Fast [140, 1302, 488, 207, 1441, 1925, 123, 294, 1660, 1238, 1026, 1270, 762, 889, 2003]. fast-pivoting [762]. faster [1243, 793]. Fault [476, 1168, 1803, 1328]. Fault-Tolerant [476, 1168, 1803, 1328]. FBP [701]. FDDI [365]. Feasibility [894, 1740, 130, 2015, 1567, 1626, 1372, 1643]. Feasible [488, 1515, 1797, 589]. Feature [148, 393, 379, 513, 168, 215, 2098, 191, 279, 642, 427, 881, 342, 1430, 2096, 1258]. Features [1761]. Federated [1924]. Feedforward [185]. Feel [193]. Fertilization [220]. Few [1743, 2055]. Few-Bit [2055]. Fiber [199]. FiberSCIP [1314]. Fiction [169]. Fidelity [1622]. Field [168]. Fields [1560]. FilMINT [926]. filter [852, 800]. Filtered [570]. Filtering [944, 1421, 1910]. Filtration [846]. Filtration-Oriented [846]. Financial [1229, 427]. Finding [335, 1002, 1906, 1993, 70, 783, 1680, 200, 1387, 1660, 1520, 1797, 589, 560, 911]. Fine [608, 1759]. Fine-Grained [1759]. Finite [299, 1960, 673, 953, 1754, 1643, 1028, 386, 349, 678, 622, 389, 407, 1072, 567, 373]. finite-buffer [678]. Finite-Horizon [953, 567]. Finite-Resource [673]. Finite-sample [1028]. finite-time [407]. finitely [667]. Firm [1949, 1641]. First [1516, 2032, 2027, 1990, 2052, 1428, 1677, 361, 961]. First-Order [2032, 2052, 1428, 1677, 961]. First-Party [1990]. Fit [866]. Fitter [343]. Fitting [368, 1090, 1511, 1451, 1672, 835, 810, 1170, 988]. Fixations [1393]. Fixed [746, 1154, 1025, 1789, 908, 1583, 5, 1317, 695, 907, 1055, 1197, 1159]. Fixed-Budget [1789]. Fixed-Charge [1025, 908, 907, 1055, 1159]. fixed-wireless [695]. Fixing [2040]. Fleet [745, 1162, 922]. flex [1262]. flex-deluge [1262]. Flexible [1425, 227, 1768, 1940, 1679, 1073, 474, 1842, 2043, 1116]. Flight [942, 1370].

Floating [1349, 1212]. floating-point [1212]. Flood [2035]. Flow [1845, 1821, 160, 313, 2021, 908, 2013, 1908, 8, 639, 1653, 316, 1013, 557, 437, 664, 479, 907, 1055, 534, 420, 334, 1034, 939, 724, 662, 507, 1148]. Flow-Based [2013]. flow-shop [507]. Flows [1143, 872, 127, 304, 2010, 1500, 1118]. Flowshop [1762, 808, 712]. Flowtime [241]. Fluid [1099]. fluxes [1276]. Fold [1327]. folksonomies [1193]. Followee [1857]. Follower [1967, 1715]. Following [1936, 1696]. Football [861]. Forcing [1590]. Ford [757]. Forecasting [1709, 2073, 1844, 185, 2045, 1999]. Forests [1397]. fork [1279]. fork/join [1279]. Forks [249]. form [481, 1183]. Formal [378, 11]. Formalism [90]. formalization [559]. Formation [1574, 2071, 1070]. Formulas [1769, 1183]. Formulation [940, 1425, 1553, 1936, 103, 1692, 1349, 974, 2013, 744, 1345, 2007, 1211, 994, 572, 847, 697, 445]. Formulations [1102, 1316, 1846, 1926, 1485, 1771, 1773, 1424, 2002, 1904, 1474, 1673, 1899, 1311, 1985, 195, 1595, 782, 424, 1106, 969, 1123, 450]. Fortification [1596, 1686, 1271]. forward [1080, 1264]. Founding [1340]. FPTAS [445]. fractal [373]. Fractional [1537]. Fractionation [1684, 1207, 1290]. fractions [438]. Fragmentation [1674]. frame [389, 667]. Framework [1478, 728, 1618, 1030, 2004, 1704, 1659, 1487, 1569, 1698, 1658, 1794, 1566, 1339, 2001, 1760, 1733, 2097, 1994, 1411, 542, 1881, 2099, 2045, 694, 668, 386, 1050, 1158, 794, 741, 1271, 1206, 809, 738, 465, 857]. Frameworks [1846, 1639, 1767]. Frank [1768]. FrankWolfe.jl [1768]. Fraud [2076, 1881, 1013]. Free [2054, 1200, 732]. Freedom [918]. fresh [777]. freshness [777]. freshness-condition-dependent [777]. Frontier [701]. Frontier-Based [701]. frvcpy [1583]. Fuel [1810]. Fuel-Constrained [1810]. Fully [2094, 941, 445]. Function [1415, 362, 1804, 1405, 1444, 1451, 132, 1366, 1672, 604, 381, 753, 978]. Functional [116, 1264]. Functions [227, 1836, 15, 102, 1453, 1509, 1545, 1644, 2009, 797, 1065, 851, 753, 859, 1194, 978, 583, 510, 920]. Funds [1086]. Further [57]. Fusion [1920, 2044, 2099]. Future [216, 162, 517, 174]. Future-Event [162]. Fuzzy [35, 434]. FX [75]. FX/8 [75].

G [2, 946]. Gallery [331]. Galls [608]. Game [2057, 1764, 1137, 1237, 1108]. Game-Theoretic [1137]. Games [1540, 1852, 2071, 1664, 1696, 1991, 1911, 1388, 1770, 1256]. GAMS [855, 1506]. Gap [571]. Gas [1253, 1401, 1333, 1937, 1368, 1462, 979]. Gathering [1730, 1594]. Gauss [75]. Gaussian [1114, 1836, 1647, 1560]. Gaussian-copula [1114]. Gaze [1393]. Gb [131]. Gd [131]. Geelen [1078]. Gene [900]. General [201, 1704, 35, 713, 1698, 266, 1483, 1635, 1509, 595, 179, 1857, 465, 1327, 1686, 1943, 1309, 641, 1091, 1266, 326, 1206, 1177]. general-purpose [1091]. Generality [616]. Generalized [1481, 1171, 117, 167, 350, 1022, 580, 1378, 995, 828, 1064, 890, 1038, 696, 1249, 1248, 638, 773, 1288, 997, 747, 852, 1034, 548, 1282, 1194, 1160, 1224, 650, 578, 988]. Generally [1613]. Generate [122]. Generated [1325, 667]. Generating [570, 1749, 42, 65, 566, 1336, 429, 100, 1956]. Generation [1846, 425, 369, 1678, 942, 1344, 1524, 78, 1361, 1861, 312, 1335, 101, 640, 1982, 1748, 1345, 1963, 1523, 1667, 1338, 1066, 411, 495, 1158, 1101, 1067, 670, 1221, 991, 534, 839, 1274, 862, 889, 818, 1170, 864, 631, 450, 502]. Generator [1834, 1135]. Generators [1498, 802, 1426, 290, 1495, 1496, 1497, 1029, 945, 339, 1126, 592]. Generic [1363, 568, 1023, 1051]. Genetic [343, 209, 299, 208, 426, 298, 275, 781, 344, 102, 345, 99, 297, 342, 347, 346, 210, 348, 898, 552, 532, 567, 1058, 592]. genetic/optimization [567]. Geometric [140, 2093, 1378, 123, 970]. Geometry [1629]. Ghost [277]. GI [1959, 1959, 89, 289, 413]. GI-distributed [413]. GI/GI/1 [1959]. GI/PH/1 [89, 289]. Gittins [762]. given [1114]. Global [1845, 672, 887, 1589, 601, 51, 2015, 2077, 1647, 1953, 814, 1129, 527, 753, 859, 970, 738, 980]. Globalized [1285, 1888]. Globally [1420, 821, 749]. Gomory [876, 558]. Gondzio [483]. Good [1864, 1886, 346, 592]. GPU [1246, 1762]. GPU-Accelerated [1762]. Gradient [230, 1362, 115, 1571, 1754, 2009, 1858, 1867, 1323, 873, 1128, 959, 497]. Gradient-Based [1858, 1323]. Gradient-Boosted [1571]. Gradients [1768, 337, 1105]. Grained [1759]. Grammar [1067]. Grammar-based [1067]. Grammars [85]. Grand [1950]. granular [555]. Graph [133, 1692, 1785, 311, 43, 85, 1137, 312, 1466, 267, 200, 88, 616, 264, 2014, 370, 1292, 1039, 991, 456, 1052, 1142, 1214, 402, 1096]. Graph-Based [43, 85, 2014]. Graph-Grammars [85]. graph-theoretical [1039]. Graphical [177, 806]. Graphics [1412, 1305]. Graphs [1855, 359, 35, 473, 1569, 944, 788, 1803, 1406, 11, 1727, 1327, 1917, 1161, 643, 885, 559, 1018, 518]. GRASP [635, 731, 237, 408, 455, 518, 619]. Gray [1946]. Greedy [18, 1551]. Greeks [1225]. Green [1817]. Greenberg [1340]. GRG [213, 106]. GRG2 [186]. GRG2-Based [186]. Grid [779, 855, 1747, 948, 1320, 981, 858]. Grid-Enabled [855]. Grid-Level [1320]. grooming [759]. Group [1409, 1476, 1140, 521, 1945, 1935, 1430, 349, 1110]. Group-Based [1945]. Grouping [1549, 2098, 1469]. Groups [1822, 1158]. Growth [1330, 1196]. GTH [340]. Guarantee [730, 445]. Guaranteed [1943]. Guarantees [1398, 1874, 1188, 1076, 711]. Guest [671, 522, 539, 596, 683, 136]. Guide [170, 637]. Guided [746, 549, 1422, 1055, 784]. Guidelines [150]. Guillotine [936, 135, 1246, 1039, 1257, 1259]. guillotine-cutting [1039]. Gunbarrel [1401, 1462].

Hamiltonian [2024, 1255]. Hamiltonians [2023]. Hand [1670, 1605, 1703]. Handling [448, 1964]. Handover [1533]. Hankel [2073]. haplotype [600]. Haplotyping [599, 899, 839]. Hard [1762, 422, 850]. Hardware [673]. Hardware/Software [673]. Harmonic [230, 497]. Harris [461]. Harvey [1340]. Hashing [74]. Having [228]. Hazards [1778]. HDR [1371]. HDR-Brachytherapy [1371]. Head [639]. Head-Body-Tail [639]. Health [1476, 1618]. heavy [988]. heavy-tailed [988]. Hedging [1986, 1080, 1205]. Height [1186]. Helpfulness [1518]. Helping [218]. Hessians [844]. Heterogeneity [1988]. Heterogeneous [1620, 1755, 1878, 1767, 2076, 1460]. Heuristic [201, 1309, 992, 845, 1957, 887, 1485, 866, 1484, 755, 1980, 798, 1182, 135, 638, 357, 2089, 863, 189, 435, 1974, 1372, 595, 294, 1166, 276, 1057, 1932, 708, 890, 1139, 623, 696, 905, 1107, 907, 653, 418, 747, 487, 1199, 665, 548, 714, 1226]. Heuristics [1590, 488, 443, 328, 1352, 255, 157, 1904, 1925, 34, 429, 62, 123, 1380, 1853, 542, 1633, 1546, 617, 801, 391, 453, 848, 1288, 373, 969, 925]. Hidden [2051, 1407]. Hierarchical [197, 1095, 90, 1677, 366, 1162]. hierarchy [1003]. High [1768, 854, 1344, 861, 1506, 881, 544, 1622, 1971, 860, 1778, 2096, 806, 1495, 1084, 1043]. High-Dimensional [881, 544, 1778, 2096, 1495, 1043]. High-Fidelity [1622]. High-Performance [1768, 1506, 1971]. High-Throughput [854, 861, 860]. high-volume [1084]. Higher [2061, 1725, 1016, 971]. higher-dimensional [1016, 971]. Higher-Order [1725]. Highly [1333]. hill [650]. hill-climbing [650]. Historical [1083]. Homing [1095]. homogeneous [697]. Homophily [1728, 2058]. Homotopy [1911]. Hop [122, 1035, 1172, 377, 1052]. Hop-Constrained [122, 1035]. Hopfield [434]. Horizon [396, 953, 1241, 567]. Horn [3]. Hose [940]. Hospital [1611, 1679, 863, 1844]. Hot [1899]. hotel [1265]. hour [1191]. hour-ahead [1191]. Hub [1880, 1661, 1992, 1903, 695, 938, 905, 375, 924]. hub-and-spoke [905]. hull [1261, 389]. Hulls [242]. Human [1384, 2092, 1705]. Human-Robot [2092]. Humanitarian [1655]. Hungarian [176]. Hybrid [992, 790, 1856, 528, 428, 1629, 292, 1820, 1008, 1138, 492, 567, 518, 784]. hybridized [460]. Hypatia.jl [1773]. hyperbox [1043]. Hypercube [269, 69]. Hyperexponential [1516]. hyperexponentials [988]. hypergraph [1117]. Hyperlocal [1658]. hyperplanes [1066]. Hypothesis [425].

I-SMOOTH [1065]. ICD [1135]. ICU [1842]. ICU-Downstream [1842]. Identical [254, 798, 837]. Identification [257, 1408, 412, 600]. Identify [1362]. Identifying [1991, 33, 2044, 1638, 315, 1393, 1708, 1098, 389]. Idle [318]. II [65, 30, 648, 85, 1974, 591, 297]. ILP [744]. Image [1949, 1722, 277, 81, 775]. Image-Based [1722]. Imbalance [1867]. immunization [1108]. Impact [1440, 1997]. Imperfect [1625]. implantable [1135]. Implementation [10, 40, 340, 167, 69, 351, 85, 2019, 317, 27, 9, 339]. Implementations [272, 98, 115, 603]. Implementing [129, 720, 576, 818]. implication [923]. Implications [141, 1802]. Implicit [1492, 95]. Importance [1180, 1225]. Improve [1977]. Improved [1590, 421, 274, 1843, 2092, 313, 1423, 1206, 1668, 285, 1570, 669, 1188, 1007, 976, 1160]. Improvement [1551, 409, 211, 1227]. Improving [1640, 1242, 1709, 1765, 754, 1291, 719, 77, 861, 800, 919, 1667, 1141]. IMRT [1694, 1021]. Incorporating [1684, 1371, 2093, 1289, 260]. Incorporation [910]. increase [1280]. Incremental [1322, 923, 95, 387]. Independent [394, 241, 2075, 625]. Index [470, 471, 562, 563, 611, 612, 656, 657, 704, 705, 768, 769, 826, 827, 878, 879, 933, 934, 983, 984, 2015, 1086, 953, 635, 762, 793]. Indexed [1553, 782, 1211, 450]. Individual [1640, 1521, 1440]. Individualized [1478]. Induced [1316, 358, 1408]. Induction [686, 1788, 774, 1112]. Inductive [1727]. Industrial [535, 439, 498, 1192]. industrial-scale [1192]. Industry [2093, 1265]. Inequalities [1381, 1960, 1670, 33, 414, 1965, 1975, 965, 1064, 1285, 894, 740, 902, 633, 380, 1221, 401, 366, 964]. Inequality [1566, 1081]. Inequity [1692]. Inexact [893]. Infeasibilities [287]. infeasibility [1310]. Infeasible [80, 33, 91, 335, 410]. Infer [386]. Inference [1794, 244, 1132]. Infinite [178, 373, 727]. Infinite-State [178]. Infinitely [1882]. Infinity [1604]. Influence [1728, 1987, 1439, 1997, 44, 1654, 1693]. Influencer [1824]. Influenza [1337]. Information [1835, 407, 1761, 1625, 1558, 429, 1823, 1435, 1833, 2099, 1621, 1013, 888, 771, 388, 1304, 1251]. informational [663]. Informative [1520]. INFORMS [1553, 498, 483, 398, 1864, 1610, 1851, 1340]. infrastructure [1948]. InfrastructureModels [1948]. Infrastructures [1441]. Ingot [306]. inhomogeneous [914]. Initial [1046, 129]. Initialization [495, 2026]. Inner [2040, 1393, 1097]. Innovation [1958]. Input [14, 1750, 1751, 1544, 1977, 1100, 810, 349]. Input-Model [1544]. Inscribed [1960, 1315]. Insights [799, 1700, 2026, 151]. Inspection [1579, 452]. Inspired [2027, 1827, 1885]. Instance [82, 2026]. Instances [1793, 293, 1336, 284, 457, 373]. instantiation [1019]. Institute [531, 546, 554, 565, 575, 585, 613, 615, 628, 645, 654, 659]. Instrumentation [1916]. Insurance [1575]. Integer [201, 1450, 1334, 1665, 1918, 1316, 1769, 1415, 1242, 672, 1744, 1896, 1485, 1700, 1195, 1204, 1601, 246, 1168, 1980, 1749, 167, 876, 1487, 558, 831, 535, 1631, 1889, 1344, 1901, 1537, 414, 2008, 1843, 252, 1350, 1904, 1474, 1937, 1073, 1925, 1626, 861, 1629, 1699, 2082, 1571, 1433, 1871, 1372, 571, 1414, 1419, 1449, 384, 1671, 1451, 1526, 756, 677, 235, 1873, 5, 1722, 2014, 912, 1345, 525, 1672, 1595, 477, 1811, 1546, 1763, 1420, 1947, 1605, 1578, 1585, 962, 926, 1499, 325, 890, 1124, 582, 386, 1268, 1133, 538, 385, 902, 439, 629, 1259]. integer [498, 401, 410, 1055, 1097, 442, 850, 566, 419, 1078, 605, 527, 964, 1247, 911, 1277, 1053, 662, 1160, 807, 383, 857, 864]. integer-programming-based [1124]. Integral [3, 1748, 621]. Integrals [1740]. Integrated [1478, 1659, 359, 2039, 872, 1278, 1673, 1741, 1994, 1718, 1321, 922, 1395, 1806, 1148, 1619, 1907, 772, 1237, 1075, 333]. Integrating [514, 1521, 2076, 1857, 210]. Integration [76, 2040, 1483]. Intelligence [179, 543]. Intelligent [642, 226, 2045, 1280, 735, 376, 526]. Intensity [1684, 833, 2088, 927, 870]. Intensity-Modulated [1684, 833, 927, 870]. Intensive [1767, 1622]. Intentions [1642]. Interacting [1539, 1129]. Interaction [1408]. Interarrival [462]. Interconnecting [365]. Interconnection [229]. Interconnections [96]. Interdicting [2079].

Interdiction [1442, 1232, 1664, 1388, 1828, 1245, 1770, 1595, 1686, 1271, 724]. Interested [2036]. interface [815]. Interfaces [177]. Interhub [1880]. Interim [721]. Interior [165, 57, 278, 1696, 2038, 109, 202, 2019, 266, 191, 196, 27, 9, 194, 195, 1885, 329, 432, 1266, 1301, 483, 371, 697]. Interior-Point [278, 1696, 2019, 194, 195, 1301, 697]. Intermediate [1973, 784]. Intermittent [1924]. International [892]. Internet [462]. Internet-Type [462]. Interpretable [1957, 1879, 1857]. Intersection [1579, 1604]. Interval [2037, 388, 267, 954, 1091, 1118, 1182, 440]. Interval-Valued [2037]. Intervals [2080, 1100]. Intra [464]. Intra-Ring [464]. Intranets [746]. Introduction [259, 2018, 1, 7, 13, 20, 28, 39, 48, 56, 64, 73, 84, 94, 105, 114, 43, 85, 1970, 136, 2085, 804, 383]. Intrusion [780, 800]. Intrusion-Detection [780]. Inventory [992, 307, 1900, 1373, 232, 1909, 1346, 1545, 1633, 1827, 1102, 708, 1287, 777, 681, 1250, 848, 1278, 1150, 990]. inventory-routing [848]. Inverse [227, 1700, 212, 1775, 1950, 1536]. Inversion [239, 1396, 1354, 438, 336, 717]. Inverting [319, 739, 694]. Investigation [908]. Irrational [1707]. irregular [583]. irt [1816]. Isomorphic [288]. Isotonic [652]. Issue [1864, 1610, 761, 522, 1851, 2018, 1501, 596, 1502]. Issues [268, 148, 131, 967, 1251]. Item [2070, 2006, 1816, 1963, 1329, 1634, 1706]. Items [1558, 1125]. Iterated [1974, 1763, 503]. Iteration [97, 1062, 134, 1194]. Iterative [268, 1957, 2072, 1240, 1518, 1657, 2091, 1668, 1247]. iteratively [1065].

J [498, 1340, 483]. Jackknifing [126]. Jacobi [75]. JANOS [1659]. Java [474]. Jeroslow [2]. Job [205, 79, 1759, 295, 276, 204, 309, 935, 479, 1175]. Job-Shop [79, 935]. Jobs [1962, 241]. Join [249, 1279]. Joins [261, 473]. Joint [306, 1358, 1586, 1908, 995, 681]. Journal [1864, 1553, 1610, 1851, 1340]. judgment [533]. Julia [1948, 1724, 1505, 1933, 1173]. Jump [1516, 802].

K-Cuts [558]. Karmarkar [10]. Kernel [2070, 1966]. Kernels [1587]. Kernighan [536]. Kestrel [815]. Key [471, 563, 612, 657, 705, 769, 827, 879, 934, 984]. Keys [208]. Keyword [1860]. Kidney [1557, 1681]. King [790]. KKT [1301]. Knapsack [246, 415, 1232, 1740, 1388, 1886, 965, 1800, 1486, 1561, 1231, 446, 1588, 1064, 996, 325, 1292, 582, 623, 1074, 361, 1107, 670, 1182, 1012, 733, 903]. Knowledge [259, 58, 3, 428, 1229, 1362, 55, 257, 910, 264, 543, 1641, 1575, 774, 873, 959, 367]. Knowledge-Driven [1641]. Knowledge-Enabled [1229]. knowledge-gradient [873, 959]. Kolmogorov [1264]. Krige [1017]. Kriging [1646, 1841, 1331, 1252]. Kriging-Assisted [1841]. Kronecker [458, 594].

L [498, 1242, 1925]. L-Shaped [1242, 1925]. Label [1616, 1932, 1112].

Label-Correcting [1616]. label-dependent [1112]. Labels [348]. Lagrangean [941, 905, 502]. Lagrangian [1568, 1437, 1749, 1465, 1288, 1256, 729, 2052, 1545, 707, 1792, 1582]. Lagrangian-Type [1792]. Lags [1745]. Laguerre [319]. Landing [790]. Language [253, 44, 524, 405]. Languages [378, 884, 836]. LANs [365]. Laplace [239, 319, 438, 694, 739, 403, 717, 1354]. Large [745, 766, 1957, 2087, 746, 293, 1664, 1412, 1498, 1029, 213, 974, 1789, 223, 26, 1383, 1981, 2042, 1974, 155, 1577, 123, 106, 1464, 59, 1167, 1827, 284, 1548, 1874, 1645, 536, 632, 337, 376, 938, 1496, 1497, 1248, 1288, 487, 382, 494, 1189, 733, 663, 818, 1127]. large-margin [1127]. Large-Order [1498, 1029]. Large-Scale [745, 766, 2087, 746, 1664, 213, 1789, 2042, 1974, 155, 1577, 59, 1827, 284, 1548, 1874, 1645, 632, 337, 376, 938, 1248, 1288, 487, 382, 1189, 663, 818]. Largest [1960]. Laser [1999]. Last [196]. Late [258, 1112]. Latency [2025, 1654, 1479]. Latency-Constrained [2025]. Lateness [1347]. Latent [1759]. lattice [327, 339, 1126]. Lattices [1487]. laws [1072]. Layer [408, 646]. Layered [1052, 210]. Laying [643]. Layout [476, 1186, 357, 668, 821, 1109, 1148]. lead [990]. Learning [1640, 1384, 1949, 1478, 352, 1896, 1852, 1610, 1765, 1639, 1728, 1825, 2037, 501, 1203, 1883, 1614, 1468, 2062, 2056, 1904, 1861, 1925, 1193, 2098, 2042, 1798, 1946, 1539, 1733, 1705, 1868, 2058, 1966, 2051, 1638, 184, 1956, 1406, 1377, 1542, 812, 1998, 210, 1982, 1727, 1548, 1642, 1677, 1697, 1575, 1867, 1778, 1467, 1463, 1582, 1275, 1069, 1295, 842, 1258, 795, 367, 510, 1127]. Learning-Based [1542, 1697, 1275]. Learning-Supported [1861]. Least [917, 1439, 2080, 81]. Least-Cost [1439]. Least-Squares [917, 2080, 81]. Leg [1370]. Lemke [69]. Length [299, 1622, 718, 1306]. Let [1999]. Level [308, 1412, 1989, 1320, 1317, 1156, 1305, 999, 726]. level-2 [1305]. level-3 [999]. liars [1280]. Libraries [55, 646]. Library [1972, 1901, 2050, 1816, 104]. Lies [216]. lifespan [772]. Lifetime [1730, 1772, 1238]. LIFO [764]. Lift [1068, 1470]. Lift-and-Project [1068, 1470]. Lifted [401, 414, 1965, 964, 807]. lifting [838]. Light [1134, 17]. Lightpath [1774]. Like [1764, 946]. Likelihood [1392, 2060, 1750, 1648]. Likelihood-Ratio [1750]. Limited [1399, 1468]. Limits [1565, 1854]. Lin [536]. Line [209, 822, 2092, 223, 408, 1419, 353, 2075, 1630, 504, 1014]. Linear [1665, 1198, 1415, 475, 1119, 672, 1836, 2055, 192, 1485, 16, 1601, 937, 80, 103, 1143, 278, 831, 2038, 838, 713, 2072, 137, 107, 130, 1507, 1240, 1901, 212, 1902, 91, 82, 1775, 802, 138, 6, 77, 15, 1421, 1566, 102, 1453, 1383, 1270, 1514, 191, 196, 27, 9, 115, 1372, 1335, 24, 127, 1776, 1131, 1956, 1451, 71, 1799, 95, 59, 315, 161, 110, 1905, 292, 1587, 1672, 1870, 1420, 1885, 412, 593, 797, 1243, 896, 894, 335, 526, 508, 979, 976, 1200, 629, 437, 329, 432, 410, 1258, 381, 442, 467, 961, 327, 339, 533, 851]. linear [487, 382, 483, 715, 1282, 384, 371, 707, 785, 1053, 697, 1103, 807, 920, 383]. Linear-Programming [1143, 1335]. Linear-programming-based [838]. Linear-Quadratic [127, 785]. linear-time [961]. Linearization [1412, 1305, 999]. Linearized [1573]. linearly [961]. Lines [354, 390]. Link [2078, 1880, 849, 1591, 1174]. Linkage [719]. Links [1387]. Lion [146]. Lipschitz [1357]. Lists [162, 1807, 315]. Liver [1435, 1003]. Load [269, 1614, 1703, 1593]. loaded [1054]. Loading [940, 1510, 1742, 1697, 764, 901, 698, 805]. Local [205, 1261, 1048, 391, 549, 96, 1974, 1647, 1372, 309, 1587, 1660, 1057, 928, 935, 824, 623, 778, 460, 794, 760, 405, 853, 990, 792, 711, 784, 738]. local-search-based [623, 778]. localized [1295]. LOCALIZER [405]. Localizing [287]. Locating [80]. Location [308, 1769, 430, 1880, 1171, 1661, 1556, 1902, 141, 1679, 690, 2054, 1635, 1673, 2001, 1779, 1541, 1509, 86, 1992, 1947, 1603, 1087, 1309, 449, 1273, 928, 915, 695, 1101, 938, 1011, 1249, 1248, 1026, 1237, 758, 1288, 606, 924, 1052, 1311, 834, 969]. Location-Aware [1603]. location-routing [1101]. location/allocation [1249, 1248]. Locational [1855]. Locations [2057, 1679]. Locking [1370]. Locks [1390]. Logarithmic [1911, 1081]. Logic [1957, 1856, 1758, 501, 1600, 233, 523, 78, 1673, 1741, 1630, 1499, 385, 1011, 776, 383]. Logic-Based [1856, 1758, 1600, 1673, 1741, 1630, 1011]. Logistic [1884, 1194]. Logistics [1657, 1655]. Logit [1471, 1267, 1179]. Łojasiewicz [2009]. Long [2043, 1495]. long-cycle [1495]. Long-Tail [2043]. lookahead [390]. looping [757]. Loosely [1585]. Lost [2070]. Lot [781, 1909, 1146, 1671, 1801, 1345, 1329, 1634, 1261, 1241, 837, 1296, 1185]. Lot-Sizing [781, 1909, 1671, 1345, 1261, 1241, 837, 1296]. Low [570, 1589, 1803, 429, 1479]. Low-Cost [429]. Low-Diameter [1803]. Low-Discrepancy [570]. Low-Latency [1479]. Low-Rank [1589]. Lower [146, 937, 1484, 1747, 1989, 639, 1297, 377, 468]. Lower-Level [1989]. LP [558, 2072, 1413, 77, 626]. LP-based [626]. LP-Representations [77]. LPFML [629]. LPFORM [317]. LSTM [1844]. LU [1200, 576, 59]. lunch [732].

M [461, 1059, 1090, 121, 241, 2081, 131, 1059, 946, 121]. M/G/1 [946]. M/G/1-Type [946]. M/Gb/1 [131]. M/Gd/1 [131]. M/M/c [1059]. Machine [1640, 352, 1896, 1553, 180, 1825, 1324, 917, 158, 272, 1203, 258, 318, 2056, 1363, 2013, 1861, 61, 1673, 1733, 2058, 1655, 1423, 1702, 1956, 1377, 1308, 1796, 1982, 41, 1356, 1570, 316, 828, 1275, 632, 1211, 411, 503, 391, 1069, 798, 479, 619, 1219, 468, 840, 1215, 423, 450, 502]. Machines [895, 2070, 241, 1165, 639, 1329, 166, 837, 504, 511]. Macro [392]. Macromolecular [609]. MAD [1711]. Made [1865]. Main [262]. Maintaining [1163]. Maintenance [1409, 1994, 1559]. Major [193]. Majority [260]. Make [1395]. Make-to-Order [1395]. Makes [1711]. Makespan [255, 241]. Making [2068, 916, 2094, 521, 1597]. Malicious [1685]. Man [146]. Management [531, 546, 554, 565, 575, 585, 613, 615, 628, 645, 654, 659, 1573, 1030, 259, 393, 307, 400, 1599, 1187, 232, 451, 1346, 1321, 177, 644, 777, 663, 1123, 1103, 820, 1041, 1265]. Managerial [292]. Managing [1683, 1781, 1844]. Manipulation [47]. Manufacturing [1337, 823, 735]. Many [1099, 1837, 1811, 1002, 1306]. Many-Server [1099, 1306]. many-to-many [1002]. Many-to-One [1811]. MAP [1286, 1286]. Maps [348, 1109]. Margin [1404, 1127]. Marginals [1525, 835, 495]. maritime [1162, 1150]. markers [898]. Market [457, 779, 1229, 1946, 1999, 1907, 981, 1191, 1170]. Marketing [426, 1824]. Markets [1718, 1176]. Markov [268, 914, 458, 2037, 42, 65, 178, 1802, 1313, 6, 245, 946, 723, 326, 795, 594, 567, 1733, 2051, 248, 211, 762, 598, 1180, 12, 727, 1560, 66, 134]. Markov-Chain [42, 65]. Markov-Modulated [66]. Markovian [1040, 560, 1089]. Masking [348]. Mass [68]. Massively [120, 41]. Match [74]. Matching [1316, 139, 257, 236, 265, 234, 200, 2041, 1777, 1723, 284, 1811, 1976, 835, 775, 1002]. Matchings [416, 788]. Materials [1098]. Mathematical [817, 1030, 154, 379, 424, 1359, 686, 522, 103, 253, 1448, 1652, 183, 152, 1106, 603, 668, 322, 1501]. mathematical-programming [668]. Matheuristic [1373, 1542, 1287]. MathOptInterface [1652]. Matlab [1612]. Matrices [47, 1905]. Matrix [1676, 1879, 2073, 1927, 432, 1566, 36, 41, 894, 909, 455, 371]. Matroid [112]. Matroids [1468]. Max [509, 160, 830, 1763, 1621, 1036, 1182, 1962, 1938, 1352]. Max- [1036]. Max-Cut [830, 1352]. Max-Flow [160]. MAX-SAT [1938]. maximal [805]. maximal-space [805]. Maximally [699]. Maximization [1987, 1439, 1997, 1654, 1693, 1736, 1808, 1935, 1602, 785]. maximize [695, 888]. Maximizing [1804, 1120]. Maximum [1316, 1836, 1392, 1730, 1840, 488, 1670, 1465, 1592, 1649, 1322, 1699, 290, 8, 1167, 1315, 718, 1273, 1007, 1049, 724, 1238, 763]. Maximum-Entropy [1840]. maximum-flow [724]. MDD [1437]. MDD-Based [1437]. Mean [1442, 1360, 1065, 729]. mean-constrained [1065]. Mean-Risk [1442]. mean-variance [729]. Means [286, 351, 493, 956]. Measurable [15]. Measure [1455, 1711]. Measurement [1518, 2096, 1252]. Measures [148, 1344, 1598, 2061, 1572, 17, 1269, 987, 1184]. Measuring [156]. Mechanism [910, 260]. Mechanisms [720, 1459]. Media [1824, 1658, 1997]. Median [537, 2033, 1456, 142, 974, 1686, 449, 651, 375, 1293, 1255, 834, 1106]. medians [1115]. Mediation [1952]. Medical [1504]. Meet [1355]. Meet-in-the-Middle [1355]. Memetic [1917, 696, 1172]. Memorial [1340]. Memoriam [2, 461, 1403]. Memory [262, 1314, 458, 421, 588]. memory-efficient [458]. Merging [522, 547, 881, 1501]. Meritorious [2046]. Mesh [288]. Meshes [267]. Message [950, 1365, 1033]. Meta [617]. Meta-heuristics [617]. metabolism [1276]. metaconstraints [1210]. Metaheuristic [360, 929, 734, 435, 796]. Metaheuristics [2031, 794, 871]. Metamodeling [229]. Metamodels [1986, 2080, 1331]. Metastatic [1330]. metastrategy [663]. Method [319, 1918, 1242, 227, 1676, 1568, 1119, 352, 570, 154, 129, 444, 16, 1195, 1204, 2074, 1936, 165, 57, 1487, 1791, 1343, 2068, 92, 1696, 2038, 713, 158, 1911, 916, 2060, 351, 1738, 2061, 1843, 1787, 1515, 1937, 223, 1941, 496, 863, 2003, 1788, 279, 9, 155, 1648, 1419, 1958, 1857, 1799, 1411, 640, 493, 1945, 292, 1992, 1953, 1098, 1520, 1725, 675, 387, 810, 1283, 954, 1091, 1050, 589, 459, 1003, 633, 1007, 572, 870, 329, 648, 742, 1266, 1193, 851, 1081, 1078, 483, 753, 978, 697, 820, 920, 1296, 1036]. Method-Based [1411]. Methodologies [688]. Methodology [746, 1291, 311]. Methods [1684, 1360, 176, 1700, 368, 1825, 298, 278, 272, 130, 1333, 462, 1584, 273, 109, 78, 202, 2019, 1737, 1928, 1539, 1506, 191, 196, 115, 100, 281, 1380, 231, 5, 194, 1398, 195, 1706, 1947, 1686, 1582, 1317, 1302, 380, 1249, 1234, 726, 1179, 371, 1308, 1215, 865, 864]. Metric [1455, 2098]. Metrics [1371, 152]. Metropolitan [1661]. Microfluidic [972]. Microgrid [1585, 2049]. Middle [1355]. Migration [2034, 1856]. Migratory [1522]. Military [1504]. Milliseconds [1744]. MILP [1957, 492, 1457]. MILP-Based [1957]. MILP/CP [492]. Min [509, 1743, 190, 1698, 1763, 1621, 437, 664, 1182, 1117]. min-cost [437, 664]. min-cut [1117]. Min-Max [509, 1621, 1182]. Min-Sum [190]. Min-Sup-Min [1743]. mines [1139]. Minimal [1186, 80, 965, 446, 786]. Minimal-Height [1186]. Minimality [1485]. Minimally [33]. Minimax [505, 1577, 834]. Minimization [1442, 1347, 2034, 1980, 1324, 1401, 2015, 255, 1535, 2006, 408, 797, 709]. Minimize [258, 763]. Minimizing [1962, 61, 118, 510, 423]. Minimum [1821, 1804, 937, 416, 2031, 1785, 1674, 234, 200, 1739, 1166, 1595, 277, 1660, 1327, 1328, 2081, 1593, 1066, 797, 643, 1004, 870, 689, 1138, 638, 377, 1015, 1230, 852, 871, 1141, 1235]. Minimum-Cost [1821]. minimum-energy [1015]. Minimum-Weight [416, 2081]. Mining [425, 1291, 682, 1795, 540, 543, 544, 541, 1196, 424, 1006, 1110, 1092, 1502, 1000]. Mining-Based [1291]. MINLP [1530]. MINLPLib [538]. MINLPs [1333]. Minmax [449, 651, 1118]. Minor [1636]. MinRLE [1488]. MINSAT [501]. minsum [482]. MIP [571]. MIPLIB [850]. MIQP [1141]. MIR [902]. missile [1237]. Missing [1491]. Missions [1823]. Mitigating [1435]. Mitigation [2035]. Mixed [1450, 1334, 1665, 1918, 1769, 1415, 1744, 1896, 1700, 1195, 1601, 1980, 876, 558, 1631, 2037, 1537, 2008, 252, 1421, 1350, 1904, 1474, 1937, 1073, 1626, 1699, 1571, 1433, 1871, 1457, 1419, 1451, 756, 1799, 235, 5, 1722, 1345, 1672, 1546, 1947, 1585, 962, 926, 582, 1133, 538, 902, 1130, 410, 1097, 997, 850, 566, 419, 964, 1247, 1053, 1160, 807, 864]. mixed-binary [997]. Mixed-Integer [1665, 1918, 1769, 1744, 1896, 1601, 1980, 876, 1631, 1537, 2008, 252, 1350, 1904, 1474, 1937, 1073, 1626, 1571, 1433, 1871, 1451, 5, 1345, 1672, 1947, 1585, 962, 926, 1133, 538, 410, 1097, 850, 964, 1247, 1053, 1160, 807, 864]. MIXMAX [1426]. mixtures [1170]. ML [1959]. Mobile [746]. Mobility [1705, 1844]. mod [699]. mod- [699]. Mode [1483]. Model [745, 790, 122, 67, 1940, 76, 400, 359, 2059, 232, 262, 1586, 944, 1535, 2021, 365, 1325, 2077, 2054, 1362, 1635, 1923, 1471, 1071, 1522, 55, 1759, 1751, 1544, 1844, 101, 464, 177, 1183, 1622, 1869, 1428, 2045, 777, 960, 1133, 899, 898, 1158, 1039, 1267, 1237, 636, 1281, 930, 1092, 1031, 1179, 1044, 1108, 1132, 1104, 785, 434, 1152]. Model-Free [2054]. Modeled [1712]. Modeling [790, 931, 1659, 154, 378, 426, 1601, 58, 678, 359, 485, 1210, 1632, 253, 846, 1259, 1933, 1531, 2025, 43, 85, 1386, 1868, 1244, 22, 1579, 101, 177, 1040, 815, 1217, 524, 1269, 794, 660, 405, 646, 776, 333]. Modelling [1384, 1793, 836]. Models [2035, 672, 1669, 505, 1557, 2034, 54, 1396, 686, 103, 369, 1512, 2031, 6, 1707, 486, 1646, 1448, 1892, 26, 1071, 2051, 1674, 1647, 1457, 2088, 948, 616, 1985, 474, 1998, 1796, 1982, 1942, 185, 1881, 370, 1575, 1623, 1820, 2096, 1089, 1087, 603, 806, 458, 538, 1295, 1128, 1130, 492, 581, 1274, 1310, 1277, 800, 1177, 1170]. Modern [1531]. Modifications [82]. Modified [1062, 1266]. Modular [90, 1288]. Modularity [486]. Modularization [58]. Modulated [1684, 833, 66, 927, 870]. modulus [1029]. molecular [596]. Molecule [958]. Molecules [1362]. Moment [1716, 1902, 2061, 622]. Moments [1396, 520, 1725]. Mond [1989]. Monge [1551]. Monitoring [771, 1443, 1685]. monochromatic [1018]. monomial [1007]. Monotone [145, 273, 380, 1266, 510, 1183]. Monotonicity [1388, 1331]. Monotropic [158, 467]. Monte [1122, 1151, 1714, 1928, 2065, 1180, 865]. Morris [1411]. Morse [601]. Mortar [1603]. Mothership [1436]. motivated [1193]. Movement [1727]. Moves [891]. MP [850]. MR1478041 [483]. MR2549129 [920]. MR2663501 [884, 883, 882]. MR2663502 [884]. MR2663503 [884]. MRSA [931]. MS [398, 394, 396]. Much [323, 38]. Multi [1755, 1878, 1948, 761, 2070, 2006, 302, 1963, 1329, 1634, 1858, 708, 774, 1067, 1228]. multi-activity [1067]. Multi-Bidder [2006]. multi-cloud [1228]. multi-decision-tree [774]. Multi-infrastructure [1948]. Multi-Issue [761]. Multi-Item [2070, 2006, 1963, 1329, 1634]. multi-period [708]. Multi-Resolution [1858]. Multi-resource [1755, 1878]. Multi-Tier [302]. multiactivity [1233]. Multiattribute [15]. multicategory [618]. Multiclass [1286, 1059, 591, 1227]. Multicommodity [1437, 2021, 127, 1542, 304, 2010, 557, 363, 437, 1055, 534, 334, 662, 1159]. multicommodity-flow [662]. Multicompartment [1819]. Multicomponent [1559]. Multicover [1550]. Multicriteria [831]. Multidepot [34]. Multidimensional [1313, 1486, 623, 1012, 903]. Multidrop [199]. Multidrug [1477]. Multidrug-Resistant [1477]. Multiechelon [1801]. Multifaceted [101]. multifacility [1300]. Multigraph [2066]. Multigrid [1706]. multigroup [1133]. Multiknapsack [1394, 1562]. multilabel [1023]. multilateral [1251]. Multileaf [845]. Multilevel [1079, 1217, 781, 1909, 1644, 1706, 1826, 689, 1311]. multimachine [677]. Multimodal [2002, 1934, 804]. Multimode [691]. Multinomial [1471, 1884]. Multiobjective [1353, 1669, 2055, 1488, 1929, 1455, 2090, 1584, 1616, 1541, 1129, 1449, 1830, 1508, 1907, 1023, 847, 808, 715, 1299, 911, 809, 1082]. multipass [390]. Multiperiod [1735]. Multiple [1821, 1620, 1498, 1134, 765, 68, 1547, 1421, 606, 1800, 1539, 1486, 1037, 598, 24, 468, 1606, 568, 1196, 1496, 1497, 1029, 1295, 670, 1125, 356, 339, 1126, 1136, 1019, 649, 592, 1197, 1251]. Multiple-Choice [1421]. Multiple-Cost-Row [24]. Multiple-machine [468]. multiple-objective [649]. Multiple-Product [568].

Multiple-Query [68]. Multiplicative [1699]. Multiplier [128]. Multipliers [2068, 2038]. Multipliers-Based [2038]. Multiprecision [1901]. Multiprocessor [255, 639, 711]. Multiproduct [1578, 785]. multiresolution [660]. Multiresource [1619]. multiresponse [989]. Multiretailer [2070]. Multirow [1604]. Multiscale [1726, 1031]. Multiserver [49, 485, 1054]. Multisource [2076]. Multisourcing [1281]. Multistage [1427, 846, 2094, 1524, 1552, 1247, 1958, 1653, 2086, 1069, 1301, 729, 1236]. multistart [421, 738]. Multitechnology [490]. multitrip [1047]. Multitype [1619]. Multiunit [799]. Multivariable [1588]. Multivariate [1836, 1946, 1170, 1428, 835, 960]. multivehicle [1102, 1287]. Multiversion [1921]. Multiview [1642]. Music [683, 684, 687, 685, 688]. Music-Retrieval [688]. Must [2078]. Must-Link [2078]. myopically [888].

n [241]. Names [76]. Narrating [1999]. Nash [1540, 1991, 1889, 1761]. Nash-Bargaining [1761]. National [400]. Natural [1425, 1773, 1401, 1368, 1462]. Navigation [2036]. Near [1918, 632, 1492, 1587, 995, 1256]. Near-Linear [1587]. Near-Optimal [1918, 1492, 632, 995, 1256]. Nearest [222, 570, 1400, 1071, 749, 594]. nearest-neighbor [749]. Nearly [1438, 248]. Need [1793, 323]. needs [981]. negative [1092]. negotiation [1251]. Negotiations [761]. Neighbor [570, 1538, 1071, 749]. Neighborhood [745, 766, 2089, 1379, 556, 764, 1248, 758, 1000, 1159]. Neighborhoods [1464]. Neighbors [1400]. Neonatal [1622]. NEOS [815]. Nested [331, 2060, 1726, 2003, 1576, 1267, 930, 729, 1179]. Net [184, 326]. Nets [185]. Network [940, 187, 1425, 1442, 122, 40, 1669, 2087, 305, 394, 1771, 1728, 2071, 1856, 916, 682, 52, 1286, 1766, 141, 2021, 2027, 908, 1375, 2042, 1657, 2058, 44, 291, 591, 2091, 127, 182, 519, 1828, 1542, 229, 464, 177, 1942, 23, 1374, 60, 1820, 2079, 669, 1156, 1500, 449, 1139, 915, 941, 1038, 1035, 634, 459, 363, 907, 1055, 534, 417, 1174, 1075, 1281, 334, 576, 630, 366, 1034, 1190, 724, 792, 649, 1123, 1103, 820, 1159, 1041, 1265]. Network-Based [23]. network-flow [534]. network-interdiction [724]. Networked [1430]. Networks [476, 1615, 1574, 2055, 1900, 748, 1730, 490, 287, 1692, 117, 1154, 1774, 1401, 199, 942, 96, 687, 1439, 608, 1886, 1854, 313, 690, 1095, 223, 26, 128, 1099, 2065, 183, 1767, 22, 1090, 62, 87, 1648, 198, 1654, 1693, 186, 1978, 304, 1167, 1408, 25, 240, 292, 1827, 1859, 32, 1653, 1462, 1463, 1594, 1300, 651, 364, 695, 1040, 634, 376, 759, 1027, 998, 979, 453, 1269, 580, 581, 966, 594, 420, 1015, 1070, 803, 1264, 834, 939, 968, 452, 440, 406, 1279, 791, 465, 1238]. Networks-Inspired [1827]. Neural [40, 2055, 1900, 687, 1886, 1658, 1975, 2065, 183, 291, 62, 1648, 184, 182, 406, 186, 185, 23, 292, 1827, 1859, 2045, 669, 453]. Neural-Net [184]. neural-network [669]. News [1658]. Newsvendor [1963, 1578]. Newton [222, 231]. NLP [738]. No [1390, 347, 498, 732, 483, 939]. no-free-lunch [732]. No-Wait [1390]. Node [788, 2065, 590, 1691, 966, 1220].

Node-Weighted [788]. Nodes [1917]. Noise [139, 1326, 1444, 2022, 2020]. Noise-Aware [2022]. noisy [600]. Non [1877, 482, 273, 1386, 1040, 715, 583]. Non-approximability [482]. non-linear [715]. non-Markovian [1040]. Non-Monotone [273]. Non-Order-Associated [1877]. Non-Poisson [1386]. non-preemptive [583]. Nonadaptability [1987]. Nonconvex [886, 1890, 1694, 1870, 1420]. Nonconvexity [1589]. nondeterministic [436]. nondifferentiability [707]. Nondominated [315, 1508, 734, 911]. Nonhomogeneous [228]. Nonidentical [1401, 1329]. noniterative [785]. Nonlinear [887, 1927, 2008, 159, 1794, 157, 1737, 2024, 266, 116, 99, 1571, 1577, 1379, 1848, 106, 292, 926, 1285, 538, 589, 1158, 979, 1097, 674, 618, 980]. Nonmonotonic [274]. nonnegative [1065]. Nonparametric [1343, 1868, 1759, 1656, 1587]. Nonpreemptive [1553, 49, 1211, 497]. nonprobability [336]. nonrenewal [349]. Nonstandard [99]. Nonstationary [1386, 875, 726, 1284, 1264]. nonsuccessive [327]. Nonsymmetric [1612]. Norm [242, 1511, 1604]. Normal [1428, 835, 873, 1183, 1170]. normal-copula [835]. normalization [322]. Normalized [1098]. NORTA [495, 1005]. Note [1996, 92, 1286, 124, 1389, 1417, 1432, 1473, 1503, 1529, 1555, 1581, 1609, 1651, 1689, 1720, 1753, 1784, 1814, 1832, 1850, 1863, 1876, 1894, 1914, 1931, 1955, 1969, 1984, 2017, 2030, 2048, 2064, 2084, 688, 81, 498, 920]. Note-Based [688]. Novel [778, 2022, 1673]. Nuclear [1098]. Number [1960, 1498, 802, 1962, 1465, 1426, 1466, 1495, 1153, 339, 1126, 402, 592, 1197, 763]. Numeric [1640]. Numerical [239, 154, 368, 131, 6, 202, 289, 223, 134, 621, 36, 4, 438, 914, 336, 717]. Numerically [319, 1922, 876, 1297, 694, 954]. Nurse [1063, 1981].

OAR [1371]. Obfuscation [1835]. Object [1949, 680]. Objective [1405, 1164, 1372, 1366, 1837, 1413, 381, 1414, 724, 649]. Objectives [241, 1577]. Objects [226]. Oblique [618]. Observability [2037]. Observable [1733]. observational [1132]. Observations [1797]. Observed [1444, 567]. obstacle [1121]. Obstetric [1683]. obtained [743]. obviate [732]. obviating [707]. Odd [1592]. Odd-Cycle [1592]. Off [1461, 2012, 1021]. offer [981]. Offerings [1471]. office [760]. office-space-allocation [760]. Offline [1443, 1395, 1602, 1591]. offs [1006]. Old [273]. On-Demand [2041, 2002]. On/Off [1461]. One [201, 2070, 77, 14, 45, 1056, 1564, 1447, 1811, 1570, 709, 874, 840]. one-center [874]. One-Dimensional [45, 1447, 709]. One-Machine [1570, 840]. One-Warehouse [2070]. Online [1369, 799, 1744, 480, 1188, 2039, 1835, 1518, 1105, 1854, 1443, 1761, 390, 1872, 1958, 474, 1395, 1602, 1723, 1520, 1463, 1591, 1302, 952, 1045, 1226]. Only [2076, 1860, 978]. onto [1961]. ontologies [1193]. Ontology [900]. Open [180, 1865, 1973, 1583, 1336, 2049, 709, 1139, 479, 1001, 465, 763]. open-pit [1139]. open-shop [1001]. Open-Source [1865, 1973, 1583, 1336, 2049]. open-stacks [709]. Operating [951, 1600, 1298, 1606, 1239, 1084].

Operation [310]. Operational [1625]. Operations [1949, 531, 546, 554, 565, 575, 585, 613, 615, 628, 645, 654, 671, 659, 1573, 393, 1661, 1851, 2018, 2097, 1173, 1994, 1833, 342, 347, 474, 458, 504]. Operations-Research [474]. Operator [380]. Operator-splitting [380]. Opinions [148, 1226]. OPL [525]. Opportunities [393, 1815, 586, 2062]. OPT [2089]. Optic [199]. Optical [464, 759, 939]. Optima [831, 712]. Optimal [1381, 1918, 1845, 1476, 1620, 1326, 1461, 819, 2057, 439, 535, 1991, 254, 679, 1187, 1551, 163, 50, 498, 1492, 1258, 1367, 1921, 1072, 74, 1191, 2021, 1135, 1552, 356, 144, 1920, 1788, 2082, 70, 605, 803, 1638, 1767, 1337, 1718, 972, 1517, 1908, 119, 1873, 624, 1978, 1479, 1953, 1176, 1653, 1885, 1621, 1898, 1594, 1585, 995, 821, 632, 749, 380, 1237, 1256, 1001, 762, 1290, 1276, 832, 1127]. Optimality [1845, 2008, 255, 1985, 1398, 989, 1227]. Optimisation [1353]. Optimising [1892]. Optimization [1684, 1120, 1918, 1939, 833, 209, 514, 1573, 1662, 1743, 1568, 1371, 2004, 208, 216, 1418, 352, 1948, 1113, 1213, 1527, 1669, 269, 1048, 1446, 1744, 1756, 1879, 1896, 1940, 1897, 887, 1700, 1207, 1855, 748, 1341, 855, 1746, 1589, 1865, 1716, 854, 1734, 1154, 298, 1487, 1488, 1791, 1343, 1631, 35, 2038, 1512, 1632, 601, 680, 1839, 1847, 2072, 1357, 1944, 2023, 395, 513, 517, 1358, 160, 1933, 1248, 1901, 515, 1537, 261, 270, 1775, 1455, 2090, 68, 1584, 1795, 1782, 523, 1367, 1787, 2032, 1794, 1572, 215, 1421, 2056, 1592, 975, 2027, 2026, 219, 516, 1737, 2024, 2077]. Optimization [1919, 1652, 1941, 1717, 1471, 1657, 1950, 2097, 1966, 1541, 1514, 44, 1694, 1368, 1993, 1624, 1647, 1638, 1571, 1407, 929, 155, 1655, 1612, 1511, 2052, 881, 1732, 1848, 1526, 1489, 1882, 1469, 231, 1985, 1830, 1736, 1560, 188, 1636, 1998, 1656, 1460, 1508, 1393, 1056, 1786, 23, 1644, 1827, 1837, 1811, 1430, 2007, 1858, 1953, 1924, 1523, 1459, 1490, 4, 860, 1677, 1792, 1757, 1623, 1820, 1323, 1559, 1338, 1192, 814, 725, 843, 989, 1500, 1196, 1285, 886, 1243, 681, 824, 337, 1050, 898, 385, 1027, 1017, 815, 979, 1217, 586, 1301, 1105, 463, 1072, 492, 1023, 581, 732]. optimization [966, 467, 847, 741, 567, 1244, 949, 1129, 1284, 715, 1299, 1289, 1247, 884, 1310, 753, 859, 862, 1021, 1132, 707, 1219, 1308, 406, 818, 882, 738, 650, 1082, 1043, 1226]. Optimization-based [1120, 1248, 1459]. Optimization-Driven [1469]. Optimize [1415, 417]. Optimized [1740, 2020, 1874]. Optimizer [2022, 161]. Optimizing [845, 1330, 1836, 1670, 1889, 96, 604, 1524, 1109, 429, 1859, 548]. Optimum [51]. option [1225, 865]. Options [1838, 1354]. OR/MS [398, 394, 396]. OR/SM [333]. Order [1769, 1877, 1498, 2032, 244, 1637, 2052, 5, 1395, 912, 2009, 1428, 1677, 1725, 1623, 1496, 1497, 1029, 508, 661, 961, 1192]. Order-Preserving [912]. Ordered [2033, 1456, 834]. Ordering [18, 1698, 896, 460, 371]. Orderings [1765, 1079]. Orders [765, 2011]. Ordinal [1656]. Organisms [1477]. organized [1070]. Organizing [292]. Orientation [833, 1534]. Oriented [1734, 846, 160]. Orienteering [119, 1805, 1627, 374]. Origin [917]. Original [212]. Orthogonal [866]. Oscillation [1422]. Other [597]. Out-of-Hospital [1611]. Outbreaks [1685]. outcome [911]. Outer [2040, 926, 1097]. outer-inner [1097]. Outlook [1871]. Outpatient [1325, 1302]. Output [493]. outsourcing [1190]. OVA [1071]. Overbooking [916]. Overhead [97]. Overheads [249]. Overlapping [1706, 737]. overlay [941]. Overload [1906, 1189]. Overload-Checking [1906]. Overnight [2066]. Overtime [1410]. Overview [43, 383].

P2P [1269]. Package [1865, 1505, 1973, 1612, 1777, 1971]. Packet [26, 420, 803]. Packet-Switched [26]. Packing [1922, 1846, 1355, 1563, 1424, 549, 1866, 1745, 551, 1674, 1144, 1606, 1447, 1057, 1454, 814, 1016, 641, 801, 963, 1257, 1076, 435, 665, 913, 710, 1051]. packings [743]. Pages [719]. Pairs [1897, 1438, 1002]. Pairwise [139, 1347, 2071]. Pairwise-Disjoint [139]. Palindromes [597]. Panacea [347]. Pandemic [2037]. Papers [1864]. PaPILO [1901]. Parallel [21, 169, 268, 430, 170, 936, 148, 153, 176, 1829, 117, 340, 328, 1412, 69, 254, 1319, 97, 249, 272, 323, 271, 168, 174, 163, 1901, 270, 957, 781, 1854, 98, 1363, 2013, 1673, 2042, 171, 150, 152, 1284, 45, 120, 1423, 1449, 24, 127, 859, 1848, 1165, 1796, 204, 173, 639, 1905, 4, 41, 1645, 166, 603, 951, 411, 798, 1497, 380, 837, 504, 566, 487, 619, 1215, 511, 857]. Parallelization [856, 1314, 1667, 557]. Parallelizing [444]. Parameter [2026, 481, 1127]. Parameterized [1615]. Parameters [2037, 157]. Parametric [352, 241]. Parapint [1848]. Pareto [1334, 1060, 831, 971, 712]. parity [740]. parsimony [899, 599, 839]. Part [43, 85, 296, 297, 535, 19, 30, 591, 590]. Partial [1819, 74, 1550, 1273]. Partial-Match [74]. Partially [1733, 1338, 567]. particle [1129]. Particular [35]. Partition [1427, 141, 676, 1317, 1096]. Partition-Based [1317]. Partitioning [642, 399, 264, 1748, 418, 487, 1019, 1117]. Parts [1346]. Party [1835, 1990]. Passage [1516]. Passenger [1806]. Passing [950]. Passive [1997]. Patch [1187]. Path [476, 1926, 1438, 787, 1936, 1696, 1740, 1025, 53, 261, 2002, 1441, 1565, 1675, 1434, 906, 486, 1616, 408, 124, 1245, 1422, 635, 731, 817, 967, 692, 420, 968, 452]. Path-Based [1926, 2002, 452]. Path-Following [1936, 1696]. Path-reduced [906]. Paths [755, 1348, 243, 1710, 376, 375, 534]. Patient [1619, 1254]. Patrol [1899]. Pattern [1485, 684, 1376, 1291, 71]. Patterns [2043, 1010, 1045]. Paulson [1645]. Pay [799]. payoffs [1225]. PDE [1937]. PDE-Constrained [1937]. Peaceman [2074, 1738]. pediatric [1108]. Peer [1728, 941, 1070]. peer-to-peer [941, 1070]. Penalties [1565]. Penalty [1121, 298, 238, 1333, 1515, 127, 1799, 132, 1235]. Penalty-based [1121]. Percent [1772]. Percentile [823]. Perfect [416, 265, 234, 200]. Performance [268, 1768, 426, 199, 1291, 262, 149, 313, 75, 280, 722, 1506, 152, 124, 1377, 711, 1613, 956, 1971, 678, 385, 987, 373, 1028, 507, 1141]. Period [1344, 1185, 708, 1261, 718]. Periodic [1687, 320, 355, 1318, 356, 1010]. perishability [708]. PERLE [1488]. Permutation [1762, 625, 507]. Permutations [1402, 489]. Persistent [251].

Personal [1777]. Personalization [540]. Personalized [1597, 1869, 1067, 1233]. Personnel [863]. Perspective [394, 1482, 182, 177, 1578, 1757, 1513, 1127]. Perspectives [515, 1838, 225, 882, 500, 1181]. PERT [968]. Perturbation [293, 1866]. Perturbation-Based [1866]. perturbations [518]. Pessimistic [1902, 1490]. Peta [1762]. Peta-Scale [1762]. Petri [326]. PH [89, 289, 1286, 1090]. Phantom [497]. Phase [1397, 157, 945]. phase-type [945]. phases [1283]. Pht [591, 590, 590]. Pht/Pht/ [591, 590]. Phylogenetic [608]. Physical [68]. Physician [1410, 1351]. Picker [1531, 2000, 1637]. Picking [1637, 1742]. Pickup [1437, 764]. Pickup-and-Delivery [1437]. Pickups [1818]. Pieces [145]. Piecewise [107, 1146, 1453, 1451, 1526, 1672, 1065, 467, 851, 920]. piecewise-constant [1065]. Piecewise-Linear [107, 467]. Pipeline [1949, 1368, 1462]. piracy [1269]. pit [1139]. Pivot [201]. pivoting [762]. Placement [1892, 1923, 852, 452]. Placing [348]. plan [1021]. Planar [1273, 647, 648, 32]. Plane [1676, 303, 243, 234, 1582, 363, 633, 838, 1224, 1159]. Planes [1756, 1815, 14, 1776, 821, 620, 566]. planner [1254, 2049]. Planning [1684, 833, 1878, 1253, 1480, 1825, 1599, 1399, 1758, 1410, 53, 2067, 2093, 1625, 1694, 1741, 595, 124, 1732, 1298, 1830, 1398, 1374, 1606, 2086, 1619, 1338, 1907, 927, 1196, 1239, 870, 1205, 741]. Plant [357, 1717, 1011, 758]. plant-location [758]. Plants [1532]. Platform [394, 2049]. Platforms [2036, 1377]. Platoon [1897]. Play [1764]. POI [1869]. Point [222, 165, 57, 278, 1696, 2038, 466, 109, 202, 2019, 266, 191, 196, 27, 9, 2052, 194, 195, 1885, 1212, 775, 801, 329, 1266, 1301, 483, 371, 697]. point-based [801]. point-matching [775]. pointed [389]. Points [1334, 139, 1515, 234, 589, 911]. Poisson [1396, 228, 1386, 889, 66]. Policies [679, 1187, 1551, 1504, 780, 702, 995, 1227, 624]. Policy [1551, 1062, 1362, 2065, 211, 1898, 1820, 1069, 873, 723, 1227, 1284, 1031, 1194]. Polishing [756]. politeness [771]. Polling [1396]. Polygon [144, 1115]. Polygonal [466]. Polygons [331]. Polyhedra [3, 1109]. Polyhedral [940, 305, 1700, 1810, 1671, 1510, 1978, 324, 922]. polyhedron [667]. Polynomial [2094, 1587, 1424, 1270, 445]. Polyphonic [685]. Polytope [1361]. Polytopic [1315]. Pool [861]. Pooling [51, 256, 116, 951]. pools [747]. Population [1844, 1710, 600]. Populations [1620, 599]. portable [1495, 1496]. Portfolio [1701, 1791, 1343, 1344, 1953, 1151]. portfolios [1049]. Pos [222]. Positioning [2074, 602]. Positive [755, 1092, 1693, 1860]. Positive-versus-negative [1092]. Positivity [1676]. Possibilistic [15]. Post [883]. Post-decision [883]. Posterior [1713]. Posterior-Based [1713]. Postman [1579, 409]. posynomial [970]. Potable [1978]. Potential [601, 1266]. Power [1845, 739, 1615, 286, 1829, 1143, 1401, 462, 1524, 1726, 1717, 1532, 1994, 1908, 2007, 1653, 1338, 326]. Power-Efficiently [1615]. Power-Series [286]. Power-Tailed [462]. powerful [1058, 883]. Practical [1111, 467, 1656, 173, 1490]. Practice [514, 513, 225, 194, 1078].

Practitioner [345]. Pre [1998]. Pre-Trained [1998]. Precedence [61, 1570]. Precision [2072, 1028]. Preconditioned [115]. preconditioner [594, 791]. Predefined [1063]. Predicate [477]. Predict [1622]. Predicting [799, 1625, 1482, 962]. Prediction [1477, 917, 849, 1861, 1760, 1657, 1733, 1872, 1641, 1591, 1174]. Prediction-and-Optimization [1657]. Prediction-Based [1872]. Prediction-Driven [1477]. Predictions [1640, 395, 1071, 1727]. Predictive [1659, 1440, 1994, 1998, 1634, 1491, 1092]. preemption [828]. Preemptive [50, 1165, 356, 583]. Preference [1384, 1539, 1868, 1591, 734]. Preference-Learning [1868]. preference-nondominated [734]. Preferences [1857]. Preferential [1985]. Prefetching [682]. Preprocessing [116, 968, 235, 25, 110, 240, 541, 580]. prerelease [1269]. Prescriptive [1659, 1896]. Presence [2076, 1207, 716]. Presence-Only [2076]. Preserving [912, 1342]. Presolve [1450, 329, 483, 1310]. Presolving [1901]. Pretrained [2044]. prevention [1013, 800]. Price [1456, 1569, 1819, 1679, 975, 1363, 1800, 1981, 1817, 1532, 1509, 1423, 1790, 1380, 1447, 1742, 1454, 1697, 2010, 2095, 681, 489, 938, 1239, 963, 1055, 1255, 803, 1214, 939, 1179, 1233, 1051]. Prices [1790, 820]. Pricing [1369, 1771, 1692, 1267, 1838, 1986, 1981, 569, 1354, 730, 1080, 636, 1214, 776, 1179, 1108, 865, 1041, 1265]. Primal [165, 57, 758, 109, 266, 9, 70, 630, 265, 200, 1380, 838, 976, 334, 1206]. Primal- [70]. Primal-Dual [165, 57, 266, 9, 265, 200, 758]. Primal-dual-based [630]. Primary [198]. Principal [67, 2053]. Principle [1355]. printed [675]. printed-circuit [675]. Prior [910, 329, 483]. Priorities [1712]. Prioritization [1923]. Priority [520, 1059, 49, 497]. Privacy [918, 1342, 1013]. Privacy-Preserving [1342]. Private [1468, 1467, 1251]. Prize [1915, 1324, 1164, 1339, 1663]. Prize-Collecting [1324, 1164, 1339, 1663]. Probabilistic [336, 108, 78, 1429, 399, 824, 1230, 1044]. Probabilities [178, 245, 244, 49, 89, 1272, 560]. Probability [239, 1518, 1444, 1428, 742]. Probing [235]. Problem [308, 21, 1381, 1316, 1769, 766, 1216, 1952, 940, 209, 79, 509, 992, 1425, 2012, 790, 936, 1922, 1846, 1394, 176, 76, 1979, 1915, 1438, 1771, 866, 2074, 1617, 415, 537, 1484, 1557, 490, 1171, 1840, 2033, 360, 488, 133, 1692, 1556, 1670, 1324, 1563, 1534, 1856, 1456, 238, 2031, 142, 942, 1025, 1972, 1551, 549, 190, 51, 1358, 1349, 135, 1022, 212, 256, 957, 303, 255, 1628, 275, 1886, 1819, 2000, 478, 781, 1782, 1843, 1465, 1434, 849, 2092, 1747, 1679, 357, 2027, 98, 486, 908, 1616, 1361, 1474, 237, 1448, 1861, 1583, 1649]. Problem [1891, 2077, 1800, 301, 872, 1164, 1809, 1745, 1322, 1635, 1673, 189, 2001, 861, 1817, 116, 1532, 1899, 1522, 1486, 551, 830, 1037, 291, 1974, 1674, 257, 1731, 1561, 99, 1332, 302, 1562, 87, 1729, 1231, 446, 1385, 1436, 182, 296, 297, 285, 1671, 1693, 260, 119, 1550, 104, 294, 1545, 86, 1579, 1691, 1777, 290, 118, 8, 812, 464, 31, 1464, 1510, 1166, 276, 1945, 204, 912, 1345, 639, 23, 1606, 1564, 568, 1329, 1947, 2066, 1588, 1459, 1805, 1490, 1057, 1570, 1578, 1772, 1686, 1697, 2095, 1903, 691, 1593, 1582, 316, 708, 1121, 1309, 1156, 731, 1066, 412]. problem [1287, 651, 1009, 1273, 1124, 387, 1008, 1292, 782, 582, 824, 915, 941, 1246, 1035, 556, 896, 623, 967, 1074, 1115, 764, 1004, 361, 973, 1039, 503, 1101, 740, 696, 775, 1260, 391, 994, 798, 620, 552, 1002, 963, 572, 870, 901, 1293, 1011, 1118, 374, 1257, 421, 1107, 670, 1241, 1172, 1297, 1182, 460, 689, 772, 1138, 1130, 1249, 638, 1305, 1026, 1175, 999, 1076, 758, 409, 907, 1055, 789, 534, 757, 692, 1023, 606, 418, 874, 795, 760, 1174, 1015, 747, 1052, 1230, 796, 1001, 436, 1012, 1255, 630, 605, 808, 1047, 852, 698]. problem [1034, 1142, 1214, 373, 913, 1058, 548, 1254, 714, 324, 1162, 805, 1228, 1294, 804, 710, 455, 384, 903, 862, 699, 518, 871, 724, 677, 792, 1303, 1014, 1307, 969, 1096, 583, 809, 784, 869, 712, 555, 1215, 434, 1296, 1159, 578, 1148, 1185, 1181, 1235, 502]. Problems [1918, 222, 1821, 1573, 1662, 314, 430, 1347, 1568, 140, 269, 331, 1926, 1485, 217, 16, 1553, 1601, 1855, 246, 1715, 1516, 1980, 2070, 298, 1791, 1355, 1412, 18, 535, 1631, 35, 1424, 713, 1632, 1344, 1944, 1388, 846, 974, 1933, 1762, 1531, 270, 218, 1766, 2008, 1698, 1565, 1675, 2061, 2090, 1584, 1795, 1789, 143, 1794, 311, 1421, 46, 893, 1973, 1515, 1937, 1649, 1625, 1652, 2054, 1981, 1339, 1375, 1941, 34, 1626, 1378, 1629, 1741, 236, 1407, 929, 1871, 45, 267, 1423, 1577, 1391, 1536, 595, 1908, 1776, 2052, 881, 1663, 123, 1545, 1828, 1799, 235, 1165]. Problems [1144, 1460, 1366, 1748, 1508, 1633, 1695, 1167, 1837, 1963, 1447, 1811, 1763, 1634, 1548, 2099, 1667, 2075, 1867, 1645, 1102, 669, 828, 1261, 817, 821, 536, 829, 593, 372, 632, 928, 364, 886, 1211, 634, 625, 557, 411, 938, 363, 385, 1267, 439, 375, 437, 664, 1259, 1248, 498, 560, 991, 482, 1266, 848, 492, 837, 1023, 456, 600, 924, 588, 850, 566, 847, 1075, 1081, 1278, 487, 435, 1271, 382, 1199, 334, 665, 1299, 494, 729, 733, 1274, 834, 1116, 1179, 968, 1184, 626, 1177, 468, 818, 1103, 662, 820, 970, 650, 980, 507, 1043, 577]. problems [450]. Procedure [165, 1456, 1911, 228, 128, 88, 353, 956, 1645, 691, 412, 996, 1302, 1005, 463, 660, 722, 366, 715]. Procedures [1713, 2050, 908, 2089, 700, 1517, 1056, 973, 500, 409, 581, 436]. Process [392, 1291, 1325, 1073, 1733, 1741, 1647, 1759, 1090, 1013, 1006, 987]. Processes [268, 1516, 368, 2037, 250, 228, 1386, 211, 1712, 66, 2069, 1820, 1198, 810, 875, 723, 326, 660, 567, 727]. Processing [1949, 29, 1412, 685, 431, 277, 1356, 951, 479, 1305, 440, 1219, 624]. Processors [254, 50, 356]. produce [777]. produced [327]. Product [754, 1781, 1941, 1946, 90, 1958, 568, 1642, 772, 594, 1150, 1276]. Production [306, 2039, 1373, 68, 1687, 2067, 1146, 872, 1483, 1741, 164, 595, 1395, 1992, 1907, 1102, 708, 772, 1278, 785]. Products [306, 2050, 1395, 1196, 773]. profile [773]. profile-association [773]. Profiles [1620]. Profit [1627, 785]. profit-maximization [785]. Profits [1805]. Program [1618, 82, 161, 1578, 335, 432, 1141]. programme [911]. Programming [1450, 1334, 1665, 10, 1316, 1769, 1540, 1119, 672, 1353, 1890, 1724, 332, 192, 1485, 16, 1195, 1204, 379, 1168, 165, 1359, 1895, 522, 1143, 1670, 167, 1951, 1401, 1505, 1889, 158, 942, 137, 916, 271, 1927, 1929, 253, 130, 160, 1240, 212, 159, 138, 1838, 523, 15, 1504, 3, 157, 252, 2093, 1781, 1552, 1474, 1073, 1809, 496, 1717, 861, 1378, 191, 196, 1629, 183, 27, 9, 856, 152, 1433, 1871, 1680, 45, 1372, 571, 1414, 1419, 1335, 1449, 24, 1776, 881, 1131, 1671, 1451, 756, 71, 1320, 235, 5, 59, 1801, 912, 1345, 525]. Programming [25, 110, 240, 1595, 477, 1546, 1763, 1420, 1947, 1653, 1976, 2086, 2010, 2075, 1623, 1559, 943, 1198, 412, 817, 668, 821, 1499, 593, 607, 322, 475, 890, 935, 1124, 1008, 582, 386, 424, 1133, 1145, 538, 602, 1501, 361, 385, 526, 1069, 838, 1239, 1217, 1107, 524, 629, 1172, 1259, 991, 1266, 1191, 442, 456, 850, 533, 566, 1270, 419, 487, 1001, 1078, 382, 904, 605, 646, 527, 852, 776, 1180, 710, 384, 371, 883, 1184, 677, 1106, 1177, 1096, 1053, 618, 649, 869, 1103, 662, 970, 836, 807, 980, 383, 445, 857, 1181, 763]. Programming-Based [942, 1133, 442, 487]. Programs [201, 1415, 1427, 1601, 1589, 1749, 80, 278, 831, 1319, 1549, 107, 1507, 1902, 91, 414, 1965, 1492, 77, 2040, 1566, 1350, 266, 1383, 1989, 1699, 120, 1956, 1681, 106, 11, 1722, 2014, 110, 1672, 1422, 1420, 1885, 1605, 1703, 1874, 1513, 1585, 962, 1317, 1964, 926, 1268, 589, 422, 902, 976, 329, 1234, 401, 410, 1055, 381, 1097, 1122, 997, 483, 964, 1162, 707, 649, 697, 1160, 1224, 807, 925, 864, 1236, 1141]. Progress [192, 442]. Progressive [1664, 1296, 1205]. Project [1068, 1470, 1599, 451, 1795, 2062, 691, 620]. Project-Scheduling [691]. Projection [1315, 674]. projection-based [674]. Projections [1961, 1196]. Projective [130, 1776]. Promise [218]. Promising [1357]. Promotional [2093]. Proofs [14]. Propagation [1596, 295, 440, 1964, 1212, 620]. Proper [1936]. Properties [481, 1476, 385, 1991, 1551, 493, 909, 795, 1170]. proportional [1109]. Propositional [233]. Propositionalization [958]. Pros [1764]. Prospect [2068]. Prostate [1371]. Protection [1880, 755, 918, 388]. Protective [1777]. Protein [603, 2074, 1408, 1145, 604]. Proteins [1408]. prototype [749, 333]. Prototyping [1506]. Proud [193]. Provably [1492, 1886, 643]. Providing [1226]. provisioning [1228]. Proxy [702]. Pruning [891, 701, 496, 88, 832]. Pseudo [1424, 1498]. Pseudo-polynomial [1424]. Pseudo-Random [1498]. PTSD [1942]. Public [1476, 1618]. Publishing [1342]. Pump [1372, 1978]. Pure [1889, 1372, 899, 599]. purpose [641, 1091]. Pursuit [2042]. py [1816]. py-irt [1816]. pyJedAI [2050]. PyMOSO [1488]. PyNumero [1848]. Python [1865, 1972, 1973, 1816]. PyVRP [1971].

Q [795]. Q-learning [795]. QAP [1738]. QHDOPT [2024]. QoS [746]. QoS-Capable [746]. Quad [315]. Quad-Trees [315]. Quadratic [766, 1442, 2004, 1394, 1461, 246, 1341, 415, 165, 1589, 1412, 1434, 1592, 1624, 1332, 1433, 1562, 1231, 127, 2088, 1985, 31, 1564, 1422, 1420, 1513, 1582, 481, 1064, 582, 1283, 896, 696, 633, 552, 1118, 421, 1107, 1305, 999, 436, 1199, 733, 785, 807, 1141, 1181]. quadratic-cut [633]. quadratic-form [481]. Quadratically [1527, 1535, 1513]. Quality [1030, 1874, 1267, 1006, 1277, 1021]. Quantifying [1100, 1021]. Quantile [1543, 1980, 1787, 1517, 1656, 1644]. Quantile-Based [1656]. Quantization [362]. Quantum [1851, 2018, 2023, 2025, 2021, 2027, 2019, 2026, 2024, 2022, 2085, 1833, 1636, 2020]. Quantum-Inspired [2027]. Quasi [946, 1714, 1466, 1712, 231, 865]. Quasi-Birth [1712]. Quasi-Birth-Death [946]. Quasi-Clique [1466]. quasi-Monte [865]. Quasi-Newton [231]. Quay [1945]. QUBO [1352]. Query [68, 188, 1157, 510]. Queue [1959, 244, 49, 121, 1711, 718, 349, 678, 622, 813, 1306, 579]. Queueing [1769, 54, 131, 1286, 986, 1090, 591, 590, 1712, 1272, 560, 1076, 726, 1264, 869, 497, 465]. Queueing-Time [131]. Queues [131, 520, 1059, 462, 89, 289, 1099, 1318, 1613, 485, 1227, 1054, 1306, 717]. Quickly [42, 65].

R [1488]. R-MinRLE [1488]. R-PERLE [1488]. Rachford [2074, 1738]. radial [753, 859]. Radiation [833, 845, 1732, 1398, 927, 1207, 741]. radio [803]. Radiosurgery [1830]. Radiotherapy [1684, 1478, 1872, 870, 1290]. Radius [222, 974, 1626]. Ramping [1361]. RANDMOD [82]. Random [208, 1498, 1944, 1507, 802, 944, 1426, 1841, 1560, 640, 1667, 1826, 495, 1495, 661, 723, 339, 1126, 889, 592, 1170, 507]. Random-Variate [640]. Randomization [1441]. Randomize [1754]. Randomized [1631, 1332, 1321, 1828, 958, 1057]. Randomizing [82]. Rank [1589, 1670, 1785, 14, 1406, 1564, 1932, 1114]. Rank-One [1564]. Ranking [2033, 1713, 2059, 500, 1789, 1597, 1932, 1218]. Ranking-and-Selection [1713]. Rapid [1693, 1560, 1056, 1189]. Rapidly [448]. Rare [1481, 1318, 742]. Rare-Event [1318, 742]. ratcatcher [647]. Rates [2009, 1136]. Rating [1440, 1482]. Ratio [2060, 1750, 1572, 1648, 1943, 1049]. Rational [1040]. Ray [1567]. ]ary [397]. Reactive [206, 360, 455, 812, 556]. Real [1369, 448, 1986, 431, 1398, 1742, 1603, 1191]. Real-Time [1369, 448, 1986, 431, 1398, 1742, 1603, 1191]. Reasoning [55, 399]. Rebalancing [2097]. Recognition [2046]. Recognizing [138, 1905, 909]. Recommendation [1521, 1658, 944, 1857, 1869, 2043, 1120, 1218]. recommendations [1125]. Recommender [1440, 993, 1887, 1603, 1289]. Reconfigurable [673, 267]. reconstructing [1158]. Reconstruction [542, 898]. Recourse [1743, 1944, 846, 1681, 1317, 1268, 1277]. recoverable [1156]. Recovery [1682, 1189]. Rectangles [1960]. Rectangular [936, 1057, 1273]. Rectilinear [242, 1016]. recurrences [327]. Recurrent [1900, 687, 1827]. Recursive [1498, 1788, 1428, 1496, 1497, 1029, 339, 1126, 911, 592]. Recycled [1797]. redefinition [377]. redeployment [904]. Redesign [1958]. Reduce [1068, 179, 1917, 1280]. Reduce-and-Split [1068]. Reduced [1926, 906]. Reducing [1477, 1751, 1544, 1489, 1081]. Reduction [222, 931, 1926, 167, 1313, 2061, 89, 1695, 1884, 1947, 1725, 1874, 1089, 457, 381, 1266, 674, 733, 865]. Reductions [1450]. Redundancy [1457, 1458]. Referees [469, 561, 655, 703, 767, 825, 877, 932, 321, 404, 529, 610]. Reference [1837]. Refinement [1967, 2072, 1240, 1152]. Reflected [1516]. Reformulation [1412, 2052, 1283, 1074, 1305, 999, 1270]. reformulation-linearization [1305, 999]. Reformulations [1513, 1160].

Regenerator [1171]. Region [1700, 1661, 1050, 533]. Regions [139, 53, 1993, 1998]. Registration [81, 775]. Regression [1519, 1128, 2058, 1884, 1870, 1778, 1295, 1258, 1136, 1282]. Regret [2006, 1577, 1763, 449, 651, 1118, 1182, 834]. Regrets [1889]. Regularization [1890, 2082, 1873]. Regularized [2098, 1778]. Reinforcement [1610, 1765, 2037, 842, 1582]. rejection [1005]. Rejoinder [153, 398, 174, 220, 196, 283, 884, 347]. Related [2050, 1649, 1339, 928]. Relating [2014]. Relational [261, 473]. Relations [1533]. Relationship [1406, 544, 263]. Relationship-Based [544]. Relationships [1706]. Relative [1446, 2019, 651]. Relaxation [1413, 1738, 1165, 1256, 1096, 1181, 502]. Relaxations [1461, 1979, 1253, 1840, 1978, 1723, 1009]. Relaxed [1569, 1382]. Relay [1747]. Relays [1375]. Release [1921, 677, 1307, 1219]. Reliability [21, 1640, 1346, 1457, 32, 60, 817, 1038]. Reliable [1154, 1635, 1087, 987, 966, 969, 1163]. Relinking [408, 1422, 635, 731]. relocation [1235]. Remarks [2044]. remedy [1122]. remember [1014]. Reneging [1712]. renewal [875, 1162, 621]. renewal-type [621]. renewals [1153]. Reoptimization [187, 664]. Repeating [178]. repetitive [727]. replacement [1135]. Replenishment [1915, 995, 784]. Replication [1046, 917, 1117]. Replication-Dependent [1046]. repopulation [1207]. Reporting [148, 150]. Reports [1641]. Representation [1728, 103, 212, 55, 474, 1727, 851, 920]. Representations [77, 1946, 464, 477, 1811, 1212, 786]. Representative [310, 899]. Reproducible [2096]. Request [2036]. Requests [1755, 1683]. Requirements [448, 1766, 1654]. Rerouting [1774]. Resampling [1977, 1867]. Rescheduling [765, 2011, 1806]. Research [514, 531, 546, 554, 565, 575, 585, 613, 615, 628, 645, 654, 671, 659, 153, 393, 1851, 2018, 174, 225, 171, 1173, 892, 1833, 214, 474, 406]. Researcher [342, 347]. Reserve [1790]. reserved [533]. reserved-judgment [533]. Resident [262]. Resilient [1480]. Resistant [1477]. Resolution [338, 2050, 1834, 1858, 644, 572]. Resolution-Related [2050]. Resolving [1071]. Resource [2012, 1682, 1926, 1203, 1795, 673, 950, 1577, 1116, 1510, 1695, 1576, 691, 372, 1755, 1878, 620, 692, 356, 1274, 663]. Resource-Constrained [1795, 691, 1116, 620, 1274]. Resource-Window [1926]. resources [981, 1031]. Respondent [918]. Respondent-Defined [918]. Responder [2027]. Response [1475, 1478, 833, 1816, 1680, 1710, 1050, 463, 674]. Response-Adaptive [1475, 1680]. response-surface [1050]. Responsible [1864]. responsive [617]. restart [1049]. Restless [1409]. Restoration [1347, 1829, 486]. Restricted [1738, 475]. Restrictions [486, 420]. result [874, 961]. Results [535, 108, 230, 1514, 1871, 290, 1510, 1331, 407, 482, 1281, 324, 804, 1132, 1104]. Retail [2093]. Retailers [1603]. retailing [681]. Retrial [121, 718]. Retrieval [688]. Retrospective [1487, 1690]. revenue [1123, 1103, 820, 1041, 1265]. revenues [1172]. Reversal [537]. reversals [489]. Review [2039, 37, 637, 93, 83, 113, 1520, 808, 868, 406]. Reviewers [982, 1032, 1094, 1155, 1208, 1263, 1312, 1364, 1416, 1493, 1607, 1812, 1912, 2028, 2046]. Reviews [72, 1518, 1958, 63, 1077, 1169, 1520, 1226]. Revised [248, 463]. revisited [1221]. Revisiting [1838]. Revolution [517]. Reward [1572]. Reward-Risk [1572]. Rich [1972]. Ride [2041, 1008, 577]. Right [1670, 1605, 1703]. Right-Hand [1670, 1605, 1703]. Ring [464, 304]. Ripple [1821]. Risk [1620, 1442, 797, 1348, 1343, 1319, 1344, 1598, 2061, 1572, 1443, 2080, 1086, 1717, 1561, 1544, 427, 1828, 1245, 1870, 2086, 1703, 1080, 1217, 1122, 1151, 1284, 1184, 1236]. Risk-Adjustable [1703]. Risk-Averse [1319, 1086, 1717, 1561, 1828, 1245, 2086, 1217, 1284, 1236]. Risk-Based [1870]. Risks [1575, 1826]. RNA [1362, 1760]. road [998]. Robert [2]. Robot [2092]. Robust [940, 1821, 1476, 1573, 1662, 1743, 1371, 1967, 2055, 1392, 1446, 1601, 1557, 1865, 1716, 1734, 1599, 1631, 1017, 1512, 1906, 1944, 2094, 1750, 2032, 1794, 2040, 2056, 1062, 1737, 1919, 2042, 1626, 1717, 1779, 1888, 1966, 1993, 1624, 1433, 1220, 1655, 1190, 1776, 2052, 1732, 1545, 1526, 1489, 1882, 1828, 1736, 1019, 1460, 1808, 1685, 1602, 1786, 1458, 1963, 1523, 1805, 1356, 1703, 1903, 1826, 1757, 1623, 1156, 1285, 886, 1118, 1247, 1104]. Robustness [54]. ROC [1786]. Role [152, 1152, 644, 527]. Rolling [1806]. Room [1600, 1298, 1606, 951, 1239, 1084]. Rootfinding [54]. Rostering [1063]. rotamers [1145]. Rotating [1704]. Rotation [175]. Rounded [1975]. Rounding [46, 1321, 602]. Roundoff [1200]. Roundoff-error-free [1200]. Route [1583, 1887, 198, 118]. routed [759]. routines [649]. Routing [509, 992, 672, 1926, 1915, 217, 505, 1149, 360, 1373, 1534, 1972, 1358, 1349, 1531, 957, 1405, 1819, 2000, 1843, 1854, 1973, 237, 1970, 223, 1891, 34, 1818, 128, 1817, 1637, 1336, 1532, 1522, 288, 1974, 236, 62, 1721, 1436, 296, 297, 1853, 118, 1460, 1633, 1742, 2066, 1667, 1697, 1102, 817, 1287, 364, 556, 1101, 998, 1297, 848, 794, 906, 418, 420, 747, 1199, 1047, 1254, 1150, 1294, 868, 699, 1307, 784, 555]. RoutingBlocks [1973]. Row [1111, 24, 821, 977]. Rows [1678, 178]. RSOME [1865]. Rule [1957, 1944, 1902, 1788, 1807, 101, 1708, 1280, 653]. Rule-Based [101, 1280, 653]. Rules [1713, 1526, 210, 1125, 773, 1028]. Rumor [1429]. Rumor-Spreading [1429]. Run [1445]. Rural [1579, 409]. RWA [1774].

s [1272, 1090, 121]. Saddle [2052]. Saddle-Point [2052]. Safe [876, 1083, 1297]. Salary [1759]. Sales [1709, 307, 2070, 2093]. Salesman [140, 1979, 1915, 1484, 1022, 2089, 1809, 1037, 1385, 182, 104, 123, 294, 536, 1009, 1124, 824, 764, 740, 1260, 994, 1023, 373, 1058, 714, 434]. SALOME [353]. Sample [1441, 1492, 1908, 1028]. Sample-Based [1492]. Samples [2091]. Sampling [1427, 1326, 1481, 1840, 1775, 1264, 1223, 1517, 1643, 1656, 1898, 888, 716, 1005, 1072, 1271, 1180, 1225]. sandwich [971]. Sarcasm [2044]. SARS [597]. SAT [1938]. Satisfaction [477]. Satisfiability [233, 387, 923]. Satisficing [2041]. Saves [111]. Say [76]. Scalable [1684, 1701, 1610, 1661, 1497, 1441, 1747, 1816, 1848, 1320, 925, 1289]. Scale [745, 766, 2087, 746, 1664, 213, 1762, 1789, 2042, 1974, 155, 1577, 59, 1827, 284, 1870, 1548, 1874, 1645, 632, 337, 376, 938, 1248, 1288, 487, 382, 494, 1189, 663, 818, 1192]. Scaling [1766, 115, 336, 1308]. Scaling-up [1766]. Scatter [587, 775, 1541, 830, 738, 625, 715]. Scattered [1915, 2000]. Scenario [1427, 1069, 1319, 1549, 1344, 2061, 848, 1526, 1469, 697, 1725, 1874, 1184]. Scenarios [1512]. scene [1121]. Schedule [1921, 922]. Scheduled [1374]. Scheduler [392]. schedules [1207]. Scheduling [205, 79, 2012, 1967, 1704, 1926, 217, 111, 1553, 180, 2039, 1324, 254, 1758, 1906, 1410, 479, 1762, 50, 1600, 255, 258, 1687, 275, 1795, 2062, 1782, 318, 1524, 1679, 320, 355, 1363, 2013, 1981, 863, 1673, 189, 673, 569, 241, 573, 783, 1871, 1655, 164, 1994, 1423, 1718, 1390, 295, 1702, 1872, 179, 1934, 1298, 1796, 276, 1945, 204, 1395, 1978, 309, 1988, 2066, 1356, 1570, 1772, 1619, 691, 943, 166, 669, 828, 632, 935, 1302, 1139, 1211, 411, 503, 1067, 391, 798, 620, 1239, 1042, 858, 772, 1175, 453, 330, 482, 906, 504, 949, 803, 808, 390]. scheduling [1233, 619, 677, 626, 711, 1219, 1084, 468, 583, 840, 818, 712, 1215, 511, 507, 450]. schema [629]. Scheme [2087, 1890, 1589, 569, 1466, 1780, 1963, 1490, 818, 423, 445]. Schemes [1492, 2094]. Schrage [498]. Science [1864, 393, 1833]. Sciences [531, 546, 554, 565, 575, 585, 613, 615, 628, 645, 654, 659]. SCIP [1314]. Scope [1624]. scores [600]. scoring [604]. Screening [1476, 1620, 1411, 1942, 1056, 919]. SDDP.jl [1505]. SDP [1739, 1582]. SDP-Based [1582]. Seamless [1934]. Search [205, 201, 745, 766, 206, 1048, 425, 587, 1195, 1204, 1063, 360, 754, 2057, 328, 92, 1498, 108, 1357, 549, 720, 51, 19, 30, 274, 270, 1628, 357, 102, 1448, 1866, 2077, 863, 2065, 1541, 830, 1974, 1647, 1372, 1668, 1379, 296, 297, 1958, 812, 31, 276, 204, 309, 1837, 1422, 1660, 1634, 2045, 1057, 1323, 1917, 1932, 843, 806, 928, 935, 1008, 824, 556, 625, 623, 778, 764, 775, 459, 391, 1027, 901, 1172, 460, 1248, 638, 1175, 758, 1055, 723, 463, 794, 588, 760, 419, 436, 405, 1284, 715, 990, 1000, 711, 784, 555, 738, 1082, 507, 1176]. Search-Based [2077, 794]. search-enhanced [806]. search-space [507]. Searching [598, 592]. Seasonal [1337]. Seating [1394]. Second [1769, 892, 1790, 1623]. Second-Order [1769, 1623]. Second-Price [1790]. Secondary [1760, 198]. Section [1970, 82]. Sector [1830]. secure [1163]. Security [1187, 780, 1548, 1881, 1303]. Security-Constrained [1548, 1303]. See [1835]. Segment [143]. Segmental [684]. Segmentation [686]. Seidel [75]. Selecting [306, 1543, 819, 453, 157, 1898]. Selection [1360, 1701, 379, 1661, 2071, 1791, 1713, 1344, 1883, 1537, 1178, 1789, 944, 1761, 2098, 2065, 642, 1223, 1517, 427, 198, 1780, 1597, 1884, 1430, 1329, 860, 1520, 1822, 2096, 1907, 1145, 749, 1069, 905, 500, 1258, 1220, 1132, 1296]. Selective [62]. Self [1614, 2036, 1070, 1718, 292, 2069, 697]. self-dual [697]. Self-Exciting [2069]. Self-Interested [2036]. Self-Learning [1614]. Self-organized [1070]. Self-Organizing [292]. Self-Scheduling [1718]. Selling [1825]. Semantic [1073, 1210, 1127]. Semantics [378, 431]. Semantics-Based [431]. Semicontinuous [1513]. Semidefinite [1540, 1838, 1623, 821, 1145, 602, 456, 1096, 1141, 1181]. semismooth [494].

Semisupervised [1860]. Sensitive [358, 1112, 1193]. Sensitivities [1598, 1151]. Sensitivity [1779, 17, 432, 674]. Sensor [1615, 1730, 448, 1923, 1767, 1594, 1238]. Sensor-Intensive [1767]. Sentiment [1134]. Separability [1333]. Separable [1341, 165, 1965, 1576, 1097, 883]. Separation [1064, 652, 1678, 1592, 1975, 2006, 996]. Sepsis [1733]. Sequence [92, 1547, 252, 598, 1428, 725, 607, 606, 1010]. Sequence-Dependent [252]. Sequences [570]. Sequencing [1682, 208, 587, 1533, 252, 1362, 1130, 959]. Sequential [845, 1967, 1360, 165, 888, 1234, 98, 1471, 1887, 948, 1331, 460, 479, 1105, 919]. Series [21, 286, 849, 1946, 185, 1905, 810, 326, 1106]. Series-Parallel [21, 1905]. Server [54, 131, 1099, 1318, 815, 1042, 1306, 624]. Service [131, 238, 1766, 2036, 1679, 690, 1228, 1273, 726, 1227, 1189]. service-level-approximation [726]. Services [1524, 1374, 730, 349, 678, 636]. Session [542]. Set [1415, 1427, 1670, 1611, 1367, 1966, 1916, 929, 1655, 234, 214, 1693, 1550, 1777, 399, 1748, 1508, 1327, 1328, 2081, 996, 1138, 418, 874, 961, 839, 487, 1110, 665, 1247, 911, 1108]. Set-Covering [1916, 839]. set-covering-based [665]. set-mining [1110]. set-partitioning-based [418]. Sets [1918, 1957, 1960, 1601, 1168, 242, 80, 466, 2015, 1413, 2090, 1638, 1882, 1708, 1660, 894, 1072, 971, 650]. Setting [892, 1790]. Settings [1331]. Settlement [1718]. Setup [709, 678, 1241, 1185]. Setups [252, 2075, 1630, 1215]. Several [1024, 1088]. Shabbir [1403]. Shaped [1242, 1843, 1925]. Shared [1314, 1116]. Sharing [1835, 2097]. SHEDR [1658]. Sheet [306]. shelf [777]. shelf-space [777]. Shift [1410, 1067]. Shilling [993]. Shipping [1483]. Shop [205, 79, 180, 1871, 276, 204, 309, 639, 316, 935, 1175, 1001, 468, 507]. shop-scheduling [468]. shopping [1045]. Shops [295, 479]. short [772]. Shortening [162]. Shortest [1438, 787, 755, 1348, 1025, 1565, 1675, 1434, 1616, 1245, 376, 375, 692, 968]. Shortest-Path [1438, 787, 1025, 692, 968]. Shortest-Paths [755, 375]. Shortfall [1525]. show [862]. Showcasing [1793]. Shrinkage [1977]. shutdown [949]. sibling [898, 1158]. Side [2074, 1670, 1748, 1703, 1621, 602, 534]. Side-Chain [2074]. Sides [1605]. Significance [541]. Silver [172]. Similarity [210, 1722, 1158]. Similarity-Based [210]. SimOpt [1847]. Simple [154, 822, 1485, 1711, 915, 1026, 758, 440, 1014]. Simplex [129, 444, 175, 1961, 272, 107, 1402, 459, 976, 463, 334, 576, 1034, 1206]. simplex-based [459]. Simulated [1326, 866, 96, 46, 87, 132, 407]. Simulating [368, 2080]. Simulation [169, 514, 1046, 227, 170, 931, 67, 746, 819, 1487, 1488, 338, 1839, 1847, 1357, 2060, 517, 716, 168, 174, 515, 1782, 1367, 1750, 1787, 1443, 516, 225, 637, 280, 2054, 1841, 171, 1400, 1386, 1318, 279, 283, 1751, 1544, 281, 226, 214, 90, 2088, 231, 1194, 1560, 156, 1411, 1643, 11, 493, 956, 1056, 173, 1644, 1827, 1858, 823, 370, 1323, 737, 843, 989, 1100, 810, 1050, 1017, 513, 875, 500, 463, 1072, 559, 722, 919, 1043].

Simulation-Based [1787, 1194]. Simulation-Optimization [1839, 1847]. Simulations [17, 481]. Simultaneous [858, 650, 1884]. Simultaneously [860]. Single [828, 1404, 1553, 54, 131, 1324, 1445, 1344, 1531, 258, 2000, 318, 61, 1818, 1318, 590, 1702, 86, 1370, 1907, 708, 821, 632, 996, 1211, 774, 503, 938, 391, 1150, 834, 1219, 423, 502]. single- [774]. Single-Commodity [1818]. single-facility [834]. Single-Machine [1324, 318, 828, 632, 503, 502]. Single-Period [1344]. Single-Picker [1531, 2000]. single-row [821]. Single-Run [1445]. Single-Server [54, 131, 1318]. single-sourcing [708]. Singly [120]. Size [1766, 1666, 1886, 666]. Sizes [306]. Sizing [781, 1909, 1146, 1671, 1801, 1345, 1329, 1634, 1261, 1241, 837, 1296, 1185]. Skart [956]. Skewness [956]. Skewness- [956]. Slack [355]. Slicing [686, 1342]. SM [333]. Small [1670, 422]. SMART [1031]. SMOOTH [1065]. Smoothing [354, 1065, 785]. SNP [600]. SOCEMO [1299]. Social [1864, 1574, 1692, 1439, 1824, 1586, 1658, 1997, 2058, 1654, 1693, 1463, 1591, 1027]. Socially [1991]. Soft [448, 1742]. Software [1895, 1488, 1921, 1970, 2024, 673, 2049, 1777, 1019, 857, 82]. Solar [1532]. Solution [509, 268, 269, 415, 1980, 293, 1664, 18, 713, 1972, 1349, 2000, 6, 1266, 365, 1761, 1779, 1899, 45, 1729, 1391, 1536, 733, 1663, 1873, 1686, 962, 457, 325, 941, 458, 375, 1249, 432, 813, 991, 794, 581, 1294]. Solutions [1918, 1415, 1476, 1743, 535, 1631, 1886, 1535, 289, 1625, 115, 1391, 756, 1374, 1309, 821, 829, 632, 439, 1002, 498, 734, 605, 1201, 727, 621]. Solvable [1766]. Solve [218, 46, 1981, 881, 1548, 1917, 1011, 1305, 502]. solved [534]. Solver [1583, 1332, 571, 1739, 1971, 926, 1301]. Solvers [2031, 252, 738, 1141]. Solving [1845, 1959, 1485, 400, 1133, 1438, 411, 133, 1773, 278, 1412, 2031, 374, 107, 974, 1933, 1333, 1762, 1175, 1965, 1455, 275, 1789, 1434, 1794, 837, 2040, 1566, 456, 850, 1383, 2054, 1989, 334, 1214, 929, 548, 1162, 1379, 1671, 1956, 1579, 724, 1691, 792, 1938, 106, 95, 912, 1167, 1420, 860, 1792, 1874, 1630, 943, 1317, 1087, 1009, 582, 1246, 1200, 901, 723, 1081, 1255, 710, 677, 649, 869, 1764]. Some [1840, 35, 108, 1313, 290, 66, 1331, 385, 327]. Sorting [489, 1539]. SOS [1739]. SOS-SDP [1739]. Source [1865, 1973, 1583, 1336, 2049, 771]. Sources [1134]. Sourcing [1900, 1826, 708]. Space [1979, 1195, 1204, 369, 1405, 1313, 1202, 2026, 908, 1699, 1419, 276, 1323, 777, 1009, 760, 805, 507]. Space- [1405]. Spanning [1381, 303, 1397, 285, 1166, 1595, 1593, 689, 638, 377, 1230, 871, 1177, 809]. Spare [486, 1346, 417]. Sparse [47, 2055, 1519, 1701, 2073, 1664, 1791, 1965, 1535, 1362, 2053, 45, 1536, 371, 106, 59, 1167, 1917, 643, 844, 402]. Sparsity [1961, 1677]. Sparsity-Exploiting [1961]. Spatial [1359, 1883, 1244]. Spatio [2069]. Spatio-Temporal [2069]. Spatiobiologically [1478]. Spatiotemporal [1684, 1611]. Spatiotemporally [1290]. Special [671, 1864, 1610, 683, 522, 1501, 854, 1851, 2018, 253, 109, 1970, 1502, 1428, 1098, 596, 1126]. specialties [1084]. specified [495]. spectra [1183]. Spectral [250, 1426, 339, 722]. Speed [1358, 68, 1308]. Speeding [787, 1645]. Speeds [316]. speedup [1049]. Speedups [151]. Spheres [1866].

Spiral [686]. Splicing [1822]. Split [1068, 1093, 1891, 1818, 457]. Splitting [2004, 1195, 1481, 2074, 278, 1738, 464, 1038, 380, 1247]. spoke [905]. Sponsored [720]. Spot [1899]. Spreading [1429]. SQLMP [103]. square [814]. Squares [917, 2080, 1739, 1938, 81]. SSPMO [715]. SSS [214]. Stability [2071, 251, 1811, 1218, 402]. Stabilization [1335, 973]. Stabilized [1678, 1861, 1221]. Stabilizing [1950]. Stable [1670, 954, 1002, 1256]. Stacking [2070]. Stacks [1037, 709, 763]. Staffing [1351, 1842]. Stage [1573, 1662, 2035, 1749, 278, 1512, 1632, 1944, 1507, 2025, 2056, 1395, 1757, 1317, 731, 1243, 1257, 1234, 997, 1281, 747, 1150, 1277, 1104, 864]. stage-unrestricted [1257]. Staged [265, 200, 789, 662]. StAMPL [846]. Star [1617, 301]. Stars [1408]. start [649]. State [1046, 332, 1979, 131, 369, 1632, 271, 178, 159, 49, 89, 637, 191, 182, 868, 1462, 1009, 500, 722, 674]. State-of-the [868]. State-of-the-Art [159, 637, 182]. State-of-the-Art-Survey [332]. State-Space [1979, 1009]. State-Variable [1632]. States [2073, 883]. Static [1038, 2097, 1561, 23]. Station [1556, 690, 1892]. Stationary [1696, 1911, 245, 1428, 1272]. stations [452]. Statistical [755, 443, 1646, 1157, 1677]. Statistically [187]. Statistics [244, 1400, 508, 661]. Status [4]. Stay [1622]. Steady [1046, 131, 178, 49, 89, 500, 722, 674]. Steady-State [1046, 131, 178, 49, 500, 722, 674]. Steiner [314, 133, 1172, 377, 301, 1164, 1339, 87, 1663, 518, 609]. Step [158, 902]. stepwise [1226]. Stereotactic [1830]. Stiff [12]. Stochastic [1427, 1878, 269, 1756, 2087, 1890, 1724, 332, 1050, 1749, 1865, 1895, 1538, 369, 1445, 278, 1696, 1774, 1319, 1632, 1505, 1911, 1758, 271, 1551, 846, 1507, 262, 1405, 1600, 451, 1492, 1843, 1787, 1524, 2032, 313, 311, 1552, 223, 1841, 1717, 1779, 248, 1561, 1718, 571, 881, 1671, 1320, 1542, 1084, 1801, 1345, 25, 110, 240, 1331, 1374, 1947, 1653, 1976, 2086, 1605, 1578, 1792, 1874, 1323, 1559, 1338, 1317, 1121, 593, 981, 824, 1268, 1252, 1069, 1217, 1042, 1128, 1234, 381, 1301, 463, 848, 997, 594, 1284, 1162, 729, 1180, 884, 1031, 753, 859, 968, 990, 1184, 1177, 697]. stochastic [662, 820, 791, 882, 836, 925, 413, 864, 1236]. stochastically [1072]. Stock [866, 535, 1424, 135, 893, 431, 1861, 1776, 1727, 1806, 1999, 731, 709, 973, 439, 498, 1296]. Stock-Cutting [866]. Stockout [681]. Stopping [299, 1713, 762, 1028]. Stops [1973]. Storage [1334, 1915, 682, 2000, 1320, 1191]. stores [1196]. Straight [408]. Strategic [1923, 1422]. Strategically [1491]. Strategies [2057, 1795, 429, 382, 1828, 1853, 1656, 276, 602, 421, 419, 1127]. Strategy [1369, 1477, 51, 1378, 952, 1142, 832]. stratification [600]. streams [960]. Strength [1111, 1802]. Strengthened [1268, 1738]. string [605]. Strip [936, 551, 1057]. Strip-Packing [551, 1057]. Strong [1165, 1275, 1050]. stronger [1243]. Structural [1476, 1027, 1722, 795, 1174]. Structure [305, 1512, 608, 109, 1350, 1652, 1362, 1798, 1760, 1677, 412, 604, 1126, 1304, 1199, 903, 727, 1159]. Structured [359, 232, 2051, 11, 1602, 507, 333]. Structures [10, 253, 2050, 101, 1016, 327, 852]. Studies [175, 1132]. Study [205, 728, 79, 746, 1348, 1791, 278, 160, 233, 1921, 428, 1164, 1332, 1810, 1872, 639, 1588, 475, 824, 1080, 664, 742, 1006, 850, 419, 964, 793, 626]. subgradient [818]. Subgraph [1649]. Subgroups [1708]. Subject [471, 1219]. Subjective [1797]. Sublanguage [103]. Submatrix [912]. Submodular [1808, 1935, 1770, 1602, 1219, 1308]. Submodularity [2054]. Subnetwork [314]. Suboptimality [2005]. Subproblem [1715]. Subproblems [2082, 1873]. Subquadratic [2089]. Subrouting [214]. Subset [1755, 819, 1791, 1537, 1520, 1822, 749, 335, 1304, 1132]. Substations [2035]. Substitution [1346, 1801, 1763, 1578]. Substitution-Dependent [1346]. Substitutions [1825]. Subsystem [488]. Subsystems [33]. Successive [1442]. Sum [2031, 190, 1698, 1764, 1739, 1938, 791]. Sum-of-Squares [1739]. Summary [1313]. Sup [1743]. Supercomputers [1762]. superior [851, 920]. Supervised [1959, 1639, 1538, 1925, 1098]. Supply [2067, 975, 1321, 1826, 1281]. Support [895, 917, 448, 1901, 184, 524, 961, 949]. Supported [1861]. suppressed [1214]. Suppression [2020]. Surface [833, 1050, 463]. surfaces [1060]. Surgery [1682, 1967, 1734, 1399, 951, 1205]. surgical [1254]. Surrogate [1793, 1357, 1407, 1837, 1299]. Surrogate-Based [1357]. Surveillance [849, 1552, 1229]. Survey [332, 1746, 159, 1584, 182, 550, 842, 726]. Survivability [1095]. survivable [1035, 363]. Survival [170]. Survive [168, 171]. Switched [26]. Switching [1461, 793, 1303]. symbol [1109]. Symbolic [520, 1904]. Symmetric [1438, 557, 294]. Symmetries [1964]. Symmetry [175, 1534, 1161, 837]. symmetry-breaking [837]. Synchronization [679, 97]. Synchronized [1547]. Synchronous [464]. Synergies [1833]. Synergistic [1449]. Synthesis [2021]. Synthetic [862]. System [790, 1360, 746, 400, 359, 1829, 278, 1286, 232, 82, 1726, 993, 1923, 1887, 591, 590, 164, 1457, 101, 399, 5, 177, 186, 568, 1858, 927, 750, 1272, 1495, 905, 460, 1237, 1244]. Systematic [1352]. Systems [1030, 154, 1900, 1480, 1440, 338, 682, 262, 2002, 1687, 43, 85, 975, 1761, 2097, 1637, 1368, 429, 115, 1994, 1379, 1712, 221, 263, 688, 2007, 823, 1603, 1797, 1280, 735, 376, 617, 815, 1200, 1250, 726, 1289, 1228, 1189, 1163, 497, 1192].

Table [1186]. Tableau [558, 1111]. tableaux [977, 964]. Tables [755, 1235]. Taboo [204]. Tabu [201, 806, 206, 587, 360, 328, 92, 108, 19, 30, 274, 1690, 357, 2077, 296, 812, 31, 1422, 1008, 625, 459, 588, 715, 555]. Tabu-Search [357]. Tactical [1625]. Tagging [1558]. Taguchi [1017]. Tail [1344, 639, 2043, 1005]. Tailed [462, 988]. Talent [943]. Tandem [1090]. Tank [434]. Tar [149]. Tardiness [1324, 1059, 318, 61, 1702, 1165, 782, 503, 391, 330, 840, 1165]. Tardy [1962]. Target [1330, 1371, 1734, 150, 1382, 23, 2099]. Target-Oriented [1734]. targeted [773]. Targeting [1477]. Targets [1349]. Tariff [748]. Task [673, 156, 1081]. Tasks [254, 320, 355, 356]. Taxi [1661, 1887]. Taxonomy [328]. Tchebycheff [649]. TDMA [455]. Team [1574, 812, 1805, 1627]. Team-Learning [812]. Teamwork [1291]. Technical [92, 1286, 498, 124, 81]. Technique [293, 1412, 350, 68, 1305, 999]. Technique-Based [1412, 999].

Techniques [10, 1980, 1335, 235, 204, 1420, 1964, 1060, 527, 710, 865]. Technologies [395, 571]. technology [1031]. Telecommunication [26, 364, 695, 420]. Telecommunications [1095, 417, 636]. Televised [1586]. Temperature [132]. Template [1636, 384]. Template-Based [1636]. Temporal [2073, 2069, 1074]. Tensor [1709, 1376, 791]. Terminal [21]. Terminals [1533]. Terminator [2040]. Terms [900]. Terrains [124]. Territory [1024, 1088]. terrorism [930]. TES [250, 100]. Test [1793, 2034, 900, 290, 989, 593, 538]. Testbed [1847]. Testing [310, 1476, 153, 1140, 38, 594, 972, 996]. Tests [1620, 443, 1646, 339]. Text [348, 1641]. Text-Embedding [1641]. Textual [1229]. TGVx [1869]. Their [462, 1857, 385, 1947]. theorems [732]. Theoretic [1137, 370, 1237]. Theoretical [1478, 1785, 290, 725, 1039, 1206]. Theoretically [1943]. Theory [1615, 2068, 197, 513, 141, 223, 1816, 17, 264, 194, 407, 1105, 580]. Theory-Based [264]. Therapist [1679]. Therapy [833, 1732, 927, 1207]. Third [1835]. Third-Party [1835]. threading [603]. threat [1157]. Three [180, 137, 549, 635, 641, 801]. Three-Dimensional [137, 549, 641, 801]. three-index [635]. Three-Machine [180]. Threshold [1614, 273, 800, 1141]. Threshold-Based [1614]. Thresholding [274, 1866]. Through-Fleet-Assignment [745]. Throughput [854, 861, 860, 1279]. Tied [1071]. Tier [302]. Tight [1908]. Tightening [1735]. Time [1369, 2012, 782, 1063, 131, 360, 1534, 448, 249, 2037, 942, 1405, 2094, 1565, 1675, 318, 849, 311, 1986, 1679, 908, 431, 237, 1809, 1745, 1946, 1099, 1351, 1817, 1429, 358, 1090, 2091, 296, 297, 118, 1613, 185, 1398, 1905, 1587, 1742, 823, 1806, 1627, 1603, 1697, 962, 450, 914, 475, 1009, 810, 556, 485, 994, 998, 407, 1191, 1054, 961, 747, 714, 1294, 1306, 743, 619, 990, 1106, 579, 717, 662, 882, 413, 445]. Time-Correlated [1405]. Time-Dependent [1565, 1675, 1679, 1809, 1817, 1090, 1613, 998]. Time-Expanded [942, 2091]. Time-indexed [782, 450]. Time-Resource [2012]. Time-Sensitive [358]. Time-Series [849, 810, 1106]. Time-Space [908]. time-staged [662]. Time-Variant [1806]. Time-Varying [2037, 1099, 1351, 1627, 1054, 1306]. Times [341, 1405, 462, 1356, 678, 1241, 1219, 1185]. Timetable [1806]. Timetabling [291, 892, 1262, 778]. Together [216]. Tolerance [1988]. Tolerant [476, 1168, 1803, 1328]. Tomahto [76]. Tomato [76]. Tonal [686, 685]. Tool [846, 347, 582]. Toolbox [1768]. Toolkit [1721]. Tools [1839, 52, 219, 1970, 883]. Top [1521, 816, 1898, 1218]. Top- [1898, 1521, 816]. top-K [1218]. Topic [1586]. Topologies [22, 915, 366]. topology [507]. Toroidal [288]. Total [1324, 258, 61, 2082, 1702, 1873, 782, 503, 391]. totally [1277]. Tour [547, 119, 1233]. Tracing [1911]. Tracking [2099]. tractable [1281]. Trade [2012, 1006, 1021]. Trade-off [2012, 1021]. trade-offs [1006]. Tradeoffs [52]. Trading [431]. Traffic [400, 2073, 17, 759, 455]. traffic-grooming [759]. Trained [1998, 1859]. Training [2055, 1648, 186]. Trajectories [790, 274]. Transaction [944, 431]. Transaction-Based [944]. Transactional [341, 1342]. Transfer [2037]. Transferable [1404]. Transform [1396].

Transformation [350]. Transformations [1940, 601, 985]. Transforming [875]. Transforms [239, 319, 739, 403, 438, 694, 336]. Transient [1441, 986, 1462, 1089, 979, 813]. Transient-State [1462]. Transients [1046]. Transit [1376, 2002, 617]. Transition [2051]. Transitions [42, 65, 1781]. translation [581]. transmembrane [604]. Transmission [931, 1477, 1461, 1747, 2007, 1462, 1303]. Transparent [365, 856]. Transport [1333, 1937, 1885]. Transportation [1551, 1483, 99, 2091, 1934, 1374, 376, 772]. Transshipment [1533]. Travel [1405]. Traveler [1216]. Traveling [140, 1979, 1915, 1484, 2089, 1809, 1037, 1385, 182, 104, 123, 294, 536, 1009, 1124, 824, 764, 740, 1260, 994, 1023, 373, 1058, 714, 434]. Treatment [833, 845, 1872, 1732, 1830, 1398, 1708, 870, 741, 1021]. Tree [1381, 1730, 133, 1815, 303, 946, 701, 301, 1164, 1339, 1788, 2065, 302, 285, 1663, 1166, 1595, 1594, 1593, 960, 774, 361, 901, 885, 1172, 689, 638, 848, 1112, 1110, 930, 436, 852, 871, 1177, 809]. tree-based [848, 1110]. Tree-Like [946]. Tree-Pruning [701]. tree-search [436]. Tree-Star [301]. TreeHugger [900]. Trees [197, 1666, 496, 1571, 358, 609, 315, 666, 1069, 532, 377, 832, 618, 1238]. Trend [1910]. Trial [1475]. Trials [1680, 1621]. Triangle [1195]. Triangular [946]. Truck [1897]. Trunk [223]. Trust [1700, 1998, 1050]. trust-region [1050]. Trustworthy [993]. TSP [480, 1437, 293, 740, 1197]. TSPLIB [104]. TSPs [475]. TSPTW [528]. Tumor [1478, 1330, 1207]. turnaround [949]. tutorial [842]. TV [1586]. Two [308, 21, 1573, 1662, 2035, 1846, 1836, 1749, 1563, 278, 902, 1111, 1512, 1632, 1944, 1507, 50, 2025, 1397, 2056, 2082, 87, 1718, 1150, 200, 86, 1395, 1672, 1697, 1621, 1757, 1317, 316, 1261, 1156, 731, 817, 372, 977, 1243, 1246, 622, 1257, 1259, 479, 1234, 560, 789, 997, 966, 1281, 747, 435, 1018, 710, 1277, 1104, 712, 864]. Two-Armed [1621]. Two-Dimensional [1563, 1397, 1697, 731, 1246, 1257, 1259, 560, 789, 435, 710]. Two-Level [308, 1156]. Two-Machine [316, 479]. Two-Matching [200]. two-moment [622]. two-node [966]. two-path [817]. two-period [1261]. two-resource [372]. Two-Row [1111, 977]. Two-Settlement [1718]. Two-Stage [1573, 1662, 2035, 1749, 278, 1512, 1632, 1944, 1507, 2025, 2056, 1395, 1757, 1317, 1150, 731, 1243, 1234, 997, 1281, 747, 1277, 1104, 864]. two-staged [789]. Two-step [902]. Two-Terminal [21]. Type [462, 946, 929, 1792, 1630, 437, 945, 621]. Type-2 [1630]. Typed [378]. Types [89]. typing [1210].

UAV [1556]. Unbiased [1445]. unbounded [1114, 667]. Uncapacitated [308, 908, 1671, 25, 1311]. Uncertain [2023, 1635, 1823, 519, 1806, 1356, 1623, 1285, 1042, 1190]. Uncertainties [1371, 1782, 1726, 1514]. Uncertainty [1478, 940, 1601, 1855, 1596, 2067, 1750, 1919, 1567, 1966, 1433, 1994, 1671, 474, 1796, 1977, 1460, 1703, 449, 1100, 951, 1205, 1247, 1236, 1148].

Uncommon [1074]. Unconstrained [781]. Uncovering [1710]. Underestimators [1341]. Underground [1795]. underlying [795]. Understanding [1440, 1482]. Undirected [200, 409]. Unforeseen [1690]. Unified [1373, 2001, 694, 794, 465]. Uniform [241, 100, 1495, 508]. Unifying [1591, 386]. unimodular [1277]. Unique [76]. Unit [1412, 1361, 1474, 1567, 1810, 1982, 1622, 1548, 1305, 1303]. Units [1401]. Univariate [1525, 350, 978]. university [1262, 778]. UNIX [392]. Unknown [124]. Unlabeled [1860]. Unmanned [1823]. Unobserved [1790]. Unrelated [50, 1165, 166, 511]. unrestricted [1257]. Unspecified [1366]. Unsupervised [1705]. Unwrapping [1397]. Updates [1927]. Upper [146, 1442, 1484, 2015, 1322, 1064, 997]. Upper-Bounding [1442]. Urban [2073]. Usage [540, 542, 541]. Use [310, 165, 1090, 1982, 1145, 560, 326, 1127]. useful [335]. User [1440, 1482, 1857, 1706, 637]. User-Item [1706]. Using [201, 1618, 845, 122, 931, 153, 426, 582, 1283, 1246, 1900, 1341, 1765, 359, 761, 1895, 2073, 686, 133, 1487, 1991, 1376, 1011, 96, 2015, 377, 1965, 1441, 403, 1658, 1443, 365, 89, 46, 2093, 1363, 2013, 676, 1073, 1892, 2054, 1362, 1173, 115, 127, 710, 881, 1844, 71, 1938, 106, 1408, 2043, 1083, 1703, 1575, 1645, 1797, 1630, 631, 1089, 821, 824, 337, 526, 1069, 858, 421, 844, 1305, 453, 1151, 1266, 837, 1191, 636, 1214, 1289, 1058, 1018, 324, 859, 1021, 978, 618, 348]. usually [1026]. UTVPI [923].

vacation [1272]. Vaccine [1337, 1108]. Valid [1381, 1735, 1053, 366]. Validation [725, 944, 1646]. Valuable [347]. Valuation [35, 221]. Valuation-Based [221]. Value [1343, 1631, 1410, 15, 2080, 1539, 358, 1828, 1374, 2086, 888, 1122, 978, 1041]. Value-at-Risk [2080, 1122]. Value-Driven [1539]. Valued [2037]. values [327, 440]. valve [915]. VaR [1252]. Variable [1063, 1765, 764, 1735, 1632, 350, 2040, 1379, 1884, 675, 996, 556, 377, 758]. Variables [271, 1344, 1861, 1210, 661, 1081]. Variance [1360, 1445, 2025, 381, 1947, 481, 737, 729, 423, 1010]. variance-based [1010]. Variances [1802]. Variant [1806]. Variants [175, 1703, 1175]. Variate [640]. Variates [100]. Variation [558, 2082, 1873]. Variational [2025, 2022, 633, 380]. Varying [2037, 1099, 1351, 1627, 1054, 1306]. Vector [362, 895, 917, 1837, 961, 1170]. Vectors [315, 495, 327, 889]. Vehicle [509, 217, 1533, 360, 1534, 1972, 1358, 1349, 957, 1405, 1819, 1843, 1973, 237, 1583, 1970, 1891, 1892, 34, 1818, 1817, 1974, 236, 1823, 62, 1721, 296, 297, 1853, 118, 1460, 1742, 2066, 1667, 1697, 2095, 556, 1297, 794, 418, 747, 1199, 1047, 1294, 868, 699, 1307, 784, 555]. Vehicle-Routing [1819, 556, 699, 784, 555]. Vehicles [1024, 1235]. Verification [2015]. VeRoViz [1721]. Version [486]. Versions [75]. versus [1092]. Vertex [1167, 1660, 796]. Vertical [1079]. Vertiport [1661]. Very [745, 766, 293, 1167, 1248]. Veterans [1942]. Via [1678, 438, 1100, 1940, 1038, 1481, 379, 165, 1252, 1670, 547, 1889, 1987, 1200, 2060, 1259, 991, 1962, 1397, 1886, 1122, 2040, 2026, 223, 1714, 2080, 2098, 1400, 1946, 1950, 1993, 605, 803, 950, 1544, 1511, 1321, 1391, 1247, 2052, 1956, 1451, 1526, 231, 1736, 1560, 592, 912, 1393, 1056, 1398, 1420, 1043, 1858, 823, 988, 1667, 860, 1642, 1917, 1582]. Victory [1404]. View [345]. Viewers [1997]. violated [699]. Violations [76, 1131]. Viral [917]. Virtual [476, 398, 1441, 1717, 128, 1400, 1479, 364]. Visualization [47, 216, 217, 1538, 215, 219, 1721, 544, 36]. VNS [1379]. Volatility [1641, 1999]. Volume [471, 1371, 1960, 1694, 1315, 1084]. Voronoi [141, 138]. Vote [1404]. Voting [1518]. Voting-Based [1518]. VRP [1971]. VRPSolverEasy [1972]. vs [774, 513, 1717, 1871]. Vulnerability [1425, 1844].

W [690]. W-CDMA [690]. W3C [629]. Wait [1390]. Waiting [341, 1565, 717, 579]. Waiting-time [717]. walks [885]. Warehouse [2070]. Warehouses [1915, 679, 1531]. Warm [649]. Warm-start [649]. Warmstart [1982]. Wasserstein [1716, 2040]. Water [1978, 1236]. Wavelength [581, 759]. wavelength-routed [759]. wavelet [722]. wavelet-based [722]. way [1117]. Ways [686]. WDM [1774, 581, 939]. Weapon [23]. Weapon-Target [23]. web [1502, 624, 393, 754, 720, 395, 719, 573, 540, 542, 543, 397, 541]. web-based [1502]. Web-Catalog [754]. Web-Usage [540, 542, 541]. websites [1045]. Wedelin [890]. Weight [416, 1465, 234, 1328, 2081]. Weighted [2071, 53, 258, 1962, 788, 1702, 1165, 782, 503, 391, 874, 1108, 649, 423]. Weighting [1660]. Weighting-Based [1660]. Weir [1989]. while [1018]. Wide [393, 397, 1142]. width [698]. Will [168, 171]. Willingness [799]. Window [1926, 1534]. Windows [1809, 1817, 1697, 1009, 556, 994, 747, 714, 1294, 619, 2012, 360, 237, 296, 297, 118, 1742]. winner [653]. Wireless [1615, 1730, 1154, 1594, 695, 1015, 1238]. without [1272, 1157]. Wolfe [1768, 1074]. Word [471, 563, 612, 657, 705, 769, 827, 879, 934, 984, 196]. Words [598]. Work [258, 944, 344]. Workers [1915]. Workflow [644]. Workforce [1704]. Workload [354]. Works [1352]. World [393, 637, 397]. Worst [1525, 1747, 34, 1514, 1188]. Worst-Case [1525, 1747, 34, 1514, 1188]. Would [193].

XML [629]. Xpress [850].

Yield [1826].

Z [498]. Zag [1140]. ZDDs [1363]. Zero [201, 1590, 1889, 1764, 77, 1056, 1214]. Zero-One [201, 77, 1056]. Zero-Sum [1764]. zero-suppressed [1214]. Zeroth [2009]. Zig [1140]. Zig-Zag [1140]. Zigzag [1082]. zones [1273]. zoom [5].

References

Greenberg:1989:EIa

[1] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 1(1):1, Winter 1989. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1.1.1.

Anonymous:1989:MRG

[2] Anonymous. In memoriam: Robert G. Jeroslow (1942–1988). ORSA Journal on Computing, 1(1):2–6, Winter 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1.1.2.

Jeroslow:1989:DPI

[3] Robert G. Jeroslow, (deceased) and Jinchang Wang. Dynamic programming, integral polyhedra and Horn clause knowledge base. ORSA Journal on Computing, 1(1):7–19, Winter 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1.1.7.

Zenios:1989:PNO

[4] Stavros A. Zenios. Parallel numerical optimization: Current status and an annotated bibliography. ORSA Journal on Computing, 1(1):20–43, Winter 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1.1.20.

Singhal:1989:FOB

[5] Jaya Singhal, Roy E. Marsten, and Thomas L. Morin. Fixed order branch-and-bound methods for mixed-integer programming: The zoom system. ORSA Journal on Computing, 1(1):44–51, Winter 1989. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1.1.44.

Heyman:1989:NSL

[6] Daniel P. Heyman and Alyson Reeves. Numerical solution of linear equations arising in Markov chain models. ORSA Journal on Computing, 1 (1):52–60, Winter 1989. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1.1.52.

Greenberg:1989:EIb

[7] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 1(2):61, Spring 1989. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1.2.61.

Shahrokhi:1989:AAM

[8] Farhad Shahrokhi. Approximation algorithms for the maximum concurrent flow problem. ORSA Journal on Computing, 1(2):62–69, Spring 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.1.2.62.

McShane:1989:IPD

[9] Kevin A. McShane, Clyde L. Monma, and David Shanno. An implementation of a primal-dual interior point method for linear programming. ORSA Journal on Computing, 1(2):70–83, Spring 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1.2.70.

Adler:1989:DSP

[10] Ilan Adler, Narendra Karmarkar, Mauricio G. C. Resende, and Geraldo Veiga. Data structures and programming techniques for the implementation of Karmarkar’s algorithm. ORSA Journal on Computing, 1(2): 84–106, Spring 1989. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1.2.84.

Som:1989:FDE

[11] Tapas K. Som and Robert G. Sargent. A formal development of event graphs as an aid to structured and efficient simulation programs. ORSA Journal on Computing, 1(2):107–125, Spring 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1.2.107.

Reibman:1989:ASM

[12] Andrew Reibman, Kishor Trivedi, Sanjaya Kumar, and Gianfranco Ciardo. Analysis of stiff Markov chains. ORSA Journal on Computing, 1 (2):126–133, Spring 1989. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1.2.126.

Greenberg:1989:EIc

[13] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 1(3):135–136, Summer 1989. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1.3.135.

Hooker:1989:IPR

[14] John N. Hooker, Jr. Input proofs and rank one cutting planes. ORSA Journal on Computing, 1(3):137–145, Summer 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1.3.137.

Inuiguchi:1989:PLP

[15] Masahiro Inuiguchi, Hidetomo Ichihashi, and Hideo Tanaka. Possibilistic linear programming with measurable multiattribute value functions. ORSA Journal on Computing, 1(3):146–158, Summer 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1.3.146.

Boggs:1989:AEM

[16] Paul T. Boggs, Paul D. Domich, Janet R. Donaldson, and Christoph Witzgall. Algorithmic enhancements to the method of centers for linear programming problems. ORSA Journal on Computing, 1(3):159–171, Summer 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1.3.159.

Simon:1989:NES

[17] Burton Simon. A new estimator of sensitivity measures for simulations based on light traffic theory. ORSA Journal on Computing, 1(3):172–180, Summer 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1.3.172.

Dechter:1989:GSO

[18] Avi Dechter and Rina Dechter. On the greedy solution of ordering problems. ORSA Journal on Computing, 1(3):181–189, Summer 1989. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1.3.181.

Glover:1989:TSP

[19] Fred Glover. Tabu search — Part I. ORSA Journal on Computing, 1(3):190–206, Summer 1989. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1.3.190.

Greenberg:1989:EId

[20] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 1(4):207–208, Fall 1989. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1.4.207.

Aboelfotoh:1989:SPB

[21] Hosam M. Aboelfotoh and Charles J. Colbourn. Series-parallel bounds for the two-terminal reliability problem. ORSA Journal on Computing, 1(4):209–222, Fall 1989. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1.4.209.

Moose:1989:MND

[22] Robert L. Moose, Jr. Modeling networks with dynamic topologies. ORSA Journal on Computing, 1(4):223–231, Fall 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1.4.223.

Wacholder:1989:NNB

[23] Eitan Wacholder. A neural network-based optimization algorithm for the static weapon-target assignment problem. ORSA Journal on Computing, 1(4):232–246, Fall 1989. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1.4.232.

Phillips:1989:AAP

[24] Andrew T. Phillips and J. Ben Rosen. Anomalous acceleration in parallel multiple-cost-row linear programming. ORSA Journal on Computing, 1 (4):247–251, Fall 1989. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1.4.247.

Wallace:1989:PSP

[25] Stein W. Wallace and Roger J-B Wets. Preprocessing in stochastic programming: The case of uncapacitated networks. ORSA Journal on Computing, 1(4):252–270, Fall 1989. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1.4.252.

LeBlanc:1989:CMC

[26] Larry J. LeBlanc and Ronald V. Simmons. Continuous models for capacity design of large packet-switched telecommunication networks. ORSA Journal on Computing, 1(4):271–286, Fall 1989. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1.4.271.

Marsten:1989:IDA

[27] Roy E. Marsten, Matthew J. Saltzman, David F. Shanno, George S. Pierce, and J. F. Ballintijn. Implementation of a dual affine interior point algorithm for linear programming. ORSA Journal on Computing, 1(4):287–297, Fall 1989. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1.4.287.

Greenberg:1990:EIa

[28] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 2(1):1–2, Winter 1990. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2.1.1.

Anonymous:1990:EPW

[29] Anonymous. Expediting processing: What can the author do? ORSA Journal on Computing, 2(1):3, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2.1.3.

Glover:1990:TSP

[30] Fred Glover. Tabu search — Part II. ORSA Journal on Computing, 2 (1):4–32, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2.1.4.

Skorin-Kapov:1990:TSA

[31] Jadranka Skorin-Kapov. Tabu search applied to the quadratic assignment problem. ORSA Journal on Computing, 2(1):33–45, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.2.1.33.

Whited:1990:RCP

[32] David E. Whited, Douglas R. Shier, and James P. Jarvis. Reliability computations for planar networks. ORSA Journal on Computing, 2(1):46–60, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2.1.46.

Gleeson:1990:IMI

[33] John Gleeson and Jennifer Ryan. Identifying minimally infeasible subsystems of inequalities. ORSA Journal on Computing, 2(1):61–63, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.2.1.61.

Li:1990:WCA

[34] Chung-Lun Li and David Simchi-Levi. Worst-case analysis of heuristics for multidepot capacitated vehicle routing problems. ORSA Journal on Computing, 2(1):64–73, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2.1.64.

Delgado:1990:VOP

[35] Miguel Delgado, Jose L. Verdegay, and Maria-Amparo Vila. On valuation and optimization problems in fuzzy graphs: A general approach and some particular cases. ORSA Journal on Computing, 2(1):74–83, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.2.1.74.

Tuchman:1990:MVD

[36] Allan M. Tuchman and Michael W. Berry. Matrix visualization in the design of numerical algorithms. ORSA Journal on Computing, 2(1):84–92, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2.1.84.

Jones:1990:BR

[37] Richard H. Jones. Book review. ORSA Journal on Computing, 2(1):93, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2.1.93.

Greenberg:1990:CTW

[38] Harvey J. Greenberg. Computational testing: Why, how and how much. ORSA Journal on Computing, 2(1):94–97, Winter 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2.1.94.

Greenberg:1990:EIb

[39] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 2(2):99, Spring 1990. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2.2.99.

Barker:1990:NNI

[40] Allen L. Barker, Donald E. Brown, and Worthy N. Martin. A neural network implementation of a data association algorithm. ORSA Journal on Computing, 2(2):100–111, Spring 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2.2.100.

Zenios:1990:MBM

[41] Stavros A. Zenios. Matrix balancing on a massively parallel connection machine. ORSA Journal on Computing, 2(2):112–125, Spring 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.2.2.112.

Fox:1990:GMC

[42] Bennett L. Fox. Generating Markov-chain transitions quickly: I. ORSA Journal on Computing, 2(2):126–135, Spring 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2.2.126.

Jones:1990:IGB

[43] Christopher V. Jones. An introduction to graph-based modeling systems, part i: Overview. ORSA Journal on Computing, 2(2):136–151, Spring 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.2.2.136.

Lustig:1990:ICL

[44] Irvin J. Lustig. The influence of computer language on computational comparisons: An example from network optimization. ORSA Journal on Computing, 2(2):152–161, Spring 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2.2.152.

Nicol:1990:PSS

[45] David M. Nicol. Parallel solution of sparse one-dimensional dynamic programming problems. ORSA Journal on Computing, 2(2):162–173, Spring 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.2.2.162.

Kelly:1990:USA

[46] James Kelly, Bruce Golden, and Arjang Assad. Using simulated annealing to solve controlled rounding problems. ORSA Journal on Computing, 2(2):174–185, Spring 1990. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2.2.174.

Alvarado:1990:MVS

[47] Fernando L. Alvarado. Manipulation and visualization of sparse matrices. ORSA Journal on Computing, 2(2):186–207, Spring 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2.2.186.

Greenberg:1990:EIc

[48] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 2(3):209–210, Summer 1990. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2.3.209.

Kao:1990:CSS

[49] Edward P. C. Kao and Kumar S. Narayanan. Computing steady-state probabilities of a nonpreemptive priority multiserver queue. ORSA Journal on Computing, 2(3):211–218, Summer 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2.3.211.

Gonzalez:1990:OPS

[50] Teofilo Gonzalez, Eugene L. Lawler, and Sartaj Sahni. Optimal preemptive scheduling of two unrelated processors. ORSA Journal on Computing, 2(3):219–224, Summer 1990. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2.3.219.

Floudas:1990:DSG

[51] Christodoulos A. Floudas and Avanish Aggarwal. A decomposition strategy for global optimum search in the pooling problem. ORSA Journal on Computing, 2(3):225–235, Summer 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2.3.225.

Gavish:1990:BND

[52] Bezalel Gavish and Kemal Altinkemer. Backbone network design tools with economic tradeoffs. ORSA Journal on Computing, 2(3):236–252, Summer 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2.3.236.

Gewali:1990:PPW

[53] Laxmi P. Gewali, Alex C. Meng, Joseph S. B. Mitchell, and Simeon Ntafos. Path planning in 0/1/∞ weighted regions with applications. ORSA Journal on Computing, 2(3):253–272, Summer 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2.3.253.

Chaudhry:1990:RRS

[54] Mohan L. Chaudhry, Carl M. Harris, and William G. Marchal. Robustness of rootfinding in single-server queueing models. ORSA Journal on Computing, 2(3):273–286, Summer 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2.3.273.

Mannino:1990:MLK

[55] Michael V. Mannino, Betsy S. Greenberg, and Sa Neung Hong. Model libraries: Knowledge representation and reasoning. ORSA Journal on Computing, 2(3):287–301, Summer 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2.3.287.

Greenberg:1990:EId

[56] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 2(4):303, Fall 1990. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2.4.303.

Choi:1990:FDP

[57] In Chan Choi, Clyde L. Monma, and David F. Shanno. Further development of a primal-dual interior point method. ORSA Journal on Computing, 2(4):304–311, Fall 1990. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2.4.304.

Chakravarty:1990:KMA

[58] Amiya K. Chakravarty and Diptendu Sinha. Knowledge modularization for adaptive decision modeling. ORSA Journal on Computing, 2 (4):312–324, Fall 1990. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2.4.312.

Suhl:1990:CSL

[59] Uwe H. Suhl and Leena M. Suhl. Computing sparse LU factorizations for large-scale linear programming bases. ORSA Journal on Computing, 2(4):325–335, Fall 1990. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2.4.325.

Yang:1990:ANR

[60] Che-Liang Yang and Peter Kubat. An algorithm for network reliability bounds. ORSA Journal on Computing, 2(4):336–345, Fall 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2.4.336.

Leung:1990:MTT

[61] Joseph Y.-T. Leung and Gilbert H. Young. Minimizing total tardiness on a single machine with precedence constraints. ORSA Journal on Computing, 2(4):346–352, Fall 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2.4.346.

Nygard:1990:NNS

[62] Kendall E. Nygard, Paul Juell, and Nagesh Kadaba. Neural networks for selective vehicle routing heuristics. ORSA Journal on Computing, 2 (4):353–364, Fall 1990. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2.4.353.

Saltzman:1990:BR

[63] Matthew J. Saltzman and Kevin S. McCurley. Book reviews. ORSA Journal on Computing, 2(4):365–368, Fall 1990. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2.4.365.

Greenberg:1991:EIa

[64] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 3(1):1–2, Winter 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.1.1.

Fox:1991:GMC

[65] Bennett L. Fox and Andrew R. Young. Generating Markov-chain transitions quickly: II. ORSA Journal on Computing, 3(1):3–11, Winter 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.3.1.3.

Sitaraman:1991:ASM

[66] H. Sitaraman. Approximation of some Markov-modulated Poisson processes. ORSA Journal on Computing, 3(1):12–22, Winter 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.3.1.12.

Bauer:1991:DES

[67] Kenneth W. Bauer, Jr., Bipin Kochar, and Joseph J. Talavage. Discrete event simulation model decomposition by principal components analysis. ORSA Journal on Computing, 3(1):23–32, Winter 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.3.1.23.

Helman:1991:MPT

[68] Paul Helman and Arnon Rosenthal. A mass production technique to speed multiple-query optimization and physical database design. ORSA Journal on Computing, 3(1):33–55, Winter 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.1.33.

DeLeone:1991:PIL

[69] Renato De Leone and Terence T.-H. Ow. Parallel implementation of Lemke’s algorithm on the hypercube. ORSA Journal on Computing, 3 (1):56–62, Winter 1991. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.3.1.56.

Megiddo:1991:FPD

[70] Nimrod Megiddo. On finding primal- and dual-optimal bases. ORSA Journal on Computing, 3(1):63–65, Winter 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.1.63.

Roy:1991:PCU

[71] Asim Roy and Somnath Mukhopadhyay. Pattern classification using linear programming. ORSA Journal on Computing, 3(1):66–80, Winter 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.3.1.66.

Greenberg:1991:BR

[72] Harvey J. Greenberg, John W. Chinneck, and Gerald M. Karam. Book reviews. ORSA Journal on Computing, 3(1):81–83, Winter 1991. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.3.1.81.

Greenberg:1991:EIb

[73] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 3(2):85, Spring 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.2.85.

Jagannathan:1991:OPM

[74] Raj Jagannathan. Optimal partial-match hashing design. ORSA Journal on Computing, 3(2):86–91, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.3.2.86.

Kempka:1991:PCJ

[75] David N. Kempka, Jeffery L. Kennington, and Hossam A. Zaki. Performance characteristics of the Jacobi and the Gauss–Seidel versions of the auction algorithm on the Alliant FX/8. ORSA Journal on Computing, 3(2):92–106, Spring 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.2.92.

Bhargava:1991:UNV

[76] Hemant K. Bhargava, Steven O. Kimbrough, and Ramayya Krishnan. Unique names violations, a problem for model integration or you say tomato, i say tomahto. ORSA Journal on Computing, 3(2):107–120, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.3.2.107.

Hoffman:1991:ILR

[77] Karla L. Hoffman and Manfred Padberg. Improving LP-representations of zero-one linear programs for branch-and-cut. ORSA Journal on Computing, 3(2):121–134, Spring 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.2.121.

Jaumard:1991:CGM

[78] Brigitte Jaumard, Pierre Hansen, and Marcus Poggi de Aragão. Column generation methods for probabilistic logic. ORSA Journal on Computing, 3(2):135–148, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.3.2.135.

Applegate:1991:CSJ

[79] David Applegate and William Cook. A computational study of the jobshop scheduling problem. ORSA Journal on Computing, 3(2):149–156, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.3.2.149.

Chinneck:1991:LMI

[80] John W. Chinneck and Erik W. Dravnieks. Locating minimal infeasible constraint sets in linear programs. ORSA Journal on Computing, 3(2): 157–168, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.3.2.157.

Zikan:1991:TNL

[81] Karel Zikan. Technical note — least-squares image registration. ORSA Journal on Computing, 3(2):169–172, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.2.169.

Greenberg:1991:SSR

[82] Harvey J. Greenberg. SOFTWARE SECTION — RANDMOD: A system for randomizing modifications to an instance of a linear program. ORSA Journal on Computing, 3(2):173–175, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.2.173.

Ryan:1991:BR

[83] Jennifer Ryan. Book review. ORSA Journal on Computing, 3(2):176, Spring 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.3.2.176.

Greenberg:1991:EIc

[84] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 3(3):179, Summer 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.3.179.

Jones:1991:IGB

[85] Christopher V. Jones. An introduction to graph-based modeling systems, part II: Graph-grammars and the implementation. ORSA Journal on Computing, 3(3):180–206, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.3.3.180.

Rosen:1991:CCT

[86] J. Ben Rosen and Guo-Liang Xue. Computational comparison of two algorithms for the Euclidean single facility location problem. ORSA Journal on Computing, 3(3):207–212, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.3.207.

Osborne:1991:CTS

[87] Lawrence J. Osborne and Billy E. Gillett. A comparison of two simulated annealing algorithms applied to the directed Steiner problem on networks. ORSA Journal on Computing, 3(3):213–225, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.3.3.213.

Sager:1991:PPE

[88] Thomas J. Sager and Shi-Jen Lin. A pruning procedure for exact graph coloring. ORSA Journal on Computing, 3(3):226–230, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.3.3.226.

Kao:1991:USR

[89] Edward P. C. Kao. Using state reduction for computing steady state probabilities of queues of GI/PH/1 types. ORSA Journal on Computing, 3(3):231–240, Summer 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.3.231.

Portier:1991:PAF

[90] Frederick J. Portier and Robert E. Haymond. A product automaton formalism for hierarchical, modular, discrete simulation. ORSA Journal on Computing, 3(3):241–252, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.3.241.

Greenberg:1991:ADI

[91] Harvey J. Greenberg and Frederic H. Murphy. Approaches to diagnosing infeasible linear programs. ORSA Journal on Computing, 3(3):253–261, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.3.3.253.

Dammeyer:1991:TNC

[92] Frank Dammeyer, Peter Forst, and Stefan Voss. Technical note — on the cancellation sequence method of tabu search. ORSA Journal on Computing, 3(3):262–265, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.3.3.262.

Lodwick:1991:BR

[93] Weldon A. Lodwick. Book review. ORSA Journal on Computing, 3 (3):266, Summer 1991. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.3.3.266.

Greenberg:1991:EId

[94] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 3(4):267–268, Fall 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.4.267.

Stuckey:1991:ILC

[95] Peter J. Stuckey. Incremental linear constraint solving and detection of implicit equalities. ORSA Journal on Computing, 3(4):269–274, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.3.4.269.

Fetterolf:1991:OIL

[96] Peter C. Fetterolf and G. Anandalingam. Optimizing interconnections of local area networks: An approach using simulated annealing. ORSA Journal on Computing, 3(4):275–287, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.4.275.

Downey:1991:BSO

[97] Peter J. Downey. Bounding synchronization overhead for parallel iteration. ORSA Journal on Computing, 3(4):288–298, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.3.4.288.

Kennington:1991:EAD

[98] Jeffery L. Kennington and Zhiming Wang. An empirical analysis of the dense assignment problem: Sequential and parallel implementations. ORSA Journal on Computing, 3(4):299–306, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.3.4.299.

Michalewicz:1991:NGA

[99] Zbigniew Michalewicz, George A. Vignaux, and Matthew Hobbs. A nonstandard genetic algorithm for the nonlinear transportation problem. ORSA Journal on Computing, 3(4):307–316, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.3.4.307.

Melamed:1991:TCM

[100] Benjamin Melamed. TES: A class of methods for generating autocorrelated uniform variates. ORSA Journal on Computing, 3(4):317–329, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.3.4.317.

Rozenblit:1991:RBG

[101] Jerzy W. Rozenblit and Yueh M. Huang. Rule-based generation of model structures in multifaceted modeling and system design. ORSA Journal on Computing, 3(4):330–344, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.3.4.330.

Koehler:1991:LDF

[102] Gary J. Koehler. Linear discriminant functions determined by genetic search. ORSA Journal on Computing, 3(4):345–357, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.3.4.345.

Choobineh:1991:SDS

[103] Joobin Choobineh. SQLMP: A data sublanguage for representation and formulation of linear mathematical models. ORSA Journal on Computing, 3(4):358–375, Fall 1991. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.3.4.358.

Reinelt:1991:TTS

[104] Gerhard Reinelt. TSPLIB — a traveling salesman problem library. ORSA Journal on Computing, 3(4):376–384, Fall 1991. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.3.4.376.

Greenberg:1992:EIa

[105] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 4(1):1, Winter 1992. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.4.1.1.

Smith:1992:SLS

[106] Stuart Smith and Leon Lasdon. Solving large sparse nonlinear programs using GRG. ORSA Journal on Computing, 4(1):2–15, Winter 1992. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.4.1.2.

Fourer:1992:SPL

[107] Robert Fourer and Roy E. Marsten. Solving piecewise-linear programs: Experiments with a simplex approach. ORSA Journal on Computing, 4 (1):16–31, Winter 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.1.16.

Faigle:1992:SCR

[108] Ulrich Faigle and Walter Kern. Some convergence results for probabilistic tabu search. ORSA Journal on Computing, 4(1):32–37, Winter 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.4.1.32.

Hurd:1992:ESS

[109] James K. Hurd and Frederic H. Murphy. Exploiting special structure in primal dual interior point methods. ORSA Journal on Computing, 4 (1):38–44, Winter 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.1.38.

Wallace:1992:PSP

[110] Stein W. Wallace and Roger J-B. Wets. Preprocessing in stochastic programming: The case of linear programs. ORSA Journal on Computing, 4(1):45–59, Winter 1992. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.4.1.45.

Boguslavsky:1992:SCS

[111] Leonid B. Boguslavsky, Jr., Edward G. Coffman, Edgar N. Gilbert, and Alexander Y. Kreinin. Scheduling checks and saves. ORSA Journal on Computing, 4(1):60–69, Winter 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.4.1.60.

Lee:1992:MAA

[112] Jon Lee and Jennifer Ryan. Matroid applications and algorithms. ORSA Journal on Computing, 4(1):70–98, Winter 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.4.1.70.

Vohra:1992:BR

[113] Rakesh V. Vohra. Book review. ORSA Journal on Computing, 4(1):99, Winter 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.4.1.99.

Greenberg:1992:EIb

[114] Harvey J. Greenberg. Editor’s introduction. ORSA Journal on Computing, 4(2):101–102, Spring 1992. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.4.2.101.

Mehrotra:1992:IAS

[115] Sanjay Mehrotra. Implementations of affine scaling methods: Approximate solutions of systems of linear equations using preconditioned conjugate gradient methods. ORSA Journal on Computing, 4(2):103–118, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.4.2.103.

Lodwick:1992:PNF

[116] Weldon A. Lodwick. Preprocessing nonlinear functional constraints with applications to the pooling problem. ORSA Journal on Computing, 4 (2):119–131, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.2.119.

Clark:1992:GNP

[117] Robert H. Clark, Jeffery L. Kennington, Robert R. Meyer, and Muthukrishnan Ramamurti. Generalized networks: Parallel algorithms and an empirical analysis. ORSA Journal on Computing, 4(2):132–145, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.4.2.132.

Savelsbergh:1992:VRP

[118] Martin W. P. Savelsbergh. The vehicle routing problem with time Windows: Minimizing route duration. ORSA Journal on Computing, 4(2): 146–154, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.2.146.

Ramesh:1992:OAO

[119] R. Ramesh, Yong-Seok Yoon, and Mark H. Karwan. An optimal algorithm for the orienteering tour problem. ORSA Journal on Computing, 4(2):155–165, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.2.155.

Nielsen:1992:MPA

[120] Soren S. Nielsen and Stavros A. Zenios. Massively parallel algorithms for singly constrained convex programs. ORSA Journal on Computing, 4(2):166–181, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.2.166.

Yang:1992:CMM

[121] Tao Yang, Morton J. M. Posner, and James G. C. Templeton. The Ca/M/s/m retrial queue: A computational approach. ORSA Journal on Computing, 4(2):182–191, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.4.2.182.

Balakrishnan:1992:UHC

[122] Anantaram Balakrishnan and Kemal Altinkemer. Using a hopconstrained model to generate alternative communication network design. ORSA Journal on Computing, 4(2):192–205, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.2.192.

Reinelt:1992:FHL

[123] Gerhard Reinelt. Fast heuristics for large geometric traveling salesman problems. ORSA Journal on Computing, 4(2):206–217, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.4.2.206.

Rao:1992:TNP

[124] Nageswara S. V. Rao. Technical note — on performance of path planning algorithms in unknown terrains. ORSA Journal on Computing, 4(2): 218–224, Spring 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.2.218.

Anonymous:1992:E

[125] Anonymous. From the Editor. ORSA Journal on Computing, 4(3):225, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.4.3.225.

Glynn:1992:JUB

[126] Peter W. Glynn and Philip Heidelberger. Jackknifing under a budget constraint. ORSA Journal on Computing, 4(3):226–234, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.4.3.226.

Pinar:1992:PDM

[127] Mustafa C. Pinar and Stavros A. Zenios. Parallel decomposition of multicommodity network flows using a linear-quadratic penalty algorithm. ORSA Journal on Computing, 4(3):235–249, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.3.235.

Lin:1992:NMA

[128] Frank Y. S. Lin and James R. Yee. A new multiplier adjustment procedure for the distributed computation of routing assignments in virtual circuit data networks. ORSA Journal on Computing, 4(3):250–266, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.4.3.250.

Bixby:1992:ISM

[129] Robert E. Bixby. Implementing the simplex method: The initial basis. ORSA Journal on Computing, 4(3):267–284, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.3.267.

Fraley:1992:AAF

[130] Chris Fraley and Jean-Philippe Vial. Alternative approaches to feasibility in projective methods for linear programming. ORSA Journal on Computing, 4(3):285–299, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.4.3.285.

Chaudhry:1992:NIC

[131] Mohan L. Chaudhry. Numerical issues in computing steady-state queueing-time distributions of single-server bulk-service queues: M/Gb/1 and M/Gd/1. ORSA Journal on Computing, 4(3):300–310, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.4.3.300.

Stern:1992:SAT

[132] Julio M. Stern. Simulated annealing with a temperature dependent penalty function. ORSA Journal on Computing, 4(3):311–319, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.4.3.311.

Chopra:1992:SST

[133] Sunil Chopra, Edgar R. Gorres, and M. R. Rao. Solving the Steiner tree problem on a graph using branch and cut. ORSA Journal on Computing, 4(3):320–335, Summer 1992. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.4.3.320.

Stewart:1992:NEI

[134] William J. Stewart and Wei Wu. Numerical experiments with iteration and aggregation for Markov chains. ORSA Journal on Computing, 4(3): 336–350, Summer 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.3.336.

Ghandforoush:1992:HAG

[135] Parviz Ghandforoush and John J. Daniels. A heuristic algorithm for the guillotine constrained cutting stock problem. ORSA Journal on Computing, 4(3):351–356, Summer 1992. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.4.3.351.

Mitchell:1992:GEI

[136] Joseph S. B. Mitchell and Jan Karel Lenstra. Guest Editors’ introduction. ORSA Journal on Computing, 4(4):357–359, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.4.357.

Eppstein:1992:DTD

[137] David Eppstein. Dynamic three-dimensional linear programming. ORSA Journal on Computing, 4(4):360–368, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.4.4.360.

Hartvigsen:1992:RVD

[138] David Hartvigsen. Recognizing Voronoi diagrams with linear programming. ORSA Journal on Computing, 4(4):369–374, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.4.369.

Arkin:1992:MPP

[139] Esther M. Arkin, Klara Kedem, Joseph S. B. Mitchell, Josef Sprinzak, and Michael Werman. Matching points into pairwise-disjoint noise regions: Combinatorial bounds and algorithms. ORSA Journal on Computing, 4(4):375–386, Fall 1992. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.4.4.375.

Bentley:1992:FAG

[140] Jon Jouis Bentley. Fast algorithms for geometric traveling salesman problems. ORSA Journal on Computing, 4(4):387–411, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.4.387.

Hakimi:1992:VPN

[141] S. Louis Hakimi, Martine Labbé Labbé, and Edward Schmeichel. The Voronoi partition of a network and its implications in location theory. ORSA Journal on Computing, 4(4):412–417, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.4.412.

ElGindy:1992:EAC

[142] Hossam ElGindy and J. Mark Keil. Efficient algorithms for the capacitated 1-median problem. ORSA Journal on Computing, 4(4):418–425, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.4.4.418.

Imai:1992:SCP

[143] Hiroshi Imai, D. T. Lee, and Chung-Do Yang. 1-segment center problems. ORSA Journal on Computing, 4(4):426–434, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.4.426.

Koutsoupias:1992:OBP

[144] Elias Koutsoupias, Christos H. Papadimitriou, and Martha Sideri. On the optimal bisection of a polygon. ORSA Journal on Computing, 4 (4):435–438, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.4.435.

Chandru:1992:MPC

[145] V. Chandru, V. T. Rajan, and R. Swaminathan. Monotone pieces of chains. ORSA Journal on Computing, 4(4):439–446, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.4.4.439.

Alonso:1992:LMU

[146] Laurent Alonso, Arthur S. Goldstein, and Edward M. Reingold. “lion and man”: Upper and lower bounds. ORSA Journal on Computing, 4 (4):447–452, Fall 1992. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.4.4.447.

Golden:1993:E

[147] Bruce L. Golden. From the Editor. ORSA Journal on Computing, 5(1): 1, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.5.1.1.

Barr:1993:FAR

[148] Richard S. Barr and Betty L. Hickman. Feature article — reporting computational experiments with parallel algorithms: Issues, measures, and experts’ opinions. ORSA Journal on Computing, 5(1):2–18, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.1.2.

Gustafson:1993:CTB

[149] John L. Gustafson. Commentary — the “Tar Baby” of computing: Performance analysis. ORSA Journal on Computing, 5(1):19–21, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.1.19.

Loshin:1993:CPR

[150] H. David Loshin and Alex Vasilevsky. Commentary — parallel reporting guidelines are on target. ORSA Journal on Computing, 5(1):22–23, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.5.1.22.

Lusk:1993:CSI

[151] Ewing L. Lusk. Commentary — speedups and insights. ORSA Journal on Computing, 5(1):24–25, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.5.1.24.

Miller:1993:CRP

[152] Donald L. Miller and Joseph F. Pekny. Commentary — the role of performance metrics for parallel mathematical programming algorithms. ORSA Journal on Computing, 5(1):26–28, Winter 1993. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.5.1.26.

Barr:1993:RUP

[153] Richard S. Barr and Betty L. Hickman. Rejoinder — using parallel empirical testing to advance algorithm research. ORSA Journal on Computing, 5(1):29–32, Winter 1993. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.5.1.29.

Bhargava:1993:DAM

[154] Hemant K. Bhargava. Dimensional analysis in mathematical modeling systems: A simple numerical method. ORSA Journal on Computing, 5 (1):33–39, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.5.1.33.

Nash:1993:BML

[155] Stephen G. Nash and Ariela Sofer. A barrier method for large-scale constrained optimization. ORSA Journal on Computing, 5(1):40–53, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.1.40.

Shanker:1993:MCD

[156] Murali S. Shanker, W. David Kelton, and Rema Padman. Measuring congestion for dynamic task allocation in distributed simulation. ORSA Journal on Computing, 5(1):54–68, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.5.1.54.

Johnson:1993:SPP

[157] Mary A. Johnson. Selecting parameters of phase distributions: Combining nonlinear programming, heuristics, and Erlang distributions. ORSA Journal on Computing, 5(1):69–83, Winter 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.5.1.69.

Eckstein:1993:ASM

[158] Jonathan Eckstein. The alternating step method for monotropic programming on the connection machine CM-2. ORSA Journal on Computing, 5(1):84–96, Winter 1993. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.5.1.84.

Hansen:1993:SAS

[159] Pierre Hansen, Brigitte Jaumard, and Vincent Mathon. State-of-the-art survey — constrained nonlinear 0–1 programming. ORSA Journal on Computing, 5(2):97–119, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.5.2.97.

Gallo:1993:TPE

[160] Giorgio Gallo and Maria Grazia Scutellà. Toward a programming environment for combinatorial optimization: A case study oriented to maxflow computations. ORSA Journal on Computing, 5(2):120–133, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.2.120.

Vanderbei:1993:AAL

[161] Robert J. Vanderbei. ALPO: Another linear program optimizer. ORSA Journal on Computing, 5(2):134–146, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.5.2.134.

Fox:1993:SFE

[162] Bennett L. Fox. Shortening future-event lists. ORSA Journal on Computing, 5(2):147–150, Spring 1993. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.5.2.147.

Ghandeharizadeh:1993:OBA

[163] Shahram Ghandeharizadeh, Robert R. Meyer, Gary L. Schultz, and Jonathan Yackel. Optimal balanced assignments and a parallel database application. ORSA Journal on Computing, 5(2):151–167, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.2.151.

Nussbaum:1993:PSS

[164] Miguel Nussbaum and Eduardo A. Parra. A production scheduling system. ORSA Journal on Computing, 5(2):168–181, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.5.2.168.

Carpenter:1993:SQP

[165] Tamra J. Carpenter, Irvin J. Lustig, John M. Mulvey, and David F. Shanno. Separable quadratic programming via a primal-dual interior point method and its use in a sequential procedure. ORSA Journal on Computing, 5(2):182–191, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.5.2.182.

vandeVelde:1993:DBA

[166] S. L. van de Velde. Duality-based algorithms for scheduling unrelated parallel machines. ORSA Journal on Computing, 5(2):192–205, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.2.192.

Cook:1993:IGB

[167] William Cook, Thomas Rutherford, Herbert E. Scarf, and David Shallcross. An implementation of the generalized basis reduction algorithm for integer programming. ORSA Journal on Computing, 5(2):206–212, Spring 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.5.2.206.

Fujimoto:1993:FAP

[168] Richard M. Fujimoto. Feature article — parallel discrete event simulation: Will the field survive? ORSA Journal on Computing, 5(3):213–230, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.5.3.213.

Abrams:1993:CPD

[169] Marc Abrams. Commentary — parallel discrete event simulation: Fact or fiction? ORSA Journal on Computing, 5(3):231–233, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.3.231.

Bagrodia:1993:CSG

[170] Rajive Bagrodia. Commentary — a survival guide for parallel simulation. ORSA Journal on Computing, 5(3):234–235, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.5.3.234.

Lin:1993:CWP

[171] Yi-Bing Lin. Commentary — will parallel simulation research survive? ORSA Journal on Computing, 5(3):236–238, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.5.3.236.

Reynolds:1993:CSB

[172] Paul F. Reynolds, Jr. Commentary — the Silver bullet. ORSA Journal on Computing, 5(3):239–241, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.5.3.239.

Unger:1993:CPP

[173] Brian W. Unger and John G. Cleary. Commentary — practical parallel discrete event simulation. ORSA Journal on Computing, 5(3):242–244, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.5.3.242.

Fujimoto:1993:RFD

[174] Richard M. Fujimoto. Rejoinder — future directions in parallel simulation research. ORSA Journal on Computing, 5(3):245–248, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.3.245.

Borgwardt:1993:ESA

[175] Karl Heinz Borgwardt, Renate Damm, Rudolf Donig, and Gabriele Joas. Empirical studies on the average efficiency of simplex variants under rotation symmetry. ORSA Journal on Computing, 5(3):249–260, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.5.3.249.

Bertsekas:1993:PAH

[176] Dimitri P. Bertsekas and David A. Castañon. Parallel asynchronous Hungarian methods for the assignment problem. ORSA Journal on Computing, 5(3):261–274, Summer 1993. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.5.3.261.

Steiger:1993:GIN

[177] David Steiger, Ramesh Sharda, and Brian Leclaire. Graphical interfaces for network modeling: A model management system perspective. ORSA Journal on Computing, 5(3):275–291, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.5.3.275.

Grassmann:1993:CSS

[178] Winfried K. Grassmann and Daniel P. Heyman. Computation of steadystate probabilities for infinite-state Markov chains with repeating rows. ORSA Journal on Computing, 5(3):292–303, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.5.3.292.

Prietula:1993:AAG

[179] Michael J. Prietula, Wen-Ling Hsu, David Steier, and Allen Newell. Applying an architecture for general intelligence to reduce scheduling effort. ORSA Journal on Computing, 5(3):304–320, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.5.3.304.

Chen:1993:AAT

[180] Bo Chen and Vitaly A. Strusevich. Approximation algorithms for threemachine open shop scheduling. ORSA Journal on Computing, 5(3): 321–326, Summer 1993. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.5.3.321.

Anonymous:1993:E

[181] Anonymous. From the Editor. ORSA Journal on Computing, 5(4):327, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.5.4.327.

Potvin:1993:SAS

[182] Jean-Yves Potvin. State-of-the-art survey — the traveling salesman problem: A neural network perspective. ORSA Journal on Computing, 5 (4):328–348, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.5.4.328.

Mangasarian:1993:MPN

[183] O. L. Mangasarian. Mathematical programming in neural networks. ORSA Journal on Computing, 5(4):349–360, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.5.4.349.

Piramuthu:1993:LAN

[184] Selwyn Piramuthu, Chung-Ming Kuan, and Michael J. Shaw. Learning algorithms for neural-net decision support. ORSA Journal on Computing, 5(4):361–373, Fall 1993. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.5.4.361.

Tang:1993:FNN

[185] Zaiyong Tang and Paul A. Fishwick. Feedforward neural nets as models for time series forecasting. ORSA Journal on Computing, 5(4):374– 385, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.5.4.374.

Subramanian:1993:GBS

[186] Venkat Subramanian and Ming S. Hung. A GRG2-based system for training neural networks: Design and computational experience. ORSA Journal on Computing, 5(4):386–394, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.5.4.386.

Amini:1993:NRA

[187] Mohammad M. Amini and Richard S. Barr. Network reoptimization algorithms: A statistically designed comparison. ORSA Journal on Computing, 5(4):395–409, Fall 1993. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.5.4.395.

Seppi:1993:BAD

[188] Kevin D. Seppi, J. Wesley Barnes, and Carl N. Morris. A Bayesian approach to database query optimization. ORSA Journal on Computing, 5(4):410–419, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.5.4.410.

Liman:1993:EBH

[189] Surya Danusaputro Liman and Chung-Yee Lee. Error bound of a heuristic for the common due date scheduling problem. ORSA Journal on Computing, 5(4):420–425, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.5.4.420.

Fischetti:1993:EAM

[190] Matteo Fischetti and Paolo Toth. An efficient algorithm for the minsum arborescence problem on complete digraphs. ORSA Journal on Computing, 5(4):426–434, Fall 1993. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.5.4.426.

Lustig:1994:FAI

[191] Irvin J. Lustig, Roy E. Marsten, and David F. Shanno. Feature article — interior point methods for linear programming: Computational state of the art. ORSA Journal on Computing, 6(1):1–14, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.1.1.

Bixby:1994:CPL

[192] Robert E. Bixby. Commentary — progress in linear programming. ORSA Journal on Computing, 6(1):15–22, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.1.15.

Saunders:1994:CMC

[193] Michael A. Saunders. Commentary — major Cholesky would feel proud. ORSA Journal on Computing, 6(1):23–27, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.1.23.

Todd:1994:CTP

[194] Michael J. Todd. Commentary — theory and practice for interior-point methods. ORSA Journal on Computing, 6(1):28–31, Winter 1994. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.6.1.28.

Vanderbei:1994:CIP

[195] Robert J. Vanderbei. Commentary — interior-point methods: Algorithms and formulations. ORSA Journal on Computing, 6(1):32–34, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.6.1.32.

Lustig:1994:RLW

[196] Irvin J. Lustig, Roy E. Marsten, and David F. Shanno. Rejoinder — the last word on interior point methods for linear programming — for now. ORSA Journal on Computing, 6(1):35–36, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.1.35.

Dugat:1994:CHT

[197] Vincent Dugat and Sandra Sandri. Complexity of hierarchical trees in evidence theory. ORSA Journal on Computing, 6(1):37–49, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.6.1.37.

Pirkul:1994:PSR

[198] Hasan Pirkul and Sridhar Narasimhan. Primary and secondary route selection in backbone computer networks. ORSA Journal on Computing, 6(1):50–60, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.6.1.50.

Dewan:1994:DPC

[199] Rajiv M. Dewan. Design of performance constrained optic fiber multidrop access networks. ORSA Journal on Computing, 6(1):61–67, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.6.1.61.

Pekny:1994:SPD

[200] Joseph F. Pekny and Donald L. Miller. A staged primal-dual algorithm for finding a minimum cost perfect two-matching in an undirected graph. ORSA Journal on Computing, 6(1):68–81, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.1.68.

Aboudi:1994:TSG

[201] Ronny Aboudi and Kurt Jörnsten. Tabu search for general zero-one integer programs using the pivot and complement heuristic. ORSA Journal on Computing, 6(1):82–93, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.6.1.82.

Jung:1994:NFM

[202] Ho-Won Jung, Roy E. Marsten, and Matthew J. Saltzman. Numerical factorization methods for interior point algorithms. ORSA Journal on Computing, 6(1):94–105, Winter 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.6.1.94.

Golden:1994:E

[203] Bruce L. Golden. From the Editor. ORSA Journal on Computing, 6 (2):107, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.6.2.107.

Taillard:1994:PTS

[204] Éric D. Taillard. Parallel taboo search techniques for the job shop scheduling problem. ORSA Journal on Computing, 6(2):108–117, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.6.2.108.

Aarts:1994:CSL

[205] E. H. L. Aarts, P. J. M. van Laarhoven, J. K. Lenstra, and N. L. J. Ulder. A computational study of local search algorithms for job shop scheduling. ORSA Journal on Computing, 6(2):118–125, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.6.2.118.

Battiti:1994:RTS

[206] Roberto Battiti and Giampietro Tecchiolli. The reactive tabu search. ORSA Journal on Computing, 6(2):126–140, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.2.126.

Dorndorf:1994:FCA

[207] Ulrich Dorndorf and Erwin Pesch. Fast clustering algorithms. ORSA Journal on Computing, 6(2):141–153, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.2.141.

Bean:1994:GAR

[208] James C. Bean. Genetic algorithms and random keys for sequencing and optimization. ORSA Journal on Computing, 6(2):154–160, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.6.2.154.

Anderson:1994:GAC

[209] Edward J. Anderson and Michael C. Ferris. Genetic algorithms for combinatorial optimization: The assemble line balancing problem. ORSA Journal on Computing, 6(2):161–173, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.2.161.

Sikora:1994:DLL

[210] Riyaz Sikora and Michael Shaw. A double-layered learning approach to acquiring rules for classification: Integrating genetic algorithms with similarity-based learning. ORSA Journal on Computing, 6(2):174–187, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.6.2.174.

Melekopoglou:1994:CPI

[211] Mary Melekopoglou and Anne Condon. On the complexity of the policy improvement algorithm for Markov decision processes. ORSA Journal on Computing, 6(2):188–192, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.6.2.188.

Gondzio:1994:EOP

[212] Jacek Gondzio. On exploiting original problem data in the inverse representation of linear programming bases. ORSA Journal on Computing, 6(2):193–206, Spring 1994. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.6.2.193.

Drud:1994:CLS

[213] Arne Stolbjerg Drud. CONOPT — a large-scale GRG code. ORSA Journal on Computing, 6(2):207–216, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.2.207.

Pollatschek:1994:SSS

[214] A. M. Pollatschek. SSS: Subrouting set for experimenting in simulation research. ORSA Journal on Computing, 6(2):217–220, Spring 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.6.2.217.

Jones:1994:FAV

[215] Christopher V. Jones. Feature article — visualization and optimization. ORSA Journal on Computing, 6(3):221–257, Summer 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.3.221.

Bell:1994:CVO

[216] Peter C. Bell. Commentary — visualization and optimization: The future lies together. ORSA Journal on Computing, 6(3):258–260, Summer 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.6.3.258.

Bodin:1994:CVV

[217] Lawrence Bodin and Laurence Levy. Commentary — visualization in vehicle routing and scheduling problems. ORSA Journal on Computing, 6(3):261–269, Summer 1994. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.6.3.261.

Greenberg:1994:CPH

[218] Harvey J. Greenberg and Frederic H. Murphy. Commentary — the promise of OR is helping to solve problems. ORSA Journal on Computing, 6(3):270–272, Summer 1994. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.6.3.270.

Keiper:1994:CDT

[219] Jerry Keiper and Tom Wickham-Jones. Commentary — designing tools for visualization and optimization. ORSA Journal on Computing, 6(3): 273–277, Summer 1994. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.6.3.273.

Jones:1994:RAC

[220] Christopher V. Jones. Rejoinder — anchoring and cross-fertilization. ORSA Journal on Computing, 6(3):278–280, Summer 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.3.278.

Shenoy:1994:CVB

[221] Prakash P. Shenoy. Consistency in valuation-based systems. ORSA Journal on Computing, 6(3):281–291, Summer 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.3.281.

Al-Sultan:1994:NBR

[222] Khaled S. Al-Sultan. A Newton based radius reduction algorithm for nearest point problems in Pos cones. ORSA Journal on Computing, 6(3):292–299, Summer 1994. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.6.3.292.

Kushner:1994:NMC

[223] Harold J. Kushner and Jichuan Yang. A numerical method for controlled routing in large trunk line networks via stochastic control theory. ORSA Journal on Computing, 6(3):300–316, Summer 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.3.300.

Anonymous:1994:E

[224] Anonymous. From the Editor. ORSA Journal on Computing, 6(4):317, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.6.4.317.

Kelton:1994:PSR

[225] W. David Kelton. Perspectives on simulation research and practice. ORSA Journal on Computing, 6(4):318–328, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.4.318.

Ozden:1994:IOS

[226] Mufit Ozden. Intelligent objects in simulation. ORSA Journal on Computing, 6(4):329–341, Fall 1994. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.6.4.329.

Avramidis:1994:FME

[227] Athanassios N. Avramidis and James R. Wilson. A flexible method for estimating inverse distribution functions in simulation experiments. ORSA Journal on Computing, 6(4):342–355, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.4.342.

Johnson:1994:EEP

[228] Mary A. Johnson, Sanghoon Lee, and James R. Wilson. Experimental evaluation of a procedure for estimating nonhomogeneous Poisson processes having cyclic behavior. ORSA Journal on Computing, 6(4):356– 368, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.6.4.356.

Shaw:1994:AMI

[229] Wade H. Shaw, Jr., Nathaniel J. Davis IV, and Richard A. Raines. The application of metamodeling to interconnection network analysis. ORSA Journal on Computing, 6(4):369–380, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.6.4.369.

Jacobson:1994:CRH

[230] Sheldon H. Jacobson. Convergence results for harmonic gradient estimators. ORSA Journal on Computing, 6(4):381–397, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.4.381.

Safizadeh:1994:OSQ

[231] M. Hossein Safizadeh and Robert Signorile. Optimization of simulation via quasi-Newton methods. ORSA Journal on Computing, 6(4):398– 408, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.6.4.398.

Ghiaseddin:1994:SES

[232] Nasir Ghiaseddin, Khalil Matta, and Diptendu Sinha. A structured expert system for model management in inventory control. ORSA Journal on Computing, 6(4):409–422, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.6.4.409.

Harche:1994:CSS

[233] F. Harche, J. N. Hooker, and G. L. Thompson. A computational study of satisfiability algorithms for propositional logic. ORSA Journal on Computing, 6(4):423–435, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.6.4.423.

Osiakwan:1994:EAC

[234] Constantine K. N. Osiakwan and Selim G. Akl. An EP algorithm for computing a minimum weight perfect matching for a set of points on the plane. ORSA Journal on Computing, 6(4):436–444, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.6.4.436.

Savelsbergh:1994:PPT

[235] M. W. P. Savelsbergh. Preprocessing and probing techniques for mixed integer programming problems. ORSA Journal on Computing, 6(4):445– 454, Fall 1994. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.6.4.445.

Miller:1995:MBE

[236] Donald L. Miller. A matching based exact algorithm for capacitated vehicle routing problems. ORSA Journal on Computing, 7(1):1–9, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.1.1.

Kontoravdis:1995:GVR

[237] George Kontoravdis and Jonathan F. Bard. A GRASP for the vehicle routing problem with time Windows. ORSA Journal on Computing, 7 (1):10–23, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.1.10.

Diaby:1995:DPC

[238] Moustapha Diaby and R. Ramesh. The distribution problem with carrier service: A dual based penalty approach. ORSA Journal on Computing, 7(1):24–35, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.1.24.

Abate:1995:NIL

[239] Joseph Abate and Ward Whitt. Numerical inversion of Laplace transforms of probability distributions. ORSA Journal on Computing, 7(1): 36–43, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.1.36.

Wallace:1995:PSP

[240] Stein W. Wallace and Roger J-B Wets. Preprocessing in stochastic programming: The case of capacitated networks. ORSA Journal on Computing, 7(1):44–62, Winter 1995. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.7.1.44.

McCormick:1995:SIJ

[241] S. Thomas McCormick and Michael L. Pinedo. Scheduling n independent jobs on m uniform machines with both flowtime and makespan objectives: A parametric analysis. ORSA Journal on Computing, 7(1):63–77, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.1.63.

Chhajed:1995:HES

[242] Dilip Chhajed and Vijay Chandru. Hulls and efficient sets for the rectilinear norm. ORSA Journal on Computing, 7(1):78–83, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.1.78.

Liebling:1995:DPP

[243] Th. M. Liebling, F. Margot, D. Müller, A. Prodon, and L. Stauffer. Disjoint paths in the plane. ORSA Journal on Computing, 7(1):84–88, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.7.1.84.

Jones:1995:ECP

[244] Lee K. Jones and Richard C. Larson. Efficient computation of probabilities of events described by order statistics and applications to queue inference. ORSA Journal on Computing, 7(1):89–100, Winter 1995. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.7.1.89.

Heyman:1995:CBA

[245] Daniel P. Heyman and Meredith J. Goldsmith. Comparisons between aggregation/disaggregation and a direct algorithm for computing the stationary probabilities of a Markov chain. ORSA Journal on Computing, 7(1):101–108, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.1.101.

Bretthauer:1995:BBA

[246] Kurt M. Bretthauer, Bala Shetty, and Siddhartha Syam. A branch and bound algorithm for integer quadratic knapsack problems. ORSA Journal on Computing, 7(1):109–116, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.7.1.109.

Anonymous:1995:E

[247] Anonymous. Erratum. ORSA Journal on Computing, 7(1):116, Winter 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.1.116.

Mattingly:1995:RSC

[248] R. Bruce Mattingly. A revised stochastic complementation algorithm for nearly completely decomposable Markov chains. ORSA Journal on Computing, 7(2):117–124, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.7.2.117.

Downey:1995:BAO

[249] Peter J. Downey. Bounds and approximations for overheads in the time to join parallel forks. ORSA Journal on Computing, 7(2):125–139, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.2.125.

Jagerman:1995:SAB

[250] David L. Jagerman and Benjamin Melamed. Spectral analysis of basic TES processes. ORSA Journal on Computing, 7(2):140–148, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.2.140.

Park:1995:SPC

[251] June S. Park, Robert Bartoszyński, and Walter A. Rosenkrantz. Stability of p-persistent CSMA/CD. ORSA Journal on Computing, 7(2): 149–159, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.2.149.

Jordan:1995:CCM

[252] Carsten Jordan and Andreas Drexl. A comparison of constraint and mixed-integer programming solvers for batch sequencing with sequencedependent setups. ORSA Journal on Computing, 7(2):160–165, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.2.160.

Fourer:1995:ESS

[253] Robert Fourer and David M. Gay. Expressing special structures in an algebraic modeling language for mathematical programming. ORSA Journal on Computing, 7(2):166–190, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.7.2.166.

DellAmico:1995:OST

[254] Mauro Dell’Amico and Silvano Martello. Optimal scheduling of tasks on identical parallel processors. ORSA Journal on Computing, 7(2): 191–200, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.2.191.

Han:1995:AOM

[255] Sunan Han, Dawei Hong, and Joseph Y.-T. Leung. On the asymptotic optimality of multiprocessor scheduling heuristics for the makespan minimization problem. ORSA Journal on Computing, 7(2):201–204, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.2.201.

Greenberg:1995:APP

[256] Harvey J. Greenberg. Analyzing the pooling problem. ORSA Journal on Computing, 7(2):205–217, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.7.2.205.

Mehrez:1995:MIP

[257] Abraham Mehrez and Geoffrey Steinberg. A matching identification problem: A knowledge acquisition approach. ORSA Journal on Computing, 7(2):218–231, Spring 1995. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.7.2.218.

Hariri:1995:SMS

[258] A. M. A. Hariri, C. N. Potts, and L. N. Van Wassenhove. Single machine scheduling to minimize total weighted late work. ORSA Journal on Computing, 7(2):232–242, Spring 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.7.2.232.

Basu:1995:ICK

[259] Amit Basu. Introduction to the cluster on knowledge and data management. ORSA Journal on Computing, 7(3):243, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.7.3.243.

Ram:1995:IMC

[260] Sudha Ram and Sridhar Narasimhan. Incorporating the majority consensus concurrency control mechanism into the database allocation problem. ORSA Journal on Computing, 7(3):244–256, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.7.3.244.

Gopal:1995:APO

[261] Ram D. Gopal, R. Ramesh, and Stanley Zionts. Access path optimization in relational joins. ORSA Journal on Computing, 7(3):257–268, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.3.257.

Goes:1995:SMP

[262] Paulo B. Goes. A stochastic model for performance evaluation of main memory resident database systems. ORSA Journal on Computing, 7(3): 269–282, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.3.269.

Triantaphyllou:1995:RBC

[263] Evangelos Triantaphyllou and Allen L. Soyster. A relationship between CNF and DNF systems derivable from examples. ORSA Journal on Computing, 7(3):283–285, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.7.3.283.

Srikanth:1995:GTB

[264] Rajan Srikanth. A graph theory-based approach for partitioning knowledge bases. ORSA Journal on Computing, 7(3):286–297, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.3.286.

Miller:1995:SPD

[265] Donald L. Miller and Joseph F. Pekny. A staged primal-dual algorithm for perfect b-matching with edge capacities. ORSA Journal on Computing, 7(3):298–320, Summer 1995. CODEN ???? ISSN 0899-1499 (print),

2326-3245 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.7.3.298.

Lasdon:1995:PDP

[266] Leon S. Lasdon, John Plummer, and Gang Yu. Primal-dual and primal interior point algorithms for general nonlinear programs. ORSA Journal on Computing, 7(3):321–332, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.7.3.321.

Olariu:1995:IGP

[267] S. Olariu, J. L. Schwing, and J. Zhang. Interval graph problems on reconfigurable meshes. ORSA Journal on Computing, 7(3):333–348, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.3.333.

Archibald:1995:PII

[268] T. W. Archibald, K. I. M. McKinnon, and L. C. Thomas. Performance issues for the iterative solution of Markov decision processes on parallel computers. ORSA Journal on Computing, 7(3):349–357, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.3.349.

Berland:1995:LBH

[269] Nils Jacob Berland. Load balancing in hypercube solution of stochastic optimization problems. ORSA Journal on Computing, 7(3):358–364, Summer 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.7.3.358.

Grama:1995:PSA

[270] Ananth Grama and Vipin Kumar. Parallel search algorithms for discrete optimization problems. ORSA Journal on Computing, 7(4):365–385, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.4.365.

Eschenbach:1995:PAS

[271] Elizabeth A. Eschenbach, Christine A. Shoemaker, and Hugh M. Caffey. Parallel algorithms for stochastic dynamic programming with continuous state and control variables. ORSA Journal on Computing, 7 (4):386–401, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.4.386.

Eckstein:1995:DPI

[272] Jonathan Eckstein, İ. İlkay Boduroğlu, Lazaros C. Polymenakos, and Donald Goldfarb. Data-parallel implementations of dense simplex methods on the connection machine CM-2. ORSA Journal on Computing, 7 (4):402–416, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.4.402.

Hu:1995:OBA

[273] T. C. Hu, Andrew B. Kahng, and Chung-Wen Albert Tsao. Old bachelor acceptance: A new class of non-monotone threshold accepting methods. ORSA Journal on Computing, 7(4):417–425, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.7.4.417.

Glover:1995:TTI

[274] Fred Glover. Tabu thresholding: Improved search by nonmonotonic trajectories. ORSA Journal on Computing, 7(4):426–442, Fall 1995. CO-DEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.7.4.426.

Herrmann:1995:SCS

[275] Jeffrey W. Herrmann and Chung-Yee Lee. Solving a class scheduling problem with a genetic algorithm. ORSA Journal on Computing, 7 (4):443–452, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-

3245 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.7.4.443.

Storer:1995:PHS

[276] Robert H. Storer, S. David Wu, and Renzo Vaccari. Problem and heuristic space search strategies for job shop scheduling. ORSA Journal on Computing, 7(4):453–467, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.7.4.453.

Woodruff:1995:GIP

[277] David L. Woodruff. Ghost image processing for minimum covariance determinants. ORSA Journal on Computing, 7(4):468–473, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.7.4.468.

Czyzyk:1995:SAS

[278] Joseph Czyzyk, Robert Fourer, and Sanjay Mehrotra. A study of the augmented system and column-splitting approaches for solving two-stage stochastic linear programs by interior-point methods. ORSA Journal on Computing, 7(4):474–490, Fall 1995. CODEN ???? ISSN 0899-1499 (print), 2326-3245 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.7.4.474.

McGeoch:1996:FAT

[279] Catherine C. McGeoch. Feature article — toward an experimental method for algorithm simulation. INFORMS Journal on Computing, 8(1):1–15, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.1.1.

LEcuyer:1996:CSA

[280] Pierre L’Ecuyer. Commentary — simulation of algorithms for performance analysis. INFORMS Journal on Computing, 8(1):16–20, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.8.1.16.

Orlin:1996:CEM

[281] James B. Orlin. Commentary — on experimental methods for algorithm simulation. INFORMS Journal on Computing, 8(1):21–23, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.8.1.21.

Shier:1996:CAA

[282] Douglas R. Shier. Commentary — on algorithm analysis. INFORMS Journal on Computing, 8(1):24–26, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.8.1.24.

McGeoch:1996:RCA

[283] Catherine C. McGeoch. Rejoinder — challenges in algorithm simulation. INFORMS Journal on Computing, 8(1):27–28, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.8.1.27.

Williamson:1996:CEA

[284] David P. Williamson and Michel X. Goemans. Computational experience with an approximation algorithm on large-scale Euclidean matching instances. INFORMS Journal on Computing, 8(1):29–40, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.8.1.29.

Punnen:1996:IAC

[285] Abraham P. Punnen and K. P. K. Nair. An improved algorithm for the constrained bottleneck spanning tree problem. INFORMS Journal on Computing, 8(1):41–44, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.8.1.41.

Blanc:1996:CDM

[286] J. P. C. Blanc and R. D. Van der Mei. Computation of derivatives by means of the power-series algorithm. INFORMS Journal on Computing, 8(1):45–54, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.1.45.

Chinneck:1996:LDI

[287] John W. Chinneck. Localizing and diagnosing infeasibilities in networks. INFORMS Journal on Computing, 8(1):55–62, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.8.1.55.

Mao:1996:IRT

[288] Weizhen Mao and David M. Nicol. Isomorphic routing on a toroidal mesh. INFORMS Journal on Computing, 8(1):63–73, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.8.1.63.

Kao:1996:CAA

[289] Edward P. C. Kao. A comparison of alternative approaches for numerical solutions of GI/PH/1 queues. INFORMS Journal on Computing, 8 (1):74–85, Winter 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.1.74.

Sanchis:1996:SET

[290] Laura A. Sanchis and Arun Jagota. Some experimental and theoretical results on test case generators for the maximum clique problem. INFORMS Journal on Computing, 8(2):87–102, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.8.2.87.

Mausser:1996:AAN

[291] Helmut E. Mausser, Michael J. Magazine, and John B. Moore. Application of an annealed neural network to a timetabling problem. INFORMS Journal on Computing, 8(2):103–117, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.8.2.103.

Wang:1996:SOA

[292] Shouhong Wang. A self-organizing approach to managerial nonlinear discriminant analysis: A hybrid method of linear discriminant analysis and neural networks. INFORMS Journal on Computing, 8(2):118–124, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.8.2.118.

Codenotti:1996:PET

[293] Bruno Codenotti, Giovanni Manzini, Luciano Margara, and Giovanni Resta. Perturbation: An efficient technique for the solution of very large instances of the Euclidean TSP. INFORMS Journal on Computing, 8 (2):125–133, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.2.125.

Renaud:1996:FCH

[294] Jacques Renaud, Fayez F. Boctor, and Gilbert Laporte. A fast composite heuristic for the symmetric traveling salesman problem. INFORMS Journal on Computing, 8(2):134–143, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.8.2.134.

Pesch:1996:CPB

[295] Erwin Pesch and Ulrich A. W. Tetzlaff. Constraint propagation based scheduling of job shops. INFORMS Journal on Computing, 8(2):144–157, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.8.2.144.

Potvin:1996:VRPa

[296] Jean-Yves Potvin, Tanguy Kervahut, Bruno-Laurent Garcia, and Jean-Marc Rousseau. The vehicle routing problem with time Windows part i: Tabu search. INFORMS Journal on Computing, 8(2):158–164, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.8.2.158.

Potvin:1996:VRPb

[297] Jean-Yves Potvin and Samy Bengio. The vehicle routing problem with time Windows part II: Genetic search. INFORMS Journal on Computing, 8(2):165–172, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.2.165.

Coit:1996:APM

[298] David W. Coit, Alice E. Smith, and David M. Tate. Adaptive penalty methods for genetic optimization of constrained combinatorial problems. INFORMS Journal on Computing, 8(2):173–182, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.8.2.173.

Aytug:1996:SCF

[299] Haldun Aytug and Gary J. Koehler. Stopping criteria for finite length genetic algorithms. INFORMS Journal on Computing, 8(2):183–191, Spring 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.8.2.183.

Anonymous:1996:E

[300] Anonymous. From the Editor. INFORMS Journal on Computing, 8 (3):193, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.3.193.

Lee:1996:BCA

[301] Youngho Lee, Steve Y. Chiu, and Jennifer Ryan. A branch and cut algorithm for a Steiner tree-star problem. INFORMS Journal on Computing, 8(3):194–201, Summer 1996. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.8.3.194.

Mirchandani:1996:MTT

[302] Prakash Mirchandani. The multi-tier tree problem. INFORMS Journal on Computing, 8(3):202–218, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.8.3.202.

Hall:1996:ECP

[303] Leslie Hall. Experience with a cutting plane algorithm for the capacitated spanning tree problem. INFORMS Journal on Computing, 8(3): 219–234, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.3.219.

Vachani:1996:MFR

[304] Rita Vachani, Alexander Shulman, Peter Kubat, and Julie Ward. Multicommodity flows in ring networks. INFORMS Journal on Computing, 8(3):235–242, Summer 1996. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.8.3.235.

Bienstock:1996:CND

[305] Daniel Bienstock and Oktay Günlük. Capacitated network design — polyhedral structure and computation. INFORMS Journal on Computing, 8(3):243–259, Summer 1996. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.8.3.243.

Balakrishnan:1996:SIS

[306] Anantaram Balakrishnan and Srimathy Gopalakrishnan. Selecting ingot sizes for joint production of sheet products. INFORMS Journal on Computing, 8(3):260–273, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.8.3.260.

Bitran:1996:IMC

[307] Gabriel R. Bitran and Susana V. Mondschein. Inventory management in catalog sales companies. INFORMS Journal on Computing, 8(3): 274–288, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.8.3.274.

Aardal:1996:TLU

[308] Karen Aardal, Martine Labbé, Janny Leung, and Maurice Queyranne. On the two-level uncapacitated facility location problem. INFORMS Journal on Computing, 8(3):289–301, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.8.3.289.

Vaessens:1996:JSS

[309] R. J. M. Vaessens, E. H. L. Aarts, and J. K. Lenstra. Job shop scheduling by local search. INFORMS Journal on Computing, 8(3):302–317, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.8.3.302.

Ahuja:1996:URO

[310] Ravindra K. Ahuja and James B. Orlin. Use of representative operation counts in computational testing of algorithms. INFORMS Journal on Computing, 8(3):318–330, Summer 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.8.3.318.

Jayawardena:1996:MSG

[311] T. Jayawardena and S. Yakowitz. Methodology for stochastic graph completion-time problems. INFORMS Journal on Computing, 8(4):331– 343, Fall 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.8.4.331.

Mehrotra:1996:CGA

[312] Anuj Mehrotra and Michael A. Trick. A column generation approach for graph coloring. INFORMS Journal on Computing, 8(4):344–354, Fall 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.8.4.344.

Jarvis:1996:IAA

[313] James P. Jarvis and Douglas R. Shier. An improved algorithm for approximating the performance of stochastic flow networks. INFORMS Journal on Computing, 8(4):355–360, Fall 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.8.4.355.

Averbakh:1996:BSS

[314] Igor Averbakh and Oded Berman. Bottleneck Steiner subnetwork problems with k-connectivity constraints. INFORMS Journal on Computing, 8(4):361–366, Fall 1996. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.8.4.361.

Sun:1996:QTL

[315] Minghe Sun and Ralph E. Steuer. Quad-trees and linear lists for identifying nondominated criterion vectors. INFORMS Journal on Computing, 8(4):367–375, Fall 1996. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.8.4.367.

vanHoesel:1996:ATM

[316] C. P. M. van Hoesel, A. P. M. Wagelmans, and M. van Vliet. An O(n log n) algorithm for the two-machine flow shop problem with controllable machine speeds. INFORMS Journal on Computing, 8(4):376– 382, Fall 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.8.4.376.

Ma:1996:IL

[317] Pai-Chun Ma, Frederic H. Murphy, and Edward A. Stohr. An implementation of LPFORM. INFORMS Journal on Computing, 8(4):383– 401, Fall 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.8.4.383.

Hoogeveen:1996:BBA

[318] J. A. Hoogeveen and S. L. van de Velde. A branch-and-bound algorithm for single-machine earliness–tardiness scheduling with idle time. INFORMS Journal on Computing, 8(4):402–412, Fall 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.8.4.402.

Abate:1996:LMN

[319] Joseph Abate, Gagan L. Choudhury, and Ward Whitt. On the Laguerre method for numerically inverting Laplace transforms. INFORMS Journal on Computing, 8(4):413–427, Fall 1996. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.8.4.413.

Korst:1996:SPT

[320] Jan Korst, Emile Aarts, and Jan Karel Lenstra. Scheduling periodic tasks. INFORMS Journal on Computing, 8(4):428–435, Fall 1996. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.8.4.428.

Golden:1996:AR

[321] Bruce Golden. Appreciation to referees. INFORMS Journal on Computing, 8(4):436–437, Fall 1996. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.8.4.436.

Bala:1997:MPA

[322] Mohan Bala and Kipp Martin. A mathematical programming approach to data base normalization. INFORMS Journal on Computing, 9(1):1–14, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.9.1.1.

Eckstein:1997:HMC

[323] Jonathan Eckstein. How much communication does parallel branch and bound need? INFORMS Journal on Computing, 9(1):15–29, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.9.1.15.

Palubeckis:1997:BBA

[324] Gintaras Palubeckis. A branch-and-bound approach using polyhedral results for a clustering problem. INFORMS Journal on Computing, 9 (1):30–42, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.9.1.30.

Babayev:1997:NKS

[325] Djangir A. Babayev, Fred Glover, and Jennifer Ryan. A new knapsack solution approach by integer equivalent aggregation and consistency determination. INFORMS Journal on Computing, 9(1):43–50, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.9.1.43.

Koole:1997:UPS

[326] Ger Koole. On the use of the power series algorithm for general Markov processes, with an application to a Petri net. INFORMS Journal on Computing, 9(1):51–56, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.1.51.

LEcuyer:1997:BLS

[327] Pierre L’Ecuyer. Bad lattice structures for vectors of nonsuccessive values produced by some linear recurrences. INFORMS Journal on Computing, 9(1):57–60, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.9.1.57.

Crainic:1997:TTP

[328] Teodor Gabriel Crainic, Michel Toulouse, and Michel Gendreau. Toward a taxonomy of parallel tabu search heuristics. INFORMS Journal on Computing, 9(1):61–72, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.1.61.

Gondzio:1997:PAL

[329] Jacek Gondzio. Presolve analysis of linear programs prior to applying an interior point method. INFORMS Journal on Computing, 9(1):73–91, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/

ijoc.9.1.73. See addendum [483].

Hoogeveen:1997:ETS

[330] J. A. Hoogeveen and S. L. van de Velde. Earliness-tardiness scheduling around almost equal due dates. INFORMS Journal on Computing, 9 (1):92–99, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.9.1.92.

Bhadury:1997:AGP

[331] J. Bhadury, V. Chandru, A. Maheshwari, and R. Chandrasekaran. Art gallery problems for convex nested polygons. INFORMS Journal on Computing, 9(1):100–110, Winter 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.1.100.

Birge:1997:SAS

[332] John R. Birge. State-of-the-art-survey — stochastic programming: Computation and applications. INFORMS Journal on Computing, 9(2):111– 133, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.9.2.111.

Wright:1997:SPI

[333] Gordon P. Wright, N. Dan Worobetz, Myong Kang, Radha V. Mookerjee, and Radha Chandrasekharan. OR/SM: a prototype integrated modeling environment based on structured modeling. INFORMS Journal on Computing, 9(2):134–153, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.2.134.

McBride:1997:SMF

[334] Richard D. McBride and John W. Mamer. Solving multicommodity flow problems with a primal embedded network simplex algorithm. INFORMS Journal on Computing, 9(2):154–163, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.9.2.154.

Chinneck:1997:FUS

[335] John W. Chinneck. Finding a useful subset of constraints for analysis in an infeasible linear program. INFORMS Journal on Computing, 9 (2):164–174, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.9.2.164.

Choudhury:1997:PSN

[336] Gagan L. Choudhury and Ward Whitt. Probabilistic scaling for the numerical inversion of nonprobability transforms. INFORMS Journal on Computing, 9(2):175–184, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.2.175.

Bischof:1997:CGL

[337] Christian H. Bischof, Ali Bouaricha, Peyvand M. Khademi, and Jorge J. Moré. Computing gradients in large-scale optimization using automatic differentiation. INFORMS Journal on Computing, 9(2):185–194, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.9.2.185.

Deuermeyer:1997:AAD

[338] Bryan L. Deuermeyer, Guy L. Curry, Andrew T. Duchowski, and Srilakshmi Venkatesh. An automatic approach to deadlock detection and resolution in discrete simulation systems. INFORMS Journal on Computing, 9(2):195–205, Spring 1997. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.9.2.195.

LEcuyer:1997:ILS

[339] Pierre L’Ecuyer and Raymond Couture. An implementation of the lattice and spectral tests for multiple recursive linear random number generators. INFORMS Journal on Computing, 9(2):206–217, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.9.2.206.

Cohen:1997:PIG

[340] David M. Cohen, Daniel P. Heyman, Asya Rabinovitch, and Danit Brown. A parallel implementation of the GTH algorithm. INFORMS Journal on Computing, 9(2):218–223, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.9.2.218.

Daley:1997:EWT

[341] D. J. Daley and L. D. Servi. Estimating waiting times from transactional data. INFORMS Journal on Computing, 9(2):224–229, Spring 1997. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.9.2.224.

Reeves:1997:FAG

[342] Colin R. Reeves. Feature article — genetic algorithms for the operations researcher. INFORMS Journal on Computing, 9(3):231–250, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.9.3.231.

Ahuja:1997:CDF

[343] Ravindra K. Ahuja and James B. Orlin. Commentary — developing fitter genetic algorithms. INFORMS Journal on Computing, 9(3):251–253, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.9.3.251.

Kershenbaum:1997:CWG

[344] Aaron Kershenbaum. Commentary — when genetic algorithms work best. INFORMS Journal on Computing, 9(3):254–255, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.9.3.254.

Levine:1997:CGA

[345] David Levine. Commentary — genetic algorithms: A Practitioner’s view. INFORMS Journal on Computing, 9(3):256–259, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.9.3.256.

Ross:1997:CWG

[346] Peter Ross. Commentary — what are genetic algorithms good at? IN-FORMS Journal on Computing, 9(3):260–262, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.9.3.260.

Reeves:1997:RGA

[347] Colin R. Reeves. Rejoinder — genetic algorithms: No panacea, but a valuable tool for the operations researcher. INFORMS Journal on Computing, 9(3):263–265, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.3.263.

Verner:1997:PTL

[348] Oleg V. Verner, Roger L. Wainwright, and Dale A. Schoenefeld. Placing text labels on maps and diagrams using genetic algorithms with masking. INFORMS Journal on Computing, 9(3):266–275, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.9.3.266.

Chakravarthy:1997:FCQ

[349] Srinivas R. Chakravarthy and Lan Bin. A finite capacity queue with nonrenewal input and exponential dynamic group services. INFORMS Journal on Computing, 9(3):276–287, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.9.3.276.

Glen:1997:GUC

[350] Andrew G. Glen, Lawrence M. Leemis, and John H. Drew. A generalized univariate change-of-variable transformation technique. INFORMS Journal on Computing, 9(3):288–295, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.9.3.288.

Fishman:1997:IBM

[351] George S. Fishman and L. Stephen Yarberry. An implementation of the batch means method. INFORMS Journal on Computing, 9(3):296–310, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.9.3.296.

Bennett:1997:POM

[352] Kristin P. Bennett and Erin J. Bredensteiner. A parametric optimization method for machine learning. INFORMS Journal on Computing, 9(3): 311–318, Summer 1997. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.9.3.311.

Scholl:1997:SBB

[353] Armin Scholl and Robert Klein. SALOME: A bidirectional branch-and-bound procedure for assembly line balancing. INFORMS Journal on Computing, 9(4):319–334, Fall 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.4.319.

Pinnoi:1997:BCA

[354] Anulark Pinnoi and Wilbert F. Wilhelm. A branch and cut approach for workload smoothing on assembly lines. INFORMS Journal on Computing, 9(4):335–350, Fall 1997. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.9.4.335.

Korst:1997:SPT

[355] Jan Korst, Emile Aarts, and Jan Karel Lenstra. Scheduling periodic tasks with slack. INFORMS Journal on Computing, 9(4):351–362, Fall 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.9.4.351.

Kijowski:1997:ORA

[356] B. A. Kijowski and U. S. Palekar. Optimal resource assignment of preemptive periodic tasks on multiple processors. INFORMS Journal on Computing, 9(4):363–373, Fall 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.9.4.363.

Kaku:1997:TSH

[357] Bharat K. Kaku and Joseph B. Mazzola. A tabu-search heuristic for the dynamic plant layout problem. INFORMS Journal on Computing, 9(4):374–384, Fall 1997. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.9.4.374.

Murphy:1997:AVI

[358] Catherine K. Murphy and Michel Benaroch. Adding value to induced decision trees for time-sensitive data. INFORMS Journal on Computing, 9(4):385–396, Fall 1997. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.9.4.385.

Chari:1997:IMS

[359] Kaushal Chari and Tarun K. Sen. An integrated modeling system for structured modeling using model graphs. INFORMS Journal on Computing, 9(4):397–416, Fall 1997. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.9.4.397.

Chiang:1997:RTS

[360] Wen-Chyuan Chiang and Robert A. Russell. A reactive tabu search metaheuristic for the vehicle routing problem with time Windows. INFORMS Journal on Computing, 9(4):417–430, Fall 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.9.4.417.

Cho:1997:DFD

[361] Geon Cho and Dong X. Shaw. A depth-first dynamic programming algorithm for the tree knapsack problem. INFORMS Journal on Computing, 9(4):431–438, Fall 1997. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.9.4.431.

Altinel:1997:VQA

[362] İ. Kuban Altınel, John Oommen, and Necati Aras. Vector quantization for arbitrary distance function estimation. INFORMS Journal on Computing, 9(4):439–451, Fall 1997. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.9.4.439.

Dahl:1998:CPA

[363] Geir Dahl and Mechthild Stoer. A cutting plane algorithm for multicommodity survivable network design problems. INFORMS Journal on Computing, 10(1):1–11, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.1.1.

Bartolacci:1998:VCA

[364] Michael R. Bartolacci and S. David Wu. A virtual clustering approach for routing problems in telecommunication networks. INFORMS Journal on Computing, 10(1):12–24, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.1.12.

Kaefer:1998:ILF

[365] Frederick Kaefer and June S. Park. Interconnecting LANs and a FDDI backbone using transparent bridges: A model and solution algorithms. INFORMS Journal on Computing, 10(1):25–39, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.10.1.25.

Mitra:1998:DAP

[366] Sabyasachi Mitra and Ishwar Murthy. A dual ascent procedure with valid inequalities for designing hierarchical network topologies. INFORMS Journal on Computing, 10(1):40–55, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.10.1.40.

Realff:1998:AEB

[367] Matthew J. Realff and George Stephanopoulos. On the application of explanation-based learning to acquire control knowledge for branch and bound algorithms. INFORMS Journal on Computing, 10(1):56–71, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
1.56.

Cario:1998:NMF

[368] Marne C. Cario and Barry L. Nelson. Numerical methods for fitting and simulating autoregressive-to-anything processes. INFORMS Journal on Computing, 10(1):72–81, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.1.72.

Ciardo:1998:DSS

[369] Gianfranco Ciardo, Joshua Gluckman, and David Nicol. Distributed state space generation of discrete-state stochastic models. INFORMS Journal on Computing, 10(1):82–93, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.10.1.82.

Yucesan:1998:CSM

[370] Enver Yücesan and Lee Schruben. Complexity of simulation models: A graph theoretic approach. INFORMS Journal on Computing, 10(1): 94–106, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.10.1.94.

Rothberg:1998:SMO

[371] Edward Rothberg and Bruce Hendrickson. Sparse matrix ordering methods for interior point linear programming. INFORMS Journal on Computing, 10(1):107–113, Winter 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.1.107.

Armstrong:1998:EAC

[372] R. Armstrong, S. Gu, and L. Lei. An efficient algorithm for a class of two-resource allocation problems. INFORMS Journal on Computing, 10(1):114–120, Winter 1998. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.10.1.114.

Moscato:1998:PHF

[373] Pablo Moscato and Michael G. Norman. On the performance of heuristics on finite and infinite fractal instances of the Euclidean traveling salesman problem. INFORMS Journal on Computing, 10(2):121–132, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
2.121.

Fischetti:1998:SOP

[374] Matteo Fischetti, Juan José Salazar González, and Paolo Toth. Solving the orienteering problem through branch-and-cut. INFORMS Journal on Computing, 10(2):133–148, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.2.133.

Ernst:1998:ESA

[375] A. T. Ernst and M. Krishnamoorthy. An exact solution approach based on shortest-paths for p-hub median problems. INFORMS Journal on Computing, 10(2):149–162, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.2.149.

Chou:1998:ASP

[376] Yu-Li Chou, H. Edwin Romeijn, and Robert L. Smith. Approximating shortest paths in large-scale networks with an application to intelligent transportation systems. INFORMS Journal on Computing, 10(2): 163–179, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.10.2.163.

Gouveia:1998:UVR

[377] Luis Gouveia. Using variable redefinition for computing lower bounds for minimum spanning and Steiner trees with hop constraints. INFORMS Journal on Computing, 10(2):180–188, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.10.2.180.

Bhargava:1998:FSA

[378] Hemant K. Bhargava, Ramayya Krishnan, and Peter Piela. On formal semantics and analysis of typed modeling languages: An analysis of ascend. INFORMS Journal on Computing, 10(2):189–208, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.10.2.189.

Bradley:1998:FSM

[379] P. S. Bradley, O. L. Mangasarian, and W. N. Street. Feature selection via mathematical programming. INFORMS Journal on Computing, 10 (2):209–217, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.10.2.209.

Eckstein:1998:OSM

[380] Jonathan Eckstein and Michael C. Ferris. Operator-splitting methods for monotone affine variational inequalities, with a parallel application to optimal control. INFORMS Journal on Computing, 10(2):218–235, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
2.218.

Higle:1998:VRO

[381] Julia L. Higle. Variance reduction and objective function evaluation in stochastic linear programs. INFORMS Journal on Computing, 10(2): 236–247, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.10.2.236.

Maros:1998:SCA

[382] István Maros and Gautam Mitra. Strategies for creating advanced bases for large-scale linear programming problems. INFORMS Journal on Computing, 10(2):248–260, Spring 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.2.248.

Williams:1998:CBI

[383] H. Paul Williams and John M. Wilson. Connections between integer linear programming and constraint logic programming—an overview and introduction to the cluster of articles. INFORMS Journal on Computing, 10(3):261–264, Summer 1998. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.10.3.261.

Proll:1998:ILP

[384] Les Proll and Barbara Smith. Integer linear programming and constraint programming approaches to a template design problem. INFORMS Journal on Computing, 10(3):265–275, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.10.3.265.

Darby-Dowman:1998:PSC

[385] Ken Darby-Dowman and James Little. Properties of some combinatorial optimization problems and their effect on the performance of integer programming and constraint logic programming. INFORMS Journal on Computing, 10(3):276–286, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.3.276.

Bockmayr:1998:BIU

[386] Alexander Bockmayr and Thomas Kasper. Branch and Infer: a unifying framework for integer and finite domain constraint programming. IN-FORMS Journal on Computing, 10(3):287–300, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.10.3.287.

Bennaceur:1998:IBB

[387] Hachemi Bennaceur, Idir Gouachi, and Gérard Plateau. An incremental branch-and-bound method for the satisfiability problem. IN-FORMS Journal on Computing, 10(3):301–308, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.10.3.301.

Gopal:1998:IPC

[388] Ram D. Gopal, Paulo B. Goes, and Robert S. Garfinkel. Interval protection of confidential information in a database. INFORMS Journal on Computing, 10(3):309–322, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.3.309.

Dula:1998:AIF

[389] J. H. Dulá, R. V. Helgason, and N. Venugopal. An algorithm for identifying the frame of a pointed finite conical hull. INFORMS Journal on Computing, 10(3):323–330, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.10.3.323.

Motwani:1998:OSL

[390] Rajeev Motwani, Vijay Saraswat, and Eric Torng. Online scheduling with lookahead: multipass assembly lines. INFORMS Journal on Computing, 10(3):331–340, Summer 1998. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.10.3.331.

Crauwels:1998:LSH

[391] H. A. J. Crauwels, C. N. Potts, and L. N. Van Wassenhove. Local search heuristics for the single machine total weighted tardiness scheduling problem. INFORMS Journal on Computing, 10(3):341–350, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
3.341.

Benhenni:1998:AAU

[392] Robert A. Benhenni. An approximate analysis of a UNIX macro process scheduler. INFORMS Journal on Computing, 10(3):351–357, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
3.351.

Bhargava:1998:FAW

[393] Hemant K. Bhargava and Ramayya Krishnan. Feature article — the world wide Web: Opportunities for operations research and management science. INFORMS Journal on Computing, 10(4):359–383, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
4.359.

Bradley:1998:CDD

[394] Gordon H. Bradley and Arnold H. Buss. Commentary — dynamic, distributed, platform independent OR/MS applications — a network perspective. INFORMS Journal on Computing, 10(4):384–387, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
4.384.

Fourer:1998:CPW

[395] Robert Fourer. Commentary — predictions for Web technologies in optimization. INFORMS Journal on Computing, 10(4):388–389, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
4.388.

Geoffrion:1998:CNH

[396] Arthur M. Geoffrion. Commentary — a new horizon for OR/MS. IN-FORMS Journal on Computing, 10(4):390–392, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.10.4.390.

Trick:1998:RAW

[397] Michael Trick. ]ary — the world wide Web: It’s the customers. IN-FORMS Journal on Computing, 10(4):393–395, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.10.4.393.

Bhargava:1998:RME

[398] Hemant K. Bhargava and Ramayya Krishnan. Rejoinder — OR/MS, electronic commerce, and the virtual INFORMS community. INFORMS Journal on Computing, 10(4):396–398, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.10.4.396.

Sarkar:1998:PAS

[399] Sumit Sarkar and Deb Ghosh. Partitioning the attribute set for a probabilistic reasoning system. INFORMS Journal on Computing, 10(4):399– 416, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.10.4.399.

Boyd:1998:EAS

[400] E. Andrew Boyd, Rusty Burlingame, and Kenneth Lindsay. An efficient algorithm for solving an air traffic management model of the national airspace system. INFORMS Journal on Computing, 10(4):417–426, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
4.417.

Gu:1998:LCI

[401] Zonghao Gu, George L. Nemhauser, and Martin W. P. Savelsbergh. Lifted cover inequalities for 0–1 integer programs: computation. IN-FORMS Journal on Computing, 10(4):427–437, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.10.4.427.

Sewell:1998:BBA

[402] E. C. Sewell. A branch and bound algorithm for the stability number of a sparse graph. INFORMS Journal on Computing, 10(4):438–447, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
4.438.

Harris:1998:DEU

[403] Carl M. Harris and William G. Marchal. Distribution estimation using Laplace transforms. INFORMS Journal on Computing, 10(4):448–458, Fall 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.10.
4.448.

Golden:1998:AR

[404] Bruce L. Golden. Appreciation to referees. INFORMS Journal on Computing, 10(4):459–460, Fall 1998. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.10.4.459.

Michel:1999:LML

[405] Laurent Michel and Pascal Van Hentenryck. LOCALIZER: a modeling language for local search. INFORMS Journal on Computing, 11(1):1–14, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.11.1.1.

Smith:1999:NNC

[406] Kate A. Smith. Neural networks for combinatorial optimization: a review of more than a decade of research. INFORMS Journal on Computing, 11(1):15–34, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.1.15.

Fleischer:1999:ITF

[407] Mark Fleischer and Sheldon H. Jacobson. Information theory and the finite-time behavior of the simulated annealing algorithm: experimental results. INFORMS Journal on Computing, 11(1):35–43, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.org/doi/abs/10.1287/ijoc.11.1.35.

Laguna:1999:GPR

[408] Manuel Laguna and Rafael Marti. GRASP and path relinking for 2-layer straight line crossing minimization. INFORMS Journal on Computing, 11(1):44–52, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.1.44.

Hertz:1999:IPU

[409] Alain Hertz, Gilbert Laporte, and Pierrette Nanchen Hugo. Improvement procedures for the undirected rural postman problem. INFORMS Journal on Computing, 11(1):53–62, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.11.1.53.

Guieu:1999:AIM

[410] Olivier Guieu and John W. Chinneck. Analyzing infeasible mixed-integer and integer linear programs. INFORMS Journal on Computing, 11(1): 63–77, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.1.63.

Chen:1999:SPM

[411] Zhi-Long Chen and Warren B. Powell. Solving parallel machine scheduling problems by column generation. INFORMS Journal on Computing, 11(1):78–94, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.1.78.

Andersen:1999:EPS

[412] Erling D. Andersen. On exploiting problem structure in a basis identification procedure for linear programming. INFORMS Journal on Computing, 11(1):95–103, Winter 1999. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.11.1.95.

Wilhelm:1999:ASA

[413] W. E. Wilhelm and Pradip Som. Analysis of stochastic assembly with GI-distributed assembly time. INFORMS Journal on Computing, 11(1): 104–116, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.1.104.

Gu:1999:LCI

[414] Zonghao Gu, George L. Nemhauser, and Martin W. P. Savelsbergh. Lifted cover inequalities for 0–1 integer programs: Complexity. IN-FORMS Journal on Computing, 11(1):117–123, Winter 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.11.1.117.

Caprara:1999:ESQ

[415] Alberto Caprara, David Pisinger, and Paolo Toth. Exact solution of the quadratic knapsack problem. INFORMS Journal on Computing, 11(2): 125–137, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.2.125.

Cook:1999:CMW

[416] William Cook and André Rohe. Computing minimum-weight perfect matchings. INFORMS Journal on Computing, 11(2):138–148, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.11.
2.138.

Kennington:1999:EDA

[417] Jeffery L. Kennington and Jason E. Whitler. An efficient decomposition algorithm to optimize spare capacity in a telecommunications network. INFORMS Journal on Computing, 11(2):149–160, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.11.2.149.

Kelly:1999:SPB

[418] James P. Kelly and Jiefeng Xu. A set-partitioning-based heuristic for the vehicle routing problem. INFORMS Journal on Computing, 11(2): 161–172, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.2.161.

Linderoth:1999:CSS

[419] J. T. Linderoth and M. W. P. Savelsbergh. A computational study of search strategies for mixed integer programming. INFORMS Journal on Computing, 11(2):173–187, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.11.2.173.

LeBlanc:1999:PRT

[420] Larry J. LeBlanc, Jerome Chifflet, and Philippe Mahey. Packet routing in telecommunication networks with path and flow restrictions. INFORMS Journal on Computing, 11(2):188–197, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.11.2.188.

Fleurent:1999:ICM

[421] Charles Fleurent and Fred Glover. Improved constructive multistart strategies for the quadratic assignment problem using adaptive memory. INFORMS Journal on Computing, 11(2):198–204, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.11.2.198.

Cornuejols:1999:CHS

[422] Gérard Cornuéjols and Milind Dawande. A class of hard small 0–1 programs. INFORMS Journal on Computing, 11(2):205–210, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.11.
2.205.

Woeginger:1999:ASM

[423] Gerhard J. Woeginger. An approximation scheme for minimizing agreeably weighted variance on a single machine. INFORMS Journal on Computing, 11(2):211–216, Spring 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.11.2.211.

Bradley:1999:MPD

[424] P. S. Bradley, Usama M. Fayyad, and O. L. Mangasarian. Mathematical programming for data mining: formulations and challenges. IN-FORMS Journal on Computing, 11(3):217–238, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.11.3.217.

Bhargava:1999:DMD

[425] Hemant K. Bhargava. Data mining by decomposition: Adaptive search for hypothesis generation. INFORMS Journal on Computing, 11(3): 239–247, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.3.239.

Bhattacharyya:1999:DMP

[426] Siddhartha Bhattacharyya. Direct marketing performance modeling using genetic algorithms. INFORMS Journal on Computing, 11(3):248– 257, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.3.248.

Piramuthu:1999:FSF

[427] Selwyn Piramuthu. Feature selection for financial credit-risk evaluation decisions. INFORMS Journal on Computing, 11(3):258–266, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.11.
3.258.

Kumar:1999:SCH

[428] Akhil Kumar and Ignacio Olmeda. A study of composite or hybrid classifiers for knowledge discovery. INFORMS Journal on Computing, 11(3): 267–277, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.3.267.

Mannino:1999:OES

[429] Michael V. Mannino and Vijay S. Mookerjee. Optimizing expert systems: Heuristics for efficiently generating low-cost information acquisition strategies. INFORMS Journal on Computing, 11(3):278–291, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.11.
3.278.

Averbakh:1999:PCA

[430] Igor Averbakh and Oded Berman. Parallel complexity of additive location problems. INFORMS Journal on Computing, 11(3):292–298, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.11.
3.292.

Konana:1999:SBT

[431] Prabhudev Konana and Sudha Ram. Semantics-based transaction processing for real-time databases: The case of automated stock trading. INFORMS Journal on Computing, 11(3):299–315, Summer 1999. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.11.3.299.

Greenberg:1999:MSA

[432] Harvey J. Greenberg. Matrix sensitivity analysis from an interior solution of a linear program. INFORMS Journal on Computing, 11(3): 316–327, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.3.316.

Anonymous:1999:A

[433] Anonymous. APMOD 2000. INFORMS Journal on Computing, 11 (3):328, Summer 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.3.328.

Wolfe:1999:FHT

[434] William J. Wolfe. A fuzzy Hopfield–Tank traveling salesman problem model. INFORMS Journal on Computing, 11(4):329–344, Fall 1999. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.11.4.329.

Lodi:1999:HMA

[435] Andrea Lodi, Silvano Martello, and Daniele Vigo. Heuristic and metaheuristic approaches for a class of two-dimensional bin packing problems. INFORMS Journal on Computing, 11(4):345–357, Fall 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.11.4.345.

Maniezzo:1999:EAN

[436] Vittorio Maniezzo. Exact and approximate nondeterministic tree-search procedures for the quadratic assignment problem. INFORMS Journal on Computing, 11(4):358–369, Fall 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.11.4.358.

Frangioni:1999:BTD

[437] Antonio Frangioni and Giorgio Gallo. A bundle type dual-ascent approach to linear multicommodity min-cost flow problems. INFORMS Journal on Computing, 11(4):370–393, Fall 1999. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.11.4.370.

Abate:1999:CLT

[438] Joseph Abate and Ward Whitt. Computing Laplace transforms for numerical inversion via continued fractions. INFORMS Journal on Computing, 11(4):394–405, Fall 1999. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.11.4.394.

Degraeve:1999:OIS

[439] Zeger Degraeve and Linus Schrage. Optimal integer solutions to industrial cutting stock problems. INFORMS Journal on Computing, 11 (4):406–419, Fall 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.4.406. See note [498].

Schug:1999:PIV

[440] Brett W. Schug and Matthew J. Realff. Propagation of interval values in simple processing networks. INFORMS Journal on Computing, 11 (4):420–431, Fall 1999. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.11.4.420.

Kelton:2000:Ea

[441] W. David Kelton. From the Editor. INFORMS Journal on Computing, 12(1):1, Winter 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.1.1.

Johnson:2000:PLP

[442] Ellis L. Johnson, George L. Nemhauser, and Martin W. P. Savelsbergh. Progress in linear programming-based algorithms for integer programming: an exposition. INFORMS Journal on Computing, 12(1):2–23, Winter 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.12.1.2.11900.

Coffin:2000:SAC

[443] Marie Coffin and Matthew J. Saltzman. Statistical analysis of computational tests of algorithms and heuristics. INFORMS Journal on Computing, 12(1):24–44, Winter 2000. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.12.1.24.11899.

Bixby:2000:PDS

[444] Robert E. Bixby and Alexander Martin. Parallelizing the dual simplex method. INFORMS Journal on Computing, 12(1):45–56, Winter 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.12.
1.45.11902.

Woeginger:2000:WDD

[445] Gerhard J. Woeginger. When does a dynamic programming formulation guarantee the existence of a fully polynomial time approximation scheme (FPTAS)? INFORMS Journal on Computing, 12(1):57–74, Winter 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.12.
1.57.11901.

Pisinger:2000:MAB

[446] David Pisinger. A minimal algorithm for the bounded knapsack problem. INFORMS Journal on Computing, 12(1):75–82, Winter 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.12.1.75.11898.

Kelton:2000:Eb

[447] W. David Kelton. From the Editor. INFORMS Journal on Computing, 12(2):83, Spring 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.2.83.

Datta:2000:HSD

[448] Anindya Datta and Igor R. Viguier. Handling sensor data in rapidly changing environments to support soft real-time requirements. IN-FORMS Journal on Computing, 12(2):84–103, Spring 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.12.2.84.11893.

Averbakh:2000:MRM

[449] Igor Averbakh and Oded Berman. Minmax regret median location on a network under uncertainty. INFORMS Journal on Computing, 12(2): 104–110, Spring 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.2.104.11897.

vandenAkker:2000:TIF

[450] J. M. van den Akker, C. A. J. Hurkens, and M. W. P. Savelsbergh. Time-indexed formulations for machine scheduling problems: column generation. INFORMS Journal on Computing, 12(2):111–124, Spring 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.12.
2.111.11896.

Gutjahr:2000:SBB

[451] W. J. Gutjahr, C. Strauss, and E. Wagner. A stochastic branch-and-bound approach to activity crashing in project management. INFORMS Journal on Computing, 12(2):125–135, Spring 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.12.2.125.11894.

Rosenkrantz:2000:APB

[452] Daniel J. Rosenkrantz, Giri K. Tayi, and S. S. Ravi. Algorithms for path-based placement of inspection stations on networks. INFORMS Journal on Computing, 12(2):136–149, Spring 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.12.2.136.11895.

Gupta:2000:SSH

[453] Jatinder N. D. Gupta, Randall S. Sexton, and Enar A. Tunc. Selecting scheduling heuristics using neural networks. INFORMS Journal on Computing, 12(2):150–162, Spring 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.12.2.150.11893.

Kelton:2000:Ec

[454] W. David Kelton. From the Editor. INFORMS Journal on Computing, 12(3):163, Summer 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.3.163.

Prais:2000:RGA

[455] Marcelo Prais and Celso C. Ribeiro. Reactive GRASP: an application to a matrix decomposition problem in TDMA traffic assignment. IN-FORMS Journal on Computing, 12(3):164–176, Summer 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.12.3.164.12639.

Karisch:2000:SGB

[456] Stefan E. Karisch, Franz Rendl, and Jens Clausen. Solving graph bisection problems with semidefinite programming. INFORMS Journal on Computing, 12(3):177–191, Summer 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.12.3.177.12637.

Aardal:2000:MSB

[457] K. Aardal, R. E. Bixby, C. A. J. Hurkens, A. K. Lenstra, and J. W. Smeltink. Market split and basis reduction: towards a solution of the Cornuéjols–Dawande instances. INFORMS Journal on Computing, 12 (3):192–202, Summer 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.3.192.12635.

Buchholz:2000:CME

[458] Peter Buchholz, Gianfranco Ciardo, Susanna Donatelli, and Peter Kemper. Complexity of memory-efficient Kronecker operations with applications to the solution of Markov models. INFORMS Journal on Computing, 12(3):203–222, Summer 2000. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.12.3.203.12634.

Crainic:2000:SBT

[459] Teodor Gabriel Crainic, Michel Gendreau, and Judith M. Farvolden. A simplex-based tabu search method for capacitated network design. IN-FORMS Journal on Computing, 12(3):223–236, Summer 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.12.3.223.12638.

Gambardella:2000:ACS

[460] Luca Maria Gambardella and Marco Dorigo. An ant colony system hybridized with a new local search for the sequential ordering problem. IN-FORMS Journal on Computing, 12(3):237–255, Summer 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.12.3.237.12636.

Gass:2000:MCM

[461] Saul I. Gass and Donald Gross. In memoriam: Carl M. Harris. INFORMS Journal on Computing, 12(4):257–260, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.12.4.257.11883.

Harris:2000:ITQ

[462] Carl M. Harris, Percy H. Brill, and Martin J. Fischer. Internet-type queues with power-tailed interarrival times and computational methods for their analysis. INFORMS Journal on Computing, 12(4):261–271, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.12.
4.261.11882.

Humphrey:2000:RSS

[463] David G. Humphrey and James R. Wilson. A revised simplex search procedure for stochastic simulation response surface optimization. IN-FORMS Journal on Computing, 12(4):272–283, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.12.4.272.11879.

Sherali:2000:EMR

[464] Hanif D. Sherali, J. Cole Smith, and Youngho Lee. Enhanced model representations for an intra-ring synchronous optical network design problem allowing demand splitting. INFORMS Journal on Computing, 12 (4):284–298, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.4.284.11884.

Vroblefski:2000:GOC

[465] Mark Vroblefski, R. Ramesh, and Stanley Zionts. General open and closed queueing networks with blocking: a unified framework for approximation. INFORMS Journal on Computing, 12(4):299–316, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.12.
4.299.11878.

Diaz-Banez:2000:APS

[466] José-Miguel Díaz-Báñez, Francisco Gómez, and Ferran Hurtado. Approximation of point sets by 1-corner polygonal chains. INFORMS Journal on Computing, 12(4):317–323, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.12.4.317.11880.

Kontogiorgis:2000:PPL

[467] Spyros Kontogiorgis. Practical piecewise-linear approximation for monotropic optimization. INFORMS Journal on Computing, 12(4):324– 340, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.12.4.324.11877.

Sourd:2000:MML

[468] Francis Sourd and Wim Nuijten. Multiple-machine lower bounds for shop-scheduling problems. INFORMS Journal on Computing, 12(4):341– 352, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.12.4.341.11881.

Anonymous:2000:AR

[469] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 12(4):353–355, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.4.353.

Anonymous:2000:AI

[470] Anonymous. Author index. INFORMS Journal on Computing, 12(4): 356–357, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.12.4.356.

Anonymous:2000:SCK

[471] Anonymous. Subject classification and key word index to volume 12. INFORMS Journal on Computing, 12(4):358–360, Fall 2000. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.12.4.358.11885.

Kelton:2001:Ea

[472] W. David Kelton. From the Editor. INFORMS Journal on Computing, 13(1):1, Winter 2001. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.13.1.1.9747.

Gopal:2001:CGD

[473] Ram D. Gopal, R. Ramesh, and Stanley Zionts. Cascade graphs: Design, analysis and algorithms for relational joins. INFORMS Journal on Computing, 13(1):2–28, Winter 2001. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.13.1.2.9746.

Shapiro:2001:FJR

[474] Joel A. Shapiro, Warren B. Powell, and David Bernstein. A flexible Java representation for uncertainty in online operations-research models. INFORMS Journal on Computing, 13(1):29–55, Winter 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.1.29.9749.

Balas:2001:LTD

[475] Egon Balas and Neil Simonetti. Linear time dynamic-programming algorithms for new classes of restricted TSPs: a computational study. IN-FORMS Journal on Computing, 13(1):56–75, Winter 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.1.56.9748.

Ball:2001:FTV

[476] Michael O. Ball and Andrew Vakhutinsky. Fault-tolerant virtual path layout in ATM networks. INFORMS Journal on Computing, 13(1):76–94, Winter 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.13.1.76.9750.

Williams:2001:RAP

[477] H. P. Williams and Hong Yan. Representations of the all different predicate of constraint satisfaction in integer programming. INFORMS Journal on Computing, 13(2):96–103, Spring 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.13.2.96.10515.

Hochbaum:2001:BCC

[478] Dorit S. Hochbaum and Eli V. Olinick. The bounded cycle-cover problem. INFORMS Journal on Computing, 13(2):104–119, Spring 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.2.104.10516.

Glass:2001:SBS

[479] C. A. Glass, C. N. Potts, and V. A. Strusevich. Scheduling batches with sequential job processing for two-machine flow and open shops. IN-FORMS Journal on Computing, 13(2):120–137, Spring 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.2.120.10521.

Blom:2001:OTA

[480] Michiel Blom, Sven O. Krumke, Willem E. de Paepe, and Leen Stougie. The online TSP against fair adversaries. INFORMS Journal on Computing, 13(2):138–148, Spring 2001. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.13.2.138.10517.

Alexopoulos:2001:PBQ

[481] Christos Alexopoulos, David Goldsman, and Gamze Tokol. Properties of batched quadratic-form variance parameter estimators for simulations. INFORMS Journal on Computing, 13(2):149–156, Spring 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.2.149.10518.

Hoogeveen:2001:NAR

[482] Han Hoogeveen, Petra Schuurman, and Gerhard J. Woeginger. Nonapproximability results for scheduling problems with minsum criteria. INFORMS Journal on Computing, 13(2):157–168, Spring 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.2.157.10520.

Meszaros:2001:APA

[483] Csaba Mészáros and Jacek Gondzio. Addendum to: “Presolve analysis of linear programs prior to applying an interior point method” [IN-FORMS J. Comput. 9 (1997), no. 1, 73–91; MR1478041 (99a:90120)] by Gondzio. INFORMS Journal on Computing, 13(2):169–170, Spring 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.13.
2.169.10519. See [329].

Kelton:2001:Eb

[484] W. David Kelton. From the Editor. INFORMS Journal on Computing, 13(3):171, Summer 2001. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.13.3.171.12628.

Chaudhry:2001:MAD

[485] M. L. Chaudhry, U. C. Gupta, and V. Goswami. Modeling and analysis of discrete-time multiserver queues with batch arrivals: GI^{X}/Geom/m. IN-FORMS Journal on Computing, 13(3):172–180, Summer 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.3.172.12627.

Kennington:2001:PRV

[486] Jeffery L. Kennington and Mark W. Lewis. The path restoration version of the spare capacity allocation problem with modularity restrictions: Models, algorithms, and an empirical analysis. INFORMS Journal on Computing, 13(3):181–190, Summer 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.13.3.181.12629.

Linderoth:2001:PLP

[487] Jeff T. Linderoth, Eva K. Lee, and Martin W. P. Savelsbergh. A parallel, linear programming-based heuristic for large-scale set partitioning problems. INFORMS Journal on Computing, 13(3):191–209, Summer 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.13.
3.191.12630.

Chinneck:2001:FHM

[488] John W. Chinneck. Fast heuristics for the maximum feasible subsystem problem. INFORMS Journal on Computing, 13(3):210–223, Summer 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.13.
3.210.12632.

Caprara:2001:SPR

[489] Alberto Caprara, Giuseppe Lancia, and See-Kiong Ng. Sorting permutations by reversals through branch-and-price. INFORMS Journal on Computing, 13(3):224–244, Summer 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.13.3.224.12631.

Chamberland:2001:DPM

[490] Steven Chamberland and Brunilde Sansò. On the design problem of multitechnology networks. INFORMS Journal on Computing, 13(3): 245–256, Summer 2001. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.13.3.245.12633.

Kelton:2001:Ec

[491] W. David Kelton. From the Editor. INFORMS Journal on Computing, 13(4):257, Fall 2001. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.13.4.257.9735.

Jain:2001:AHM

[492] Vipul Jain and Ignacio E. Grossmann. Algorithms for hybrid MILP/CP models for a class of optimization problems. INFORMS Journal on Computing, 13(4):258–276, Fall 2001. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.13.4.258.9733.

Steiger:2001:CPB

[493] Natalie M. Steiger and James R. Wilson. Convergence properties of the batch means method for simulation output analysis. INFORMS Journal on Computing, 13(4):277–293, Fall 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.13.4.277.9737.

Munson:2001:SAL

[494] Todd S. Munson, Francisco Facchinei, Michael C. Ferris, Andreas Fischer, and Christian Kanzow. The semismooth algorithm for large scale complementarity problems. INFORMS Journal on Computing, 13(4): 294–311, Fall 2001. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.13.4.294.9734.

Chen:2001:ING

[495] Huifen Chen. Initialization for NORTA: generation of random vectors with specified marginals and correlations. INFORMS Journal on Computing, 13(4):312–331, Fall 2001. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.13.4.312.9736.

Li:2001:DPB

[496] Xiao-Bai Li, James Sweigart, James Teng, Joan Donohue, and Lori Thombs. A dynamic programming based pruning method for decision trees. INFORMS Journal on Computing, 13(4):332–344, Fall 2001. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.13.4.332.9732.

Vazquez-Abad:2001:PHG

[497] Felisa J. Vázquez-Abad and Sheldon H. Jacobson. Phantom harmonic gradient estimators for nonpreemptive priority queueing systems. IN-FORMS Journal on Computing, 13(4):345–359, Fall 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.13.4.345.9731.

Goulimis:2001:TNO

[498] Constantine N. Goulimis. Technical note on: “Optimal integer solutions to industrial cutting stock problems” [INFORMS J. Comput. 11 (1999), no. 4, 406–419; 1758626] by Z. Degraeve and L. Schrage. INFORMS Journal on Computing, 13(4):360, Fall 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.13.4.360.9738. See [439].

Kelton:2002:Ea

[499] W. David Kelton. From the Editor. INFORMS Journal on Computing, 14(1):1, Winter 2002. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.14.1.1.7711.

Goldsman:2002:RSS

[500] David Goldsman, Seong-Hee Kim, William S. Marshall, and Barry L. Nelson. Ranking and selection for steady-state simulation: procedures and perspectives. INFORMS Journal on Computing, 14(1):2–19, Winter 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
1.2.7710.

Felici:2002:MAL

[501] Giovanni Felici and Klaus Truemper. A MINSAT approach for learning in logic domains. INFORMS Journal on Computing, 14(1):20–36, Winter 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
1.20.7709.

vandenAkker:2002:CCG

[502] Marjan van den Akker, Han Hoogeveen, and Steef van de Velde. Combining column generation and Lagrangean relaxation to solve a singlemachine common due date problem. INFORMS Journal on Computing, 14(1):37–51, Winter 2002. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.14.1.37.7706.

Congram:2002:IDA

[503] Richard K. Congram, Chris N. Potts, and Steef L. van de Velde. An iterated dynasearch algorithm for the single-machine total weighted tardiness scheduling problem. INFORMS Journal on Computing, 14(1):52–67, Winter 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.14.1.52.7712.

Kyparisis:2002:ALS

[504] George J. Kyparisis and Christos Koulamas. Assembly-line scheduling with concurrent operations and parallel machines. INFORMS Journal on Computing, 14(1):68–80, Winter 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.14.1.68.7708.

Brumbaugh-Smith:2002:MMD

[505] James P. Brumbaugh-Smith and Douglas R. Shier. Minimax models for diverse routing. INFORMS Journal on Computing, 14(1):81–95, Winter 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
1.81.7707.

Kelton:2002:Eb

[506] W. David Kelton. From the Editor. INFORMS Journal on Computing, 14(2):97, Spring 2002. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.14.2.97.122.

Watson:2002:CSR

[507] Jean-Paul Watson, Laura Barbulescu, L. Darrell Whitley, and Adele E. Howe. Contrasting structured and random permutation flow-shop scheduling problems: search-space topology and algorithm performance. INFORMS Journal on Computing, 14(2):98–123, Spring 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.14.2.98.120.

Diniz:2002:CDL

[508] Morganna Carmem Diniz, Edmundo de Souza e Silva, and H. Richard Gail. Calculating the distribution of a linear combination of uniform order statistics. INFORMS Journal on Computing, 14(2):124–131, Spring 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
2.124.121.

Applegate:2002:SMM

[509] David Applegate, William Cook, Sanjeeb Dash, and André Rohe. Solution of a min-max vehicle routing problem. INFORMS Journal on Computing, 14(2):132–143, Spring 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.14.2.132.118.

Torvik:2002:MAQ

[510] Vetle I. Torvik and Evangelos Triantaphyllou. Minimizing the average query complexity of learning monotone Boolean functions. INFORMS Journal on Computing, 14(2):144–174, Spring 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.14.2.144.117.

Vredeveld:2002:ECA

[511] Tjark Vredeveld and Cor Hurkens. Experimental comparison of approximation algorithms for scheduling unrelated parallel machines. INFORMS Journal on Computing, 14(2):175–189, Spring 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.14.2.175.119.

Kelton:2002:Ec

[512] W. David Kelton. From the Editor. INFORMS Journal on Computing, 14(3):191, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.14.3.191.114.

Fu:2002:FAO

[513] Michael C. Fu. Feature article: Optimization for simulation: Theory vs. practice. INFORMS Journal on Computing, 14(3):192–215, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
3.192.113.

Andradottir:2002:SOI

[514] Sigrún Andradóttir. Simulation optimization: Integrating research and practice. INFORMS Journal on Computing, 14(3):216–219, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
3.216.112.

Glynn:2002:APS

[515] Peter W. Glynn. Additional perspectives on simulation for optimization. INFORMS Journal on Computing, 14(3):220–222, Summer 2002. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.14.3.220.110.

Kelly:2002:SOE

[516] James P. Kelly. Simulation optimization is evolving. INFORMS Journal on Computing, 14(3):223–225, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.14.3.223.108.

Fu:2002:SOF

[517] Michael C. Fu. Simulation optimization in the future: Evolution or revolution? INFORMS Journal on Computing, 14(3):226–227, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
3.226.107.

Ribeiro:2002:HGP

[518] Celso C. Ribeiro, Eduardo Uchoa, and Renato F. Werneck. A hybrid GRASP with perturbations for the Steiner problem in graphs. IN-FORMS Journal on Computing, 14(3):228–246, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.14.3.228.116.

Riis:2002:CND

[519] Morten Riis and Kim Allan Andersen. Capacitated network design with uncertain demand. INFORMS Journal on Computing, 14(3):247–260, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.14.3.247.111.

Derkic:2002:SCM

[520] Steve Derkic and James E. Stafford. Symbolic computation of moments in priority queues. INFORMS Journal on Computing, 14(3):261–277, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.14.3.261.115.

Rees:2002:EAG

[521] Jackie Rees and Gary J. Koehler. An evolutionary approach to group decision making. INFORMS Journal on Computing, 14(3):278–292, Summer 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
3.278.109.

Chinneck:2002:GES

[522] John W. Chinneck. From the guest Editor: Special issue on the merging of mathematical programming and constraint programming. INFORMS Journal on Computing, 14(4):293–294, Fall 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.14.4.293.2829.

Hooker:2002:LOC

[523] John N. Hooker. Logic, optimization, and constraint programming. INFORMS Journal on Computing, 14(4):295–321, Fall 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.14.4.295.2828.

Fourer:2002:EAM

[524] Robert Fourer and David M. Gay. Extending an algebraic modeling language to support constraint programming. INFORMS Journal on Computing, 14(4):322–344, Fall 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.14.4.322.2825.

VanHentenryck:2002:CIP

[525] Pascal Van Hentenryck. Constraint and integer programming in OPL. INFORMS Journal on Computing, 14(4):345–372, Fall 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.14.4.345.2826.

Davey:2002:EIB

[526] Bruce Davey, Natashia Boland, and Peter J. Stuckey. Efficient intelligent backtracking using linear programming. INFORMS Journal on Computing, 14(4):373–386, Fall 2002. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.14.4.373.2823.

Milano:2002:RIP

[527] Michela Milano, Greger Ottosson, Philippe Refalo, and Erlendur S. Thorsteinsson. The role of integer programming techniques in constraint programming’s global constraints. INFORMS Journal on Computing, 14 (4):387–402, Fall 2002. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.14.4.387.2830.

Focacci:2002:HEA

[528] Filippo Focacci, Andrea Lodi, and Michela Milano. A hybrid exact algorithm for the TSPTW. INFORMS Journal on Computing, 14(4):403–417, Fall 2002. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.14.
4.403.2827.

Kelton:2002:AR

[529] W. David Kelton. Appreciation to referees. INFORMS Journal on Computing, 14(4):418–420, Fall 2002. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.14.4.418.2824.

Kelton:2003:Ea

[530] W. David Kelton. From the Editor. INFORMS Journal on Computing, 15(1):1, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.1.1.15154.

Anonymous:2003:IORa

[531] Anonymous. Institute for Operations Research and the Management Sciences. INFORMS Journal on Computing, 15(1):2, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.15.1.2.15153.

Fu:2003:GAB

[532] Zhiwei Fu, Bruce L. Golden, Shreevardhan Lele, S. Raghavan, and Edward A. Wasil. A genetic algorithm-based approach for building accurate decision trees. INFORMS Journal on Computing, 15(1):3–22, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
1.3.15152.

Lee:2003:LPA

[533] Eva K. Lee, Richard J. Gallagher, and David A. Patterson. A linear programming approach to discriminant analysis with a reservedjudgment region. INFORMS Journal on Computing, 15(1):23–41, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
1.23.15158.

Holmberg:2003:MNF

[534] Kaj Holmberg and Di Yuan. A multicommodity network-flow problem with side constraints on paths solved by column generation. INFORMS Journal on Computing, 15(1):42–57, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.15.1.42.15151.

Degraeve:2003:OIS

[535] Zeger Degraeve and Marc Peeters. Optimal integer solutions to industrial cutting-stock problems: Part 2, benchmark results. INFORMS Journal on Computing, 15(1):58–81, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.15.1.58.15156.

Applegate:2003:CLK

[536] David Applegate, William Cook, and André Rohe. Chained Lin– Kernighan for large traveling salesman problems. INFORMS Journal on Computing, 15(1):82–92, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.15.1.82.15157.

Caprara:2003:RMP

[537] Alberto Caprara. The reversal median problem. INFORMS Journal on Computing, 15(1):93–113, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.15.1.93.15155.

Bussieck:2003:MCT

[538] Michael R. Bussieck, Arne Stolbjerg Drud, and Alexander Meeraus. MINLPLib—a collection of test models for mixed-integer nonlinear programming. INFORMS Journal on Computing, 15(1):114–119, Winter 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
1.114.15159.

Tuzhilin:2003:GCE

[539] Alexander Tuzhilin and Louiqa Raschid. From the guest Co-Editors. INFORMS Journal on Computing, 15(2):121–122, Spring 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.15.2.121.14446.

Shahabi:2003:EAW

[540] Cyrus Shahabi and Farnoush Banaei-Kashani. Efficient and anonymous Web-usage mining for Web personalization. INFORMS Journal on Computing, 15(2):123–147, Spring 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.15.2.123.14444.

Zheng:2003:ESD

[541] Zhiqiang Zheng, Balaji Padmanabhan, and Steven O. Kimbrough. On the existence and significance of data preprocessing biases in Webusage mining. INFORMS Journal on Computing, 15(2):148–170, Spring 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
2.148.14449.

Spiliopoulou:2003:FES

[542] Myra Spiliopoulou, Bamshad Mobasher, Bettina Berendt, and Miki Nakagawa. A framework for the evaluation of session reconstruction heuristics in Web-usage analysis. INFORMS Journal on Computing, 15(2): 171–190, Spring 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.2.171.14445.

Srivastava:2003:WBI

[543] Jaideep Srivastava and Robert Cooley. Web business intelligence: Mining the Web for actionable knowledge. INFORMS Journal on Computing, 15(2):191–207, Spring 2003. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.15.2.191.14447.

Strehl:2003:RBC

[544] Alexander Strehl and Joydeep Ghosh. Relationship-based clustering and visualization for high-dimensional data mining. INFORMS Journal on Computing, 15(2):208–230, Spring 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.15.2.208.14448.

Kelton:2003:Eb

[545] David Kelton. From the Editor. INFORMS Journal on Computing, 15 (3):231, Summer 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.3.231.16081.

Anonymous:2003:IORb

[546] Anonymous. Institute for Operations Research and the Management Sciences. INFORMS Journal on Computing, 15(3):232, Summer 2003. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:/

/pubsonline.informs.org/doi/abs/10.1287/ijoc.15.3.232.16079.

Cook:2003:TMB

[547] William Cook and Paul Seymour. Tour merging via branchdecomposition. INFORMS Journal on Computing, 15(3):233–248, Summer 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
3.233.16078.

Nauss:2003:SGA

[548] Robert M. Nauss. Solving the generalized assignment problem: an optimizing and heuristic approach. INFORMS Journal on Computing, 15 (3):249–266, Summer 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.3.249.16075.

Faroe:2003:GLS

[549] Oluf Faroe, David Pisinger, and Martin Zachariasen. Guided local search for the three-dimensional bin-packing problem. INFORMS Journal on Computing, 15(3):267–283, Summer 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.15.3.267.16080.

deVries:2003:CAS

[550] Sven de Vries and Rakesh V. Vohra. Combinatorial auctions: A survey. INFORMS Journal on Computing, 15(3):284–309, Summer 2003. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:/

/pubsonline.informs.org/doi/abs/10.1287/ijoc.15.3.284.16077.

Martello:2003:EAS

[551] Silvano Martello, Michele Monaci, and Daniele Vigo. An exact approach to the strip-packing problem. INFORMS Journal on Computing, 15(3): 310–319, Summer 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.3.310.16082.

Drezner:2003:NGA

[552] Zvi Drezner. A new genetic algorithm for the quadratic assignment problem. INFORMS Journal on Computing, 15(3):320–330, Summer 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
3.320.16076.

Anonymous:2003:E

[553] Anonymous. From the Editor. INFORMS Journal on Computing, 15(4): 331, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.15.4.331.24892.

Anonymous:2003:IORc

[554] Anonymous. Institute for Operations Research and the Management Sciences. INFORMS Journal on Computing, 15(4):332, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.15.4.332.24891.

Toth:2003:GTS

[555] Paolo Toth and Daniele Vigo. The granular tabu search and its application to the vehicle-routing problem. INFORMS Journal on Computing, 15(4):333–346, Fall 2003. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.15.4.333.24890.

Braysy:2003:RVN

[556] Olli Bräysy. A reactive variable neighborhood search for the vehiclerouting problem with time windows. INFORMS Journal on Computing, 15(4):347–368, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.4.347.24896.

Cappanera:2003:SAP

[557] Paola Cappanera and Antonio Frangioni. Symmetric and asymmetric parallelization of a cost-decomposition algorithm for multicommodity flow problems. INFORMS Journal on Computing, 15(4):369–384, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
4.369.24887.

Cornuejols:2003:KCV

[558] Gérard Cornuéjols, Yanjun Li, and Dieter Vandenbussche. K-cuts: A variation of Gomory mixed integer cuts from the LP tableau. INFORMS Journal on Computing, 15(4):385–396, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.15.4.385.24893.

Ingalls:2003:ECF

[559] Ricki G. Ingalls, Douglas J. Morrice, Enver Yücesan, and Andrew B. Whinston. Execution conditions: a formalization of event cancellation in simulation graphs. INFORMS Journal on Computing, 15(4):397–411, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic).

URL https://pubsonline.informs.org/doi/abs/10.1287/ijoc.15.
4.397.24888.

Grassmann:2003:UEF

[560] Winfried K. Grassmann. The use of eigenvalues for finding equilibrium probabilities of certain Markovian two-dimensional queueing problems.

INFORMS Journal on Computing, 15(4):412–421, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.15.4.412.24889.

Anonymous:2003:AR

[561] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 15(4):422–425, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.4.422.24897.

Anonymous:2003:AI

[562] Anonymous. Author index. INFORMS Journal on Computing, 15(4): 426–427, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.4.426.24895.

Anonymous:2003:KWI

[563] Anonymous. Key word index. INFORMS Journal on Computing, 15 (4):428–429, Fall 2003. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.15.4.428.24894.

Kelton:2004:Ea

[564] W. David Kelton. From the Editor. INFORMS Journal on Computing, 16(1):1, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.16.1.1.27671.

Anonymous:2004:IORa

[565] Anonymous. Institute for Operations Research and the Management Sciences. INFORMS Journal on Computing, 16(1):2, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.16.1.2.27670.

Lee:2004:GCP

[566] Eva K. Lee. Generating cutting planes for mixed integer programming problems in a parallel computing environment. INFORMS Journal on Computing, 16(1):3–26, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1030.0027.

Lin:2004:HGO

[567] Zong-Zhi Lin, James C. Bean, and Chelsea C. White. A hybrid genetic/ optimization algorithm for finite-horizon, partially observed Markov decision processes. INFORMS Journal on Computing, 16(1):27–38, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1020.0024.

Wilhelm:2004:BCA

[568] Wilbert E. Wilhelm and Radu Gadidov. A branch-and-cut approach for a generic multiple-product, assembly-system design problem. INFORMS Journal on Computing, 16(1):39–55, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1020.0025.

Makri:2004:NPS

[569] Alexandra Makri and Diego Klabjan. A new pricing scheme for airline crew scheduling. INFORMS Journal on Computing, 16(1):56–67, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1020.0026.

Bennett:2004:FNN

[570] Mihoko V. Bennett and Thomas R. Willemain. The filtered nearest neighbor method for generating low-discrepancy sequences. INFORMS Journal on Computing, 16(1):68–72, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1020.0004.

Parija:2004:BGB

[571] Gyana R. Parija, Shabbir Ahmed, and Alan J. King. On bridging the gap between stochastic integer programming and MIP solver technologies. INFORMS Journal on Computing, 16(1):73–83, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1020.0005.

Elloumi:2004:NFR

[572] Sourour Elloumi, Martine Labbé, and Yves Pochet. A new formulation and resolution method for the p-center problem. INFORMS Journal on Computing, 16(1):84–94, Winter 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1030.0028.

Menon:2004:SBA

[573] Syam Menon and Ali Amiri. Scheduling banner advertisements on the Web. INFORMS Journal on Computing, 16(1):95–105, Winter 2004. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1020.0003.

Kelton:2004:Eb

[574] W. David Kelton. From the Editor. INFORMS Journal on Computing, 16(2):107, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.16.2.107.33159.

Anonymous:2004:IORb

[575] Anonymous. Institute for Operations Research and the Management Sciences. INFORMS Journal on Computing, 16(2):108, Spring 2004. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:/

/pubsonline.informs.org/doi/abs/10.1287/ijoc.16.2.108.33158.

McBride:2004:ILF

[576] Richard D. McBride and John W. Mamer. Implementing an LU factorization for the embedded network simplex algorithm. INFORMS Journal on Computing, 16(2):109–119, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0086.

dePaepe:2004:CAC

[577] Willem E. de Paepe, Jan Karel Lenstra, Jiri Sgall, René A. Sitters, and Leen Stougie. Computer-aided complexity classification of diala-ride problems. INFORMS Journal on Computing, 16(2):120–132, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0052.

Yagiura:2004:ECA

[578] Mutsunori Yagiura, Toshihide Ibaraki, and Fred Glover. An ejection chain approach for the generalized assignment problem. INFORMS Journal on Computing, 16(2):133–151, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1030.0036.

Shortle:2004:ACW

[579] John F. Shortle, Percy H. Brill, Martin J. Fischer, Donald Gross, and Denise M. B. Masi. An algorithm to compute the waiting time distribution for the M/G/1 queue. INFORMS Journal on Computing, 16(2): 152–161, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1030.0045.

Kennington:2004:GNT

[580] Jeffery L. Kennington and Karen R. Lewis. Generalized networks: the theory of preprocessing and an empirical analysis. INFORMS Journal on Computing, 16(2):162–173, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1030.0031.

Kennington:2004:WTW

[581] Jeffery L. Kennington and Eli V. Olinick. Wavelength translation in WDM networks: optimization models and solution procedures. IN-FORMS Journal on Computing, 16(2):174–187, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0038.

Billionnet:2004:UMI

[582] Alain Billionnet and Éric Soutif. Using a mixed integer programming tool for solving the 0-1 quadratic knapsack problem. INFORMS Journal on Computing, 16(2):188–197, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1030.0029.

Sourd:2004:CAP

[583] Francis Sourd. The continuous assignment problem and its application to preemptive and non-preemptive scheduling with irregular cost functions. INFORMS Journal on Computing, 16(2):198–208, Spring 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0034.

Kelton:2004:Ec

[584] W. David Kelton. From the Editor. INFORMS Journal on Computing, 16(3):209, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.16.3.209.40143.

Anonymous:2004:IORc

[585] Anonymous. Institute for Operations Research and the Management Sciences. INFORMS Journal on Computing, 16(3):210, Summer 2004. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:/

/pubsonline.informs.org/doi/abs/10.1287/ijoc.16.3.210.40145.

Greenberg:2004:OCO

[586] Harvey J. Greenberg, William E. Hart, and Giuseppe Lancia. Opportunities for combinatorial optimization in computational biology. IN-FORMS Journal on Computing, 16(3):211–231, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0073.

Blazewicz:2004:DST

[587] Jacek Błażewicz, Fred Glover, and Marta Kasprzak. DNA sequencing — tabu and scatter search combined. INFORMS Journal on Computing, 16(3):232–240, Summer 2004. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1030.0049.

Kulturel-Konak:2004:ETS

[588] Sadan Kulturel-Konak, Bryan A. Norman, David W. Coit, and Alice E. Smith. Exploiting tabu search memory in constrained problems. IN-FORMS Journal on Computing, 16(3):241–254, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0040.

Chinneck:2004:CCM

[589] John W. Chinneck. The constraint consensus method for finding approximately feasible points in nonlinear programs. INFORMS Journal on Computing, 16(3):255–265, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1030.0046.

Nelson:2004:PPQ

[590] Barry L. Nelson and Michael R. Taaffe. The Pht/Pht/∞ queueing system: Part I — the single node. INFORMS Journal on Computing, 16 (3):266–274, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0070.

Nelson:2004:PPK

[591] Barry L. Nelson and Michael R. Taaffe. The [Pht/Pht/∞]K queueing system: Part II — the multiclass network. INFORMS Journal on Computing, 16(3):275–283, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0071.

Tang:2004:SGM

[592] Hui-Chin Tang and Chiang Kao. Searching for good multiple recursive random number generators via a genetic algorithm. INFORMS Journal on Computing, 16(3):284–290, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1030.0039.

Ariyawansa:2004:NCS

[593] K. A. Ariyawansa and Andrew J. Felt. On a new collection of stochastic linear programming test problems. INFORMS Journal on Computing, 16(3):291–299, Summer 2004. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1030.0037.

Langville:2004:TNK

[594] Amy N. Langville and William J. Stewart. Testing the nearest Kronecker product preconditioner on Markov chains and stochastic automata networks. INFORMS Journal on Computing, 16(3):300–315, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0041.

Pochet:2004:GHP

[595] Yves Pochet and Mathieu Van Vyve. A general heuristic for production planning problems. INFORMS Journal on Computing, 16(3):316–327, Summer 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0042.

Greenberg:2004:GES

[596] Harvey J. Greenberg. From the guest editor for the special issue on computational molecular biology/bioinformatics. INFORMS Journal on Computing, 16(4):329–330, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0100.

Chew:2004:PSO

[597] David S. H. Chew, Kwok Pui Choi, Hans Heidner, and Ming-Ying Leung. Palindromes in SARS and other coronaviruses. INFORMS Journal on Computing, 16(4):331–340, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0087.

Park:2004:SMW

[598] Yonil Park and John L. Spouge. Searching for multiple words in a Markov sequence. INFORMS Journal on Computing, 16(4):341–347, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0095.

Lancia:2004:HPP

[599] Giuseppe Lancia, Maria Cristina Pinotti, and Romeo Rizzi. Haplotyping populations by pure parsimony: complexity of exact and approximation algorithms. INFORMS Journal on Computing, 16(4):348–359, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0085.

Kimmel:2004:CPN

[600] Gad Kimmel, Roded Sharan, and Ron Shamir. Computational problems in noisy SNP and haplotype analysis: block scores, block identification, and population stratification. INFORMS Journal on Computing, 16(4):360–370, Fall 2004. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0088.

Doye:2004:GOM

[601] Jonathan P. K. Doye, Robert H. Leary, Marco Locatelli, and Fabio Schoen. Global optimization of Morse clusters by potential energy transformations. INFORMS Journal on Computing, 16(4):371–379, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0084.

Chazelle:2004:SPA

[602] Bernard Chazelle, Carl Kingsford, and Mona Singh. A semidefinite programming approach to side chain positioning with new rounding strategies. INFORMS Journal on Computing, 16(4):380–392, Fall 2004. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0096.

Andonov:2004:PTM

[603] Rumen Andonov, Stefan Balev, and Nicola Yanev. Protein threading: from mathematical models to parallel implementations. INFORMS Journal on Computing, 16(4):393–405, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0092.

Gray:2004:OES

[604] Genetha Anne Gray, Tamara G. Kolda, Ken Sale, and Malin M. Young. Optimizing an empirical scoring function for transmembrane protein structure determination. INFORMS Journal on Computing, 16(4):406– 418, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0102.

Meneses:2004:OSC

[605] Cláudio N. Meneses, Zhaosong Lu, Carlos A. S. Oliveira, and Panos M. Pardalos. Optimal solutions for the closest-string problem via integer programming. INFORMS Journal on Computing, 16(4):419–429, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0090.

Just:2004:MSA

[606] Winfried Just and Gianluca Della Vedova. Multiple sequence alignment as a facility-location problem. INFORMS Journal on Computing, 16 (4):430–440, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0093.

Arslan:2004:DPB

[607] Abdullah N. Arslan and Ömer Eğecioğlu. Dynamic programming based approximation algorithms for sequence alignment with constraints. IN-FORMS Journal on Computing, 16(4):441–458, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0097.

Gusfield:2004:FSG

[608] Dan Gusfield, Satish Eddhu, and Charles Langley. The fine structure of galls in phylogenetic networks. INFORMS Journal on Computing, 16 (4):459–469, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0099.

Stanton:2004:STD

[609] Courtney Stanton and J. MacGregor Smith. Steiner trees and 3-D macromolecular conformation. INFORMS Journal on Computing, 16(4):470– 485, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0101.

Kelton:2004:AR

[610] W. David Kelton. Appreciation to referees. INFORMS Journal on Computing, 16(4):486–489, Fall 2004. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0114.

Anonymous:2004:AI

[611] Anonymous. Author index. INFORMS Journal on Computing, 16(4): 490–491, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0115.

Anonymous:2004:KWI

[612] Anonymous. Key word index. INFORMS Journal on Computing, 16 (4):492–493, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0116.

Anonymous:2004:IORd

[613] Anonymous. Institute for Operations Research and the Management Sciences: Board of Directors and Editors for 2004. INFORMS Journal on Computing, 16(4):494, Fall 2004. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.16.4.494.

Anonymous:2005:Ea

[614] Anonymous. From the Editor. INFORMS Journal on Computing, 17(1): 1, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0130.

Anonymous:2005:IORa

[615] Anonymous. Institute for Operations Research and the Management Sciences: Board of Directors and Editors for 2005. INFORMS Journal on Computing, 17(1):2, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.17.1.2.

Savage:2005:GEG

[616] Eric L. Savage, Lee W. Schruben, and Enver Yücesan. On the generality of event-graph models. INFORMS Journal on Computing, 17(1):3–9, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0053.

Crainic:2005:MHC

[617] Teodor Gabriel Crainic, Federico Malucelli, Maddalena Nonato, and François Guertin. Meta-heuristics for a class of demand-responsive transit systems. INFORMS Journal on Computing, 17(1):10–24, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0051.

Street:2005:OMD

[618] W. Nick Street. Oblique multicategory decision trees using nonlinear programming. INFORMS Journal on Computing, 17(1):25–31, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0047.

Rojanasoonthon:2005:GPM

[619] Siwate Rojanasoonthon and Jonathan Bard. A GRASP for parallel machine scheduling with time windows. INFORMS Journal on Computing, 17(1):32–51, Winter 2005. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1030.0048.

Demassey:2005:CPB

[620] Sophie Demassey, Christian Artigues, and Philippe Michelon. Constraintpropagation-based cutting planes: an application to the resourceconstrained project scheduling problem. INFORMS Journal on Computing, 17(1):52–65, Winter 2005. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1030.0043.

Tortorella:2005:NSR

[621] Michael Tortorella. Numerical solutions of renewal-type integral equations. INFORMS Journal on Computing, 17(1):66–74, Winter 2005. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0054.

Choi:2005:TMA

[622] Dae W. Choi, Nam K. Kim, and Kyung C. Chae. A two-moment approximation for the GI/G/c queue with finite capacity. INFORMS Journal on Computing, 17(1):75–81, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1030.0058.

Cappanera:2005:LSB

[623] Paola Cappanera and Marco Trubian. A local-search-based heuristic for the demand-constrained multidimensional knapsack problem. INFORMS Journal on Computing, 17(1):82–98, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1030.0050.

Tan:2005:OPP

[624] Yong Tan, Kamran Moinzadeh, and Vijay S. Mookerjee. Optimal processing policies for an e-commerce web server. INFORMS Journal on Computing, 17(1):99–110, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1030.0044.

Campos:2005:CIS

[625] Vicente Campos, Manuel Laguna, and Rafael Martí. Contextindependent scatter and tabu search for permutation problems. IN-FORMS Journal on Computing, 17(1):111–122, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0057.

Savelsbergh:2005:ESL

[626] Martin W. P. Savelsbergh, R. N. Uma, and Joel Wein. An experimental study of LP-based approximation algorithms for scheduling problems. INFORMS Journal on Computing, 17(1):123–136, Winter 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0055.

Anonymous:2005:Eb

[627] Anonymous. From the Editor. INFORMS Journal on Computing, 17 (2):137, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0144.

Anonymous:2005:IORb

[628] Anonymous. Institute for Operations Research and the Management Sciences: Board of Directors and Editors for 2005. INFORMS Journal on Computing, 17(2):138, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.17.2.138.

Fourer:2005:LWX

[629] Robert Fourer, Leo Lopes, and Kipp Martin. LPFML: a W3C XML schema for linear and integer programming. INFORMS Journal on Computing, 17(2):139–158, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0120.

Melkonian:2005:PDB

[630] Vardges Melkonian and Éva Tardos. Primal-dual-based algorithms for a directed network design problem. INFORMS Journal on Computing, 17(2):159–174, Spring 2005. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0066.

deCarvalho:2005:UED

[631] José M. Valério de Carvalho. Using extra dual cuts to accelerate column generation. INFORMS Journal on Computing, 17(2):175–182, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0060.

Avella:2005:NOS

[632] Pasquale Avella, Maurizio Boccia, and Bernardo D’Auria. Near-optimal solutions of large-scale single-machine scheduling problems. INFORMS Journal on Computing, 17(2):183–191, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0069.

Denault:2005:ACC

[633] Michel Denault and Jean-Louis Goffin. The analytic-center cutting-plane method for variational inequalities: a quadratic-cut approach. INFORMS Journal on Computing, 17(2):192–206, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1030.0065.

Camponogara:2005:DCN

[634] Eduardo Camponogara and Sarosh N. Talukdar. Designing communication networks to decompose network control problems. INFORMS Journal on Computing, 17(2):207–223, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1030.0056.

Aiex:2005:GPR

[635] Renata M. Aiex, Mauricio G. C. Resende, Panos M. Pardalos, and Gerardo Toraldo. GRASP with path relinking for three-index assignment. INFORMS Journal on Computing, 17(2):224–247, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0059.

Keon:2005:NPM

[636] Neil Keon and G. “Anand” Anandalingam. A new pricing model for competitive telecommunications services using congestion discounts. IN-FORMS Journal on Computing, 17(2):248–262, Spring 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1030.0062.

Kleijnen:2005:SAR

[637] Jack P. C. Kleijnen, Susan M. Sanchez, Thomas W. Lucas, and Thomas M. Cioppa. State-of-the-art review: A User’s guide to the brave new world of designing simulation experiments. INFORMS Journal on Computing, 17(3):263–289, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1050.0136.

Golden:2005:HSG

[638] Bruce Golden, S. Raghavan, and Daliborka Stanojević. Heuristic search for the generalized minimum spanning tree problem. INFORMS Journal on Computing, 17(3):290–304, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0077.

Vandevelde:2005:LBH

[639] Ann Vandevelde, Han Hoogeveen, Cor Hurkens, and Jan Karel Lenstra. Lower bounds for the head-body-tail problem on parallel machines: A computational study of the multiprocessor flow shop. INFORMS Journal on Computing, 17(3):305–320, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0082.

Smith:2005:AAM

[640] J. Cole Smith and Sheldon H. Jacobson. An analysis of the alias method for discrete random-variate generation. INFORMS Journal on Computing, 17(3):321–327, Summer 2005. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1030.0063.

Brunetta:2005:GPA

[641] Lorenzo Brunetta and Philippe Grégoire. A general purpose algorithm for three-dimensional packing. INFORMS Journal on Computing, 17(3): 328–338, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1030.0068.

Olafsson:2005:IPF

[642] Sigurdur Ólafsson and Jaekyung Yang. Intelligent partitioning for feature selection. INFORMS Journal on Computing, 17(3):339–355, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0104.

Caprara:2005:LSG

[643] Alberto Caprara and Juan-José Salazar-González. Laying out sparse graphs with provably minimum bandwidth. INFORMS Journal on Computing, 17(3):356–373, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0083.

Zeng:2005:ERR

[644] Daniel D. Zeng and J. Leon Zhao. Effective role resolution in workflow management. INFORMS Journal on Computing, 17(3):374–387, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0067.

Anonymous:2005:IORc

[645] Anonymous. Institute for Operations Research and the Management Sciences: Board of Directors and Editors for 2005. INFORMS Journal on Computing, 17(3):388, Summer 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.17.3.388.

Michel:2005:MLC

[646] Laurent Michel and Pascal Van Hentenryck. A modeling layer for constraint-programming libraries. INFORMS Journal on Computing, 17 (4):389–401, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0076.

Hicks:2005:PBDa

[647] Illya V. Hicks. Planar branch decompositions. I. The ratcatcher. IN-FORMS Journal on Computing, 17(4):402–412, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0075.

Hicks:2005:PBDb

[648] Illya V. Hicks. Planar branch decompositions. II. The cycle method. INFORMS Journal on Computing, 17(4):413–421, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0074.

Sun:2005:WSR

[649] Minghe Sun. Warm-start routines for solving augmented weighted Tchebycheff network programs in multiple-objective network programming. INFORMS Journal on Computing, 17(4):422–437, Fall 2005. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0140.

Vaughan:2005:SGH

[650] Diane E. Vaughan, Sheldon H. Jacobson, Shane N. Hall, and Laura A. McLay. Simultaneous generalized hill-climbing algorithms for addressing sets of discrete optimization problems. INFORMS Journal on Computing, 17(4):438–450, Fall 2005. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0064.

Averbakh:2005:MRR

[651] Igor Averbakh. The minmax relative regret median problem on networks. INFORMS Journal on Computing, 17(4):451–461, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0080.

Chandrasekaran:2005:IS

[652] R. Chandrasekaran, Young U. Ryu, Varghese S. Jacob, and Sungchul Hong. Isotonic separation. INFORMS Journal on Computing, 17(4):462– 474, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1030.0061.

Jones:2005:HWD

[653] Joni L. Jones and Gary J. Koehler. A heuristic for winner determination in rule-based combinatorial auctions. INFORMS Journal on Computing, 17(4):475–489, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0072.

Anonymous:2005:IORd

[654] Anonymous. Institute for Operations Research and the Management Sciences: Board of Directors and Editors for 2005. INFORMS Journal on Computing, 17(4):490, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.17.4.490.

Anonymous:2005:AR

[655] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 17(4):491–493, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0164.

Anonymous:2005:AI

[656] Anonymous. Author index. INFORMS Journal on Computing, 17(4): 494–495, Fall 2005. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0165.

Anonymous:2005:KWI

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0166.

Anonymous:2006:E

[658] Anonymous. From the Editor. INFORMS Journal on Computing, 18(1): 1, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.18.1.1.

Anonymous:2006:IOR

[659] Anonymous. Institute for Operations Research and the Management Sciences — Board of Directors and Editors for 2006. INFORMS Journal on Computing, 18(1):2, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.18.1.2.

Kuhl:2006:AMP

[660] Michael E. Kuhl, Sachin G. Sumant, and James R. Wilson. An automated multiresolution procedure for modeling complex arrival processes. INFORMS Journal on Computing, 18(1):3–18, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0113.

Evans:2006:DOS

[661] Diane L. Evans, Lawrence M. Leemis, and John H. Drew. The distribution of order statistics for discrete random variables with applications to bootstrapping. INFORMS Journal on Computing, 18(1):19–30, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0105.

Topaloglu:2006:DPA

[662] Huseyin Topaloglu and Warren B. Powell. Dynamic-programming approximations for stochastic time-staged integer multicommodity-flow problems. INFORMS Journal on Computing, 18(1):31–42, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0079.

Shapiro:2006:MLS

[663] Joel A. Shapiro and Warren B. Powell. A metastrategy for large-scale resource management based on informational decomposition. INFORMS Journal on Computing, 18(1):43–60, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0111.

Frangioni:2006:CSC

[664] Antonio Frangioni and Antonio Manca. A computational study of cost reoptimization for min-cost flow problems. INFORMS Journal on Computing, 18(1):61–70, Winter 2006. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0081.

Monaci:2006:SCB

[665] Michele Monaci and Paolo Toth. A set-covering-based heuristic approach for bin-packing problems. INFORMS Journal on Computing, 18(1):71– 85, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0089.

Cornuejols:2006:EES

[666] Gérard Cornuéjols, Miroslav Karamanov, and Yanjun Li. Early estimates of the size of branch-and-bound trees. INFORMS Journal on Computing, 18(1):86–96, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0107.

Dula:2006:AFF

[667] José H. Dulá and Francisco J. López. Algorithms for the frame of a finitely generated unbounded polyhedron. INFORMS Journal on Computing, 18(1):97–110, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0109.

Anjos:2006:NMP

[668] Miguel F. Anjos and Anthony Vannelli. A new mathematicalprogramming framework for facility-layout design. INFORMS Journal on Computing, 18(1):111–118, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0103.

Agarwal:2006:IAN

[669] Anurag Agarwal, Varghese S. Jacob, and Hasan Pirkul. An improved augmented neural-network approach for scheduling problems. INFORMS Journal on Computing, 18(1):119–128, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0108.

Forrest:2006:CGA

[670] John J. H. Forrest, Jayant Kalagnanam, and Laszlo Ladanyi. A column-generation approach to the multiple knapsack problem with color constraints. INFORMS Journal on Computing, 18(1):129–134, Winter 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0078.

Anonymous:2006:GES

[671] Anonymous. From the guest Editor — special cluster on operations research in electrical and computer engineering. INFORMS Journal on Computing, 18(2):135–136, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0188.

Behjat:2006:ILP

[672] Laleh Behjat, Anthony Vannelli, and William Rosehart. Integer linear programming models for global routing. INFORMS Journal on Computing, 18(2):137–150, Spring 2006. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0127.

Loo:2006:TSF

[673] Sin Ming Loo and B. Earl Wells. Task scheduling in a finite-resource, reconfigurable hardware/software codesign environment. INFORMS Journal on Computing, 18(2):151–172, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0106.

Pai:2006:PBR

[674] Praveen Pai, Emad Gad, Ramachandra Achar, Michel Nakhla, and Roni Khazaka. A projection-based reduction approach to computing sensitivity of steady-state response of nonlinear circuits. INFORMS Journal on Computing, 18(2):173–185, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1050.0150.

Ali:2006:DAV

[675] Shirook M. Ali, Natalia K. Nikolova, and Mohamed H. Bakr. A discrete adjoint variable method for printed-circuit board computer-aided design.

INFORMS Journal on Computing, 18(2):186–196, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0138.

Kucar:2006:UEP

[676] Dorothy Kucar and Anthony Vannelli. Using eigenvectors to partition circuits. INFORMS Journal on Computing, 18(2):197–208, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0154.

Sadykov:2006:IPC

[677] Ruslan Sadykov and Laurence A. Wolsey. Integer programming and constraint programming in solving a multimachine assignment scheduling problem with deadlines and release dates. INFORMS Journal on Computing, 18(2):209–217, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1040.0110.

Chang:2006:MPA

[678] Seok Ho Chang and Dae Won Choi. Modeling and performance analysis of a finite-buffer queue with batch arrivals, batch services, and setup times: the M^{X}/G^{Y} /1/K + B queue with setup times. INFORMS Journal on Computing, 18(2):218–228, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0098.

Dey:2006:OSP

[679] Debabrata Dey, Zhongju Zhang, and Prabuddha De. Optimal synchronization policies for data warehouses. INFORMS Journal on Computing, 18(2):229–242, Spring 2006. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0094.

Dutta:2006:OOC

[680] Kaushik Dutta, Samit Soni, Sridhar Narasimhan, and Anindya Datta. Optimization in object caching. INFORMS Journal on Computing, 18 (2):243–254, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1040.0119.

Bhargava:2006:SCJ

[681] Hemant K. Bhargava, Daewon Sun, and Susan H. Xu. Stockout compensation: joint inventory and price optimization in electronic retailing. INFORMS Journal on Computing, 18(2):255–266, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0091.

Fang:2006:DMB

[682] Xiao Fang, Olivia R. Liu Sheng, Wei Gao, and Balakrishna R. Iyer. A data-mining-based prefetching approach to caching for network storage systems. INFORMS Journal on Computing, 18(2):267–282, Spring 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0142.

Chew:2006:GES

[683] Elaine Chew. From the Guest Editor — special cluster on computation in music. INFORMS Journal on Computing, 18(3):283–284, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0199.

Conklin:2006:SPD

[684] Darrell Conklin and Christina Anagnostopoulou. Segmental pattern discovery in music. INFORMS Journal on Computing, 18(3):285–293, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0122.

Gomez:2006:TDP

[685] Emilia Gómez. Tonal description of polyphonic audio for music content processing. INFORMS Journal on Computing, 18(3):294–304, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0126.

Chew:2006:SIA

[686] Elaine Chew. Slicing it all ways: Mathematical models for tonal induction, approximation, and segmentation using the spiral array. IN-FORMS Journal on Computing, 18(3):305–320, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0134.

Franklin:2006:RNN

[687] Judy A. Franklin. Recurrent neural networks for music computation. IN-FORMS Journal on Computing, 18(3):321–338, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0131.

Uitdenbogerd:2006:MEN

[688] Alexandra L. Uitdenbogerd, Abhijit Chattaraj, and Justin Zobel. Methodologies for evaluation of note-based music-retrieval systems. IN-FORMS Journal on Computing, 18(3):339–347, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0139.

Gamvros:2006:MCM

[689] Ioannis Gamvros, Bruce Golden, and S. Raghavan. The multilevel capacitated minimum spanning tree problem. INFORMS Journal on Computing, 18(3):348–365, Summer 2006. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0123.

Kalvenes:2006:BSL

[690] Joakim Kalvenes, Jeffery Kennington, and Eli Olinick. Base station location and service assignments in W-CDMA networks. INFORMS Journal on Computing, 18(3):366–376, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0129.

Zhu:2006:BCP

[691] Guidong Zhu, Jonathan F. Bard, and Gang Yu. A branch-and-cut procedure for the multimode resource-constrained project-scheduling problem. INFORMS Journal on Computing, 18(3):377–390, Summer 2006. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1040.0121.

Irnich:2006:SPP

[692] Stefan Irnich and Daniel Villeneuve. The shortest-path problem with resource constraints and k-cycle elimination for k ≥ 3. INFORMS Journal on Computing, 18(3):391–406, Summer 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0117.

Kelton:2006:E

[693] W. David Kelton. From the Editor. INFORMS Journal on Computing, 18(4):407, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0211.

Abate:2006:UFN

[694] Joseph Abate and Ward Whitt. A unified framework for numerically inverting Laplace transforms. INFORMS Journal on Computing, 18 (4):408–421, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0137.

Bollapragada:2006:BCC

[695] Ramesh Bollapragada, Yanjun Li, and Uday S. Rao. Budget-constrained, capacitated hub location to maximize expected demand coverage in fixedwireless telecommunication networks. INFORMS Journal on Computing, 18(4):422–432, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0143.

Cordeau:2006:MHG

[696] Jean-François Cordeau, Manlio Gaudioso, Gilbert Laporte, and Luigi Moccia. A memetic heuristic for the generalized quadratic assignment problem. INFORMS Journal on Computing, 18(4):433–443, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1040.0128.

Sun:2006:SFS

[697] Jie Sun and Xinwei Liu. Scenario formulation of stochastic linear programs and the homogeneous self-dual interior-point method. INFORMS Journal on Computing, 18(4):444–454, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0112.

Moonen:2006:EAL

[698] Linda S. Moonen and Frits C. R. Spieksma. Exact algorithms for a loading problem with bounded clique width. INFORMS Journal on Computing, 18(4):455–465, Fall 2006. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0124.

Reinelt:2006:MVM

[699] Gerhard Reinelt and Klaus M. Wenger. Maximally violated mod-p cuts for the capacitated vehicle-routing problem. INFORMS Journal on Computing, 18(4):466–479, Fall 2006. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1040.0125.

Lodi:2006:DBA

[700] Andrea Lodi, Michela Milano, and Louis-Martin Rousseau. Discrepancybased additive bounding procedures. INFORMS Journal on Computing, 18(4):480–493, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0168.

Huo:2006:FFB

[701] Xiaoming Huo, Seoung Bum Kim, Kwok-Leung Tsui, and Shuchun Wang. FBP: A frontier-based tree-pruning algorithm. INFORMS Journal on Computing, 18(4):494–505, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1050.0133.

Tan:2006:ADD

[702] Yong Tan, Yonghua Ji, and Vijay S. Mookerjee. Analyzing documentduplication effects on policies for browser and proxy caching. INFORMS Journal on Computing, 18(4):506–522, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1040.0118.

Anonymous:2006:AR

[703] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 18(4):523–526, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0214.

Anonymous:2006:AI

[704] Anonymous. Author index. INFORMS Journal on Computing, 18(4): 527–528, Fall 2006. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0213.

Anonymous:2006:KWI

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0212.

Mirchandani:2007:E

[706] Prakash Mirchandani. From the Editor. INFORMS Journal on Computing, 19(1):1, Winter 2007. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0223.

Sherali:2007:ELD

[707] Hanif D. Sherali and Churlzu Lim. Enhancing Lagrangian dual optimization for linear programs by obviating nondifferentiability. INFORMS Journal on Computing, 19(1):3–13, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1050.0158.

Ahuja:2007:HAM

[708] Ravindra K. Ahuja, Wei Huang, H. Edwin Romeijn, and Dolores Romero Morales. A heuristic approach to the multi-period single-sourcing problem with production and inventory capacities and perishability constraints. INFORMS Journal on Computing, 19(1):14–26, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0151.

Belov:2007:SOS

[709] Gleb Belov and Guntram Scheithauer. Setup and open-stacks minimization in one-dimensional stock cutting. INFORMS Journal on Computing, 19(1):27–35, Winter 2007. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1050.0132.

Pisinger:2007:UDT

[710] David Pisinger and Mikkel Sigurd. Using decomposition techniques and constraint programming for solving the two-dimensional bin-packing problem. INFORMS Journal on Computing, 19(1):36–51, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0181.

Schuurman:2007:PGL

[711] Petra Schuurman and Tjark Vredeveld. Performance guarantees of local search for multiprocessor scheduling. INFORMS Journal on Computing, 19(1):52–63, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0152.

Tkindt:2007:EPO

[712] Vincent T’kindt, Federico Della Croce, and Jean-Louis Bouquard. Enumeration of Pareto optima for a flowshop scheduling problem with two criteria. INFORMS Journal on Computing, 19(1):64–72, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0167.

DiGiacomo:2007:LCG

[713] Laura Di Giacomo, Giacomo Patrizi, and Emanuele Argento. Linear complementarity as a general solution method to combinatorial problems. INFORMS Journal on Computing, 19(1):73–79, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0146.

Ohlmann:2007:CAH

[714] Jeffrey W. Ohlmann and Barrett W. Thomas. A compressed-annealing heuristic for the traveling salesman problem with time windows. IN-FORMS Journal on Computing, 19(1):80–90, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0145.

Molina:2007:SST

[715] Julian Molina, Manuel Laguna, Rafael Martí, and Rafael Caballero. SSPMO: a scatter tabu search procedure for non-linear multiobjective optimization. INFORMS Journal on Computing, 19(1):91–100, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0149.

Fu:2007:SAD

[716] Michael C. Fu, Jian-Qiang Hu, Chun-Hung Chen, and Xiaoping Xiong. Simulation allocation for determining the best design in the presence of correlated sampling. INFORMS Journal on Computing, 19(1):101–111, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0141.

Shortle:2007:WTD

[717] John F. Shortle, Martin J. Fischer, and Percy H. Brill. Waiting-time distribution of M/D_{N}/1 queues through numerical Laplace inversion. INFORMS Journal on Computing, 19(1):112–120, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0148.

Artalejo:2007:AAM

[718] Jesus R. Artalejo, Antonis Economou, and M. J. Lopez-Herrero. Algorithmic analysis of the maximum queue length in a busy period for the M/M/c retrial queue. INFORMS Journal on Computing, 19(1):121–126, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0156.

Gupta:2007:ILW

[719] Rakesh Gupta, Amitava Bagchi, and Sumit Sarkar. Improving linkage of Web pages. INFORMS Journal on Computing, 19(1):127–136, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0147.

Feng:2007:ISS

[720] Juan Feng, Hemant K. Bhargava, and David M. Pennock. Implementing sponsored search in Web search engines: Computational evaluation of alternative mechanisms. INFORMS Journal on Computing, 19(1): 137–148, Winter 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0135.

Kelton:2007:IE

[721] W. David Kelton. From the interim Editor. INFORMS Journal on Computing, 19(2):149, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1070.0229.

Lada:2007:PWB

[722] Emily K. Lada, James R. Wilson, Natalie M. Steiger, and Jeffrey A. Joines. Performance of a wavelet-based spectral procedure for steadystate simulation analysis. INFORMS Journal on Computing, 19(2):150– 160, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0161.

Hu:2007:ERP

[723] Jiaqiao Hu, Michael C. Fu, Vahid R. Ramezani, and Steven I. Marcus. An evolutionary random policy search algorithm for solving Markov decision processes. INFORMS Journal on Computing, 19(2):161–174, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0155.

Royset:2007:SBO

[724] Johannes O. Royset and R. Kevin Wood. Solving the bi-objective maximum-flow network-interdiction problem. INFORMS Journal on Computing, 19(2):175–184, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0191.

Adomavicius:2007:VSO

[725] Gediminas Adomavicius and Alexander Tuzhilin. Validation sequence optimization: a theoretical approach. INFORMS Journal on Computing, 19(2):185–200, Spring 2007. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1050.0153.

Ingolfsson:2007:SEC

[726] Armann Ingolfsson, Elvira Akhmetshina, Susan Budge, Yongyue Li, and Xudong Wu. A survey and experimental comparison of service-level-approximation methods for nonstationary M(t)/M/s(t) queueing systems with exhaustive discipline. INFORMS Journal on Computing, 19 (2):201–214, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0157.

Riska:2007:ESI

[727] Alma Riska and Evgenia Smirni. ETAQA solutions for infinite Markov processes with repetitive structure. INFORMS Journal on Computing, 19(2):215–228, Spring 2007. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1050.0160.

Andreello:2007:ECB

[728] Giuseppe Andreello, Alberto Caprara, and Matteo Fischetti. Embedding {0, 1/2}-cuts in a branch-and-cut framework: A computational study. INFORMS Journal on Computing, 19(2):229–238, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0162.

Parpas:2007:CAN

[729] Panos Parpas and Berç Rustem. Computational assessment of nested Benders and augmented Lagrangian decomposition for mean-variance multistage stochastic problems. INFORMS Journal on Computing, 19 (2):239–247, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0163.

Zhang:2007:PCS

[730] Zhongju Zhang, Debabrata Dey, and Yong Tan. Pricing communication services with delay guarantee. INFORMS Journal on Computing, 19(2): 248–260, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1050.0159.

Alvarez-Valdes:2007:GPR

[731] Ramón Alvarez-Valdes, Rafael Martí, Jose M. Tamarit, and Antonio Parajón. GRASP and path relinking for the two-dimensional two-stage cutting-stock problem. INFORMS Journal on Computing, 19(2):261– 272, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0169.

Koehler:2007:CON

[732] Gary J. Koehler. Conditions that obviate the no-free-lunch theorems for optimization. INFORMS Journal on Computing, 19(2):273–279, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0194.

Pisinger:2007:SLQ

[733] W. David Pisinger, Anders Bo Rasmussen, and Rune Sandvik. Solution of large quadratic knapsack problems through aggressive reduction. IN-FORMS Journal on Computing, 19(2):280–290, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0172.

Koksalan:2007:EMA

[734] Murat Köksalan and Selcen (Pamuk) Phelps. An evolutionary metaheuristic for approximating preference-nondominated solutions. IN-FORMS Journal on Computing, 19(2):291–301, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0170.

Chi:2007:TAI

[735] Hoi-Ming Chi, Okan K. Ersoy, Herbert Moskowitz, and Kemal Altinkemer. Toward automated intelligent manufacturing systems (AIMS). IN-FORMS Journal on Computing, 19(2):302–312, Spring 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1050.0171.

Chinneck:2007:Ea

[736] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 19(3):313, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0235.

Alexopoulos:2007:ECO

[737] Christos Alexopoulos, Nilay Tanık Argon, David Goldsman, Natalie M. Steiger, Gamze Tokol, and James R. Wilson. Efficient computation of overlapping variance estimators for simulation. INFORMS Journal on Computing, 19(3):314–327, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0198.

Ugray:2007:SSL

[738] Zsolt Ugray, Leon Lasdon, John Plummer, Fred Glover, James Kelly, and Rafael Martí. Scatter search and local NLP solvers: a multistart framework for global optimization. INFORMS Journal on Computing, 19(3):328–340, Summer 2007. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1060.0175.

Avdis:2007:PAI

[739] Efstathios Avdis and Ward Whitt. Power algorithms for inverting Laplace transforms. INFORMS Journal on Computing, 19(3):341–355, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0217.

Cook:2007:CDP

[740] William Cook, Daniel G. Espinoza, and Marcos Goycoolea. Computing with domino-parity inequalities for the traveling salesman problem (TSP). INFORMS Journal on Computing, 19(3):356–365, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0204.

Lim:2007:OFC

[741] Gino J. Lim, Michael C. Ferris, Stephen J. Wright, David M. Shepard, and Matthew A. Earl. An optimization framework for conformal radiation treatment planning. INFORMS Journal on Computing, 19(3): 366–380, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0179.

Homem-de-Mello:2007:SCE

[742] Tito Homem de Mello. A study on the cross-entropy method for rareevent probability estimation. INFORMS Journal on Computing, 19(3): 381–394, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0176.

Pisinger:2007:DPO

[743] David Pisinger. Denser packings obtained in O(n log log n) time. IN-FORMS Journal on Computing, 19(3):395–405, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1060.0192.

Lee:2007:BEI

[744] Jon Lee and François Margot. On a binary-encoded ILP coloring formulation. INFORMS Journal on Computing, 19(3):406–415, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0178.

Ahuja:2007:VLSa

[745] Ravindra K. Ahuja, Jon Goodstein, Amit Mukherjee, James B. Orlin, and Dushyant Sharma. A very large-scale neighborhood search algorithm for the combined through-fleet-assignment model. INFORMS Journal on Computing, 19(3):416–428, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0193.

Bhattacharjee:2007:DGS

[746] Sudip Bhattacharjee, Hong Zhang, R. Ramesh, and Dee H. Andrews. A decomposition and guided simulation methodology for large-scale system design: A study in QoS-capable intranets with fixed and mobile components. INFORMS Journal on Computing, 19(3):429–442, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1050.0173.

Lim:2007:TSH

[747] Andrew Lim and Xingwen Zhang. A two-stage heuristic with ejection pools and generalized ejection chains for the vehicle routing problem with time windows. INFORMS Journal on Computing, 19(3):443–457, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0186.

Bouhtou:2007:TON

[748] Mustapha Bouhtou, Stan van Hoesel, Anton F. van der Kraaij, and Jean-Luc Lutton. Tariff optimization in networks. INFORMS Journal on Computing, 19(3):458–469, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0177.

Carrizosa:2007:SGO

[749] Emilio Carrizosa, Belén Martín-Barragán, Frank Plastria, and Dolores Romero Morales. On the selection of the globally optimal prototype subset for nearest-neighbor classification. INFORMS Journal on Computing, 19(3):470–479, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0183.

Alfa:2007:ECD

[750] Attahiru Sule Alfa and Jungong Xue. Efficient computations for the discrete GI/G/1 system. INFORMS Journal on Computing, 19(3):480–484, Summer 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0190.

Chinneck:2007:Eb

[751] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 19(4):485, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0241.

Zheng:2007:CED

[752] Zhiqiang Zheng and Balaji Padmanabhan. Constructing ensembles from data envelopment analysis. INFORMS Journal on Computing, 19(4): 486–496, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0180.

Regis:2007:SRB

[753] Rommel G. Regis and Christine A. Shoemaker. A stochastic radial basis function method for the global optimization of expensive functions. INFORMS Journal on Computing, 19(4):497–509, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1060.0182.

Chiang:2007:IWC

[754] I. Robert Chiang and Manuel A. Nunez. Improving Web-catalog design for easy product search. INFORMS Journal on Computing, 19(4):510– 519, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0184.

Castro:2007:SPH

[755] Jordi Castro. A shortest-paths heuristic for statistical data protection in positive tables. INFORMS Journal on Computing, 19(4):520–533, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0185.

Rothberg:2007:EAP

[756] Edward Rothberg. An evolutionary algorithm for polishing mixed integer programming solutions. INFORMS Journal on Computing, 19 (4):534–541, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0189.

Hutson:2007:DBF

[757] Kevin R. Hutson, Terri L. Schlosser, and Douglas R. Shier. On the distributed Bellman–Ford algorithm and the looping problem. INFORMS Journal on Computing, 19(4):542–551, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1060.0195.

Hansen:2007:PDV

[758] Pierre Hansen, Jack Brimberg, Dragan Urošević, and Nenad Mladenović. Primal-dual variable neighborhood search for the simple plantlocation problem. INFORMS Journal on Computing, 19(4):552–564, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0196.

Dawande:2007:TGA

[759] Milind Dawande, Rakesh Gupta, Sanjeewa Naranpanawe, and Chelliah Sriskandarajah. A traffic-grooming algorithm for wavelength-routed optical networks. INFORMS Journal on Computing, 19(4):565–574, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0197.

Landa-Silva:2007:ACL

[760] Dario Landa-Silva and Edmund K. Burke. Asynchronous cooperative local search for the office-space-allocation problem. INFORMS Journal on Computing, 19(4):575–587, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0200.

Chari:2007:MIA

[761] Kaushal Chari and Manish Agrawal. Multi-issue automated negotiations using agents. INFORMS Journal on Computing, 19(4):588–595, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0203.

Nino-Mora:2007:FPA

[762] José Niño-Mora. A (2/3)n^{3} fast-pivoting algorithm for the Gittins index and optimal stopping of a Markov chain. INFORMS Journal on Computing, 19(4):596–606, Fall 2007. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1060.0206.

delaBanda:2007:DPM

[763] Maria Garcia de la Banda and Peter J. Stuckey. Dynamic programming to minimize the maximum number of open stacks. INFORMS Journal on Computing, 19(4):607–617, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0205.

Carrabs:2007:VNS

[764] Francesco Carrabs, Jean-François Cordeau, and Gilbert Laporte. Variable neighborhood search for the pickup and delivery traveling salesman problem with LIFO loading. INFORMS Journal on Computing, 19 (4):618–632, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0202.

Hall:2007:RMN

[765] Nicholas G. Hall, Zhixin Liu, and Chris N. Potts. Rescheduling for multiple new orders. INFORMS Journal on Computing, 19(4):633–645, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0209.

Ahuja:2007:VLSb

[766] Ravindra K. Ahuja, Krishna C. Jha, James B. Orlin, and Dushyant Sharma. Very large-scale neighborhood search for the quadratic assignment problem. INFORMS Journal on Computing, 19(4):646–657, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0201.

Anonymous:2007:AR

[767] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 19(4):658–660, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0242.

Anonymous:2007:AI

[768] Anonymous. Author index. INFORMS Journal on Computing, 19(4): 661–662, Fall 2007. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0243.

Anonymous:2007:KWI

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0244.

Anonymous:2008:E

[770] Anonymous. From the Editor. INFORMS Journal on Computing, 20(1): 1, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0257.

Eckstein:2008:MIS

[771] Jonathan Eckstein, Avigdor Gal, and Sarit Reiner. Monitoring an information source under a politeness constraint. INFORMS Journal on Computing, 20(1):3–20, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0207.

Geismar:2008:IPT

[772] H. Neil Geismar, Gilbert Laporte, Lei Lei, and Chelliah Sriskandarajah. The integrated production and transportation scheduling problem for a product with a short lifespan. INFORMS Journal on Computing, 20 (1):21–33, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1060.0208.

Hwang:2008:DGP

[773] San-Yih Hwang and Wan-Shiou Yang. Discovering generalized profileassociation rules for the targeted advertising of new products. INFORMS Journal on Computing, 20(1):34–45, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1060.0210.

Chang:2008:DTB

[774] Namsik Chang and Olivia R. Liu Sheng. Decision-tree-based knowledge discovery: single- vs. multi-decision-tree induction. INFORMS Journal on Computing, 20(1):46–54, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1060.0215.

Cordon:2008:SSP

[775] Oscar Cordón, Sergio Damas, Jose Santamaría, and Rafael Martí. Scatter search for the point-matching problem in 3D image registration. IN-FORMS Journal on Computing, 20(1):55–68, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1060.0216.

Oliveira:2008:CLP

[776] Fernando S. Oliveira. A constraint logic programming algorithm for modeling dynamic pricing. INFORMS Journal on Computing, 20(1):69–77, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1060.0218.

Bai:2008:MFP

[777] Ruibin Bai and Graham Kendall. A model for fresh produce shelfspace allocation and inventory management with freshness-condition-dependent demand. INFORMS Journal on Computing, 20(1):78–85, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0219.

Caramia:2008:NLS

[778] Massimiliano Caramia, Paolo Dell’Olmo, and Giuseppe F. Italiano. Novel local-search-based approaches to university examination timetabling. IN-FORMS Journal on Computing, 20(1):86–99, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0220.

Bapna:2008:MDG

[779] Ravi Bapna, Sanjukta Das, Robert Garfinkel, and Jan Stallaert. A market design for grid computing. INFORMS Journal on Computing, 20 (1):100–111, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0221.

Ogut:2008:IDP

[780] Hulisi Ogut, Huseyin Cavusoglu, and Srinivasan Raghunathan. Intrusiondetection policies for IT security breaches. INFORMS Journal on Computing, 20(1):112–123, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1070.0222.

Homberger:2008:PGA

[781] Jörg Homberger. A parallel genetic algorithm for the multilevel unconstrained lot-sizing problem. INFORMS Journal on Computing, 20(1):

124–132, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0224.

Bigras:2008:TIF

[782] Louis-Philippe Bigras, Michel Gamache, and Gilles Savard. Timeindexed formulations and the total weighted tardiness problem. IN-FORMS Journal on Computing, 20(1):133–142, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0225.

Mercier:2008:EFC

[783] Luc Mercier and Pascal Van Hentenryck. Edge finding for cumulative scheduling. INFORMS Journal on Computing, 20(1):143–153, Winter 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0226.

Tarantilis:2008:HGL

[784] Christos D. Tarantilis, Emmanouil E. Zachariadis, and Chris T. Kiranoudis. A hybrid guided local search for the vehicle-routing problem with intermediate replenishment facilities. INFORMS Journal on Computing, 20(1):154–168, Winter 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0230.

Singhal:2008:NAL

[785] Jaya Singhal and Kalyan Singhal. A noniterative algorithm for the linear-quadratic profit-maximization model for smoothing multiproduct production. INFORMS Journal on Computing, 20(2):169–178, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0227.

He:2008:ACM

[786] Qi-Ming He and Hanqin Zhang. An algorithm for computing minimal Coxian representations. INFORMS Journal on Computing, 20(2):179– 190, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0228.

Buriol:2008:SDS

[787] Luciana S. Buriol, Mauricio G. C. Resende, and Mikkel Thorup. Speeding up dynamic shortest-path algorithms. INFORMS Journal on Computing, 20(2):191–204, Spring 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0231.

Katriel:2008:MNW

[788] Irit Katriel. Matchings in node-weighted convex bipartite graphs. IN-FORMS Journal on Computing, 20(2):205–211, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0232.

Hifi:2008:ACT

[789] Mhand Hifi, Rym M’Hallah, and Toufik Saadi. Algorithms for the constrained two-staged two-dimensional cutting problem. INFORMS Journal on Computing, 20(2):212–221, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1070.0233.

Artiouchine:2008:KKP

[790] Konstantin Artiouchine, Philippe Baptiste, and Juliette Mattioli. The K king problem, an abstract model for computing aircraft landing trajectories: On modeling a dynamic hybrid system with constraints. INFORMS Journal on Computing, 20(2):222–233, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1070.0234.

Touzene:2008:TSP

[791] Abderezak Touzene. A tensor sum preconditioner for stochastic automata networks. INFORMS Journal on Computing, 20(2):234–242, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0236.

Salman:2008:SCL

[792] F. Sibel Salman, R. Ravi, and John N. Hooker. Solving the capacitated local access network design problem. INFORMS Journal on Computing, 20(2):243–254, Spring 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0237.

Nino-Mora:2008:FIA

[793] José Niño-Mora. A faster index algorithm and a computational study for bandits with switching costs. INFORMS Journal on Computing, 20 (2):255–269, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0238.

Irnich:2008:UMS

[794] S. Irnich. A unified modeling and solution framework for vehicle routing and local search-based metaheuristics. INFORMS Journal on Computing, 20(2):270–287, Spring 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0239.

Kunnumkal:2008:ESP

[795] Sumit Kunnumkal and Huseyin Topaloglu. Exploiting the structural properties of the underlying Markov decision problem in the Q-learning algorithm. INFORMS Journal on Computing, 20(2):288–301, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0240.

Malaguti:2008:MAV

[796] Enrico Malaguti, Michele Monaci, and Paolo Toth. A metaheuristic approach for the vertex coloring problem. INFORMS Journal on Computing, 20(2):302–316, Spring 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0245.

Aytug:2008:RMM

[797] Haldun Aytug, Gary J. Koehler, and Ling He. Risk minimization and minimum description for linear discriminant functions. INFORMS Journal on Computing, 20(2):317–331, Spring 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1070.0248.

DellAmico:2008:HEA

[798] Mauro Dell’Amico, Manuel Iori, Silvano Martello, and Michele Monaci. Heuristic and exact algorithms for the identical parallel machine scheduling problem. INFORMS Journal on Computing, 20(3):333–344, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0246.

Bapna:2008:PBW

[799] Ravi Bapna, Paulo Goes, Alok Gupta, and Gilbert Karuga. Predicting bidders’ willingness to pay in online multiunit ascending auctions: Analytical and empirical insights. INFORMS Journal on Computing, 20(3):345–355, Summer 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0247.

Ryu:2008:IIP

[800] Young U. Ryu and Hyeun-Suk Rhee. Improving intrusion prevention models: dual-threshold and dual-filter approaches. INFORMS Journal on Computing, 20(3):356–367, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1070.0249.

Crainic:2008:EPB

[801] Teodor Gabriel Crainic, Guido Perboli, and Roberto Tadei. Extreme point-based heuristics for three-dimensional bin packing. IN-FORMS Journal on Computing, 20(3):368–384, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0250.

Haramoto:2008:EJA

[802] Hiroshi Haramoto, Makoto Matsumoto, Takuji Nishimura, François Panneton, and Pierre L’Ecuyer. Efficient jump ahead for F_{2}-linear random number generators. INFORMS Journal on Computing, 20(3):385–390, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0251.

Menon:2008:OBS

[803] Syam Menon and Rakesh Gupta. Optimal broadcast scheduling in packet radio networks via branch and price. INFORMS Journal on Computing, 20(3):391–399, Summer 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1070.0252.

Pedersen:2008:BMA

[804] Christian Roed Pedersen, Lars Relund Nielsen, and Kim Allan Andersen. The bicriterion multimodal assignment problem: introduction, analysis, and experimental results. INFORMS Journal on Computing, 20(3):

400–411, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0253.

Parreno:2008:MSA

[805] F. Parreño, R. Alvarez-Valdes, J. M. Tamarit, and J. F. Oliveira. A maximal-space algorithm for the container loading problem. IN-FORMS Journal on Computing, 20(3):412–422, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0254.

Bai:2008:TSE

[806] Xue Bai, Rema Padman, Joseph Ramsey, and Peter Spirtes. Tabu searchenhanced graphical models for classification in high dimensions. IN-FORMS Journal on Computing, 20(3):423–437, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0255.

Vielma:2008:LLP

[807] Juan Pablo Vielma, Shabbir Ahmed, and George L. Nemhauser. A lifted linear programming branch-and-bound algorithm for mixed-integer conic quadratic programs. INFORMS Journal on Computing, 20(3):438–450, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1070.0256.

Minella:2008:REM

[808] Gerardo Minella, Rubén Ruiz, and Michele Ciavotta. A review and evaluation of multiobjective algorithms for the flowshop scheduling problem. INFORMS Journal on Computing, 20(3):451–471, Summer 2008. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0258.

Sourd:2008:MBB

[809] Francis Sourd and Olivier Spanjaard. A multiobjective branch-and-bound framework: application to the biobjective spanning tree problem. INFORMS Journal on Computing, 20(3):472–484, Summer 2008. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0260.

Biller:2008:EAM

[810] Bahar Biller and Barry L. Nelson. Evaluation of the ARTAFIT method for fitting time-series input processes for simulation. INFORMS Journal on Computing, 20(3):485–498, Summer 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1070.0261.

Chinneck:2008:E

[811] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 20(4):499, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1070.0299.

Shen:2008:RTS

[812] Yuelin Shen. Reactive tabu search in a team-learning problem. IN-FORMS Journal on Computing, 20(4):500–509, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1070.0259.

Griffiths:2008:ATS

[813] J. D. Griffiths, G. M. Leonenko, and J. E. Williams. Approximation to the transient solution of the M/E_{k}/1 queue. INFORMS Journal on Computing, 20(4):510–515, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1070.0262.

Addis:2008:DPS

[814] B. Addis, M. Locatelli, and F. Schoen. Disk packing in a square: a new global optimization approach. INFORMS Journal on Computing, 20(4):516–524, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0263.

Dolan:2008:KIO

[815] Elizabeth D. Dolan, Robert Fourer, Jean-Pierre Goux, Todd S. Munson, and Jason Sarich. Kestrel: an interface from optimization modeling systems to the NEOS server. INFORMS Journal on Computing, 20 (4):525–538, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0264.

Xiong:2008:TCC

[816] Hui Xiong, Wenjun Zhou, Mark Brodie, and Sheng Ma. Top-k ϕ correlation computation. INFORMS Journal on Computing, 20(4):539–552, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0265.

Andreas:2008:MPA

[817] April K. Andreas and J. Cole Smith. Mathematical programming algorithms for two-path routing problems with reliability considerations. INFORMS Journal on Computing, 20(4):553–564, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1080.0266.

Subramanian:2008:EDS

[818] Shivaram Subramanian and Hanif D. Sherali. An effective deflected subgradient optimization scheme for implementing column generation for large-scale airline crew scheduling problems. INFORMS Journal on Computing, 20(4):565–578, Fall 2008. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1080.0267.

Chen:2008:ESB

[819] Chun-Hung Chen, Donghai He, Michael Fu, and Loo Hay Lee. Efficient simulation budget allocation for selecting an optimal subset. INFORMS Journal on Computing, 20(4):579–595, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0268.

Topaloglu:2008:SAM

[820] Huseyin Topaloglu. A stochastic approximation method to compute bid prices in network revenue management problems. INFORMS Journal on Computing, 20(4):596–610, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0269.

Anjos:2008:CGO

[821] Miguel F. Anjos and Anthony Vannelli. Computing globally optimal solutions for single-row layout problems using semidefinite programming and cutting planes. INFORMS Journal on Computing, 20(4):611–617, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0270.

Blum:2008:BAS

[822] Christian Blum. Beam-ACO for simple assembly line balancing. IN-FORMS Journal on Computing, 20(4):618–627, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1080.0271.

Yang:2008:ECT

[823] Feng Yang, Bruce E. Ankenman, and Barry L. Nelson. Estimating cycle time percentile curves for manufacturing systems via simulation. INFORMS Journal on Computing, 20(4):628–643, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1080.0272.

Birattari:2008:EBL

[824] Mauro Birattari, Prasanna Balaprakash, Thomas Stützle, and Marco Dorigo. Estimation-based local search for stochastic combinatorial optimization using delta evaluations: a case study on the probabilistic traveling salesman problem. INFORMS Journal on Computing, 20(4):644– 658, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0276.

Anonymous:2008:AR

[825] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 20(4):659–661, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0301.

Anonymous:2008:AI

[826] Anonymous. Author index. INFORMS Journal on Computing, 20(4): 662–664, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0302.

Anonymous:2008:KWI

[827] Anonymous. Key word index. INFORMS Journal on Computing, 20 (4):665–666, Fall 2008. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0303.

Agnetis:2009:SMS

[828] Alessandro Agnetis, Arianna Alfieri, and Gaia Nicosia. Single-machine scheduling problems with generalized preemption. INFORMS Journal on Computing, 21(1):1–12, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0273.

Arbib:2009:EAE

[829] C. Arbib and F. Marinelli. Exact and asymptotically exact solutions for a class of assortment problems. INFORMS Journal on Computing, 21 (1):13–25, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0274.

Marti:2009:ASS

[830] Rafael Martí, Abraham Duarte, and Manuel Laguna. Advanced scatter search for the max-cut problem. INFORMS Journal on Computing, 21 (1):26–38, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0275.

DeLoera:2009:POM

[831] Jesús A. De Loera, Raymond Hemmecke, and Matthias Köppe. Pareto optima of multicriteria integer linear programs. INFORMS Journal on Computing, 21(1):39–48, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0277.

Sherali:2009:OCP

[832] Hanif D. Sherali, Antoine G. Hobeika, and Chawalit Jeenanunta. An optimal constrained pruning strategy for decision trees. INFORMS Journal on Computing, 21(1):49–61, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0278.

Aleman:2009:RSA

[833] Dionne M. Aleman, H. Edwin Romeijn, and James F. Dempsey. A response surface approach to beam orientation optimization in intensitymodulated radiation therapy treatment planning. INFORMS Journal on Computing, 21(1):62–76, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0279.

Puerto:2009:MRS

[834] J. Puerto, A. M. Rodríguez-Chía, and A. Tamir. Minimax regret singlefacility ordered median location problems on networks. INFORMS Journal on Computing, 21(1):77–87, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0280.

Avramidis:2009:ECM

[835] Athanassios N. Avramidis, Nabil Channouf, and Pierre L’Ecuyer. Efficient correlation matching for fitting discrete multivariate distributions with arbitrary marginals and normal-copula dependence. INFORMS Journal on Computing, 21(1):88–106, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0281.

Valente:2009:EAM

[836] Christian Valente, Gautam Mitra, Mustapha Sadki, and Robert Fourer. Extending algebraic modelling languages for stochastic programming. INFORMS Journal on Computing, 21(1):107–122, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1080.0282.

Jans:2009:SLS

[837] Raf Jans. Solving lot-sizing problems on parallel identical machines using symmetry-breaking constraints. INFORMS Journal on Computing, 21 (1):123–136, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0283.

Dey:2009:LPB

[838] Santanu S. Dey and Jean-Philippe Richard. Linear-programming-based lifting and its application to primal cutting-plane algorithms. INFORMS Journal on Computing, 21(1):137–150, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0284.

Lancia:2009:SCA

[839] Giuseppe Lancia and Paolo Serafini. A set-covering approach with column generation for parsimony haplotyping. INFORMS Journal on Computing, 21(1):151–166, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0285.

Sourd:2009:NEA

[840] Francis Sourd. New exact algorithms for one-machine earliness-tardiness scheduling. INFORMS Journal on Computing, 21(1):167–175, Winter 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0287.

Chinneck:2009:Ea

[841] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 21(2):177, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0331.

Gosavi:2009:RLT

[842] Abhijit Gosavi. Reinforcement learning: a tutorial survey and recent advances. INFORMS Journal on Computing, 21(2):178–192, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0305.

Andradottir:2009:BEE

[843] Sigrún Andradóttir and Andrei A. Prudius. Balanced explorative and exploitative search with estimation for simulation optimization. INFORMS Journal on Computing, 21(2):193–208, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0309.

Gebremedhin:2009:ECS

[844] Assefaw H. Gebremedhin, Arijit Tarafdar, Alex Pothen, and Andrea Walther. Efficient computation of sparse Hessians using coloring and automatic differentiation. INFORMS Journal on Computing, 21(2):209– 223, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0286.

Baatar:2009:NSE

[845] Davaatseren Baatar, Natashia Boland, Robert Johnston, and Horst W. Hamacher. A new sequential extraction heuristic for optimizing the delivery of cancer radiation treatment using multileaf collimators. INFORMS Journal on Computing, 21(2):224–241, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0288.

Fourer:2009:SFO

[846] Robert Fourer and Leo Lopes. StAMPL: A filtration-oriented modeling tool for multistage stochastic recourse problems. INFORMS Journal on Computing, 21(2):242–256, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0289.

Leyffer:2009:CCF

[847] Sven Leyffer. A complementarity constraint formulation of convex multiobjective optimization problems. INFORMS Journal on Computing, 21(2):257–267, Spring 2009. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1080.0290.

Hvattum:2009:STB

[848] Lars Magnus Hvattum, Arne Løkketangen, and Gilbert Laporte. Scenario tree-based heuristics for stochastic inventory-routing problems. IN-FORMS Journal on Computing, 21(2):268–285, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1080.0291.

Huang:2009:TSL

[849] Zan Huang and Dennis K. J. Lin. The time-series link prediction problem with applications in communication surveillance. INFORMS Journal on Computing, 21(2):286–303, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0292.

Laundy:2009:SHM

[850] Richard Laundy, Michael Perregaard, Gabriel Tavares, Horia Tipi, and Alkis Vazacopoulos. Solving hard mixed-integer programming problems with Xpress–MP: a MIPLIB 2003 case study. INFORMS Journal on Computing, 21(2):304–313, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0293.

Li:2009:SRM

[851] Han-Lin Li, Hao-Chun Lu, Chia-Hui Huang, and Nian-Ze Hu. A superior representation method for piecewise linear functions. INFORMS Journal on Computing, 21(2):314–321, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0294. See note [920].

Mofya:2009:DPA

[852] E. Chisonge Mofya and J. Cole Smith. A dynamic programming algorithm for the generalized minimum filter placement problem on tree structures. INFORMS Journal on Computing, 21(2):322–332, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0295.

Rei:2009:ABD

[853] Walter Rei, Jean-François Cordeau, Michel Gendreau, and Patrick Soriano. Accelerating Benders decomposition by local branching. INFORMS Journal on Computing, 21(2):333–345, Spring 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0296.

Chinneck:2009:ESC

[854] John W. Chinneck. From the Editor — special cluster on highthroughput optimization. INFORMS Journal on Computing, 21(3): 347–348, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0346.

Bussieck:2009:GEO

[855] Michael R. Bussieck, Michael C. Ferris, and Alexander Meeraus. Gridenabled optimization with GAMS. INFORMS Journal on Computing, 21(3):349–362, Summer 2009. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1090.0340.

Michel:2009:TPC

[856] Laurent Michel, Andrew See, and Pascal Van Hentenryck. Transparent parallelization of constraint programming. INFORMS Journal on Computing, 21(3):363–382, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0313.

Xu:2009:CES

[857] Y. Xu, T. K. Ralphs, L. Ladányi, and M. J. Saltzman. Computational experience with a software framework for parallel integer programming. INFORMS Journal on Computing, 21(3):383–397, Summer 2009. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0347.

Ferris:2009:SBS

[858] Michael C. Ferris, Christos T. Maravelias, and Arul Sundaramoorthy. Simultaneous batching and scheduling using dynamic decomposition on a grid. INFORMS Journal on Computing, 21(3):398–410, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0339.

Regis:2009:PSG

[859] Rommel G. Regis and Christine A. Shoemaker. Parallel stochastic global optimization using radial basis functions. INFORMS Journal on Computing, 21(3):411–426, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0325.

Zhang:2009:SBA

[860] Hao Howard Zhang, Leyuan Shi, Robert Meyer, Daryl Nazareth, and Warren D’Souza. Solving beam-angle selection and dose optimization simultaneously via high-throughput computing. INFORMS Journal on Computing, 21(3):427–444, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0297.

Linderoth:2009:IBF

[861] Jeff Linderoth, François Margot, and Greg Thain. Improving bounds on the football pool problem by integer programming and high-throughput computing. INFORMS Journal on Computing, 21(3):445–457, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0334.

Reilly:2009:SOP

[862] Charles H. Reilly. Synthetic optimization problem generation: show us the correlations! INFORMS Journal on Computing, 21(3):458–467, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0330.

Li:2009:CBH

[863] Jingpeng Li, Uwe Aickelin, and Edmund K. Burke. A component-based heuristic search method with evolutionary eliminations for hospital personnel scheduling. INFORMS Journal on Computing, 21(3):468–479, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0298.

Yuan:2009:ECG

[864] Yang Yuan and Suvrajeet Sen. Enhanced cut generation methods for decomposition-based branch and cut for two-stage stochastic mixedinteger programs. INFORMS Journal on Computing, 21(3):480–487, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0300.

Wang:2009:DRT

[865] Xiaoqun Wang. Dimension reduction techniques in quasi-Monte Carlo methods for option pricing. INFORMS Journal on Computing, 21(3): 488–504, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0304.

Burke:2009:SAE

[866] Edmund K. Burke, Graham Kendall, and Glenn Whitwell. A simulated annealing enhancement of the best-fit heuristic for the orthogonal stock-cutting problem. INFORMS Journal on Computing, 21(3): 505–516, Summer 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0306.

Chinneck:2009:Eb

[867] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 21(4):517, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0361.

Potvin:2009:SAR

[868] Jean-Yves Potvin. State-of-the art review—evolutionary algorithms for vehicle routing. INFORMS Journal on Computing, 21(4):518–548, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0312.

Terekhov:2009:CPA

[869] Daria Terekhov, J. Christopher Beck, and Kenneth N. Brown. A constraint programming approach for solving a queueing design and control problem. INFORMS Journal on Computing, 21(4):549–561, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1080.0307.

Ernst:2009:EMM

[870] Andreas T. Ernst, Vicky H. Mak, and Luke R. Mason. An exact method for the minimum cardinality problem in the treatment planning of intensity-modulated radiotherapy. INFORMS Journal on Computing, 21(4):562–574, Fall 2009. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1080.0308.

Rothlauf:2009:EMM

[871] Franz Rothlauf. An encoding in metaheuristics for the minimum communication spanning tree problem. INFORMS Journal on Computing, 21(4):575–584, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1080.0310.

Lei:2009:IPD

[872] Lei Lei, Hua Zhong, and W. Art Chaovalitwongse. On the integrated production and distribution problem with bidirectional flows. INFORMS Journal on Computing, 21(4):585–598, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1080.0311.

Frazier:2009:KGP

[873] Peter Frazier, Warren Powell, and Savas Dayanik. The knowledgegradient policy for correlated normal beliefs. INFORMS Journal on Computing, 21(4):599–613, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0314.

Kumar:2009:ACS

[874] Piyush Kumar and E. Alper Yıldırım. An algorithm and a core set result for the weighted Euclidean one-center problem. INFORMS Journal on Computing, 21(4):614–629, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0315.

Gerhardt:2009:TRP

[875] Ira Gerhardt and Barry L. Nelson. Transforming renewal processes for simulation of nonstationary arrival processes. INFORMS Journal on Computing, 21(4):630–640, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1080.0316.

Cook:2009:NSG

[876] William Cook, Sanjeeb Dash, Ricardo Fukasawa, and Marcos Goycoolea. Numerically safe Gomory mixed-integer cuts. INFORMS Journal on Computing, 21(4):641–649, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0324.

Anonymous:2009:AR

[877] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 21(4):650–653, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0367.

Anonymous:2009:AI

[878] Anonymous. Author index. INFORMS Journal on Computing, 21(4): 654–655, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0362.

Anonymous:2009:KWI

[879] Anonymous. Key word index. INFORMS Journal on Computing, 21(4): 656, Fall 2009. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0363.

Chinneck:2010:Ea

[880] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 22(1):1, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0371.

Powell:2010:FAM

[881] Warren B. Powell. Feature article — merging AI and OR to solve high-dimensional stochastic optimization problems using approximate dynamic programming. INFORMS Journal on Computing, 22(1):2–17, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0349.

Tsitsiklis:2010:CPS

[882] John N. Tsitsiklis. Commentary — perspectives on stochastic optimization over time [MR2663501]. INFORMS Journal on Computing, 22(1):

18–19, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0350.

Ruszczynski:2010:CPD

[883] Andrzej Ruszczyński. Commentary — post-decision states and separable approximations are powerful tools of approximate dynamic programming [MR2663501]. INFORMS Journal on Computing, 22(1):20–22, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0351.

Powell:2010:RLS

[884] Warren B. Powell. Rejoinder — the languages of stochastic optimization [MR2663501; MR2663502; MR2663503]. INFORMS Journal on Computing, 22(1):23–25, Winter 2010. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1090.0352.

Fourer:2010:CCD

[885] Robert Fourer, Chandrakant Maheshwari, Arnold Neumaier, Dominique Orban, and Hermann Schichl. Convexity and concavity detection in computational graphs: tree walks for convexity assessment. INFORMS Journal on Computing, 22(1):26–43, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1090.0321.

Bertsimas:2010:NRO

[886] Dimitris Bertsimas, Omid Nohadani, and Kwong Meng Teo. Nonconvex robust optimization for problems with constraints. INFORMS Journal on Computing, 22(1):44–58, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0319.

Bierlaire:2010:HNG

[887] M. Bierlaire, M. Thémans, and N. Zufferey. A heuristic for nonlinear global optimization. INFORMS Journal on Computing, 22(1):59–70, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0343.

Chick:2010:SSM

[888] Stephen E. Chick, Jürgen Branke, and Christian Schmidt. Sequential sampling to myopically maximize the expected value of information. IN-FORMS Journal on Computing, 22(1):71–80, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0327.

Shin:2010:AFG

[889] Kaeyoung Shin and Raghu Pasupathy. An algorithm for fast generation of bivariate Poisson random vectors. INFORMS Journal on Computing, 22(1):81–92, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0332.

Bastert:2010:GWH

[890] Oliver Bastert, Benjamin Hummel, and Sven de Vries. A generalized Wedelin heuristic for integer programming. INFORMS Journal on Computing, 22(1):93–107, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0328.

Fischetti:2010:PM

[891] Matteo Fischetti and Domenico Salvagnin. Pruning moves. INFORMS Journal on Computing, 22(1):108–119, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1090.0329.

McCollum:2010:SRA

[892] Barry McCollum, Andrea Schaerf, Ben Paechter, Paul McMullan, Rhyd Lewis, Andrew J. Parkes, Luca Di Gaspero, Rong Qu, and Edmund K. Burke. Setting the research agenda in automated timetabling: The second international timetabling competition. INFORMS Journal on Computing, 22(1):120–130, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0320.

Kiwiel:2010:IBA

[893] Krzysztof C. Kiwiel. An inexact bundle approach to cutting-stock problems. INFORMS Journal on Computing, 22(1):131–143, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0326.

Caron:2010:FCA

[894] Richard J. Caron, Tim Traynor, and Shafiu Jibrin. Feasibility and constraint analysis of sets of linear matrix inequalities. INFORMS Journal on Computing, 22(1):144–153, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1090.0323.

Carrizosa:2010:BSV

[895] Emilio Carrizosa, Belen Martin-Barragan, and Dolores Romero Morales. Binarized support vector machines. INFORMS Journal on Computing, 22(1):154–167, Winter 2010. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1090.0317.

Buchheim:2010:EAQ

[896] Christoph Buchheim, Angelika Wiegele, and Lanbo Zheng. Exact algorithms for the quadratic linear ordering problem. INFORMS Journal on Computing, 22(1):168–177, Winter 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0318.

Anonymous:2010:E

[897] Anonymous. From the Editor. INFORMS Journal on Computing, 22 (2):179, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0389.

Chaovalitwongse:2010:NOM

[898] W. Art Chaovalitwongse, Chun-An Chou, Tanya Y. Berger-Wolf, Bhaskar DasGupta, Saad Sheikh, Mary V. Ashley, and Isabel C. Caballero. New optimization model and algorithm for sibling reconstruction from genetic markers. INFORMS Journal on Computing, 22(2): 180–194, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0322.

Catanzaro:2010:CRM

[899] Daniele Catanzaro, Alessandra Godi, and Martine Labbé. A class representative model for pure parsimony haplotyping. INFORMS Journal on Computing, 22(2):195–209, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0333.

Jupiter:2010:TNT

[900] Daniel Jupiter, Jessica Şahutoğlu, and Vincent VanBuren. TreeHugger: A new test for enrichment of gene ontology terms. INFORMS Journal on Computing, 22(2):210–221, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0356.

Fanslau:2010:TSA

[901] Tobias Fanslau and Andreas Bortfeldt. A tree search algorithm for solving the container loading problem. INFORMS Journal on Computing, 22(2):222–235, Spring 2010. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1090.0338.

Dash:2010:TSM

[902] Sanjeeb Dash, Marcos Goycoolea, and Oktay Günlük. Two-step MIR inequalities for mixed integer programs. INFORMS Journal on Computing, 22(2):236–249, Spring 2010. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1090.0337.

Puchinger:2010:MKP

[903] Jakob Puchinger, Günther R. Raidl, and Ulrich Pferschy. The multidimensional knapsack problem: structure and algorithms. INFORMS Journal on Computing, 22(2):250–265, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1090.0344.

Maxwell:2010:ADP

[904] Matthew S. Maxwell, Mateo Restrepo, Shane G. Henderson, and Huseyin Topaloglu. Approximate dynamic programming for ambulance redeployment. INFORMS Journal on Computing, 22(2):266–281, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0345.

Elhedhli:2010:LHH

[905] Samir Elhedhli and Huyu Wu. A Lagrangean heuristic for hub-and-spoke system design with capacity selection and congestion. INFORMS Journal on Computing, 22(2):282–296, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1090.0335.

Irnich:2010:PRC

[906] Stefan Irnich, Guy Desaulniers, Jacques Desrosiers, and Ahmed Hadjar. Path-reduced costs for eliminating arcs in routing and scheduling. IN-FORMS Journal on Computing, 22(2):297–313, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0341.

Hewitt:2010:CEH

[907] Mike Hewitt, George L. Nemhauser, and Martin W. P. Savelsbergh. Combining exact and heuristic approaches for the capacitated fixed-charge network flow problem. INFORMS Journal on Computing, 22(2):314–325, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0348.

Kennington:2010:UTS

[908] Jeffery L. Kennington and Charles D. Nicholson. The uncapacitated time-space fixed-charge network flow problem: An empirical investigation of procedures for arc capacity assignment. INFORMS Journal on Computing, 22(2):326–337, Spring 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0354.

Goossens:2010:ARE

[909] Dries R. Goossens, Rudolf Müller, and Frits C. R. Spieksma. Algorithms for recognizing economic properties in matrix bid combinatorial auctions. INFORMS Journal on Computing, 22(3):339–352, Summer 2010. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0336.

Pardoe:2010:AAM

[910] David Pardoe, Peter Stone, Maytal Saar-Tsechansky, Tayfun Keskin, and Kerem Tomak. Adaptive auction mechanism design and the incorporation of prior knowledge. INFORMS Journal on Computing, 22(3): 353–370, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0353.

Przybylski:2010:RAF

[911] Anthony Przybylski, Xavier Gandibleux, and Matthias Ehrgott. A recursive algorithm for finding all nondominated extreme points in the outcome set of a multiobjective integer programme. INFORMS Journal on Computing, 22(3):371–386, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1090.0342.

Trapp:2010:SOP

[912] Andrew C. Trapp and Oleg A. Prokopyev. Solving the order-preserving submatrix problem via integer programming. INFORMS Journal on Computing, 22(3):387–400, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0358.

Muritiba:2010:APP

[913] Albert E. Fernandes Muritiba, Manuel Iori, Enrico Malaguti, and Paolo Toth. Algorithms for the bin packing problem with conflicts. IN-FORMS Journal on Computing, 22(3):401–415, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0355.

Arns:2010:NAI

[914] M. Arns, P. Buchholz, and A. Panchenko. On the numerical analysis of inhomogeneous continuous-time Markov chains. INFORMS Journal on Computing, 22(3):416–432, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0357.

Bodlaender:2010:VLP

[915] Hans L. Bodlaender, Albert Hendriks, Alexander Grigoriev, and Nadejda V. Grigorieva. The valve location problem in simple network topologies. INFORMS Journal on Computing, 22(3):433–442, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0365.

Erdelyi:2010:DPD

[916] Alexander Erdelyi and Huseyin Topaloglu. A dynamic programming decomposition method for making overbooking decisions over an airline network. INFORMS Journal on Computing, 22(3):443–456, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0359.

Cruz-Cano:2010:LSS

[917] Raul Cruz-Cano, David S. H. Chew, Kwok-Pui Choi, and Ming-Ying Leung. Least-squares support vector machine approach to viral replication origin prediction. INFORMS Journal on Computing, 22(3):457–470, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0360.

Kumar:2010:FPA

[918] Rajeev Kumar, Ram Gopal, and Robert Garfinkel. Freedom of privacy: Anonymous data collection with respondent-defined privacy protection. INFORMS Journal on Computing, 22(3):471–481, Summer 2010. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0364.

Wan:2010:IEE

[919] Hong Wan, Bruce E. Ankenman, and Barry L. Nelson. Improving the efficiency and efficacy of controlled sequential bifurcation for simulation factor screening. INFORMS Journal on Computing, 22(3):482–492, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0366.

Vielma:2010:NSR

[920] Juan Pablo Vielma, Shabbir Ahmed, and George Nemhauser. A note on “A superior representation method for piecewise linear functions” [MR2549129]. INFORMS Journal on Computing, 22(3):493–497, Summer 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0379. See [851].

Chinneck:2010:Eb

[921] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 22(4):499, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0423.

Sherali:2010:IAS

[922] Hanif D. Sherali, Ki-Hwan Bae, and Mohamed Haouari. Integrated airline schedule design and fleet assignment: polyhedral analysis and Benders’ decomposition approach. INFORMS Journal on Computing, 22 (4):500–513, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0368.

Schutt:2010:ISI

[923] Andreas Schutt and Peter J. Stuckey. Incremental satisfiability and implication for UTVPI constraints. INFORMS Journal on Computing, 22 (4):514–527, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0369.

Koksalan:2010:BHL

[924] Murat Köksalan and Banu Soylu. Bicriteria p-hub location problems and evolutionary algorithms. INFORMS Journal on Computing, 22(4):528– 542, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0370.

Watson:2010:SHC

[925] Jean-Paul Watson, Roger J.-B. Wets, and David L. Woodruff. Scalable heuristics for a class of chance-constrained stochastic programs. INFORMS Journal on Computing, 22(4):543–554, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0372.

Abhishek:2010:FOA

[926] Kumar Abhishek, Sven Leyffer, and Jeff Linderoth. FilMINT: an outer approximation-based solver for convex mixed-integer nonlinear programs. INFORMS Journal on Computing, 22(4):555–567, Fall 2010. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1090.0373.

Ahmed:2010:AIM

[927] Shabbir Ahmed, Ozan Gozbasi, Martin Savelsbergh, Ian Crocker, Tim Fox, and Eduard Schreibmann. An automated intensity-modulated radiation therapy planning system. INFORMS Journal on Computing, 22 (4):568–583, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1090.0374.

Bardossy:2010:DBL

[928] M. Gisela Bardossy and S. Raghavan. Dual-based local search for the connected facility location and related problems. INFORMS Journal on Computing, 22(4):584–602, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0375.

Muter:2010:CME

[929] İbrahim Muter, Ş. İlker Birbil, and Güvenç Şahin. Combination of metaheuristic and exact algorithms for solving set covering-type optimization problems. INFORMS Journal on Computing, 22(4):603–619, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1090.0376.

Lunday:2010:NET

[930] Brian J. Lunday, Hanif D. Sherali, and Theodore S. Glickman. The nested event tree model with application to combating terrorism. IN-FORMS Journal on Computing, 22(4):620–634, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0377.

Barnes:2010:MTR

[931] Sean Barnes, Bruce Golden, and Edward Wasil. MRSA transmission reduction using agent-based modeling and simulation. INFORMS Journal on Computing, 22(4):635–646, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1100.0386.

Anonymous:2010:AR

[932] Anonymous. Appreciation to referees. INFORMS Journal on Computing, 22(4):647–650, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0428.

Anonymous:2010:AI

[933] Anonymous. Author index. INFORMS Journal on Computing, 22(4): 651–652, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0426.

Anonymous:2010:KWI

[934] Anonymous. Key word index. INFORMS Journal on Computing, 22 (4):653–654, Fall 2010. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0427.

Beck:2011:CCP

[935] J. Christopher Beck, T. K. Feng, and Jean-Paul Watson. Combining constraint programming and local search for job-shop scheduling. INFORMS Journal on Computing, 23(1):1–14, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0388.

Bak:2011:PBB

[936] Sławomir Ba̧k, Jacek Błażewicz, Grzegorz Pawlak, Maciej Płaza, Edmund K. Burke, and Graham Kendall. A parallel branch-and-bound approach to the rectangular guillotine strip cutting problem. INFORMS Journal on Computing, 23(1):15–25, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0394.

Caprara:2011:DLB

[937] Alberto Caprara, Adam N. Letchford, and Juan-José Salazar-González. Decorous lower bounds for minimum linear arrangement. INFORMS Journal on Computing, 23(1):26–40, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0390.

Contreras:2011:BPL

[938] Ivan Contreras, Juan A. Díaz, and Elena Fernández. Branch and price for large-scale capacitated hub location problems with single assignment. INFORMS Journal on Computing, 23(1):41–55, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0391.

Raghavan:2011:BPW

[939] S. Raghavan and Daliborka Stanojević. Branch and price for WDM optical networks with no bifurcation of flow. INFORMS Journal on Computing, 23(1):56–74, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1100.0382.

Altin:2011:RNL

[940] Ayşegül Altın, Hande Yaman, and Mustafa Ç. Pınar. The robust network loading problem under hose demand uncertainty: Formulation, polyhedral analysis, and computations. INFORMS Journal on Computing, 23 (1):75–89, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0380.

Boschetti:2011:FDL

[941] Marco A. Boschetti, Vittorio Maniezzo, and Matteo Roffilli. A fully distributed Lagrangean solution for a peer-to-peer overlay network design problem. INFORMS Journal on Computing, 23(1):90–104, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0381.

Engineer:2011:DPB

[942] Faramroze G. Engineer, George L. Nemhauser, and Martin W. P. Savelsbergh. Dynamic programming-based column generation on timeexpanded networks: Application to the dial-a-flight problem. INFORMS Journal on Computing, 23(1):105–119, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0384.

delaBanda:2011:STS

[943] Maria Garcia de la Banda, Peter J. Stuckey, and Geoffrey Chu. Solving talent scheduling with dynamic programming. INFORMS Journal on Computing, 23(1):120–137, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1090.0378.

Huang:2011:WDC

[944] Zan Huang and Daniel Dajun Zeng. Why does collaborative filtering work? transaction-based recommendation model validation and selection by analyzing bipartite random graphs. INFORMS Journal on Computing, 23(1):138–152, Winter 2011. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1100.0385.

He:2011:ACP

[945] Qi-Ming He, Hanqin Zhang, and Jungong Xue. Algorithms for Coxianization of phase-type generators. INFORMS Journal on Computing, 23(1):153–164, Winter 2011. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1100.0383.

Houdt:2011:TMG

[946] Benny Van Houdt and Johan S. H. van Leeuwaarden. Triangular M/G/1type and tree-like quasi-birth-death Markov chains. INFORMS Journal on Computing, 23(1):165–171, Winter 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0387.

Anonymous:2011:E

[947] Anonymous. From the Editor. INFORMS Journal on Computing, 23 (2):173, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0453.

Ransbotham:2011:SGC

[948] Sam Ransbotham, Ishwar Murthy, Sabyasachi Mitra, and Sridhar Narasimhan. Sequential grid computing: Models and computational experiments. INFORMS Journal on Computing, 23(2):174–188, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0392.

Megow:2011:DSO

[949] Nicole Megow, Rolf H. Möhring, and Jens Schulz. Decision support and optimization in shutdown and turnaround scheduling. INFORMS Journal on Computing, 23(2):189–204, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0393.

Moallemi:2011:RAM

[950] Ciamac C. Moallemi and Benjamin Van Roy. Resource allocation via message passing. INFORMS Journal on Computing, 23(2):205–219, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0395.

Batun:2011:ORP

[951] Sakine Batun, Brian T. Denton, Todd R. Huschka, and Andrew J. Schaefer. Operating room pooling and parallel surgery processing under uncertainty. INFORMS Journal on Computing, 23(2):220–237, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0396.

Jank:2011:ADD

[952] Wolfgang Jank and Shu Zhang. An automated and data-driven bidding strategy for online auctions. INFORMS Journal on Computing, 23(2):

238–253, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0397.

Nino-Mora:2011:CCI

[953] José Niño-Mora. Computing a classic index for finite-horizon bandits. INFORMS Journal on Computing, 23(2):254–267, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0398.

Carrasco:2011:ENS

[954] Juan A. Carrasco. An efficient and numerically stable method for computing bounds for the interval availability distribution. INFORMS Journal on Computing, 23(2):268–283, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0399.

Dula:2011:ADE

[955] J. H. Dulá. An algorithm for data envelopment analysis. INFORMS Journal on Computing, 23(2):284–296, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0400.

Tafazzoli:2011:PSS

[956] Ali Tafazzoli, James R. Wilson, Emily K. Lada, and Natalie M. Steiger. Performance of skart: A skewness- and autoregression-adjusted batch means procedure for simulation analysis. INFORMS Journal on Computing, 23(2):297–314, Spring 2011. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1100.0401.

Groer:2011:PAV

[957] Chris Groër, Bruce Golden, and Edward Wasil. A parallel algorithm for the vehicle routing problem. INFORMS Journal on Computing, 23 (2):315–330, Spring 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0402.

Samorani:2011:REP

[958] Michele Samorani, Manuel Laguna, Robert Kirk DeLisle, and Daniel C. Weaver. A randomized exhaustive propositionalization approach for molecule classification. INFORMS Journal on Computing, 23(3):331– 345, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0404.

Negoescu:2011:KGA

[959] Diana M. Negoescu, Peter I. Frazier, and Warren B. Powell. The knowledge-gradient algorithm for sequencing experiments in drug discovery. INFORMS Journal on Computing, 23(3):346–363, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0417.

Brice:2011:CBC

[960] Pierre Brice, Wei Jiang, and Guohua Wan. A cluster-based contexttree model for multivariate data streams with applications to anomaly detection. INFORMS Journal on Computing, 23(3):364–376, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0407.

Kumar:2011:LCL

[961] Piyush Kumar and E. Alper Yıldırım. A linearly convergent lineartime first-order algorithm for support vector classification with a core set result. INFORMS Journal on Computing, 23(3):377–391, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0412.

n:2011:PST

[962] Osman Y. Özaltın, Brady Hunsaker, and Andrew J. Schaefer. Predicting the solution time of branch-and-bound algorithms for mixedinteger programs. INFORMS Journal on Computing, 23(3):392–403, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0405.

Elhedhli:2011:BPA

[963] Samir Elhedhli, Lingzi Li, Mariem Gzara, and Joe Naoum-Sawaya. A branch-and-price algorithm for the bin packing problem with conflicts. INFORMS Journal on Computing, 23(3):404–415, Summer 2011. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0406.

Narisetty:2011:LTI

[964] Amar K. Narisetty, Jean-Philippe P. Richard, and George L. Nemhauser. Lifted tableaux inequalities for 0-1 mixed-integer programs: a computational study. INFORMS Journal on Computing, 23(3):416–424, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0413.

Koehler:2011:MEB

[965] Gary J. Koehler. Minimal equivalent binary knapsack inequalities. IN-FORMS Journal on Computing, 23(3):425–429, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0408.

Konak:2011:EOR

[966] Abdullah Konak and Alice E. Smith. Efficient optimization of reliable two-node connected networks: a biobjective approach. INFORMS Journal on Computing, 23(3):430–445, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0409.

Cappanera:2011:CCA

[967] Paola Cappanera and Maria Grazia Scutellà. Color-coding algorithms to the balanced path problem: computational issues. INFORMS Journal on Computing, 23(3):446–459, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1100.0410.

Reich:2011:PSS

[968] Daniel Reich and Leo Lopes. Preprocessing stochastic shortest-path problems with application to PERT activity networks. INFORMS Journal on Computing, 23(3):460–469, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0411.

Shen:2011:RFL

[969] Zuo-Jun Max Shen, Roger Lezhou Zhan, and Jiawei Zhang. The reliable facility location problem: formulations, heuristics, and approximation algorithms. INFORMS Journal on Computing, 23(3):470–482, Summer 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0414.

Tsai:2011:EGA

[970] Jung-Fa Tsai and Ming-Hua Lin. An efficient global approach for posynomial geometric programming problems. INFORMS Journal on Computing, 23(3):483–492, Summer 2011. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1100.0403.

Rennen:2011:ESA

[971] Gijs Rennen, Edwin R. van Dam, and Dick den Hertog. Enhancement of sandwich algorithms for approximating higher-dimensional convex Pareto sets. INFORMS Journal on Computing, 23(4):493–517, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0419.

Pasaniuc:2011:OTD

[972] Bogdan Paşaniuc, Robert Garfinkel, Ion Măndoiu, and Alex Zelikovsky. Optimal testing of digital microfluidic biochips. INFORMS Journal on Computing, 23(4):518–529, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1100.0422.

Clautiaux:2011:NSP

[973] François Clautiaux, Cláudio Alves, José Valério de Carvalho, and Jürgen Rietz. New stabilization procedures for the cutting stock problem. IN-FORMS Journal on Computing, 23(4):530–545, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0415.

Garcia:2011:SLM

[974] Sergio García, Martine Labbé, and Alfredo Marín. Solving large p-median problems with a radius formulation. INFORMS Journal on Computing, 23(4):546–556, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0418.

Kaplan:2011:OSC

[975] Uğur Kaplan, Metin Türkay, Bülent Karasözen, and Lorenz T. Biegler. Optimization of supply chain systems with price elasticity of demand. INFORMS Journal on Computing, 23(4):557–568, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0421.

Elhallaoui:2011:IPS

[976] Issmail Elhallaoui, Abdelmoutalib Metrane, Guy Desaulniers, and François Soumis. An improved primal simplex algorithm for degenerate linear programs. INFORMS Journal on Computing, 23(4):569–577, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0425.

Basu:2011:ETR

[977] Amitabh Basu, Pierre Bonami, Gérard Cornuéjols, and François Margot. Experiments with two-row cuts from degenerate tableaux. INFORMS Journal on Computing, 23(4):578–590, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0437.

Siem:2011:MAU

[978] A. Y. D. Siem, D. den Hertog, and A. L. Hoffmann. A method for approximating univariate convex functions using only function value evaluations. INFORMS Journal on Computing, 23(4):591–604, Fall 2011. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0424.

Domschke:2011:CNL

[979] Pia Domschke, Bjorn Geißler, Oliver Kolb, Jens Lang, Alexander Martin, and Antonio Morsi. Combination of nonlinear and linear optimization of transient gas networks. INFORMS Journal on Computing, 23 (4):605–617, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1100.0429.

Wang:2011:NEA

[980] Yuping Wang, Hong Li, and Chuangyin Dang. A new evolutionary algorithm for a class of nonlinear bilevel programming problems and its global convergence. INFORMS Journal on Computing, 23(4):618–629, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0430.

Bapna:2011:COA

[981] Ravi Bapna, Sanjukta Das, Robert Day, Robert Garfinkel, and Jan Stallaert. A clock-and-offer auction market for grid resources when bidders face stochastic computational needs. INFORMS Journal on Computing, 23(4):630–647, Fall 2011. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1100.0431.

Anonymous:2011:AR

[982] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 23(4):648–651, Fall 2011. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1110.0486.

Anonymous:2011:AI

[983] Anonymous. Author index. INFORMS Journal on Computing, 23(4): 652–653, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0490.

Anonymous:2011:KWI

[984] Anonymous. Key word index. INFORMS Journal on Computing, 23 (4):654–655, Fall 2011. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0491.

Yang:2012:ABT

[985] Jeff X. Yang, John H. Drew, and Lawrence M. Leemis. Automating bivariate transformations. INFORMS Journal on Computing, 24(1):1–9, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0416.

Kaczynski:2012:TQA

[986] William H. Kaczynski, Lawrence M. Leemis, and John H. Drew. Transient queueing analysis. INFORMS Journal on Computing, 24(1):10–28, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0452.

Ingolfsson:2012:ERC

[987] Armann Ingolfsson and Ling Tang. Efficient and reliable computation of birth-death process performance measures. INFORMS Journal on Computing, 24(1):29–41, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1100.0435.

Yu:2012:AFH

[988] Kaiqi Yu, Mei-Ling Huang, and Percy H. Brill. An algorithm for fitting heavy-tailed distributions via generalized hyperexponentials. INFORMS Journal on Computing, 24(1):42–52, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1100.0443.

Angun:2012:ATO

[989] Ebru Angün and Jack Kleijnen. An asymptotic test of optimality conditions in multiresponse simulation optimization. INFORMS Journal on Computing, 24(1):53–65, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1100.0438.

Rossi:2012:CBL

[990] Roberto Rossi, S. Armagan Tarim, and Ramesh Bollapragada. Constraint-based local search for inventory control under stochastic demand and lead time. INFORMS Journal on Computing, 24(1):66–80, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0434.

Gualandi:2012:ESG

[991] Stefano Gualandi and Federico Malucelli. Exact solution of graph coloring problems via constraint programming and column generation. IN-FORMS Journal on Computing, 24(1):81–100, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1100.0436.

Archetti:2012:HHI

[992] Claudia Archetti, Luca Bertazzi, Alain Hertz, and M. Grazia Speranza. A hybrid heuristic for an inventory routing problem. INFORMS Journal on Computing, 24(1):101–116, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1100.0439.

Lee:2012:SAD

[993] Jong-Seok Lee and Dan Zhu. Shilling attack detection — a new approach for a trustworthy recommender system. INFORMS Journal on Computing, 24(1):117–131, Winter 2012. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1100.0440.

Dash:2012:TBF

[994] Sanjeeb Dash, Oktay Günlük, Andrea Lodi, and Andrea Tramontani. A time bucket formulation for the traveling salesman problem with time windows. INFORMS Journal on Computing, 24(1):132–147, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0432.

Adelman:2012:CNO

[995] Daniel Adelman and Diego Klabjan. Computing near-optimal policies in generalized joint replenishment. INFORMS Journal on Computing, 24(1):148–164, Winter 2012. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1100.0433.

Avella:2012:CTS

[996] Pasquale Avella, Maurizio Boccia, and Igor Vasilyev. Computational testing of a separation procedure for the knapsack set with a single continuous variable. INFORMS Journal on Computing, 24(1):165–171, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0441.

Keller:2012:DDT

[997] Brian Keller and Güzin Bayraksan. Disjunctive decomposition for twostage stochastic mixed-binary programs with generalized upper bound constraints. INFORMS Journal on Computing, 24(1):172–186, Winter 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1100.0442.

Delling:2012:CRD

[998] Daniel Delling and Giacomo Nannicini. Core routing on dynamic timedependent road networks. INFORMS Journal on Computing, 24(2):187– 201, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Hahn:2012:LRL

[999] Peter M. Hahn, Yi-Rong Zhu, Monique Guignard, William L. Hightower, and Matthew J. Saltzman. A level-3 reformulation-linearization technique-based bound for the quadratic assignment problem. INFORMS Journal on Computing, 24(2):202–209, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Samorani:2012:DMD

[1000] Michele Samorani and Manuel Laguna. Data-mining-driven neighborhood search. INFORMS Journal on Computing, 24(2):210–227, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Malapert:2012:OCP

[1001] Arnaud Malapert, Hadrien Cambazard, Christelle Guéret, Narendra Jussien, André Langevin, and Louis-Martin Rousseau. An optimal constraint programming approach to the open-shop problem. INFORMS Journal on Computing, 24(2):228–244, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Eirinakis:2012:FAS

[1002] Pavlos Eirinakis, Dimitrios Magos, Ioannis Mourtos, and Panayiotis Miliotis. Finding all stable pairs and solutions to the many-to-many stable matching problem. INFORMS Journal on Computing, 24(2):245–259, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Demirci:2012:EMB

[1003] Mehmet C. Demirci, Andrew J. Schaefer, H. Edwin Romeijn, and Mark S. Roberts. An exact method for balancing efficiency and equity in the liver allocation hierarchy. INFORMS Journal on Computing, 24(2):260–275, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Catanzaro:2012:BME

[1004] Daniele Catanzaro, Martine Labbé, Raffaele Pesenti, and Juan-José Salazar-González. The balanced minimum evolution problem. INFORMS Journal on Computing, 24(2):276–294, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Ghosh:2012:CNR

[1005] Soumyadip Ghosh and Raghu Pasupathy. C-NORTA: a rejection procedure for sampling from the tail of bivariate NORTA distributions. INFORMS Journal on Computing, 24(2):295–310, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Huang:2012:SQA

[1006] Zan Huang and Akhil Kumar. A study of quality and accuracy tradeoffs in process mining. INFORMS Journal on Computing, 24(2):311–327, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Eckstein:2012:IBB

[1007] Jonathan Eckstein and Noam Goldberg. An improved branch-and-bound method for maximum monomial agreement. INFORMS Journal on Computing, 24(2):328–341, Spring 2012. ISSN 1091-9856 (print), 1526-5528 (electronic).

Berbeglia:2012:HTS

[1008] Gerardo Berbeglia, Jean-François Cordeau, and Gilbert Laporte. A hybrid tabu search and constraint programming algorithm for the dynamic dial-a-ride problem. INFORMS Journal on Computing, 24(3):343–355, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0454.

Baldacci:2012:NSS

[1009] Roberto Baldacci, Aristide Mingozzi, and Roberto Roberti. New statespace relaxations for solving the traveling salesman problem with time windows. INFORMS Journal on Computing, 24(3):356–371, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0456.

Yang:2012:DPP

[1010] Yinghui (Catherine) Yang, Balaji Padmanabhan, Hongyan Liu, and Xiaoyu Wang. Discovery of periodic patterns in sequence data: a variancebased approach. INFORMS Journal on Computing, 24(3):372–386, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0457.

Fazel-Zarandi:2012:ULB

[1011] Mohammad M. Fazel-Zarandi and J. Christopher Beck. Using logic-based Benders decomposition to solve the capacity- and distance-constrained plant location problem. INFORMS Journal on Computing, 24(3):387– 398, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0458.

Mansini:2012:CEA

[1012] Renata Mansini and M. Grazia Speranza. CORAL: an exact algorithm for the multidimensional knapsack problem. INFORMS Journal on Computing, 24(3):399–415, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1110.0460.

Bai:2012:PFP

[1013] Xue Bai, Ram Gopal, Manuel Nunez, and Dmitry Zhdanov. On the prevention of fraud and privacy exposure in process information flow. IN-FORMS Journal on Computing, 24(3):416–432, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0461.

Sewell:2012:BBR

[1014] E. C. Sewell and S. H. Jacobson. A branch, bound, and remember algorithm for the simple assembly line balancing problem. INFORMS Journal on Computing, 24(3):433–442, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1110.0462.

Li:2012:BCA

[1015] Xiangyong Li and Y. P. Aneja. A branch-and-cut approach for the minimum-energy broadcasting problem in wireless networks. IN-FORMS Journal on Computing, 24(3):443–456, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0463.

Allen:2012:DSH

[1016] Sam D. Allen and Edmund K. Burke. Data structures for higherdimensional rectilinear packing. INFORMS Journal on Computing, 24 (3):457–470, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0464.

Dellino:2012:ROS

[1017] Gabriella Dellino, Jack P. C. Kleijnen, and Carlo Meloni. Robust optimization in simulation: Taguchi and Krige combined. INFORMS Journal on Computing, 24(3):471–484, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1110.0465.

Nobibon:2012:CGU

[1018] Fabrice Talla Nobibon, Cor A. J. Hurkens, Roel Leus, and Frits C. R. Spieksma. Coloring graphs using two colors while avoiding monochromatic cycles. INFORMS Journal on Computing, 24(3):485–499, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0466.

Spacey:2012:RSP

[1019] Simon A. Spacey, Wolfram Wiesemann, Daniel Kuhn, and Wayne Luk. Robust software partitioning with multiple instantiation. IN-FORMS Journal on Computing, 24(3):500–515, Summer 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0467.

Chinneck:2012:E

[1020] John W. Chinneck. From the Editor. INFORMS Journal on Computing, 24(4):517, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1120.0535.

Salari:2012:QTB

[1021] Ehsan Salari and H. Edwin Romeijn. Quantifying the trade-off between IMRT treatment plan quality and delivery efficiency using direct aperture optimization. INFORMS Journal on Computing, 24(4):518–533, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0474.

Golden:2012:GCS

[1022] Bruce Golden, Zahra Naji-Azimi, S. Raghavan, Majid Salari, and Paolo Toth. The generalized covering salesman problem. INFORMS Journal on Computing, 24(4):534–553, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1110.0480.

Jozefowiez:2012:GBC

[1023] Nicolas Jozefowiez, Gilbert Laporte, and Frédéric Semet. A generic branch-and-cut algorithm for multiobjective optimization problems: application to the multilabel traveling salesman problem. INFORMS Journal on Computing, 24(4):554–564, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1110.0476.

Carlsson:2012:DTA

[1024] John Gunnar Carlsson. Dividing a territory among several vehicles. INFORMS Journal on Computing, 24(4):565–577, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0479.

Engineer:2012:FCS

[1025] Faramroze G. Engineer, George L. Nemhauser, Martin W. P. Savelsbergh, and Jin-Hwa Song. The fixed-charge shortest-path problem. INFORMS Journal on Computing, 24(4):578–596, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0469.

Gortz:2012:SUF

[1026] Simon Görtz and Andreas Klose. A simple but usually fast branch-and-bound algorithm for the capacitated facility location problem. INFORMS Journal on Computing, 24(4):597–610, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1110.0468.

Dawande:2012:SSO

[1027] Milind Dawande, Vijay Mookerjee, Chelliah Sriskandarajah, and Yunxia Zhu. Structural search and optimization in social networks. INFORMS Journal on Computing, 24(4):611–623, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1110.0473.

Singham:2012:FSP

[1028] Dashi I. Singham and Lee W. Schruben. Finite-sample performance of absolute precision stopping rules. INFORMS Journal on Computing, 24 (4):624–635, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0471.

Deng:2012:LOM

[1029] Lih-Yuan Deng, Jyh-Jen Horng Shiau, and Henry Horng-Shing Lu. Large-order multiple recursive generators with modulus 2^{31} − 1. IN-FORMS Journal on Computing, 24(4):636–647, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0477.

Bai:2012:MFD

[1030] Xue Bai. A mathematical framework for data quality management in enterprise systems. INFORMS Journal on Computing, 24(4):648–664, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0475.

Powell:2012:SSM

[1031] Warren B. Powell, Abraham George, Hugo Simão, Warren Scott, Alan Lamont, and Jeffrey Stewart. SMART: a stochastic multiscale model for the analysis of energy resources, technology, and policy. INFORMS Journal on Computing, 24(4):665–682, Fall 2012. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1110.0470.

Anonymous:2012:AR

[1032] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 24(4):683–686, Fall 2012. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0532.

Woodruff:2013:ME

[1033] David L. Woodruff. Message from the Editor. INFORMS Journal on Computing, 25(1):1, Winter 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0544.

Morrison:2013:NSA

[1034] David R. Morrison, Jason J. Sauppe, and Sheldon H. Jacobson. A network simplex algorithm for the equal flow problem on a generalized network. INFORMS Journal on Computing, 25(1):2–12, Winter 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0485.

Botton:2013:BDH

[1035] Quentin Botton, Bernard Fortz, Luis Gouveia, and Michael Poss. Benders decomposition for the hop-constrained survivable network design problem. INFORMS Journal on Computing, 25(1):13–26, Winter 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0472.

Zhu:2013:MCD

[1036] Wenxing Zhu, Geng Lin, and M. M. Ali. Max-k-cut by the discrete dynamic convexized method. INFORMS Journal on Computing, 25(1):

27–40, Winter 2013. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0492.

Martinez:2013:BCA

[1037] Manuel A. Alba Martínez, Jean-François Cordeau, Mauro Dell’Amico, and Manuel Iori. A branch-and-cut algorithm for the double traveling salesman problem with multiple stacks. INFORMS Journal on Computing, 25(1):41–55, Winter 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1110.0489.

Botev:2013:SNR

[1038] Zdravko I. Botev, Pierre L’Ecuyer, Gerardo Rubino, Richard Simard, and Bruno Tuffin. Static network reliability estimation via generalized splitting. INFORMS Journal on Computing, 25(1):56–71, Winter 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0493.

Clautiaux:2013:NGT

[1039] François Clautiaux, Antoine Jouglet, and Aziz Moukrim. A new graphtheoretical model for the guillotine-cutting problem. INFORMS Journal on Computing, 25(1):72–86, Winter 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1110.0478.

Buchholz:2013:RAN

[1040] Peter Buchholz and Miklós Telek. Rational automata networks: a non-Markovian modeling approach. INFORMS Journal on Computing, 25 (1):87–101, Winter 2013. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0487.

Zhang:2013:AVD

[1041] Dan Zhang and Zhaosong Lu. Assessing the value of dynamic pricing in network revenue management. INFORMS Journal on Computing, 25(1):102–115, Winter 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1110.0488.

Erdogan:2013:DAS

[1042] S. Ayca Erdogan and Brian Denton. Dynamic appointment scheduling of a stochastic server with uncertain demand. INFORMS Journal on Computing, 25(1):116–132, Winter 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1110.0482.

Xu:2013:AHA

[1043] Jie Xu, Barry L. Nelson, and L. Jeff Hong. An adaptive hyperbox algorithm for high-dimensional discrete optimization via simulation problems. INFORMS Journal on Computing, 25(1):133–146, Winter 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1110.0481.

Rengarajan:2013:CAP

[1044] Tara Rengarajan, Nedialko Dimitrov, and David P. Morton. Convex approximations of a probabilistic bicriteria model with disruptions. IN-FORMS Journal on Computing, 25(1):147–160, Winter 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1110.0483.

Yang:2013:DOS

[1045] Yinghui (Catherine) Yang, Hongyan Liu, and Yuanjue Cai. Discovery of online shopping patterns across websites. INFORMS Journal on Computing, 25(1):161–176, Winter 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1110.0484.

Argon:2013:SSS

[1046] Nilay Tanık Argon, Sigrún Andradóttir, Christos Alexopoulos, and David Goldsman. Steady-state simulation with replication-dependent initial transients: Analysis and examples. INFORMS Journal on Computing, 25(1):177–191, Winter 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1110.0494.

Mingozzi:2013:EAM

[1047] Aristide Mingozzi, Roberto Roberti, and Paolo Toth. An exact algorithm for the multitrip vehicle routing problem. INFORMS Journal on Computing, 25(2):193–207, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1110.0495.

Bertsimas:2013:NLS

[1048] Dimitris Bertsimas, Dan A. Iancu, and Dmitriy Katz. A new local search algorithm for binary optimization. INFORMS Journal on Computing, 25 (2):208–221, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1110.0496.

Mostovyi:2013:MSR

[1049] Oleksii Mostovyi, Oleg A. Prokopyev, and Oleg V. Shylo. On maximum speedup ratio of restart algorithm portfolios. INFORMS Journal on Computing, 25(2):222–229, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1120.0497.

Chang:2013:STR

[1050] Kuo-Hao Chang, L. Jeff Hong, and Hong Wan. Stochastic trust-region response-surface method (STRONG)—a new response-surface framework for simulation optimization. INFORMS Journal on Computing, 25(2):230–243, Spring 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0498.

Sadykov:2013:PCG

[1051] Ruslan Sadykov and François Vanderbeck. Bin packing with conflicts: a generic branch-and-price algorithm. INFORMS Journal on Computing, 25(2):244–255, Spring 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0499.

Ljubic:2013:LGA

[1052] Ivana Ljubić and Stefan Gollowitzer. Layered graph approaches to the hop constrained connected facility location problem. INFORMS Journal on Computing, 25(2):256–270, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0500.

Steffy:2013:VLP

[1053] Daniel E. Steffy and Kati Wolter. Valid linear programming bounds for exact mixed-integer programming. INFORMS Journal on Computing, 25(2):271–284, Spring 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0501.

Ko:2013:CLT

[1054] Young Myoung Ko and Natarajan Gautam. Critically loaded timevarying multiserver queues: computational challenges and approximations. INFORMS Journal on Computing, 25(2):285–301, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0502.

Hewitt:2013:BPG

[1055] Mike Hewitt, George Nemhauser, and Martin W. P. Savelsbergh. Branchand-price guided search for integer programs with an application to the multicommodity fixed-charge network flow problem. INFORMS Journal on Computing, 25(2):302–316, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0503.

Tsai:2013:RSP

[1056] Shing Chih Tsai. Rapid screening procedures for zero-one optimization via simulation. INFORMS Journal on Computing, 25(2):317–331, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0504.

Zhang:2013:BSH

[1057] Defu Zhang, Lijun Wei, Stephen C. H. Leung, and Qingshan Chen. A binary search heuristic algorithm based on randomized local search for the rectangular strip-packing problem. INFORMS Journal on Computing, 25(2):332–345, Spring 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0505.

Nagata:2013:PGA

[1058] Yuichi Nagata and Shigenobu Kobayashi. A powerful genetic algorithm using edge assembly crossover for the traveling salesman problem. IN-FORMS Journal on Computing, 25(2):346–363, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0506.

Hafizoglu:2013:ETC

[1059] A. Baykal Hafızoğlu, Esma S. Gel, and Pınar Keskinocak. Expected tardiness computations in multiclass priority M/M/c queues. INFORMS Journal on Computing, 25(2):364–376, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0507.

Bokrantz:2013:AAC

[1060] Rasmus Bokrantz and Anders Forsgren. An algorithm for approximating convex Pareto surfaces based on dual techniques. INFORMS Journal on Computing, 25(2):377–393, Spring 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1120.0508.

Woodruff:2013:E

[1061] David L. Woodruff. From the Editor. INFORMS Journal on Computing, 25(3):395, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2013.0562.

Kaufman:2013:RMP

[1062] David L. Kaufman and Andrew J. Schaefer. Robust modified policy iteration. INFORMS Journal on Computing, 25(3):396–410, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0509.

Burke:2013:TPV

[1063] Edmund K. Burke, Timothy Curtois, Rong Qu, and Greet Vanden Berghe. A time predefined variable depth search for nurse rostering. IN-FORMS Journal on Computing, 25(3):411–419, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0510.

Atamturk:2013:SEC

[1064] Alper Atamtürk, Laurent Flindt Muller, and David Pisinger. Separation and extension of cover inequalities for conic quadratic knapsack constraints with generalized upper bounds. INFORMS Journal on Computing, 25(3):420–431, Summer 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0511.

Chen:2013:SIS

[1065] Huifen Chen and Bruce Schmeiser. I-SMOOTH: iteratively smoothing mean-constrained and nonnegative piecewise-constant functions. IN-FORMS Journal on Computing, 25(3):432–445, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0512.

Amaldi:2013:CGM

[1066] Edoardo Amaldi, Kanika Dhyani, and Alberto Ceselli. Column generation for the minimum hyperplanes clustering problem. INFORMS Journal on Computing, 25(3):446–460, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0513.

Cote:2013:GBC

[1067] Marie-Claude Côté, Bernard Gendron, and Louis-Martin Rousseau. Grammar-based column generation for personalized multi-activity shift scheduling. INFORMS Journal on Computing, 25(3):461–474, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0514.

Balas:2013:CLP

[1068] Egon Balas, Gérard Cornuéjols, Tamás Kis, and Giacomo Nannicini. Combining lift-and-project and reduce-and-split. INFORMS Journal on Computing, 25(3):475–487, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1120.0515.

Defourny:2013:STP

[1069] Boris Defourny, Damien Ernst, and Louis Wehenkel. Scenario trees and policy selection for multistage stochastic programming using machine learning. INFORMS Journal on Computing, 25(3):488–501, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0516.

Li:2013:SOF

[1070] Yung-Ming Li, Yong Tan, and Prabuddha De. Self-organized formation and evolution of peer-to-peer networks. INFORMS Journal on Computing, 25(3):502–516, Summer 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0517.

Lutu:2013:BMC

[1071] Patricia E. N. Lutu and Andries P. Engelbrecht. Base model combination algorithm for resolving tied predictions for K-nearest neighbor OVA ensemble models. INFORMS Journal on Computing, 25(3):517–526, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0518.

Hunter:2013:OSL

[1072] Susan R. Hunter and Raghu Pasupathy. Optimal sampling laws for stochastically constrained simulation optimization on finite sets. IN-FORMS Journal on Computing, 25(3):527–542, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0519.

Kumar:2013:FPC

[1073] Akhil Kumar, Wen Yao, and Chao-Hsien Chu. Flexible process compliance with semantic constraints using mixed-integer programming. IN-FORMS Journal on Computing, 25(3):543–559, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0520.

Caprara:2013:UDW

[1074] Alberto Caprara, (deceased), Fabio Furini, and Enrico Malaguti. Uncommon Dantzig–Wolfe reformulation for the temporal knapsack problem. INFORMS Journal on Computing, 25(3):560–571, Summer 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0521.

Li:2013:AAI

[1075] Yu Li, Jia Shu, Xi Wang, Naihua Xiu, Dachuan Xu, and Jiawei Zhang. Approximation algorithms for integrated distribution network design problems. INFORMS Journal on Computing, 25(3):572–584, Summer 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0522.

Han:2013:EAB

[1076] Jinil Han, Kyungsik Lee, Chungmok Lee, and Sungsoo Park. Exact algorithms for a bandwidth packing problem with queueing delay guarantees. INFORMS Journal on Computing, 25(3):585–596, Summer 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0523.

Saltzman:2013:BR

[1077] Matthew J. Saltzman. Book reviews. INFORMS Journal on Computing, 25(3):597–598, Summer 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2013.0566.

Margulies:2013:CGM

[1078] S. Margulies, J. Ma, and I. V. Hicks. The Cunningham–Geelen method in practice: branch-decompositions and integer programming. INFORMS Journal on Computing, 25(4):599–610, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0524.

Chimani:2013:EAM

[1079] Markus Chimani and Philipp Hungerländer. Exact approaches to multilevel vertical orderings. INFORMS Journal on Computing, 25(4):611– 624, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0525.

Du:2013:ERH

[1080] Anna Ye Du, Sanjukta Das, and R. Ramesh. Efficient risk hedging by dynamic forward pricing: a study in cloud computing. INFORMS Journal on Computing, 25(4):625–642, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1120.0526.

Li:2013:LMR

[1081] Han-Lin Li, Yao-Huei Huang, and Shu-Cherng Fang. A logarithmic method for reducing binary variables and inequality constraints in solving task assignment problems. INFORMS Journal on Computing, 25 (4):643–653, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1120.0527.

Wang:2013:ZSC

[1082] Honggang Wang. Zigzag search for continuous multiobjective optimization. INFORMS Journal on Computing, 25(4):654–665, Fall 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0528.

Yanikoglu:2013:SAA

[1083] İhsan Yanıkoğlu and Dick den Hertog. Safe approximations of ambiguous chance constraints using historical data. INFORMS Journal on Computing, 25(4):666–681, Fall 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.1120.0529.

Shylo:2013:SOR

[1084] Oleg V. Shylo, Oleg A. Prokopyev, and Andrew J. Schaefer. Stochastic operating room scheduling for high-volume specialties under block booking. INFORMS Journal on Computing, 25(4):682–692, Fall 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0530.

Fischetti:2013:BB

[1085] Matteo Fischetti and Michele Monaci. Backdoor branching. INFORMS Journal on Computing, 25(4):693–700, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0531.

Lejeune:2013:CRA

[1086] Miguel A. Lejeune and Gülay Samatlı-Paç. Construction of risk-averse enhanced index funds. INFORMS Journal on Computing, 25(4):701– 719, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0533.

Aboolian:2013:EAS

[1087] Robert Aboolian, Tingting Cui, and Zuo-Jun Max Shen. An efficient approach for solving reliable facility location models. INFORMS Journal on Computing, 25(4):720–729, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1120.0534.

Carlsson:2013:DTA

[1088] John Gunnar Carlsson and Raghuveer Devulapalli. Dividing a territory among several facilities. INFORMS Journal on Computing, 25(4):730– 742, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0536.

deSouzaeSilva:2013:ETA

[1089] Edmundo de Souza e Silva, Rosa M. M. Leão, and Raymond Marie. Efficient transient analysis of Markovian models using a block reduction approach. INFORMS Journal on Computing, 25(4):743–757, Fall 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0537.

Nasr:2013:FPM

[1090] Walid W. Nasr and Michael R. Taaffe. Fitting the Ph_{t} /M_{t} /s/c timedependent departure process for use in tandem queueing networks. IN-FORMS Journal on Computing, 25(4):758–773, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0538.

Carrasco:2013:NGP

[1091] Juan A. Carrasco. A new general-purpose method for the computation of the interval availability distribution. INFORMS Journal on Computing, 25(4):774–791, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.1120.0539.

Lutu:2013:PVN

[1092] Patricia E. N. Lutu and Andries P. Engelbrecht. Positive-versus-negative classification for model aggregation in predictive data mining. INFORMS Journal on Computing, 25(4):792–807, Fall 2013. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0540.

Fischetti:2013:ASC

[1093] Matteo Fischetti and Domenico Salvagnin. Approximating the split closure. INFORMS Journal on Computing, 25(4):808–819, Fall 2013. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.1120.0543.

Anonymous:2013:AR

[1094] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 25(4):820–822, Fall 2013. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2013.0577.

Karasan:2014:SHT

[1095] Oya Ekin Karaşan, A. Ridha Mahjoub, Onur Özkök, and Hande Yaman. Survivability in hierarchical telecommunications networks under dual homing. INFORMS Journal on Computing, 26(1):1–15, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0541.

Sotirov:2014:ESP

[1096] Renata Sotirov. An efficient semidefinite programming relaxation for the graph partition problem. INFORMS Journal on Computing, 26(1):16–30, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0542.

Hijazi:2014:OIA

[1097] Hassan Hijazi, Pierre Bonami, and Adam Ouorou. An outer-inner approximation for separable mixed-integer nonlinear programs. INFORMS Journal on Computing, 26(1):31–44, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.1120.0545.

Yang:2014:SNC

[1098] Yan T. Yang, Barak Fishbain, Dorit S. Hochbaum, Eric B. Norman, and Erik Swanberg. The supervised normalized cut method for detecting, classifying, and identifying special nuclear materials. INFORMS Journal on Computing, 26(1):45–58, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.1120.0546.

Liu:2014:ATV

[1099] Yunan Liu and Ward Whitt. Algorithms for time-varying networks of many-server fluid queues. INFORMS Journal on Computing, 26(1):59– 73, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.1120.0547.

Barton:2014:QIU

[1100] Russell R. Barton, Barry L. Nelson, and Wei Xie. Quantifying input uncertainty via simulation confidence intervals. INFORMS Journal on Computing, 26(1):74–87, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2013.0548.

Contardo:2014:EAB

[1101] Claudio Contardo, Jean-François Cordeau, and Bernard Gendron. An exact algorithm based on cut-and-column generation for the capacitated location-routing problem. INFORMS Journal on Computing, 26(1):88– 102, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0549.

Adulyasak:2014:FBC

[1102] Yossiri Adulyasak, Jean-François Cordeau, and Raf Jans. Formulations and branch-and-cut algorithms for multivehicle production and inventory routing problems. INFORMS Journal on Computing, 26(1):103–120, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0550.

Tong:2014:ALP

[1103] Chaoxu Tong and Huseyin Topaloglu. On the approximate linear programming approach for network revenue management problems. IN-FORMS Journal on Computing, 26(1):121–134, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2013.0551.

Shu:2014:DCD

[1104] Jia Shu and Miao Song. Dynamic container deployment: two-stage robust model, complexity, and computational results. INFORMS Journal on Computing, 26(1):135–149, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0552.

Huh:2014:OSO

[1105] Woonghee Tim Huh and Paat Rusmevichientong. Online sequential optimization with biased gradients: theory and applications to censored demand. INFORMS Journal on Computing, 26(1):150–159, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0553.

Seref:2014:MPF

[1106] Onur Šeref, Ya-Ju Fan, and Wanpracha Art Chaovalitwongse. Mathematical programming formulations and algorithms for discrete k-median clustering of time-series data. INFORMS Journal on Computing, 26(1): 160–172, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2013.0554.

Fomeni:2014:DPH

[1107] Franklin Djeumou Fomeni and Adam N. Letchford. A dynamic programming heuristic for the quadratic knapsack problem. INFORMS Journal on Computing, 26(1):173–182, Winter 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0555.

Robbins:2014:WSC

[1108] Matthew J. Robbins, Sheldon H. Jacobson, Uday V. Shanbhag, and Banafsheh Behzad. The weighted set covering game: a vaccine pricing model for pediatric immunization. INFORMS Journal on Computing, 26(1):183–198, Winter 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2013.0556.

Kunigami:2014:OLP

[1109] Guilherme Kunigami, Pedro J. de Rezende, Cid C. de Souza, and Tallys Yunes. Optimizing the layout of proportional symbol maps: polyhedra and computation. INFORMS Journal on Computing, 26(2):199–207, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0557.

Liu:2014:TBC

[1110] Hongyan Liu, Yinghui (Catherine) Yang, Zhuohua Chen, and Yong Zheng. A tree-based contrast set-mining approach to detecting group differences. INFORMS Journal on Computing, 26(2):208–221, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0558.

Dey:2014:PST

[1111] Santanu S. Dey, Andrea Lodi, Andrea Tramontani, and Laurence A. Wolsey. On the practical strength of two-row tableau cuts. INFORMS Journal on Computing, 26(2):222–237, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0559.

Kao:2014:CSD

[1112] Hung-Pin Kao and Kwei Tang. Cost-sensitive decision tree induction with label-dependent late constraints. INFORMS Journal on Computing, 26(2):238–252, Spring 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2013.0560.

Bergman:2014:OBB

[1113] David Bergman, Andre A. Cire, Willem-Jan van Hoeve, and J. N. Hooker. Optimization bounds from binary decision diagrams. INFORMS Journal on Computing, 26(2):253–268, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0561.

Avramidis:2014:CDU

[1114] Athanassios N. Avramidis. Constructing discrete unbounded distributions with Gaussian-copula dependence and given rank correlation. IN-FORMS Journal on Computing, 26(2):269–279, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2013.0563.

Carlsson:2014:AAC

[1115] John Gunnar Carlsson, Fan Jia, and Ying Li. An approximation algorithm for the continuous k-medians problem in a convex polygon. IN-FORMS Journal on Computing, 26(2):280–289, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2013.0564.

Rainwater:2014:RCA

[1116] Chase Rainwater, Joseph Geunes, and H. Edwin Romeijn. Resourceconstrained assignment problems with shared resource consumption and flexible demand. INFORMS Journal on Computing, 26(2):290–302, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0565.

Yazici:2014:CMC

[1117] Volkan Yazici and Cevdet Aykanat. Constrained min-cut replication for K-way hypergraph partitioning. INFORMS Journal on Computing, 26 (2):303–320, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2013.0567.

Feizollahi:2014:RMR

[1118] Mohammad Javad Feizollahi and Igor Averbakh. The robust (minmax regret) quadratic assignment problem with interval flows. INFORMS Journal on Computing, 26(2):321–335, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0568.

Basu:2014:CML

[1119] Amitabh Basu, Jesús A. De Loera, and Mark Junod. On Chubanov’s method for linear programming. INFORMS Journal on Computing, 26 (2):336–350, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2013.0569.

Adomavicius:2014:OBA

[1120] Gediminas Adomavicius and YoungOk Kwon. Optimization-based approaches for maximizing aggregate recommendation diversity. INFORMS Journal on Computing, 26(2):351–369, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0570.

Aksakalli:2014:PBA

[1121] Vural Aksakalli and Ibrahim Ari. Penalty-based algorithms for the stochastic obstacle scene problem. INFORMS Journal on Computing, 26(2):370–384, Spring 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2013.0571.

Hong:2014:CVR

[1122] L. Jeff Hong, Zhaolin Hu, and Liwei Zhang. Conditional value-at-risk approximation to value-at-risk constrained programs: a remedy via Monte Carlo. INFORMS Journal on Computing, 26(2):385–400, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0572.

Talluri:2014:NFC

[1123] Kalyan Talluri. New formulations for choice network revenue management. INFORMS Journal on Computing, 26(2):401–413, Spring 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0573.

Behdani:2014:IPB

[1124] Behnam Behdani and J. Cole Smith. An integer-programming-based approach to the close-enough traveling salesman problem. INFORMS Journal on Computing, 26(3):415–432, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0574.

Ghoshal:2014:ARR

[1125] Abhijeet Ghoshal and Sumit Sarkar. Association rules for recommendations with multiple items. INFORMS Journal on Computing, 26(3): 433–448, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2013.0575.

LEcuyer:2014:LSS

[1126] Pierre L’Ecuyer and Richard Simard. On the lattice structure of a special class of multiple recursive random number generators. INFORMS Journal on Computing, 26(3):449–460, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0576. Analysis and expo-

sure of serious lattice structure in earlier work on fast multiple recursive generators [1495, 1496, 1497, 1498, 1029].

Xu:2014:EAL

[1127] Kaiquan Xu, Stephen Shaoyi Liao, Raymond Y. K. Lau, and J. Leon Zhao. Effective active learning strategies for the use of large-margin classifiers in semantic annotation: an optimal parameter discovery perspective. INFORMS Journal on Computing, 26(3):461–483, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0578.

Fu:2014:RMA

[1128] Michael C. Fu and Huashuai Qu. Regression models augmented with direct stochastic gradient estimators. INFORMS Journal on Computing, 26(3):484–499, Summer 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2013.0579.

Mete:2014:MIP

[1129] Huseyin Onur Mete and Zelda B. Zabinsky. Multiobjective interacting particle algorithm for global optimization. INFORMS Journal on Computing, 26(3):500–513, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2013.0580.

Ghoniem:2014:EMM

[1130] Ahmed Ghoniem, Hanif D. Sherali, and Hojong Baik. Enhanced models for a mixed arrival-departure aircraft sequencing problem. IN-FORMS Journal on Computing, 26(3):514–530, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2013.0581.

Qiu:2014:CLP

[1131] Feng Qiu, Shabbir Ahmed, Santanu S. Dey, and Laurence A. Wolsey. Covering linear programming with violations. INFORMS Journal on Computing, 26(3):531–546, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2013.0582.

Sauppe:2014:CAR

[1132] Jason J. Sauppe, Sheldon H. Jacobson, and Edward C. Sewell. Complexity and approximation results for the balance optimization subset selection model for causal inference in observational studies. INFORMS Journal on Computing, 26(3):547–566, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0583.

Brooks:2014:SMM

[1133] J. Paul Brooks and Eva K. Lee. Solving a multigroup mixed-integer programming-based constrained discrimination model. INFORMS Journal on Computing, 26(3):567–585, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0584.

Fang:2014:DAS

[1134] Fang Fang, Kaushik Dutta, and Anindya Datta. Domain adaptation for sentiment classification in light of multiple sources. INFORMS Journal on Computing, 26(3):586–598, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0585.

Khojandi:2014:OIC

[1135] Anahita Khojandi, Lisa M. Maillart, Oleg A. Prokopyev, Mark S. Roberts, Timothy Brown, and William W. Barrington. Optimal implantable cardioverter defibrillator (ICD) generator replacement. IN-FORMS Journal on Computing, 26(3):599–615, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2013.0586.

Lim:2014:CRC

[1136] Eunji Lim. On convergence rates of convex regression in multiple dimensions. INFORMS Journal on Computing, 26(3):616–628, Summer 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2013.0587.

Mandala:2014:GTA

[1137] Supreet Mandala, Soundar Kumara, and Kalyan Chatterjee. A gametheoretic approach to graph clustering. INFORMS Journal on Computing, 26(3):629–643, Summer 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2013.0588.

Gendron:2014:BDB

[1138] Bernard Gendron, Abilio Lucena, Alexandre Salles da Cunha, and Luidi Simonetti. Benders decomposition, branch-and-cut, and hybrid algorithms for the minimum connected dominating set problem. INFORMS Journal on Computing, 26(4):645–657, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2013.0589.

Blom:2014:DBH

[1139] Michelle L. Blom, Christina N. Burt, Adrian R. Pearce, and Peter J. Stuckey. A decomposition-based heuristic for collaborative scheduling in a network of open-pit mines. INFORMS Journal on Computing, 26 (4):658–676, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2013.0590.

Cheng:2014:ZZA

[1140] Yongxi Cheng, Ding-Zhu Du, and Yinfeng Xu. A zig-zag approach for competitive group testing. INFORMS Journal on Computing, 26(4):677– 689, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0591.

Zheng:2014:IPM

[1141] Xiaojin Zheng, Xiaoling Sun, and Duan Li. Improving the performance of MIQP solvers for quadratic programs with cardinality and minimum threshold constraints: a semidefinite program approach. INFORMS Journal on Computing, 26(4):690–703, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2014.0592.

Morrison:2014:WBS

[1142] David R. Morrison, Jason J. Sauppe, Edward C. Sewell, and Sheldon H. Jacobson. A wide branching strategy for the graph coloring problem. INFORMS Journal on Computing, 26(4):704–717, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2014.0593.

Coffrin:2014:LPA

[1143] Carleton Coffrin and Pascal Van Hentenryck. A linear-programming approximation of AC power flows. INFORMS Journal on Computing, 26(4):718–734, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2014.0594.

Song:2014:CCB

[1144] Yongjia Song, James R. Luedtke, and Simge Küçükyavuz. Chanceconstrained binary packing problems. INFORMS Journal on Computing, 26(4):735–747, Fall 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0595.

Burkowski:2014:EUS

[1145] Forbes Burkowski, Yuen-Lam Cheung, and Henry Wolkowicz. Efficient use of semidefinite programming for selection of rotamers in protein conformations. INFORMS Journal on Computing, 26(4):748–766, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0596.

Koca:2014:LSP

[1146] Esra Koca, Hande Yaman, and M. Selim Aktürk. Lot sizing with piecewise concave production costs. INFORMS Journal on Computing, 26 (4):767–779, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2014.0597.

Dash:2014:CEC

[1147] Sanjeeb Dash, Oktay Günlük, and Juan Pablo Vielma. Computational experiments with cross and crooked cross cuts. INFORMS Journal on Computing, 26(4):780–797, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2014.0598.

Zhao:2014:IFL

[1148] Yifei Zhao and Stein W. Wallace. Integrated facility layout design and flow assignment problem under uncertainty. INFORMS Journal on Computing, 26(4):798–808, Fall 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0599.

Butsch:2014:DAR

[1149] Alexander Butsch, Jörg Kalcsics, and Gilbert Laporte. Districting for arc routing. INFORMS Journal on Computing, 26(4):809–824, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0600.

Papageorgiou:2014:TSD

[1150] Dimitri J. Papageorgiou, Ahmet B. Keha, George L. Nemhauser, and Joel Sokol. Two-stage decomposition algorithms for single product maritime inventory routing. INFORMS Journal on Computing, 26(4):825– 847, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0601.

Hong:2014:ESP

[1151] L. Jeff Hong, Sandeep Juneja, and Jun Luo. Estimating sensitivities of portfolio credit risk using Monte Carlo. INFORMS Journal on Computing, 26(4):848–865, Fall 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0602.

Xia:2014:RRA

[1152] Hao Xia, Milind Dawande, and Vijay Mookerjee. Role refinement in access control: model and analysis. INFORMS Journal on Computing, 26(4):866–884, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2014.0603.

Fisher:2014:CDN

[1153] Brent Fisher and Mohan Chaudhry. Computing the distribution for the number of renewals with bulk arrivals. INFORMS Journal on Computing, 26(4):885–892, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2014.0604.

Classen:2014:CCO

[1154] Grit Claßen, Arie M. C. A. Koster, David Coudert, and Napoleão Nepomuceno. Chance-constrained optimization of reliable fixed broadband wireless networks. INFORMS Journal on Computing, 26(4):893–909, Fall 2014. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0605.

Anonymous:2014:AR

[1155] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 26(4):910–914, Fall 2014. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0620.

Alvarez-Miranda:2015:RRT

[1156] Eduardo Álvarez-Miranda, Ivana Ljubić, S. Raghavan, and Paolo Toth. The recoverable robust two-level network design problem. INFORMS Journal on Computing, 27(1):1–19, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2014.0606.

Lu:2015:SDA

[1157] Haibing Lu, Jaideep Vaidya, Vijayalakshmi Atluri, and Yingjiu Li. Statistical database auditing without query denial threat. INFORMS Journal on Computing, 27(1):20–34, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2014.0607.

Chou:2015:CGF

[1158] Chun-An Chou, Zhe Liang, Wanpracha Art Chaovalitwongse, Tanya Y. Berger-Wolf, Bhaskar DasGupta, Saad Sheikh, Mary V. Ashley, and Isabel C. Caballero. Column-generation framework of nonlinear similarity model for reconstructing sibling groups. INFORMS Journal on Computing, 27(1):35–47, Winter 2015. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0608.

Yaghini:2015:CPN

[1159] Masoud Yaghini, Mohammad Karimi, Mohadeseh Rahbar, and Mohammad Hassan Sharifitabar. A cutting-plane neighborhood structure for fixed-charge capacitated multicommodity network design problem. IN-FORMS Journal on Computing, 27(1):48–58, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2014.0609.

Trespalacios:2015:AAI

[1160] Francisco Trespalacios and Ignacio E. Grossmann. Algorithmic approach for improved mixed-integer reformulations of convex generalized disjunctive programs. INFORMS Journal on Computing, 27(1):59–74, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0610.

vanDam:2015:BBG

[1161] E. R. van Dam and R. Sotirov. On bounding the bandwidth of graphs with symmetry. INFORMS Journal on Computing, 27(1):75–88, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0611.

Pantuso:2015:SHS

[1162] Giovanni Pantuso, Kjetil Fagerholt, and Stein W. Wallace. Solving hierarchical stochastic programs: application to the maritime fleet renewal problem. INFORMS Journal on Computing, 27(1):89–102, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0612.

Sleptchenko:2015:MSR

[1163] Andrei Sleptchenko and M. Eric Johnson. Maintaining secure and reliable distributed control systems. INFORMS Journal on Computing, 27(1): 103–117, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2014.0613.

Leitner:2015:CSE

[1164] Markus Leitner, Ivana Ljubić, and Markus Sinnl. A computational study of exact approaches for the bi-objective prize-collecting Steiner tree problem. INFORMS Journal on Computing, 27(1):118–134, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0614.

Sen:2015:SPR

[1165] Halil Şen and Kerem Bülbül. A strong preemptive relaxation for weighted tardiness and Earliness/Tardiness problems on unrelated parallel machines. INFORMS Journal on Computing, 27(1):135–150, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0615.

Steitz:2015:NHA

[1166] Wolfgang Steitz. New heuristic approaches for the bounded-diameter minimum spanning tree problem. INFORMS Journal on Computing, 27(1):151–163, Winter 2015. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0617.

Verma:2015:SMC

[1167] Anurag Verma, Austin Buchanan, and Sergiy Butenko. Solving the maximum clique and vertex coloring problems on very large sparse networks. INFORMS Journal on Computing, 27(1):164–177, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2014.0618.

Buchanan:2015:IPA

[1168] Austin Buchanan, Je Sang Sung, Sergiy Butenko, and Eduardo L. Pasiliao. An integer programming approach for fault-tolerant connected dominating sets. INFORMS Journal on Computing, 27(1):178–188, Winter 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0619.

Saltzman:2015:BR

[1169] Matthew J. Saltzman. Book reviews. INFORMS Journal on Computing, 27(1):189–192, Winter 2015. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0631.

Wang:2015:MMN

[1170] Jin Wang and Michael R. Taaffe. Multivariate mixtures of normal distributions: properties, random vector generation, fitting, and as models of market daily changes. INFORMS Journal on Computing, 27(2):193–203, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0616.

Chen:2015:GRL

[1171] Si Chen, Ivana Ljubić, and S. Raghavan. The generalized regenerator location problem. INFORMS Journal on Computing, 27(2):204–220, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0621.

Fu:2015:DPD

[1172] Zhang-Hua Fu and Jin-Kao Hao. Dynamic programming driven memetic search for the Steiner tree problem with revenues, budget, and hop constraints. INFORMS Journal on Computing, 27(2):221–237, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0622.

Lubin:2015:COR

[1173] Miles Lubin and Iain Dunning. Computing in operations research using Julia. INFORMS Journal on Computing, 27(2):238–248, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0623.

Lee:2015:NSA

[1174] Chungmok Lee, Minh Pham, Myong K. Jeong, Dohyun Kim, Dennis K. J. Lin, and Wanpracha Art Chavalitwongse. A network structural approach to the link prediction problem. INFORMS Journal on Computing, 27(2):249–267, Spring 2015. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0624.

Grimes:2015:SVJ

[1175] Diarmuid Grimes and Emmanuel Hebrard. Solving variants of the job shop scheduling problem through conflict-directed search. INFORMS Journal on Computing, 27(2):268–284, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2014.0625.

Yang:2015:OBA

[1176] Yanwu Yang, Daniel Zeng, Yinghui Yang, and Jie Zhang. Optimal budget allocation across search advertising markets. INFORMS Journal on Computing, 27(2):285–300, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2014.0626.

Shen:2015:CCP

[1177] Siqian Shen, Murat Kurt, and Jue Wang. Chance-constrained programming models and approximations for general stochastic bottleneck spanning tree problems. INFORMS Journal on Computing, 27(2):301–316, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0627.

Hong:2015:CCS

[1178] L. Jeff Hong, Jun Luo, and Barry L. Nelson. Chance constrained selection of the best. INFORMS Journal on Computing, 27(2):317–334, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0628.

Rayfield:2015:AMP

[1179] W. Zachary Rayfield, Paat Rusmevichientong, and Huseyin Topaloglu. Approximation methods for pricing problems under the nested logit model with price bounds. INFORMS Journal on Computing, 27(2): 335–357, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2014.0629.

Parpas:2015:ISS

[1180] Panos Parpas, Berk Ustun, Mort Webster, and Quang Kha Tran. Importance sampling in stochastic programming: a Markov chain Monte Carlo approach. INFORMS Journal on Computing, 27(2):358–377, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0630.

deKlerk:2015:NSP

[1181] E. de Klerk, R. Sotirov, and U. Truetsch. A new semidefinite programming relaxation for the quadratic assignment problem and its computational perspectives. INFORMS Journal on Computing, 27(2):378–391, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0634.

Furini:2015:HEA

[1182] Fabio Furini, Manuel Iori, Silvano Martello, and Mutsunori Yagiura. Heuristic and exact algorithms for the interval min-max regret knapsack problem. INFORMS Journal on Computing, 27(2):392–405, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2014.0632.

Vaisman:2015:MCM

[1183] Radislav Vaisman, Ofer Strichman, and Ilya Gertsbakh. Model counting of monotone conjunctive normal form formulas with spectra. INFORMS Journal on Computing, 27(2):406–415, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2014.0633.

Rysz:2015:SDA

[1184] Maciej Rysz, Alexander Vinel, Pavlo Krokhmal, and Eduardo L. Pasiliao. A scenario decomposition algorithm for stochastic programming problems with a class of downside risk measures. INFORMS Journal on Computing, 27(2):416–430, Spring 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2014.0635.

deAraujo:2015:PDC

[1185] Silvio Alexandre de Araujo, Bert De Reyck, Zeger Degraeve, Ioannis Fragkos, and Raf Jans. Period decompositions for the capacitated lot sizing problem with setup times. INFORMS Journal on Computing, 27(3):431–448, Summer 2015. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2014.0636.

Bilauca:2015:AMH

[1186] Mihai Bilauca, Graeme Gange, Patrick Healy, Kim Marriott, Peter Moulder, and Peter J. Stuckey. Automatic minimal-height table layout. IN-FORMS Journal on Computing, 27(3):449–461, Summer 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2014.0637.

Dey:2015:OPS

[1187] Debabrata Dey, Atanu Lahiri, and Guoying Zhang. Optimal policies for security patch management. INFORMS Journal on Computing, 27(3): 462–477, Summer 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2014.0638.

Bringmann:2015:OCI

[1188] Karl Bringmann, Benjamin Doerr, Adrian Neumann, and Jakub Sliacan. Online checkpointing with improved worst-case guarantees. IN-FORMS Journal on Computing, 27(3):478–490, Summer 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2014.0639.

Perry:2015:ARR

[1189] Ohad Perry and Ward Whitt. Achieving rapid recovery in an overload control for large-scale service systems. INFORMS Journal on Computing, 27(3):491–506, Summer 2015. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2015.0642.

Pessoa:2015:RND

[1190] Artur Alves Pessoa and Michael Poss. Robust network design with uncertain outsourcing cost. INFORMS Journal on Computing, 27(3):507–524, Summer 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2015.0643.

Jiang:2015:OHA

[1191] Daniel R. Jiang and Warren B. Powell. Optimal hour-ahead bidding in the real-time electricity market with battery storage using approximate dynamic programming. INFORMS Journal on Computing, 27(3): 525–543, Summer 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0640.

vanJaarsveld:2015:OIS

[1192] Willem van Jaarsveld and Alan Scheller-Wolf. Optimization of industrialscale assemble-to-order systems. INFORMS Journal on Computing, 27 (3):544–560, Summer 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0641.

Lau:2015:LCS

[1193] Raymond Y. K. Lau, J. Leon Zhao, Wenping Zhang, Yi Cai, and Eric W. T. Ngai. Learning context-sensitive domain ontologies from folksonomies: a cognitively motivated method. INFORMS Journal on Computing, 27(3):561–578, Summer 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0644.

Saure:2015:SBA

[1194] Antoine Sauré, Jonathan Patrick, and Martin L. Puterman. Simulationbased approximate policy iteration with generalized logistic functions. INFORMS Journal on Computing, 27(3):579–595, Summer 2015. ISSN 1091-9856 (print), 1526-5528 (electronic).

Boland:2015:CSSa

[1195] Natashia Boland, Hadi Charkhgard, and Martin Savelsbergh. A criterion space search algorithm for biobjective mixed integer programming: The triangle splitting method. INFORMS Journal on Computing, 27 (4):597–618, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0646.

Bai:2015:GPA

[1196] Xue Bai, Sudip Bhattacharjee, Fidan Boylu, and Ram Gopal. Growth projections and assortment planning of commodity products across multiple stores: a data mining and optimization approach. INFORMS Journal on Computing, 27(4):619–635, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2015.0647.

Xu:2015:AAM

[1197] Zhou Xu and Brian Rodrigues. A 3/2-approximation algorithm for the multiple TSP with a fixed number of depots. INFORMS Journal on Computing, 27(4):636–645, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0650.

Ahn:2015:LPC

[1198] Andrew Ahn and Martin Haugh. Linear programming and the control of diffusion processes. INFORMS Journal on Computing, 27(4):646–657, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2015.0651.

Martinelli:2015:EHA

[1199] Rafael Martinelli and Claudio Contardo. Exact and heuristic algorithms for capacitated vehicle routing problems with quadratic costs structure. INFORMS Journal on Computing, 27(4):658–676, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0649.

Escobedo:2015:REF

[1200] Adolfo R. Escobedo and Erick Moreno-Centeno. Roundoff-error-free algorithms for solving linear systems via Cholesky and LU factorizations. INFORMS Journal on Computing, 27(4):677–689, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0653.

Pesant:2015:ADC

[1201] Gilles Pesant. Achieving domain consistency and counting solutions for dispersion constraints. INFORMS Journal on Computing, 27(4):690– 703, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2015.0654.

Huang:2015:CCS

[1202] Jiangchuan Huang, Christoph M. Kirsch, and Raja Sengupta. Cloud computing in space. INFORMS Journal on Computing, 27(4):704–717, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2015.0652.

Gartner:2015:MLA

[1203] Daniel Gartner, Rainer Kolisch, Daniel B. Neill, and Rema Padman. Machine learning approaches for early DRG classification and resource allocation. INFORMS Journal on Computing, 27(4):718–734, Fall 2015. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0655.

Boland:2015:CSSb

[1204] Natashia Boland, Hadi Charkhgard, and Martin Savelsbergh. A criterion space search algorithm for biobjective integer programming: The balanced Box method. INFORMS Journal on Computing, 27(4):735– 754, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2015.0657.

Gul:2015:PHA

[1205] Serhat Gul, Brian T. Denton, and John W. Fowler. A progressive hedging approach for surgery planning under uncertainty. INFORMS Journal on Computing, 27(4):755–772, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0658.

Omer:2015:IPS

[1206] Jérémy Omer, Samuel Rosat, Vincent Raymond, and François Soumis. Improved primal simplex: a more general theoretical framework and an extended experimental analysis. INFORMS Journal on Computing, 27 (4):773–787, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0656.

Bortfeld:2015:ORT

[1207] Thomas Bortfeld, Jagdish Ramakrishnan, John N. Tsitsiklis, and Jan Unkelbach. Optimization of radiation therapy fractionation schedules in the presence of tumor repopulation. INFORMS Journal on Computing, 27(4):788–803, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0659.

Anonymous:2015:AR

[1208] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 27(4):804–808, Fall 2015. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2015.0685.

Anonymous:2015:EB

[1209] Anonymous. Editorial Board. INFORMS Journal on Computing, 27(4): c2–c3, Fall 2015. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc-27-04-eb.

Cire:2016:MMS

[1210] Andre A. Cire, John N. Hooker, and Tallys Yunes. Modeling with metaconstraints and semantic typing of variables. INFORMS Journal on Computing, 28(1):1–13, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0664.

Boland:2016:BIF

[1211] Natashia Boland, Riley Clement, and Hamish Waterer. A bucket indexed formulation for nonpreemptive single machine scheduling problems. INFORMS Journal on Computing, 28(1):14–30, Winter 2016. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0661. See

corrigendum [1553].

Bagnara:2016:EBF

[1212] Roberto Bagnara, Matthieu Carlier, Roberta Gori, and Arnaud Gotlieb. Exploiting binary floating-point representations for constraint propagation. INFORMS Journal on Computing, 28(1):31–46, Winter 2016. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0663.

Bergman:2016:DOD

[1213] David Bergman, Andre A. Cire, Willem-Jan van Hoeve, and J. N. Hooker. Discrete optimization with decision diagrams. INFORMS Journal on Computing, 28(1):47–66, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2015.0648.

Morrison:2016:SPP

[1214] David R. Morrison, Edward C. Sewell, and Sheldon H. Jacobson. Solving the pricing problem in a branch-and-price algorithm for graph coloring using zero-suppressed binary decision diagrams. INFORMS Journal on Computing, 28(1):67–82, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0667.

Tran:2016:DMP

[1215] Tony T. Tran, Arthur Araujo, and J. Christopher Beck. Decomposition methods for the parallel machine scheduling problem with setups. INFORMS Journal on Computing, 28(1):83–95, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0666.

Aksakalli:2016:ABE

[1216] Vural Aksakalli, O. Furkan Sahin, and Ibrahim Ari. An AO^{∗} based exact algorithm for the Canadian traveler problem. INFORMS Journal on Computing, 28(1):96–111, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0668.

Eckstein:2016:MOM

[1217] Jonathan Eckstein, Deniz Eskandani, and Jingnan Fan. Multilevel optimization modeling for risk-averse stochastic programming. INFORMS Journal on Computing, 28(1):112–128, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2015.0665.

Adomavicius:2016:CRT

[1218] Gediminas Adomavicius and Jingjing Zhang. Classification, ranking, and top-K stability of recommendation algorithms. INFORMS Journal on Computing, 28(1):129–147, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0662.

Shioura:2016:ASO

[1219] Akiyoshi Shioura, Natalia V. Shakhlevich, and Vitaly A. Strusevich. Application of submodular optimization to single machine scheduling with controllable processing times subject to release dates and deadlines. IN-FORMS Journal on Computing, 28(1):148–161, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0660.

Naoum-Sawaya:2016:RCN

[1220] Joe Naoum-Sawaya and Christoph Buchheim. Robust critical node selection by Benders decomposition. INFORMS Journal on Computing, 28(1):162–174, Winter 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2015.0671.

Gschwind:2016:DIS

[1221] Timo Gschwind and Stefan Irnich. Dual inequalities for stabilized column generation revisited. INFORMS Journal on Computing, 28(1):175–194, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2015.0670.

Anonymous:2016:EB

[1222] Anonymous. Editorial Board. INFORMS Journal on Computing, 28 (1):c2–c3, Winter 2016. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.eb.v281.

Peng:2016:DSA

[1223] Yijie Peng, Chun-Hung Chen, Michael C. Fu, and Jian-Qiang Hu. Dynamic sampling allocation and design selection. INFORMS Journal on Computing, 28(2):195–208, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0673.

Trespalacios:2016:CPA

[1224] Francisco Trespalacios and Ignacio E. Grossmann. Cutting plane algorithm for convex generalized disjunctive programs. INFORMS Journal on Computing, 28(2):209–222, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2015.0669.

Tong:2016:ISO

[1225] Shaolong Tong and Guangwu Liu. Importance sampling for option Greeks with discontinuous payoffs. INFORMS Journal on Computing, 28(2):223–235, Spring 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2015.0674.

Zhang:2016:PCO

[1226] Zunqiang Zhang, Guoqing Chen, Jin Zhang, Xunhua Guo, and Qiang Wei. Providing consistent opinions from online reviews: a heuristic stepwise optimization approach. INFORMS Journal on Computing, 28(2): 236–250, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0672.

James:2016:DES

[1227] Terry James, Kevin Glazebrook, and Kyle Lin. Developing effective service policies for multiclass queues with abandonment: asymptotic optimality and approximate policy improvement. INFORMS Journal on Computing, 28(2):251–264, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0675.

Passacantando:2016:SPP

[1228] Mauro Passacantando, Danilo Ardagna, and Anna Savi. Service provisioning problem in cloud and multi-cloud systems. INFORMS Journal on Computing, 28(2):265–277, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0681.

Li:2016:CKE

[1229] Xin Li, Kun Chen, Sherry X. Sun, Terrance Fung, Huaiqing Wang, and Daniel D. Zeng. A commonsense knowledge-enabled textual analysis approach for financial market surveillance. INFORMS Journal on Computing, 28(2):278–294, Spring 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2015.0677.

Ma:2016:MSC

[1230] Juan Ma, Foad Mahdavi Pajouh, Balabhaskar Balasundaram, and Vladimir Boginski. The minimum spanning k-core problem with bounded CVaR under probabilistic edge failures. INFORMS Journal on Computing, 28(2):295–307, Spring 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2015.0679.

Pferschy:2016:AQK

[1231] Ulrich Pferschy and Joachim Schauer. Approximation of the quadratic knapsack problem. INFORMS Journal on Computing, 28(2):308–318, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2015.0678.

Caprara:2016:BKI

[1232] Alberto Caprara, Margarida Carvalho, Andrea Lodi, and Gerhard J. Woeginger. Bilevel knapsack with interdiction constraints. INFORMS Journal on Computing, 28(2):319–333, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2015.0676.

Restrepo:2016:BPP

[1233] María I. Restrepo, Bernard Gendron, and Louis-Martin Rousseau. Branch-and-price for personalized multiactivity tour scheduling. IN-FORMS Journal on Computing, 28(2):334–350, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0683.

Gose:2016:SBM

[1234] Alexander H. Gose and Brian T. Denton. Sequential bounding methods for two-stage stochastic programs. INFORMS Journal on Computing, 28(2):351–369, Spring 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2015.0682.

vanBarneveld:2016:MEP

[1235] Thije van Barneveld. The minimum expected penalty relocation problem for the computation of compliance tables for ambulance vehicles. INFORMS Journal on Computing, 28(2):370–384, Spring 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2015.0687.

Zhang:2016:DAR

[1236] Weini Zhang, Hamed Rahimian, and Güzin Bayraksan. Decomposition algorithms for risk-averse multistage stochastic programs with application to water allocation under uncertainty. INFORMS Journal on Computing, 28(3):385–404, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2015.0684.

Han:2016:GTM

[1237] Chan Y. Han, Brian J. Lunday, and Matthew J. Robbins. A game theoretic model for the optimal location of integrated air defense system missile batteries. INFORMS Journal on Computing, 28(3):405–416, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0690.

Zhu:2016:FAA

[1238] Xiaojun Zhu, Guihai Chen, Shaojie Tang, Xiaobing Wu, and Bing Chen. Fast approximation algorithm for maximum lifetime aggregation trees in wireless sensor networks. INFORMS Journal on Computing, 28(3):

417–431, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0688.

Doulabi:2016:CPB

[1239] Seyed Hossein Hashemi Doulabi, Louis-Martin Rousseau, and Gilles Pesant. A constraint-programming-based branch-and-price-and-cut approach for operating room planning and scheduling. INFORMS Journal on Computing, 28(3):432–448, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2015.0686.

Gleixner:2016:IRL

[1240] Ambros M. Gleixner, Daniel E. Steffy, and Kati Wolter. Iterative refinement for linear programming. INFORMS Journal on Computing, 28(3):449–464, Summer 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2016.0692.

Fragkos:2016:HDA

[1241] Ioannis Fragkos, Zeger Degraeve, and Bert De Reyck. A horizon decomposition approach for the capacitated lot-sizing problem with setup times. INFORMS Journal on Computing, 28(3):465–482, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0691.

Angulo:2016:IIS

[1242] Gustavo Angulo, Shabbir Ahmed, and Santanu S. Dey. Improving the integer L-shaped method. INFORMS Journal on Computing, 28(3): 483–499, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0695.

Bertsimas:2016:DTS

[1243] Dimitris Bertsimas and Frans J. C. T. de Ruiter. Duality in two-stage adaptive linear optimization: faster computation and stronger bounds. INFORMS Journal on Computing, 28(3):500–511, Summer 2016. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0689.

Lu:2016:MOS

[1244] Fang Lu, John J. Hasenbein, and David P. Morton. Modeling and optimization of a spatial detection system. INFORMS Journal on Computing, 28(3):512–526, Summer 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2016.0694.

Song:2016:RAS

[1245] Yongjia Song and Siqian Shen. Risk-averse shortest path interdiction. IN-FORMS Journal on Computing, 28(3):527–539, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0699.

Boschetti:2016:UGC

[1246] Marco A. Boschetti, Vittorio Maniezzo, and Francesco Strappaveccia. Using GPU computing for solving the two-dimensional guillotine cutting problem. INFORMS Journal on Computing, 28(3):540–552, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0693.

Postek:2016:MAR

[1247] Krzysztof Postek and Dick den Hertog. Multistage adjustable robust mixed-integer optimization via iterative splitting of the uncertainty set. INFORMS Journal on Computing, 28(3):553–574, Summer 2016. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0696.

Ghoniem:2016:OBV

[1248] Ahmed Ghoniem, Tulay Flamand, and Mohamed Haouari. Optimizationbased very large-scale neighborhood search for generalized assignment problems with location/allocation considerations. INFORMS Journal on Computing, 28(3):575–588, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2016.0703.

Ghoniem:2016:ESM

[1249] Ahmed Ghoniem, Tulay Flamand, and Mohamed Haouari. Exact solution methods for a generalized assignment problem with location/allocation considerations. INFORMS Journal on Computing, 28 (3):589–602, Summer 2016. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0700.

Hu:2016:ISC

[1250] Jianqiang Hu, Cheng Zhang, and Chenbo Zhu. (s, S) inventory systems with correlated demands. INFORMS Journal on Computing, 28 (4):603–611, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0698.

Zheng:2016:AMN

[1251] Ronghuo Zheng, Tinglong Dai, Katia Sycara, and Nilanjan Chakraborty. Automated multilateral negotiation on multiple issues with private information. INFORMS Journal on Computing, 28(4):612–628, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0701.

Chen:2016:EVC

[1252] Xi Chen and Kyoung-Kuk Kim. Efficient VaR and CVaR measurement via stochastic kriging. INFORMS Journal on Computing, 28(4):629– 644, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0705.

Borraz-Sanchez:2016:CRG

[1253] Conrado Borraz-Sánchez, Russell Bent, Scott Backhaus, Hassan Hijazi, and Pascal Van Hentenryck. Convex relaxations for gas expansion planning. INFORMS Journal on Computing, 28(4):645–656, Fall 2016. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0697.

Nemati:2016:SPR

[1254] Sepehr Nemati, Oleg V. Shylo, Oleg A. Prokopyev, and Andrew J. Schaefer. The surgical patient routing problem: a central planner approach. INFORMS Journal on Computing, 28(4):657–673, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0706.

Marzouk:2016:BPA

[1255] Ahmed M. Marzouk, Erick Moreno-Centeno, and Halit Üster. A branchand-price algorithm for solving the Hamiltonian p-median problem. IN-FORMS Journal on Computing, 28(4):674–686, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0704.

Liu:2016:CNO

[1256] Lindong Liu, Xiangtong Qi, and Zhou Xu. Computing near-optimal stable cost allocations for cooperative games by Lagrangian relaxation. INFORMS Journal on Computing, 28(4):687–702, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0707.

Fleszar:2016:EAT

[1257] Krzysztof Fleszar. An exact algorithm for the two-dimensional stageunrestricted guillotine cutting/packing decision problem. INFORMS Journal on Computing, 28(4):703–720, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2016.0708.

Han:2016:OLL

[1258] Bin Han, Ilya O. Ryzhov, and Boris Defourny. Optimal learning in linear regression with combinatorial feature selection. INFORMS Journal on Computing, 28(4):721–735, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2016.0709.

Furini:2016:MTD

[1259] Fabio Furini, Enrico Malaguti, and Dimitri Thomopulos. Modeling twodimensional guillotine cutting problems via integer programming. IN-FORMS Journal on Computing, 28(4):736–751, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0710.

Coutinho:2016:BBA

[1260] Walton Pereira Coutinho, Roberto Quirino do Nascimento, Artur Alves Pessoa, and Anand Subramanian. A branch-and-bound algorithm for the close-enough traveling salesman problem. INFORMS Journal on Computing, 28(4):752–765, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2016.0711.

Akartunali:2016:LCT

[1261] Kerem Akartunalı, Ioannis Fragkos, Andrew J. Miller, and Tao Wu. Local cuts and two-period convex hull closures for big-bucket lot-sizing problems. INFORMS Journal on Computing, 28(4):766–780, Fall 2016. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0712.

Burke:2016:AFD

[1262] Edmund K. Burke and Yuri Bykov. An adaptive flex-deluge approach to university exam timetabling. INFORMS Journal on Computing, 28 (4):781–794, Fall 2016. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2015.0680.

Anonymous:2016:AR

[1263] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 28(4):795–799, Fall 2016. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2016.0734.

Pender:2017:SFK

[1264] Jamol Pender. Sampling the functional Kolmogorov forward equations for nonstationary queueing networks. INFORMS Journal on Computing, 29(1):1–17, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0702.

Zhang:2017:DPN

[1265] Dan Zhang and Larry Weatherford. Dynamic pricing for network revenue management: a new approach and application in the hotel industry. INFORMS Journal on Computing, 29(1):18–35, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0713.

Huang:2017:SMC

[1266] Kuo-Ling Huang and Sanjay Mehrotra. Solution of monotone complementarity and general convex programming problems using a modified potential reduction interior point method. INFORMS Journal on Computing, 29(1):36–53, Winter 2017. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2016.0715.

Davis:2017:PPU

[1267] James M. Davis, Huseyin Topaloglu, and David P. Williamson. Pricing problems under the nested logit model with a quality consistency constraint. INFORMS Journal on Computing, 29(1):54–76, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0714.

Bodur:2017:SBC

[1268] Merve Bodur, Sanjeeb Dash, Oktay Günlük, and James Luedtke. Strengthened Benders cuts for stochastic integer programs with continuous recourse. INFORMS Journal on Computing, 29(1):77–91, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0717.

Hann:2017:CPP

[1269] Il-Horn Hann and JooHee Oh. Combating prerelease piracy: modeling the effects of antipiracy measures in P2P networks. INFORMS Journal on Computing, 29(1):92–107, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2016.0719.

Li:2017:LRP

[1270] Han-Lin Li, Yao-Huei Huang, and Shu-Cherng Fang. Linear reformulation of polynomial discrete programming for fast computation. IN-FORMS Journal on Computing, 29(1):108–122, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0716.

Lozano:2017:BSF

[1271] Leonardo Lozano and J. Cole Smith. A backward sampling framework for interdiction problems with fortification. INFORMS Journal on Computing, 29(1):123–139, Winter 2017. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2016.0721.

Banik:2017:ECA

[1272] A. D. Banik and M. L. Chaudhry. Efficient computational analysis of stationary probabilities for the queueing system BMAP/G/1/N with or without vacation(s). INFORMS Journal on Computing, 29(1):140–151, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0720.

Bansal:2017:PMC

[1273] Manish Bansal and Kiavash Kianfar. Planar maximum coverage location problem with partial coverage and rectangular demand and service zones. INFORMS Journal on Computing, 29(1):152–169, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0722.

Porumbel:2017:CAC

[1274] Daniel Porumbel and François Clautiaux. Constraint aggregation in column generation models for resource-constrained covering problems. IN-FORMS Journal on Computing, 29(1):170–184, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0718.

Alvarez:2017:MLB

[1275] Alejandro Marcos Alvarez, Quentin Louveaux, and Louis Wehenkel. A machine learning-based approximation of strong branching. INFORMS Journal on Computing, 29(1):185–195, Winter 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2016.0723.

Seref:2017:ECP

[1276] Onur Şeref, J. Paul Brooks, Bernice Huang, and Stephen S. Fong. Enumeration and Cartesian product decomposition of alternate optimal fluxes in cellular metabolism. INFORMS Journal on Computing, 29(2):197–210, Spring 2017. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2016.0724.

Romeijnders:2017:AQC

[1277] Ward Romeijnders, David P. Morton, and Maarten H. van der Vlerk, (deceased). Assessing the quality of convex approximations for twostage totally unimodular integer recourse models. INFORMS Journal on Computing, 29(2):211–231, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2016.0725.

Li:2017:IPI

[1278] Feng Li, Zhi-Long Chen, and Lixin Tang. Integrated production, inventory and delivery problems: complexity and algorithms. INFORMS Journal on Computing, 29(2):232–250, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2016.0726.

Sonmez:2017:ATA

[1279] Erkut Sönmez, Alan Scheller-Wolf, and Nicola Secomandi. An analytical throughput approximation for closed fork/join networks. INFORMS Journal on Computing, 29(2):251–267, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2016.0727.

Cai:2017:HDL

[1280] Yuanfeng Cai, Zhengrui Jiang, and Vijay Mookerjee. How to deal with liars? Designing intelligent rule-based expert systems to increase accuracy or reduce cost. INFORMS Journal on Computing, 29(2):268–286, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0728.

Li:2017:MSN

[1281] Yongzhen Li, Jia Shu, Miao Song, Jiawei Zhang, and Huan Zheng. Multisourcing supply network design: two-stage chance-constrained model, tractable approximations, and computational results. INFORMS Journal on Computing, 29(2):287–300, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2016.0730.

Park:2017:AGC

[1282] Young Woong Park, Yan Jiang, Diego Klabjan, and Loren Williams. Algorithms for generalized clusterwise linear regression. INFORMS Journal on Computing, 29(2):301–317, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2016.0729.

Billionnet:2017:UCB

[1283] Alain Billionnet, Sourour Elloumi, Amélie Lambert, and Angelika Wiegele. Using a conic bundle method to accelerate both phases of a quadratic convex reformulation. INFORMS Journal on Computing, 29 (2):318–331, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0731.

Moazeni:2017:PND

[1284] Somayeh Moazeni, Warren B. Powell, Boris Defourny, and Belgacem Bouzaiene-Ayari. Parallel nonstationary direct policy search for riskaverse stochastic optimization. INFORMS Journal on Computing, 29 (2):332–349, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0733.

Ben-Tal:2017:GRO

[1285] Aharon Ben-Tal, Ruud Brekelmans, Dick den Hertog, and Jean-Philippe Vial. Globalized robust optimization for nonlinear uncertain inequalities.

INFORMS Journal on Computing, 29(2):350–366, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0735.

Gerhardt:2017:TNM

[1286] Ira Gerhardt, Barry L. Nelson, and Michael R. Taaffe. Technical note: The MAP_{t} /Ph_{t} /∞ queueing system and multiclass [MAP_{t} /Ph_{t} /∞]^{K} queueing network. INFORMS Journal on Computing, 29(2):367–376, Spring 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0736.

Archetti:2017:MMI

[1287] Claudia Archetti, Natashia Boland, and M. Grazia Speranza. A matheuristic for the multivehicle inventory routing problem. IN-FORMS Journal on Computing, 29(3):377–387, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0737.

Jena:2017:LHL

[1288] Sanjay Dominik Jena, Jean-François Cordeau, and Bernard Gendron. Lagrangian heuristics for large-scale dynamic facility location with generalized modular capacities. INFORMS Journal on Computing, 29(3): 388–404, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0738.

Muter:2017:IAD

[1289] İbrahim Muter and Tevfik Aytekin. Incorporating aggregate diversity in recommender systems using scalable optimization approaches. IN-FORMS Journal on Computing, 29(3):405–421, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0741.

Saberian:2017:SOF

[1290] Fatemeh Saberian, Archis Ghate, and Minsun Kim. Spatiotemporally optimal fractionation in radiotherapy. INFORMS Journal on Computing, 29(3):422–437, Summer 2017. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2016.0740.

Fan:2017:CPP

[1291] Shaokun Fan, Xin Li, and J. Leon Zhao. Collaboration process pattern approach to improving teamwork performance: A data mining-based methodology. INFORMS Journal on Computing, 29(3):438–456, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0739.

Bettinelli:2017:BBA

[1292] Andrea Bettinelli, Valentina Cacchiani, and Enrico Malaguti. A branchand-bound algorithm for the knapsack problem with conflict graph. IN-FORMS Journal on Computing, 29(3):457–473, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0742.

Fast:2017:BDA

[1293] Caleb C. Fast and Illya V. Hicks. A branch decomposition algorithm for the p-median problem. INFORMS Journal on Computing, 29(3): 474–488, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2016.0743.

Pecin:2017:NEE

[1294] Diego Pecin, Claudio Contardo, Guy Desaulniers, and Eduardo Uchoa. New enhancements for the exact solution of the vehicle routing problem with time windows. INFORMS Journal on Computing, 29(3):489–502, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2016.0744.

Deodhar:2017:ALM

[1295] Meghana Deodhar, Joydeep Ghosh, Maytal Saar-Tsechansky, and Vineet Keshari. Active learning with multiple localized regression models. IN-FORMS Journal on Computing, 29(3):503–522, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2016.0732.

Wu:2017:PSM

[1296] Tao Wu, Kerem Akartunalı, Raf Jans, and Zhe Liang. Progressive selection method for the coupled lot-sizing and cutting-stock problem. IN-FORMS Journal on Computing, 29(3):523–543, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0746.

Fukasawa:2017:NSL

[1297] Ricardo Fukasawa and Laurent Poirrier. Numerically safe lower bounds for the capacitated vehicle routing problem. INFORMS Journal on Computing, 29(3):544–557, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2017.0747.

Roshanaei:2017:COR

[1298] Vahid Roshanaei, Curtiss Luong, Dionne M. Aleman, and David R. Urbach. Collaborative operating room planning and scheduling. IN-FORMS Journal on Computing, 29(3):558–580, Summer 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0745.

Muller:2017:SSO

[1299] Juliane Müller. SOCEMO: surrogate optimization of computationally expensive multiobjective problems. INFORMS Journal on Computing, 29(4):581–596, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2017.0749.

Arulselvan:2017:EAD

[1300] Ashwin Arulselvan, Mohsen Rezapour, and Wolfgang A. Welz. Exact approaches for designing multifacility buy-at-bulk networks. INFORMS Journal on Computing, 29(4):597–611, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2017.0752.

Hubner:2017:DIP

[1301] Jens Hübner, Martin Schmidt, and Marc C. Steinbach. A distributed interior-point KKT solver for multistage stochastic optimization. IN-FORMS Journal on Computing, 29(4):612–630, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0748.

Berg:2017:FAM

[1302] Bjorn P. Berg and Brian T. Denton. Fast approximation methods for online scheduling of outpatient procedure centers. INFORMS Journal on Computing, 29(4):631–644, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2017.0750.

Schumacher:2017:ASC

[1303] Kathryn M. Schumacher, Amy E. M. Cohn, and Richard Li-Yang Chen. Algorithm for the N-2 security-constrained unit commitment problem with transmission switching. INFORMS Journal on Computing, 29(4): 645–659, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2017.0751.

Ma:2017:CSC

[1304] Baojun Ma, Qiang Wei, Guoqing Chen, Jin Zhang, and Xunhua Guo. Content and structure coverage: extracting a diverse information subset. INFORMS Journal on Computing, 29(4):660–675, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0753.

Goncalves:2017:GPU

[1305] Alexandre Domingues Gonçalves, Artur Alves Pessoa, Cristiana Bentes, Ricardo Farias, and Lúcia Maria de A. Drummond. A graphics processing unit algorithm to solve the quadratic assignment problem using level-2 reformulation-linearization technique. INFORMS Journal on Computing, 29(4):676–687, Fall 2017. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0754.

Pender:2017:AQL

[1306] Jamol Pender and Young Myoung Ko. Approximations for the queue length distributions of time-varying many-server queues. INFORMS Journal on Computing, 29(4):688–704, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2017.0760.

Shelbourne:2017:VRP

[1307] Benjamin C. Shelbourne, Maria Battarra, and Chris N. Potts. The vehicle routing problem with release and due dates. INFORMS Journal on Computing, 29(4):705–723, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2017.0756.

Shioura:2017:MSS

[1308] Akiyoshi Shioura, Natalia V. Shakhlevich, and Vitaly A. Strusevich. Machine speed scaling by adapting methods for convex optimization with submodular constraints. INFORMS Journal on Computing, 29(4):724– 736, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0758.

Albareda-Sambola:2017:HSF

[1309] Maria Albareda-Sambola, Elena Fernández, and Francisco Saldanha da Gama. Heuristic solutions to the facility location problem with general Bernoulli demands. INFORMS Journal on Computing, 29(4):737– 753, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0755.

Puranik:2017:DPA

[1310] Yash Puranik and Nikolaos V. Sahinidis. Deletion presolve for accelerating infeasibility diagnosis in optimization models. INFORMS Journal on Computing, 29(4):754–766, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2017.0761.

Ortiz-Astorquiza:2017:FAA

[1311] Camilo Ortiz-Astorquiza, Ivan Contreras, and Gilbert Laporte. Formulations and approximation algorithms for multilevel uncapacitated facility location. INFORMS Journal on Computing, 29(4):767–779, Fall 2017. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0757.

Anonymous:2017:AR

[1312] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 29(4):780–783, Fall 2017. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0787.

He:2018:SRC

[1313] Qi-Ming He and Attahiru Sule Alfa. Space reduction for a class of multidimensional Markov chains: A summary and some applications. INFORMS Journal on Computing, 30(1):1–10, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0759.

Shinano:2018:FSM

[1314] Yuji Shinano, Stefan Heinz, Stefan Vigerske, and Michael Winkler. Fiber-SCIP — a shared memory parallelization of SCIP. INFORMS Journal on Computing, 30(1):11–30, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2017.0762.

Zhen:2018:CMV

[1315] Jianzhe Zhen and Dick den Hertog. Computing the maximum volume inscribed ellipsoid of a polytopic projection. INFORMS Journal on Computing, 30(1):31–42, Winter 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0763.

Ahat:2018:IPF

[1316] Betül Ahat, Tınaz Ekim, and Z. Caner Taşkın. Integer programming formulations and Benders decomposition for the maximum induced matching problem. INFORMS Journal on Computing, 30(1):43–56, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0764.

vanAckooij:2018:APB

[1317] Wim van Ackooij, Welington de Oliveira, and Yongjia Song. Adaptive partition-based level decomposition methods for solving two-stage stochastic programs with fixed recourse. INFORMS Journal on Computing, 30(1):57–70, Winter 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0765.

Ma:2018:RES

[1318] Ni Ma and Ward Whitt. A rare-event simulation algorithm for periodic single-server queues. INFORMS Journal on Computing, 30(1):71–89, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0766.

Deng:2018:PSD

[1319] Yan Deng, Shabbir Ahmed, and Siqian Shen. Parallel scenario decomposition of risk-averse 0–1 stochastic programs. INFORMS Journal on Computing, 30(1):90–105, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2017.0767.

Salas:2018:BSA

[1320] Daniel F. Salas and Warren B. Powell. Benchmarking a scalable approximate dynamic programming algorithm for stochastic control of gridlevel energy storage. INFORMS Journal on Computing, 30(1):106–123, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0768.

Pedrosa:2018:ISC

[1321] Lehilton L. C. Pedrosa and Maxim Sviridenko. Integrated supply chain management via randomized rounding. INFORMS Journal on Computing, 30(1):124–136, Winter 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0769.

Li:2018:IUB

[1322] Chu-Min Li, Zhiwen Fang, Hua Jiang, and Ke Xu. Incremental upper bound for the maximum clique problem. INFORMS Journal on Computing, 30(1):137–153, Winter 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0770.

Zhou:2018:GBA

[1323] Enlu Zhou and Shalabh Bhatnagar. Gradient-based adaptive stochastic search for simulation optimization over continuous space. INFORMS Journal on Computing, 30(1):154–167, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2017.0771.

Cordone:2018:BBA

[1324] Roberto Cordone, Pierre Hosteins, and Giovanni Righini. A branchand-bound algorithm for the prize-collecting single-machine scheduling problem with deadlines and total tardiness minimization. INFORMS Journal on Computing, 30(1):168–180, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2017.0772.

Kim:2018:DDM

[1325] Song-Hee Kim, Ward Whitt, and Won Chul Cha. A data-driven model of an appointment-generated arrival process at an outpatient clinic. IN-FORMS Journal on Computing, 30(1):181–199, Winter 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0773.

Ball:2018:OSS

[1326] Robin C. Ball, Juergen Branke, and Stephan Meisel. Optimal sampling for simulated annealing under noise. INFORMS Journal on Computing, 30(1):200–215, Winter 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0774.

Zhang:2018:CMC

[1327] Zhao Zhang, Jiao Zhou, Shaojie Tang, Xiaohui Huang, and Ding-Zhu Du. Computing minimum k-connected m-fold dominating set in general graphs. INFORMS Journal on Computing, 30(2):217–224, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0776.

Zhou:2018:BBE

[1328] Jiao Zhou, Zhao Zhang, Shaojie Tang, Xiaohui Huang, and Ding-Zhu Du. Breaking the O(ln n) barrier: An enhanced approximation algorithm for fault-tolerant minimum weight connected dominating set. INFORMS Journal on Computing, 30(2):225–235, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2017.0775.

Wu:2018:ABS

[1329] Tao Wu, Zhe Liang, and Canrong Zhang. Analytics branching and selection for the capacitated multi-item lot sizing problem with nonidentical machines. INFORMS Journal on Computing, 30(2):236–258, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0777.

Badri:2018:OCT

[1330] Hamidreza Badri, Ehsan Salari, Yoichi Watanabe, and Kevin Leder. Optimizing chemoradiotherapy to target metastatic disease and tumor growth. INFORMS Journal on Computing, 30(2):259–277, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0778.

Wang:2018:SMR

[1331] Bing Wang and Jiaqiao Hu. Some monotonicity results for stochastic kriging metamodels in sequential settings. INFORMS Journal on Computing, 30(2):278–294, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2017.0779.

Mihic:2018:RDS

[1332] Krešimir Mihić, Kevin Ryan, and Alan Wood. Randomized decomposition solver with the quadratic assignment problem as a case study. INFORMS Journal on Computing, 30(2):295–308, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0781.

Geissler:2018:SHD

[1333] Björn Geißler, Antonio Morsi, Lars Schewe, and Martin Schmidt. Solving highly detailed gas transport MINLPs: Block separability and penalty alternating direction methods. INFORMS Journal on Computing, 30 (2):309–323, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2017.0780.

Adelgren:2018:ESP

[1334] Nathan Adelgren, Pietro Belotti, and Akshay Gupte. Efficient storage of Pareto points in biobjective mixed integer programming. INFORMS Journal on Computing, 30(2):324–338, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2017.0783.

Pessoa:2018:ACL

[1335] A. Pessoa, R. Sadykov, E. Uchoa, and F. Vanderbeck. Automation and combination of linear-programming based stabilization techniques in column generation. INFORMS Journal on Computing, 30(2):339–360, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2017.0784.

Lum:2018:OSD

[1336] Oliver Lum, Bruce Golden, and Edward Wasil. An open-source desktop application for generating arc-routing benchmark instances. INFORMS Journal on Computing, 30(2):361–370, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/abs/10.1287/ijoc.2017.0785.

Ozaltin:2018:ODS

[1337] Osman Y. Özaltın, Oleg A. Prokopyev, and Andrew J. Schaefer. Optimal design of the seasonal influenza vaccine with manufacturing autonomy.

INFORMS Journal on Computing, 30(2):371–387, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0786.

Zou:2018:PAS

[1338] Jikai Zou, Shabbir Ahmed, and Xu Andy Sun. Partially adaptive stochastic optimization for electric power generation expansion planning. IN-FORMS Journal on Computing, 30(2):388–401, Spring 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/abs/10.1287/ijoc.2017.0782.

Leitner:2018:DAB

[1339] Markus Leitner, Ivana Ljubić, Martin Luipersbeck, and Markus Sinnl. A dual ascent-based branch-and-bound framework for the prize-collecting Steiner tree and related problems. INFORMS Journal on Computing, 30(2):402–420, Spring 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2017.0788.

Holder:2018:MHJ

[1340] Allen Holder, Fred Murphy, and William Pierskalla. A memorial to Harvey J. Greenberg, Founding Editor of the INFORMS Journal on Computing. INFORMS Journal on Computing, 30(3):421–423, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0843.

Buchheim:2018:QCO

[1341] Christoph Buchheim and Emiliano Traversi. Quadratic combinatorial optimization using separable underestimators. INFORMS Journal on Computing, 30(3):424–437, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2017.0789.

Wang:2018:CSN

[1342] Mingzheng Wang, Zhengrui Jiang, Yu Zhang, and Haifang Yang. T - closeness slicing: A new privacy-preserving approach for transactional data publishing. INFORMS Journal on Computing, 30(3):438–453, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2017.0791.

Cui:2018:PON

[1343] Xueting Cui, Xiaoling Sun, Shushang Zhu, Rujun Jiang, and Duan Li. Portfolio optimization with nonparametric value at risk: A block coordinate descent method. INFORMS Journal on Computing, 30(3):454–471, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2017.0793.

Fairbrother:2018:SGS

[1344] Jamie Fairbrother, Amanda Turner, and Stein W. Wallace. Scenario generation for single-period portfolio selection problems with tail risk measures: Coping with high dimensions and integer variables. IN-FORMS Journal on Computing, 30(3):472–491, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2017.0790.

Tunc:2018:EMI

[1345] Huseyin Tunc, Onur A. Kilic, S. Armagan Tarim, and Roberto Rossi. An extended mixed-integer programming formulation and dynamic cut generation approach for the stochastic lot-sizing problem. INFORMS Journal on Computing, 30(3):492–506, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2017.0792.

Khademi:2018:SPI

[1346] Amin Khademi and Burak Eksioglu. Spare parts inventory management with substitution-dependent reliability. INFORMS Journal on Computing, 30(3):507–521, Summer 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2017.0794.

Averbakh:2018:LMP

[1347] Igor Averbakh and Jordi Pereira. Lateness minimization in pairwise connectivity restoration problems. INFORMS Journal on Computing, 30(3):522–538, Summer 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2017.0796.

Chicoisne:2018:RAS

[1348] Renaud Chicoisne, Fernando Ordóñez, and Daniel Espinoza. Risk averse shortest paths: A computational study. INFORMS Journal on Computing, 30(3):539–553, Summer 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2017.0795.

Gambella:2018:VRP

[1349] Claudio Gambella, Joe Naoum-Sawaya, and Bissan Ghaddar. The vehicle routing problem with floating targets: Formulation and solution approaches. INFORMS Journal on Computing, 30(3):554–569, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2017.0800.

Khaniyev:2018:SDM

[1350] Taghi Khaniyev, Samir Elhedhli, and Fatih Safa Erenay. Structure detection in mixed-integer programs. INFORMS Journal on Computing, 30(3):570–587, Summer 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2017.0797.

Liu:2018:PSE

[1351] Ran Liu and Xiaolan Xie. Physician staffing for emergency departments with time-varying demand. INFORMS Journal on Computing, 30(3): 588–607, Summer 2018. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2017.0799.

Dunning:2018:WWB

[1352] Iain Dunning, Swati Gupta, and John Silberholz. What works best when? A systematic evaluation of heuristics for Max-Cut and QUBO. INFORMS Journal on Computing, 30(3):608–624, Summer 2018. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2017.0798.

Bektas:2018:DPM

[1353] Tolga Bektaş. Disjunctive programming for multiobjective discrete optimisation. INFORMS Journal on Computing, 30(4):625–633, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2017.0804.

Song:2018:CEB

[1354] Yingda Song, Ning Cai, and Steven Kou. Computable error bounds of Laplace inversion for pricing Asian options. INFORMS Journal on Computing, 30(4):634–645, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2017.0805.

Cote:2018:MMP

[1355] Jean-François Côté and Manuel Iori. The meet-in-the-middle principle for cutting and packing problems. INFORMS Journal on Computing, 30(4):646–661, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0806.

Zhang:2018:EAD

[1356] Yuli Zhang, Zuo-Jun Max Shen, and Shiji Song. Exact algorithms for distributionally β-robust machine scheduling with uncertain processing times. INFORMS Journal on Computing, 30(4):662–676, Fall 2018. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0807.

Fan:2018:SBP

[1357] Qi Fan and Jiaqiao Hu. Surrogate-based promising area search for Lipschitz continuous simulation optimization. INFORMS Journal on Computing, 30(4):677–693, Fall 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2017.0801.

Fukasawa:2018:JVR

[1358] Ricardo Fukasawa, Qie He, Fernando Santos, and Yongjia Song. A joint vehicle routing and speed optimization problem. INFORMS Journal on Computing, 30(4):694–709, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0810.

Ceselli:2018:MPA

[1359] Alberto Ceselli, Maria Luisa Damiani, Giovanni Righini, and Diego Valorsi. Mathematical programming algorithms for spatial cloaking. IN-FORMS Journal on Computing, 30(4):710–723, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0813.

Batur:2018:MSS

[1360] Demet Batur, Lina Wang, and F. Fred Choobineh. Methods for system selection based on sequential mean–variance analysis. INFORMS Journal on Computing, 30(4):724–738, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0808.

Knueven:2018:RPC

[1361] Ben Knueven, Jim Ostrowski, and Jianhui Wang. The ramping polytope and cut generation for the unit commitment problem. INFORMS Journal on Computing, 30(4):739–749, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2017.0802.

Li:2018:KGP

[1362] Yan Li, Kristofer G. Reyes, Jorge Vazquez-Anderson, Yingfei Wang, Lydia M. Contreras, and Warren B. Powell. A knowledge gradient policy for sequencing experiments to identify the structure of RNA molecules using a sparse additive belief model. INFORMS Journal on Computing, 30(4):750–767, Fall 2018. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2017.0803.

Kowalczyk:2018:BPA

[1363] Daniel Kowalczyk and Roel Leus. A branch-and-price algorithm for parallel machine scheduling using ZDDs and generic branching. INFORMS Journal on Computing, 30(4):768–782, Fall 2018. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0809.

Anonymous:2018:AR

[1364] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 30(4):783–786, November 2018. ISSN 1091-9856 (print), 1526-5528 (electronic).

Smith:2019:ME

[1365] Alice E. Smith. Message from the Editor. INFORMS Journal on Computing, 31(1):1, Winter 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0877.

Sun:2019:DDD

[1366] Zhen Sun, Milind Dawande, Ganesh Janakiraman, and Vijay Mookerjee. Data-driven decisions for problems with an unspecified objective function. INFORMS Journal on Computing, 31(1):2–20, Winter 2019. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0818.

Hu:2019:AOS

[1367] Liujia Hu and Sigrún Andradóttir. An asymptotically optimal set approach for simulation optimization. INFORMS Journal on Computing, 31(1):21–39, Winter 2019. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0811.

Mak:2019:DCO

[1368] Terrence W. K. Mak, Pascal Van Hentenryck, Anatoly Zlotnik, and Russell Bent. Dynamic compressor optimization in natural gas pipeline systems. INFORMS Journal on Computing, 31(1):40–65, Winter 2019. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0821.

Adikari:2019:NAR

[1369] Shalinda Adikari and Kaushik Dutta. A new approach to real-time bidding in online advertisements: Auto pricing strategy. INFORMS Journal on Computing, 31(1):66–82, Winter 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0812.

Sumida:2019:AAC

[1370] Mika Sumida and Huseyin Topaloglu. An approximation algorithm for capacity allocation over a single flight leg with fare-locking. INFORMS Journal on Computing, 31(1):83–99, Winter 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0816.

Balvert:2019:ROD

[1371] Marleen Balvert, Dick den Hertog, and Aswin L. Hoffmann. Robust optimization of dose-volume metrics for prostate HDR-brachytherapy incorporating target and OAR volume delineation uncertainties. INFORMS Journal on Computing, 31(1):100–114, Winter 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0815.

Pal:2019:FPL

[1372] Aritra Pal and Hadi Charkhgard. A feasibility pump and local search based heuristic for bi-objective pure integer linear programming. IN-FORMS Journal on Computing, 31(1):115–133, Winter 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0814.

Chitsaz:2019:UDM

[1373] Masoud Chitsaz, Jean-François Cordeau, and Raf Jans. A unified decomposition matheuristic for assembly, production, and inventory routing. INFORMS Journal on Computing, 31(1):134–152, Winter 2019. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0817.

Wang:2019:SND

[1374] Xin Wang, Teodor Gabriel Crainic, and Stein W. Wallace. Stochastic network design for planning scheduled transportation services: The value of deterministic solutions. INFORMS Journal on Computing, 31 (1):153–170, Winter 2019. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0819.

Leitner:2019:EAN

[1375] Markus Leitner, Ivana Ljubić, Martin Riedler, and Mario Ruthmair. Exact approaches for network design problems with relays. INFORMS Journal on Computing, 31(1):171–192, Winter 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0820.

Du:2019:TPD

[1376] Bowen Du, Wenjun Zhou, Chuanren Liu, Yifeng Cui, and Hui Xiong. Transit pattern detection using tensor factorization. INFORMS Journal on Computing, 31(2):193–206, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0824.

Roy:2019:PCM

[1377] Asim Roy, Shiban Qureshi, Kartikeya Pande, Divitha Nair, Kartik Gairola, Pooja Jain, Suraj Singh, Kirti Sharma, Akshay Jagadale, Yi-Yang Lin, Shashank Sharma, Ramya Gotety, Yuexin Zhang, Ji Tang, Tejas Mehta, Hemanth Sindhanuru, Nonso Okafor, Santak Das, Chidambara N. Gopal, Srinivasa B. Rudraraju, and Avinash V. Kakarlapudi. Performance comparison of machine learning platforms. INFORMS Journal on Computing, 31(2):207–225, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0825.

Lu:2019:ECS

[1378] Hao-Chun Lu and Liming Yao. Efficient convexification strategy for generalized geometric programming problems. INFORMS Journal on Computing, 31(2):226–234, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0850.

Pei:2019:CVN

[1379] Jun Pei, Zorica Dražić, Milan Dražić, Nenad Mladenović, and Panos M. Pardalos. Continuous variable neighborhood search (C-VNS) for solving systems of nonlinear equations. INFORMS Journal on Computing, 31 (2):235–250, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0876.

Sadykov:2019:PHB

[1380] Ruslan Sadykov, François Vanderbeck, Artur Pessoa, Issam Tahiri, and Eduardo Uchoa. Primal heuristics for branch and price: The assets of diving methods. INFORMS Journal on Computing, 31(2):251–267, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0822.

Agarwal:2019:NVI

[1381] Yogesh Kumar Agarwal and Prahalad Venkateshan. New valid inequalities for the optimal communication spanning tree problem. INFORMS Journal on Computing, 31(2):268–284, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0827.

Tjandraatmadja:2019:TCR

[1382] Christian Tjandraatmadja and Willem-Jan van Hoeve. Target cuts from relaxed decision diagrams. INFORMS Journal on Computing, 31(2): 285–301, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0830.

Lee:2019:SLB

[1383] Ilbin Lee, Stewart Curry, and Nicoleta Serban. Solving large batches of linear programs. INFORMS Journal on Computing, 31(2):302–317, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0838.

Aggarwal:2019:MHD

[1384] Manish Aggarwal and Ali Fallah Tehrani. Modelling human decision behaviour with preference learning. INFORMS Journal on Computing, 31(2):318–334, Spring 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0823.

Poikonen:2019:BBA

[1385] Stefan Poikonen, Bruce Golden, and Edward A. Wasil. A branch-and-bound approach to the traveling salesman problem with a drone. IN-FORMS Journal on Computing, 31(2):335–346, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0826.

Liu:2019:MSN

[1386] Ran Liu, Michael E. Kuhl, Yunan Liu, and James R. Wilson. Modeling and simulation of nonstationary non-Poisson arrival processes. IN-FORMS Journal on Computing, 31(2):347–366, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0828.

Veremyev:2019:FCL

[1387] Alexander Veremyev, Oleg A. Prokopyev, and Eduardo L. Pasiliao. Finding critical links for closeness centrality. INFORMS Journal on Computing, 31(2):367–389, Spring 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0829.

Fischetti:2019:IGM

[1388] Matteo Fischetti, Ivana Ljubić, Michele Monaci, and Markus Sinnl. Interdiction games and monotonicity, with application to knapsack problems. INFORMS Journal on Computing, 31(2):390–410, Spring 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0831.

Smith:2019:NE

[1389] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 31(3):411–412, Summer 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2019.0917.

Passchyn:2019:NWS

[1390] Ward Passchyn, Dirk Briskorn, and Frits C. R. Spieksma. No-wait scheduling for locks. INFORMS Journal on Computing, 31(3):413–428, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0848.

Petit:2019:ESC

[1391] Thierry Petit and Andrew C. Trapp. Enriching solutions to combinatorial problems via solution engineering. INFORMS Journal on Computing, 31(3):429–444, Summer 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0855.

Bertsimas:2019:RML

[1392] Dimitris Bertsimas and Omid Nohadani. Robust maximum likelihood estimation. INFORMS Journal on Computing, 31(3):445–458, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0834.

Trapp:2019:IFG

[1393] Andrew C. Trapp, Wen Liu, and Soussan Djamasbi. Identifying fixations in gaze data via inner density and optimization. INFORMS Journal on Computing, 31(3):459–476, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0859.

Bergman:2019:EAQ

[1394] David Bergman. An exact algorithm for the quadratic multiknapsack problem with an application to event seating. INFORMS Journal on Computing, 31(3):477–492, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0840.

Tang:2019:ISP

[1395] Lixin Tang, Feng Li, and Zhi-Long Chen. Integrated scheduling of production and two-stage delivery of make-to-order products: Offline and online algorithms. INFORMS Journal on Computing, 31(3):493–514, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0842.

Chen:2019:CMP

[1396] Wu-Lin Chen. Computing the moments of polling models with batch Poisson arrivals by transform inversion. INFORMS Journal on Computing, 31(3):515–526, Summer 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0844.

Herszterg:2019:TDP

[1397] Ian Herszterg, Marcus Poggi, and Thibaut Vidal. Two-dimensional phase unwrapping via balanced spanning forests. INFORMS Journal on Computing, 31(3):527–543, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0832.

Ungun:2019:RTR

[1398] Baris Ungun, Lei Xing, and Stephen Boyd. Real-time radiation treatment planning with optimality guarantees via cluster and bound methods. IN-FORMS Journal on Computing, 31(3):544–558, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0841.

Deng:2019:CCS

[1399] Yan Deng, Siqian Shen, and Brian Denton. Chance-constrained surgery planning under conditions of limited and ambiguous data. IN-FORMS Journal on Computing, 31(3):559–575, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0835.

Lin:2019:VSS

[1400] Yujing Lin, Barry L. Nelson, and Linda Pei. Virtual statistics in simulation via k nearest neighbors. INFORMS Journal on Computing, 31(3): 576–592, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0839.

Deng:2019:DPA

[1401] Tianhu Deng, Yong Liang, Shixuan Zhang, Jingze Ren, and Shuyi Zheng. A dynamic programming approach to power consumption minimization in gunbarrel natural gas networks with nonidentical compressor units. INFORMS Journal on Computing, 31(3):593–611, Summer 2019. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0833.

Fukasawa:2019:PFS

[1402] Ricardo Fukasawa and Laurent Poirrier. Permutations in the factorization of simplex bases. INFORMS Journal on Computing, 31(3):612–632, Summer 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0862.

Smith:2019:MSA

[1403] J. Cole Smith. In memoriam: Shabbir Ahmed (1969–2019). INFORMS Journal on Computing, 31(4):633–635, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0931.

Blom:2019:TCM

[1404] Michelle Blom, Peter J. Stuckey, and Vanessa J. Teague. Toward computing the margin of victory in single transferable vote elections. INFORMS Journal on Computing, 31(4):636–653, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0853.

Guo:2019:VRS

[1405] Zhaoxia Guo, Stein W. Wallace, and Michal Kaut. Vehicle routing with space- and time-correlated stochastic travel times: Evaluating the objective function. INFORMS Journal on Computing, 31(4):654–670, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0906.

Raschid:2019:LRE

[1406] Louiqa Raschid, Hassan Sayyadi, and Vagelis Hristidis. Learning to rank in entity relationship graphs. INFORMS Journal on Computing, 31(4):671–688, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0837.

Muller:2019:SOC

[1407] Juliane Müller and Marcus Day. Surrogate optimization of computationally expensive black-box problems with hidden constraints. INFORMS Journal on Computing, 31(4):689–702, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0864.

Vogiatzis:2019:IEP

[1408] Chrysafis Vogiatzis and Mustafa Can Camur. Identification of essential proteins using induced stars in protein–protein interaction networks. INFORMS Journal on Computing, 31(4):703–718, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0872.

Abbou:2019:GMR

[1409] Abderrahmane Abbou and Viliam Makis. Group maintenance: A restless bandits approach. INFORMS Journal on Computing, 31(4):719– 731, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0863.

Fugener:2019:POV

[1410] Andreas Fügener and Jens O. Brunner. Planning for overtime: The value of shift extensions in physician scheduling. INFORMS Journal on Computing, 31(4):732–744, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0865.

Shi:2019:EMM

[1411] Wen Shi, Xi Chen, and Jennifer Shang. An efficient Morris methodbased framework for simulation factor screening. INFORMS Journal on Computing, 31(4):745–770, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0836.

Date:2019:LRL

[1412] Ketan Date and Rakesh Nagi. Level 2 reformulation linearization technique-based parallel algorithms for solving large quadratic assignment problems on graphics processing unit clusters. INFORMS Journal on Computing, 31(4):771–789, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0866.

Gadegaard:2019:BOB

[1413] Sune Lauth Gadegaard, Lars Relund Nielsen, and Matthias Ehrgott. Biobjective branch-and-cut algorithms based on LP relaxation and bound sets. INFORMS Journal on Computing, 31(4):790–804, Fall 2019. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0846.

Parragh:2019:BBB

[1414] Sophie N. Parragh and Fabien Tricoire. Branch-and-bound for Biobjective integer programming. INFORMS Journal on Computing, 31 (4):805–822, Fall 2019. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0856.

Altamiranda:2019:NEA

[1415] Alvaro Sierra Altamiranda and Hadi Charkhgard. A new exact algorithm to optimize a linear function over the set of efficient solutions for biobjective mixed integer linear programs. INFORMS Journal on Computing, 31(4):823–840, Fall 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0851.

Anonymous:2019:AR

[1416] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 31(4):841–845, Fall 2019. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2019.0932.

Smith:2020:NEa

[1417] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 32(1):1–2, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2019.0948.

Benade:2020:OBB

[1418] Gerdus Benadè and John N. Hooker. Optimization bounds from the branching dual. INFORMS Journal on Computing, 32(1):3–15, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0884.

Perini:2020:CSM

[1419] Tyler Perini, Natashia Boland, Diego Pecin, and Martin Savelsbergh. A criterion space method for biobjective mixed integer programming: The boxed line method. INFORMS Journal on Computing, 32(1):16–39, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0887.

Xia:2020:GSN

[1420] Wei Xia, Juan C. Vera, and Luis F. Zuluaga. Globally solving nonconvex quadratic programs via linear integer programming techniques. INFORMS Journal on Computing, 32(1):40–56, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0883.

Jornada:2020:FAB

[1421] Daniel Jornada and V. Jorge Leon. Filtering algorithms for biobjective mixed binary linear optimization problems with a multiplechoice constraint. INFORMS Journal on Computing, 32(1):57–73, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0891.

Wu:2020:ATS

[1422] Qinghua Wu, Yang Wang, and Fred Glover. Advanced tabu search algorithms for bipartite Boolean quadratic programs guided by strategic oscillation and path relinking. INFORMS Journal on Computing, 32 (1):74–89, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0871.

Oliveira:2020:IBC

[1423] Daniel Oliveira and Artur Pessoa. An improved branch-cut-and-price algorithm for parallel machine scheduling problems. INFORMS Journal on Computing, 32(1):90–100, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0854.

Delorme:2020:EPP

[1424] Maxence Delorme and Manuel Iori. Enhanced pseudo-polynomial formulations for bin packing and cutting stock problems. INFORMS Journal on Computing, 32(1):101–119, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0880.

Arslan:2020:FNF

[1425] Okan Arslan, Ola Jabali, and Gilbert Laporte. A flexible, natural formulation for the network design problem with vulnerability constraints. IN-FORMS Journal on Computing, 32(1):120–134, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0869.

LEcuyer:2020:SAM

[1426] Pierre L’Ecuyer, Paul Wambergue, and Erwan Bourceret. Spectral analysis of the MIXMAX random number generators. INFORMS Journal on Computing, 32(1):135–144, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2018.0878.

Bakir:2020:SSS

[1427] Ilke Bakir, Natashia Boland, Brian Dandurand, and Alan Erera. Sampling scenario set partition dual bounds for multistage stochastic programs. INFORMS Journal on Computing, 32(1):145–163, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0885.

Xie:2020:RCM

[1428] Jietao Xie and Juan Wu. Recursive calculation model for a special multivariate normal probability of first-order stationary sequence. INFORMS Journal on Computing, 32(1):164–171, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0852.

Mocquard:2020:PAR

[1429] Yves Mocquard, Bruno Sericola, and Emmanuelle Anceaume. Probabilistic analysis of rumor-spreading time. INFORMS Journal on Computing, 32(1):172–181, Winter 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0845.

Won:2020:COG

[1430] Daehan Won, Hasan Manzour, and Wanpracha Chaovalitwongse. Convex optimization for group feature selection in networked data. INFORMS Journal on Computing, 32(1):182–198, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0868.

Anonymous:2020:EBa

[1431] Anonymous. Editorial Board. INFORMS Journal on Computing, 32(1): ??, Winter 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.eb.v3201.

Smith:2020:NEb

[1432] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 32(2):199–200, Spring 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0968.

Mittal:2020:RQP

[1433] Areesh Mittal, Can Gokalp, and Grani A. Hanasusanto. Robust quadratic programming with mixed-integer uncertainty. INFORMS Journal on Computing, 32(2):201–218, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0901.

Hu:2020:SQS

[1434] Hao Hu and Renata Sotirov. On solving the quadratic shortest path problem. INFORMS Journal on Computing, 32(2):219–233, Spring 2020. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0861.

Nemati:2020:MIA

[1435] Sepehr Nemati, Zeynep G. Icten, Lisa M. Maillart, and Andrew J. Schaefer. Mitigating information asymmetry in liver allocation. INFORMS Journal on Computing, 32(2):234–248, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2018.0874.

Poikonen:2020:MDR

[1436] Stefan Poikonen and Bruce Golden. The mothership and drone routing problem. INFORMS Journal on Computing, 32(2):249–262, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0879.

Castro:2020:MBL

[1437] Margarita P. Castro, Andre A. Cire, and J. Christopher Beck. An MDD-based Lagrangian approach to the multicommodity pickup-and-delivery TSP. INFORMS Journal on Computing, 32(2):263–278, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0881.

Brown:2020:SNS

[1438] Gerald G. Brown and W. Matthew Carlyle. Solving the nearly symmetric all-pairs shortest-path problem. INFORMS Journal on Computing, 32 (2):279–288, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0873.

Gunnec:2020:LCI

[1439] Dilek Günneç, S. Raghavan, and Rui Zhang. Least-cost influence maximization on social networks. INFORMS Journal on Computing, 32(2): 289–302, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2019.0886.

Cheng:2020:UII

[1440] Xiaoye Cheng, Jingjing Zhang, and Lu (Lucy) Yan. Understanding the impact of individual users’ rating characteristics on the predictive accuracy of recommender systems. INFORMS Journal on Computing, 32 (2):303–320, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0882.

Guo:2020:SAF

[1441] Zhiling Guo, Jin Li, and Ram Ramesh. Scalable, adaptable, and fast estimation of transient downtime in virtual infrastructures using convex decomposition and sample path randomization. INFORMS Journal on Computing, 32(2):321–345, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0888.

Atamturk:2020:SQU

[1442] Alper Atamtürk, Carlos Deck, and Hyemin Jeon. Successive quadratic upper-bounding for discrete mean-risk minimization and network interdiction. INFORMS Journal on Computing, 32(2):346–355, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0870.

Jiang:2020:ORM

[1443] Guangxin Jiang, L. Jeff Hong, and Barry L. Nelson. Online risk monitoring using offline simulation. INFORMS Journal on Computing, 32 (2):356–375, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2019.0892.

Jian:2020:EPF

[1444] Nanjing Jian and Shane G. Henderson. Estimating the probability that a function observed with noise is convex. INFORMS Journal on Computing, 32(2):376–389, Spring 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0847.

Cui:2020:VSR

[1445] Zhenyu Cui, Michael C. Fu, Jian-Qiang Hu, Yanchu Liu, Yijie Peng, and Lingjiong Zhu. On the variance of single-run unbiased stochastic derivative estimators. INFORMS Journal on Computing, 32(2):390–407, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0897.

Bertsimas:2020:RRA

[1446] Dimitris Bertsimas and Iain Dunning. Relative robust and adaptive optimization. INFORMS Journal on Computing, 32(2):408–427, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0860.

Wei:2020:NBP

[1447] Lijun Wei, Zhixing Luo, Roberto Baldacci, and Andrew Lim. A new branch-and-price-and-cut algorithm for one-dimensional bin-packing problems. INFORMS Journal on Computing, 32(2):428–443, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2018.0867.

Kramer:2020:MMS

[1448] Raphael Kramer, Manuel Iori, and Thibaut Vidal. Mathematical models and search algorithms for the capacitated p-center problem. INFORMS Journal on Computing, 32(2):444–460, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0889.

Pettersson:2020:MIP

[1449] William Pettersson and Melih Ozlen. Multiobjective integer programming: Synergistic parallel approaches. INFORMS Journal on Computing, 32(2):461–472, Spring 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2018.0875.

Achterberg:2020:PRM

[1450] Tobias Achterberg, Robert E. Bixby, Zonghao Gu, Edward Rothberg, and Dieter Weninger. Presolve reductions in mixed integer programming. INFORMS Journal on Computing, 32(2):473–506, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2018.0857.

Rebennack:2020:PLF

[1451] Steffen Rebennack and Vitaliy Krasko. Piecewise linear function fitting via mixed-integer linear programming. INFORMS Journal on Computing, 32(2):507–530, Spring 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2019.0890.

Anonymous:2020:EBb

[1452] Anonymous. Editorial Board. INFORMS Journal on Computing, 32(2): ??, Spring 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.eb.v3202.

Kong:2020:DCP

[1453] Lingxun Kong and Christos T. Maravelias. On the derivation of continuous piecewise linear approximating functions. INFORMS Journal on Computing, 32(3):531–546, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0949.

Zhang:2020:BPC

[1454] Zheng Zhang, Brian T. Denton, and Xiaolan Xie. Branch and price for chance-constrained bin packing. INFORMS Journal on Computing, 32(3):547–564, Summer 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2019.0894.

Hale:2020:DMN

[1455] Joshua Q. Hale, Helin Zhu, and Enlu Zhou. Domination measure: A new metric for solving multiobjective optimization. INFORMS Journal on Computing, 32(3):565–581, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0920.

Deleplanque:2020:BPC

[1456] Samuel Deleplanque, Martine Labbé, Diego Ponce, and Justo Puerto. A branch-price-and-cut procedure for the discrete ordered median problem. INFORMS Journal on Computing, 32(3):582–599, Summer 2020. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0915.

Park:2020:MMC

[1457] Young Woong Park. MILP models for complex system reliability redundancy allocation with mixed components. INFORMS Journal on Computing, 32(3):600–619, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0895.

Wang:2020:DRD

[1458] Shuming Wang and Yan-Fu Li. Distributionally robust design for redundancy allocation. INFORMS Journal on Computing, 32(3):620–640, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0907.

Yekta:2020:OBM

[1459] Hoda Atef Yekta and Robert Day. Optimization-based mechanisms for the course allocation problem. INFORMS Journal on Computing, 32(3): 641–660, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2018.0849.

Subramanyam:2020:ROB

[1460] Anirudh Subramanyam, Panagiotis P. Repoussis, and Chrysanthos E. Gounaris. Robust optimization of a broad class of heterogeneous vehicle routing problems under demand uncertainty. INFORMS Journal on Computing, 32(3):661–681, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0923.

Bestuzheva:2020:CRQ

[1461] Ksenia Bestuzheva, Hassan Hijazi, and Carleton Coffrin. Convex relaxations for quadratic on/off constraints and applications to optimal transmission switching. INFORMS Journal on Computing, 32(3):682–696, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0900.

Zhang:2020:TSN

[1462] Shixuan Zhang, Sheng Liu, Tianhu Deng, and Zuo-Jun Max Shen. Transient-state natural gas transmission in gunbarrel pipeline networks. INFORMS Journal on Computing, 32(3):697–713, Summer 2020. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0904.

Zhou:2020:LCA

[1463] Fan Zhou, Kunpeng Zhang, Shuying Xie, and Xucheng Luo. Learning to correlate accounts across online social networks: an embeddingbased approach. INFORMS Journal on Computing, 32(3):714–729, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0911.

Sokol:2020:BAP

[1464] Vladyslav Sokol, Ante Ćustić, Abraham P. Punnen, and Binay Bhattacharya. Bilinear assignment problem: Large neighborhoods and experimental analysis of algorithms. INFORMS Journal on Computing, 32(3):730–746, Summer 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2019.0893.

Hosseinian:2020:LBC

[1465] Seyedmohammadhossein Hosseinian, Dalila B. M. M. Fontes, and Sergiy Butenko. A Lagrangian bound on the clique number and an exact algorithm for the maximum edge weight clique problem. INFORMS Journal on Computing, 32(3):747–762, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0898.

Miao:2020:EBS

[1466] Zhuqi Miao and Balabhaskar Balasundaram. An ellipsoidal bounding scheme for the quasi-clique number of a graph. INFORMS Journal on Computing, 32(3):763–778, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0922.

Zhou:2020:DPD

[1467] Yaqin Zhou and Shaojie Tang. Differentially private distributed learning. INFORMS Journal on Computing, 32(3):779–789, Summer 2020. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0912.

Han:2020:DPB

[1468] Kai Han, Yuntian He, Alex X. Liu, Shaojie Tang, and He Huang. Differentially private and budget-limited bandit learning over matroids. IN-FORMS Journal on Computing, 32(3):790–804, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0903.

Ryan:2020:ODS

[1469] Kevin Ryan, Shabbir Ahmed, Santanu S. Dey, Deepak Rajan, Amelia Musselman, and Jean-Paul Watson. Optimization-driven scenario grouping. INFORMS Journal on Computing, 32(3):805–821, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0924.

Balas:2020:WLP

[1470] Egon Balas and Thiago Serra. When lift-and-project cuts are different. INFORMS Journal on Computing, 32(3):822–834, Summer 2020. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0943.

Liu:2020:AOU

[1471] Nan Liu, Yuhang Ma, and Huseyin Topaloglu. Assortment optimization under the multinomial logit model with sequential offerings. IN-FORMS Journal on Computing, 32(3):835–853, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0910.

Anonymous:2020:EBc

[1472] Anonymous. Editorial Board. INFORMS Journal on Computing, 32 (3):??, Summer 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.eb.v3203.

Smith:2020:NEc

[1473] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 32(4):855–856, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.1024.

Knueven:2020:MIP

[1474] Bernard Knueven, James Ostrowski, and Jean-Paul Watson. On mixedinteger programming formulations for the unit commitment problem. INFORMS Journal on Computing, 32(4):857–876, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0944.

Ahuja:2020:AAR

[1475] Vishal Ahuja and John R. Birge. An approximation approach for response-adaptive clinical trial design. INFORMS Journal on Computing, 32(4):877–894, Fall 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0969.

Aprahamian:2020:OGT

[1476] Hrayer Aprahamian, Douglas R. Bish, and Ebru K. Bish. Optimal group testing: Structural properties and robust solutions, with application to public health screening. INFORMS Journal on Computing, 32(4):895– 911, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0942.

Barnes:2020:EPD

[1477] Sean L. Barnes, Miranda Myers, Clare Rock, Daniel J. Morgan, Lisa Pineles, Kerri A. Thom, and Anthony D. Harris. Evaluating a predictiondriven targeting strategy for reducing the transmission of multidrugresistant organisms. INFORMS Journal on Computing, 32(4):912–929, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0916.

Ajdari:2020:TFL

[1478] Ali Ajdari, Fatemeh Saberian, and Archis Ghate. A theoretical framework for learning tumor dose-response uncertainty in individualized spatiobiologically integrated radiotherapy. INFORMS Journal on Computing, 32(4):930–951, Fall 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2019.0896.

Validi:2020:ODL

[1479] Hamidreza Validi and Austin Buchanan. The optimal design of lowlatency virtual backbones. INFORMS Journal on Computing, 32(4):952– 967, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0914.

Byeon:2020:CCE

[1480] Geunyeong Byeon, Pascal Van Hentenryck, Russell Bent, and Harsha Nagarajan. Communication-constrained expansion planning for resilient distribution systems. INFORMS Journal on Computing, 32(4):968–985, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0899.

Botev:2020:SCR

[1481] Zdravko I. Botev and Pierre L’Ecuyer. Sampling conditionally on a rare event via generalized splitting. INFORMS Journal on Computing, 32 (4):986–995, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2019.0936.

Li:2020:UPU

[1482] Qiudan Li, Daniel Dajun Zeng, David Jingjun Xu, Ruoran Liu, and Riheng Yao. Understanding and predicting users’ rating behavior: A cognitive perspective. INFORMS Journal on Computing, 32(4):996–1011, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0919.

Li:2020:PTI

[1483] Feng Li, Zhou Xu, and Zhi-Long Chen. Production and transportation integration for commit-to-delivery mode with general shipping costs. IN-FORMS Journal on Computing, 32(4):1012–1029, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0935.

Carrabs:2020:AHA

[1484] Francesco Carrabs, Carmine Cerrone, Raffaele Cerulli, and Bruce Golden. An adaptive heuristic approach to compute upper and lower bounds for the close-enough traveling salesman problem. INFORMS Journal on Computing, 32(4):1030–1048, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0962.

Boccia:2020:SPM

[1485] Maurizio Boccia, Antonio Sforza, and Claudio Sterle. Simple pattern minimality problems: Integer linear programming formulations and covering-based heuristic solving approaches. INFORMS Journal on Computing, 32(4):1049–1060, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0940.

Mansini:2020:CBE

[1486] Renata Mansini and Roberto Zanotti. A core-based exact algorithm for the multidimensional multiple choice knapsack problem. INFORMS Journal on Computing, 32(4):1061–1079, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0909.

Cooper:2020:BSO

[1487] Kyle Cooper, Susan R. Hunter, and Kalyani Nagaraj. Biobjective simulation optimization on integer lattices using the epsilon-constraint method in a retrospective approximation framework. INFORMS Journal on Computing, 32(4):1080–1100, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0918.

Cooper:2020:PSM

[1488] Kyle Cooper and Susan R. Hunter. PyMOSO: Software for multiobjective simulation optimization with R-PERLE and R-MinRLE. INFORMS Journal on Computing, 32(4):1101–1108, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0902.

Roos:2020:RCR

[1489] Ernst Roos and Dick den Hertog. Reducing conservatism in robust optimization. INFORMS Journal on Computing, 32(4):1109–1127, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0913.

Zeng:2020:PSC

[1490] Bo Zeng. A practical scheme to compute the pessimistic bilevel optimization problem. INFORMS Journal on Computing, 32(4):1128–1142, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0927.

Zhang:2020:PAS

[1491] Juheng Zhang, Xiaoping Liu, and Xiao-Bai Li. Predictive analytics with strategically missing data. INFORMS Journal on Computing, 32(4): 1143–1156, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2019.0947.

Halman:2020:PNO

[1492] Nir Halman. Provably near-optimal approximation schemes for implicit stochastic and sample-based dynamic programs. INFORMS Journal on Computing, 32(4):1157–1181, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0926.

Anonymous:2020:AR

[1493] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 32(4):1182–1186, Fall 2020. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.1029.

Anonymous:2020:EBd

[1494] Anonymous. Editorial Board. INFORMS Journal on Computing, 32(4): ??, Fall 2020. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.eb.v3204.

Deng:2003:SHD

[1495] Lih-Yuan Deng and Hongquan Xu. A system of high-dimensional, efficient, long-cycle and portable uniform random number generators. ACM Transactions on Modeling and Computer Simulation, 13(4):299–309, October 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Deng:2005:EPM

[1496] Lih-Yuan Deng. Efficient and portable multiple recursive generators of large order. ACM Transactions on Modeling and Computer Simulation, 15(1):1–13, January 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Deng:2009:SPM

[1497] Lih-Yuan Deng, Huajiang Li, and Jyh-Jen Horng Shiau. Scalable parallel multiple recursive generators of large order. Parallel Computing, 35(1):29–37, January 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0167819108001099.

Deng:2012:ECS

[1498] Lih-Yuan Deng, Jyh-Jen H. Shiau, and Henry Horng-Shing Lu. Efficient computer search of large-order multiple recursive pseudo-random number generators. Journal of Computational and Applied Mathematics, 236(13): 3228–3237, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:1998:CBI

[1499] Anonymous, editor. Connections between integer programming and constraint logic programming. INFORMS (Institute for Operations Research and the Management Sciences), Catonsville, MD, USA, 1998. ISSN 1091-9856 (print), 1526-5528 (electronic). i–iv + 261–308 pp. INFORMS J. Comput. 10 (1998), no. 3.

Anonymous:1999:CON

[1500] Anonymous, editor. Combinatorial optimization and network flows. IN-FORMS (Institute for Operations Research and the Management Sciences), Catonsville, MD, USA, 1999. ISSN 1091-9856 (print), 1526-5528 (electronic). i–iv + 125–216 pp. INFORMS J. Comput. 11 (1999), no. 2.

Chinneck:2002:SIM

[1501] John W. Chinneck, editor. Special issue on the merging of mathematical programming and constraint programming. INFORMS (Institute for Operations Research and the Management Sciences), Catonsville, MD, USA, 2002. ISSN 1091-9856 (print), 1526-5528 (electronic). i–iv + 293– 417 pp. INFORMS J. Comput. 14 (2002), no. 4.

Raschid:2003:SIM

[1502] Louiqa Raschid and Alex Tuzhilin, editors. Special issue on mining webbased data for e-business applications. INFORMS (Institute for Operations Research and the Management Sciences), Catonsville, MD, USA, 2003. ISSN 1091-9856 (print), 1526-5528 (electronic). i–ii + 121–230 pp. INFORMS J. Comput. 15 (2003), no. 2.

Smith:2021:NEaa

[1503] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 33(1):1, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.1043.

Jenkins:2021:ADP

[1504] Phillip R. Jenkins, Matthew J. Robbins, and Brian J. Lunday. Approximate dynamic programming for military medical evacuation dispatching policies. INFORMS Journal on Computing, 33(1):2–26, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0930.

Dowson:2021:SJJ

[1505] Oscar Dowson and Lea Kapelevich. SDDP.jl: A Julia package for stochastic dual dynamic programming. INFORMS Journal on Computing, 33(1):27–33, Winter 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0987.

Lohmann:2021:HPP

[1506] Timo Lohmann, Michael R. Bussieck, Lutz Westermann, and Steffen Rebennack. High-performance prototyping of decomposition methods in GAMS. INFORMS Journal on Computing, 33(1):34–50, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0905.

Gangammanavar:2021:SDT

[1507] Harsha Gangammanavar, Yifan Liu, and Suvrajeet Sen. Stochastic decomposition for two-stage stochastic linear programs with random cost coefficients. INFORMS Journal on Computing, 33(1):51–71, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0929.

Tamby:2021:ENS

[1508] Satya Tamby and Daniel Vanderpooten. Enumeration of the nondominated set of multiobjective discrete optimization problems. INFORMS Journal on Computing, 33(1):72–85, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0953.

Ni:2021:BPA

[1509] Wenjun Ni, Jia Shu, Miao Song, Dachuan Xu, and Kaike Zhang. A branch-and-price algorithm for facility location with general facility cost functions. INFORMS Journal on Computing, 33(1):86–104, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0921.

Song:2021:PRB

[1510] Guopeng Song, Tamás Kis, and Roel Leus. Polyhedral results and branch-and-cut for the resource loading problem. INFORMS Journal on Computing, 33(1):105–119, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.0957.

Park:2021:ONE

[1511] Young Woong Park. Optimization for L_{1}-norm error fitting via data aggregation. INFORMS Journal on Computing, 33(1):120–142, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0908.

Doulabi:2021:EST

[1512] Hossein Hashemi Doulabi, Patrick Jaillet, Gilles Pesant, and Louis-Martin Rousseau. Exploiting the structure of two-stage robust optimization models with exponential scenarios. INFORMS Journal on Computing, 33(1):143–162, Winter 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2019.0928.

Zheng:2021:PRS

[1513] Xiaojin Zheng, Yutong Pan, and Zhaolin Hu. Perspective reformulations of semicontinuous quadratically constrained quadratic programs. IN-FORMS Journal on Computing, 33(1):163–179, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0925.

Luo:2021:CRE

[1514] Hezhi Luo, Xiaodong Ding, Jiming Peng, Rujun Jiang, and Duan Li. Complexity results and effective algorithms for worst-case linear optimization under uncertainties. INFORMS Journal on Computing, 33(1): 180–197, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2019.0941.

Kleinert:2021:CFP

[1515] Thomas Kleinert and Martin Schmidt. Computing feasible points of bilevel problems with a penalty alternating direction method. INFORMS Journal on Computing, 33(1):198–215, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0945.

Cai:2021:CAF

[1516] Ning Cai and Xuewei Yang. A computational approach to first passage problems of reflected hyperexponential jump diffusion processes. IN-FORMS Journal on Computing, 33(1):216–229, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0980.

Peng:2021:ESA

[1517] Yijie Peng, Chun-Hung Chen, Michael C. Fu, Jian-Qiang Hu, and Ilya O. Ryzhov. Efficient sampling allocation procedures for optimal quantile selection. INFORMS Journal on Computing, 33(1):230–245, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2019.0946.

Guo:2021:CVB

[1518] Xunhua Guo, Guoqing Chen, Cong Wang, Qiang Wei, and Zunqiang Zhang. Calibration of voting-based helpfulness measurement for online reviews: an iterative Bayesian probability approach. INFORMS Journal on Computing, 33(1):246–261, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0951.

Bertsimas:2021:SCR

[1519] Dimitris Bertsimas and Nishanth Mundru. Sparse convex regression. INFORMS Journal on Computing, 33(1):262–279, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0954.

Zhang:2021:RSM

[1520] Jin Zhang, Cong Wang, and Guoqing Chen. A review selection method for finding an informative subset from online reviews. INFORMS Journal on Computing, 33(1):280–299, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0950.

Canakoglu:2021:IIA

[1521] Ethem Çanakoğlu, İbrahim Muter, and Tevfik Aytekin. Integrating individual and aggregate diversity in top-N recommendation. INFORMS Journal on Computing, 33(1):300–318, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0952.

Ma:2021:MBR

[1522] Zu-Jun Ma, Fei Yang, Ying Dai, and Zuo-Jun Max Shen. The migratory beekeeping routing problem: Model and an exact algorithm. INFORMS Journal on Computing, 33(1):319–335, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0958.

Yang:2021:ROE

[1523] Haoxiang Yang, David P. Morton, Chaithanya Bandi, and Krishnamurthy Dvijotham. Robust optimization for electricity generation. IN-FORMS Journal on Computing, 33(1):336–351, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0956.

Huang:2021:MSP

[1524] Jianqiu Huang, Kai Pan, and Yongpei Guan. Multistage stochastic power generation scheduling co-optimizing energy and ancillary services. IN-FORMS Journal on Computing, 33(1):352–369, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2019.0933.

Dhara:2021:WCE

[1525] Anulekha Dhara, Bikramjit Das, and Karthik Natarajan. Worst-case expected shortfall with univariate and bivariate marginals. INFORMS Journal on Computing, 33(1):370–389, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2019.0939.

Romeijnders:2021:PCD

[1526] Ward Romeijnders and Krzysztof Postek. Piecewise constant decision rules via branch-and-bound based scenario detection for integer adjustable robust optimization. INFORMS Journal on Computing, 33(1): 390–400, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2019.0934.

Bergman:2021:DDD

[1527] David Bergman and Leonardo Lozano. Decision diagram decomposition for quadratically constrained binary optimization. INFORMS Journal on Computing, 33(1):401–418, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0938.

Anonymous:2021:EBa

[1528] Anonymous. Editorial Board. INFORMS Journal on Computing, 33(1): ??, Winter 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2021.eb.v3301.

Smith:2021:NEb

[1529] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 33(2):419–420, Spring 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2021.1076.

Berthold:2021:CAM

[1530] Timo Berthold and Jakob Witzig. Conflict analysis for MINLP. IN-FORMS Journal on Computing, 33(2):421–435, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1050.

Goeke:2021:MSP

[1531] Dominik Goeke and Michael Schneider. Modeling single-picker routing problems in classical and modern warehouses. INFORMS Journal on Computing, 33(2):436–451, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1040.

Luo:2021:BPC

[1532] Zhixing Luo, Hu Qin, T. C. E. Cheng, Qinghua Wu, and Andrew Lim. A branch-and-price-and-cut algorithm for the cable-routing problem in solar power plants. INFORMS Journal on Computing, 33(2):452–476, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0981.

Briskorn:2021:VST

[1533] Dirk Briskorn, Malte Fliedner, and Martin Tschöke. Vehicle sequencing at transshipment terminals with handover relations. INFORMS Journal on Computing, 33(2):477–494, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0964.

Dalmeijer:2021:AOS

[1534] Kevin Dalmeijer and Guy Desaulniers. Addressing orientation symmetry in the time window assignment vehicle routing problem. INFORMS Journal on Computing, 33(2):495–510, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0974.

Jiang:2021:SSQ

[1535] Shan Jiang, Shu-Cherng Fang, and Qingwei Jin. Sparse solutions by a quadratically constrained ℓ_{q}(0 < q < 1) minimization model. INFORMS Journal on Computing, 33(2):511–530, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.1004.

Phan:2021:SCS

[1536] Dzung T. Phan and Matt Menickelly. On the solution of ℓ_{0}-constrained sparse inverse covariance estimation problems. INFORMS Journal on Computing, 33(2):531–550, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.0991.

Gomez:2021:MIF

[1537] Andrés Gómez and Oleg A. Prokopyev. A mixed-integer fractional optimization approach to best subset selection. INFORMS Journal on Computing, 33(2):551–565, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1031.

Cheng:2021:SDS

[1538] Yichen Cheng, Xinlei Wang, and Yusen Xia. Supervised t-distributed stochastic neighbor embedding for data visualization and classification.

INFORMS Journal on Computing, 33(2):566–585, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0961.

Liu:2021:DDP

[1539] Jiapeng Liu, Miłosz Kadziński, Xiuwu Liao, and Xiaoxin Mao. Datadriven preference learning methods for value-driven multiple criteria sorting with interacting criteria. INFORMS Journal on Computing, 33(2): 586–606, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.0977.

Ahmadi:2021:SPN

[1540] Amir Ali Ahmadi and Jeffrey Zhang. Semidefinite programming and Nash equilibria in bimatrix games. INFORMS Journal on Computing, 33(2):607–628, Spring 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0960.

Lopez-Sanchez:2021:NSS

[1541] A. D. López-Sánchez, J. Sánchez-Oro, and M. Laguna. A new scatter search design for multiobjective combinatorial optimization with an application to facility location. INFORMS Journal on Computing, 33(2): 629–642, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.0966.

Sarayloo:2021:LBM

[1542] Fatemeh Sarayloo, Teodor Gabriel Crainic, and Walter Rei. A learningbased matheuristic for stochastic multicommodity network design. IN-FORMS Journal on Computing, 33(2):643–656, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0967.

Batur:2021:SBA

[1543] Demet Batur and F. Fred Choobineh. Selecting the best alternative based on its quantile. INFORMS Journal on Computing, 33(2):657–671, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0965.

Nelson:2021:RSI

[1544] Barry L. Nelson, Alan T. K. Wan, Guohua Zou, Xinyu Zhang, and Xi Jiang. Reducing simulation input-model risk via input model averaging. INFORMS Journal on Computing, 33(2):672–684, Spring 2021. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0994.

Rodrigues:2021:LDR

[1545] Filipe Rodrigues, Agostinho Agra, Cristina Requejo, and Erick Delage. Lagrangian duality for robust problems with decomposable functions: The case of a robust inventory problem. INFORMS Journal on Computing, 33(2):685–705, Spring 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0978.

Witzig:2021:CDH

[1546] Jakob Witzig and Ambros Gleixner. Conflict-driven heuristics for mixed integer programming. INFORMS Journal on Computing, 33(2):706–720, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0973.

Hosseininasab:2021:EMS

[1547] Amin Hosseininasab and Willem-Jan van Hoeve. Exact multiple sequence alignment by synchronized decision diagrams. INFORMS Journal on Computing, 33(2):721–738, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2019.0937.

Xavier:2021:LSL

[1548] Álinson S. Xavier, Feng Qiu, and Shabbir Ahmed. Learning to solve large-scale security-constrained unit commitment problems. INFORMS Journal on Computing, 33(2):739–756, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0976.

Deng:2021:SGD

[1549] Yan Deng, Huiwen Jia, Shabbir Ahmed, Jon Lee, and Siqian Shen. Scenario grouping and decomposition algorithms for chance-constrained programs. INFORMS Journal on Computing, 33(2):757–773, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0970.

Ran:2021:BBE

[1550] Yingli Ran, Zhao Zhang, Shaojie Tang, and Ding-Zhu Du. Breaking the r_{max} barrier: Enhanced approximation algorithms for partial set multicover problem. INFORMS Journal on Computing, 33(2):774–784, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0975.

Estes:2021:MPO

[1551] Alexander S. Estes and Michael O. Ball. Monge properties, optimal greedy policies, and policy improvement for the dynamic stochastic transportation problem. INFORMS Journal on Computing, 33(2):785–807, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0990.

Kibis:2021:MSP

[1552] Eyyüb Y. Kıbış, İ. Esra Büyüktahtakın, Robert G. Haight, Najmaddin Akhundov, Kathleen Knight, and Charles E. Flower. A multistage stochastic programming approach to the optimal surveillance and control of the emerald ash borer in cities. INFORMS Journal on Computing, 33(2):808–834, Spring 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0963.

Boland:2021:CBI

[1553] Natashia Boland, Riley Clement, and Hamish Waterer. Corrigendum to “A Bucket Indexed Formulation for Nonpreemptive Single Machine Scheduling Problems,” INFORMS Journal on Computing 28(1):14–30, 2016. INFORMS Journal on Computing, 33(2):835, Spring 2021. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0979. See

[1211].

Anonymous:2021:EBb

[1554] Anonymous. Editorial Board. INFORMS Journal on Computing, 33(2): ??, Spring 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2021.eb.v3302.

Smith:2021:NEc

[1555] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 33(3):837–838, Summer 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2021.1102.

Cicek:2021:DDU

[1556] Cihan Tugrul Cicek, Zuo-Jun Max Shen, Hakan Gultekin, and Bulent Tavli. 3-D dynamic UAV base station location problem. INFORMS Journal on Computing, 33(3):839–860, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.1034.

Carvalho:2021:RMK

[1557] Margarida Carvalho, Xenia Klimentova, Kristiaan Glorie, Ana Viana, and Miguel Constantino. Robust models for the kidney exchange problem. INFORMS Journal on Computing, 33(3):861–881, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0986.

Liu:2021:TIA

[1558] Shen Liu and Hongyan Liu. Tagging items automatically based on both content information and browsing behaviors. INFORMS Journal on Computing, 33(3):882–897, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1007.

Zhu:2021:MMO

[1559] Zhicheng Zhu, Yisha Xiang, and Bo Zeng. Multicomponent maintenance optimization: a stochastic programming approach. INFORMS Journal on Computing, 33(3):898–914, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0997.

Semelhago:2021:RDO

[1560] Mark Semelhago, Barry L. Nelson, Eunhye Song, and Andreas Wächter. Rapid discrete optimization via simulation with Gaussian Markov random fields. INFORMS Journal on Computing, 33(3):915–930, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0971.

Merzifonluoglu:2021:RAS

[1561] Yasemin Merzifonluoglu and Joseph Geunes. The risk-averse static stochastic knapsack problem. INFORMS Journal on Computing, 33(3): 931–948, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.0972.

Olivier:2021:QMP

[1562] Philippe Olivier, Andrea Lodi, and Gilles Pesant. The quadratic multiknapsack problem with conflicts and balance constraints. INFORMS Journal on Computing, 33(3):949–962, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.0983.

Cote:2021:CBD

[1563] Jean-François Côté, Mohamed Haouari, and Manuel Iori. Combinatorial Benders decomposition for the two-dimensional bin packing problem. IN-FORMS Journal on Computing, 33(3):963–978, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1014.

Wang:2021:ROQ

[1564] Yang Wang, Wei Yang, Abraham P. Punnen, Jingbo Tian, Aihua Yin, and Zhipeng Lü. The rank-one quadratic assignment problem. IN-FORMS Journal on Computing, 33(3):979–996, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1003.

He:2021:TDS

[1565] Edward He, Natashia Boland, George Nemhauser, and Martin Savelsbergh. Time-dependent shortest path problems with penalties and limits on waiting. INFORMS Journal on Computing, 33(3):997–1014, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0985.

Karimi:2021:FSC

[1566] Roya Karimi, Jianqiang Cheng, and Miguel A. Lejeune. A framework for solving chance-constrained linear matrix inequality programs. IN-FORMS Journal on Computing, 33(3):1015–1036, Summer 2021. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0982.

Li:2021:ERF

[1567] Chao Li, Muhong Zhang, and Kory Hedman. Extreme ray feasibility cuts for unit commitment with uncertainty. INFORMS Journal on Computing, 33(3):1037–1055, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.0995.

Bai:2021:ALD

[1568] Xiaodi Bai, Jie Sun, and Xiaojin Zheng. An augmented Lagrangian decomposition method for chance-constrained optimization problems. IN-FORMS Journal on Computing, 33(3):1056–1069, Summer 2021. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1001.

Gschwind:2021:BPF

[1569] Timo Gschwind, Stefan Irnich, Fabio Furini, and Roberto Wolfler Calvo. A branch-and-price framework for decomposing graphs into relaxed cliques. INFORMS Journal on Computing, 33(3):1070–1090, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0984.

Zhang:2021:IBB

[1570] Wenda Zhang, Jason J. Sauppe, and Sheldon H. Jacobson. An improved branch-and-bound algorithm for the one-machine scheduling problem with delayed precedence constraints. INFORMS Journal on Computing, 33(3):1091–1102, Summer 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0988.

Mistry:2021:MIC

[1571] Miten Mistry, Dimitrios Letsios, Gerhard Krennrich, Robert M. Lee, and Ruth Misener. Mixed-integer convex nonlinear optimization with gradient-boosted trees embedded. INFORMS Journal on Computing, 33 (3):1103–1119, Summer 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.0993.

Ji:2021:DDO

[1572] Ran Ji and Miguel A. Lejeune. Data-driven optimization of rewardrisk ratio measures. INFORMS Journal on Computing, 33(3):1120–1137, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.1002.

Ardestani-Jaafari:2021:LRC

[1573] Amir Ardestani-Jaafari and Erick Delage. Linearized robust counterparts of two-stage robust optimization problems with applications in operations management. INFORMS Journal on Computing, 33(3):1138–1161, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0959.

Berktas:2021:BBA

[1574] Nihal Berktaş and Hande Yaman. A branch-and-bound algorithm for team formation on social networks. INFORMS Journal on Computing, 33(3):1162–1176, Summer 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.1000.

Zhao:2021:KLI

[1575] Zifeng Zhao, Peng Shi, and Xiaoping Feng. Knowledge learning of insurance risks using dependence models. INFORMS Journal on Computing, 33(3):1177–1196, Summer 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.1005.

Wu:2021:NCA

[1576] Zeyang Wu, Kameng Nip, and Qie He. A new combinatorial algorithm for separable convex resource allocation with nested bound constraints. INFORMS Journal on Computing, 33(3):1197–1212, Summer 2021. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1006.

Park:2021:EAL

[1577] Jungho Park, Hadi El-Amine, and Nevin Mutlu. An exact algorithm for large-scale continuous nonlinear resource allocation problems with minimax regret objectives. INFORMS Journal on Computing, 33(3): 1213–1228, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.0999.

Zhang:2021:MNP

[1578] Jie Zhang, Weijun Xie, and Subhash C. Sarin. Multiproduct newsvendor problem with customer-driven demand substitution: a stochastic integer program perspective. INFORMS Journal on Computing, 33(3): 1229–1244, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.0996.

Roy:2021:MSI

[1579] Debdatta Sinha Roy, Adriano Masone, Bruce Golden, and Edward Wasil. Modeling and solving the intersection inspection rural postman problem. INFORMS Journal on Computing, 33(3):1245–1257, Summer 2021. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1013.

Anonymous:2021:EBc

[1580] Anonymous. Editorial Board. INFORMS Journal on Computing, 33 (3):??, Summer 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2021.eb.v3303.

Smith:2021:NEd

[1581] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 33(4):1259–1261, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2021.1126.

deMeijer:2021:SBB

[1582] Frank de Meijer and Renata Sotirov. SDP-based bounds for the quadratic cycle cover problem via cutting-plane augmented Lagrangian methods and reinforcement learning. INFORMS Journal on Computing, 33(4): 1262–1276, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2021.1075.

Kullman:2021:FOS

[1583] Nicholas D. Kullman, Aurelien Froger, Jorge E. Mendoza, and Justin C. Goodson. frvcpy: an open-source solver for the fixed route vehicle charging problem. INFORMS Journal on Computing, 33(4):1277–1283, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.1035.

Herzel:2021:AMM

[1584] Arne Herzel, Stefan Ruzika, and Clemens Thielen. Approximation methods for multiobjective optimization problems: a survey. INFORMS Journal on Computing, 33(4):1284–1299, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.1028.

Zolan:2021:DLC

[1585] Alexander J. Zolan, Michael S. Scioletti, David P. Morton, and Alexandra M. Newman. Decomposing loosely coupled mixed-integer programs for optimal microgrid design. INFORMS Journal on Computing, 33(4): 1300–1319, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.0955.

Hu:2021:CST

[1586] Yuheng Hu. Characterizing social TV activity around televised events: a joint topic model approach. INFORMS Journal on Computing, 33 (4):1320–1338, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.1038.

Wang:2021:NLT

[1587] Yining Wang, Yi Wu, and Simon S. Du. Near-linear time local polynomial nonparametric estimation with Box kernels. INFORMS Journal on Computing, 33(4):1339–1353, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1021.

Yang:2021:MBK

[1588] Yu Yang, Natashia Boland, and Martin Savelsbergh. Multivariable branching: a 0–1 knapsack problem case study. INFORMS Journal on Computing, 33(4):1354–1367, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1052.

Cen:2021:NGO

[1589] Xiaoli Cen and Yong Xia. A new global optimization scheme for quadratic programs with low-rank nonconvexity. INFORMS Journal on Computing, 33(4):1368–1383, Fall 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2020.1017.

Brimkov:2021:ICA

[1590] Boris Brimkov, Derek Mikesell, and Illya V. Hicks. Improved computational approaches and heuristics for zero forcing. INFORMS Journal on Computing, 33(4):1384–1399, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1032.

Zhou:2021:UOO

[1591] Fan Zhou, Kunpeng Zhang, Bangying Wu, Yi Yang, and Harry Jiannan Wang. Unifying online and offline preference for social link prediction. INFORMS Journal on Computing, 33(4):1400–1418, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.0989.

Junger:2021:EFO

[1592] Michael Jünger and Sven Mallach. Exact facetial odd-cycle separation for maximum cut and binary quadratic optimization. INFORMS Journal on Computing, 33(4):1419–1430, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1008.

Zhu:2021:EAM

[1593] Xiaojun Zhu and Shaojie Tang. Exact algorithms for the minimum load spanning tree problem. INFORMS Journal on Computing, 33(4):1431– 1445, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.1011.

Zhu:2021:BBA

[1594] Xiaojun Zhu and Shaojie Tang. A branch-and-bound algorithm for building optimal data gathering tree in wireless sensor networks. INFORMS Journal on Computing, 33(4):1446–1460, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.1012.

Wei:2021:IPF

[1595] Ningji Wei, Jose L. Walteros, and Foad Mahdavi Pajouh. Integer programming formulations for minimum spanning tree interdiction. IN-FORMS Journal on Computing, 33(4):1461–1480, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1018.

Gillen:2021:FAC

[1596] Colin P. Gillen, Alexander Veremyev, Oleg A. Prokopyev, and Eduardo L. Pasiliao. Fortification against cascade propagation under uncertainty. INFORMS Journal on Computing, 33(4):1481–1499, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2020.0992.

Shen:2021:RSC

[1597] Haihui Shen, L. Jeff Hong, and Xiaowei Zhang. Ranking and selection with covariates for personalized decision making. INFORMS Journal on Computing, 33(4):1500–1519, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1009.

Glynn:2021:CSD

[1598] Peter W. Glynn, Yijie Peng, Michael C. Fu, and Jian-Qiang Hu. Computing sensitivities for distortion risk measures. INFORMS Journal on Computing, 33(4):1520–1532, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1016.

Conejo:2021:RCP

[1599] Antonio J. Conejo, Nicholas G. Hall, Daniel Zhuoyu Long, and Runhao Zhang. Robust capacity planning for project management. INFORMS Journal on Computing, 33(4):1533–1550, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.1033.

Guo:2021:LBB

[1600] Cheng Guo, Merve Bodur, Dionne M. Aleman, and David R. Urbach. Logic-based Benders decomposition and binary decision diagram based approaches for stochastic distributed operating room scheduling. IN-FORMS Journal on Computing, 33(4):1551–1569, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1036.

Borrero:2021:MDA

[1601] Juan S. Borrero and Leonardo Lozano. Modeling defender-attacker problems as robust linear programs with mixed-integer uncertainty sets. IN-FORMS Journal on Computing, 33(4):1570–1589, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1041.

Torrico:2021:SRS

[1602] Alfredo Torrico, Mohit Singh, Sebastian Pokutta, Nika Haghtalab, Joseph (Seffi) Naor, and Nima Anari. Structured robust submodular maximization: Offline and online algorithms. INFORMS Journal on Computing, 33(4):1590–1607, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.0998.

Zeng:2021:LAR

[1603] Daniel Zeng, Yong Liu, Ping Yan, and Yanwu Yang. Location-aware realtime recommender systems for brick-and-mortar retailers. INFORMS Journal on Computing, 33(4):1608–1623, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/pdf/10.1287/ijoc.2020.1020.

Xavier:2021:MIC

[1604] Álinson S. Xavier, Ricardo Fukasawa, and Laurent Poirrier. Multirow intersection cuts based on the infinity norm. INFORMS Journal on Computing, 33(4):1624–1643, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/pdf/10.1287/ijoc.2020.1027.

Zhang:2021:BIP

[1605] Junlong Zhang and Osman Y. Özaltın. Bilevel integer programs with stochastic right-hand sides. INFORMS Journal on Computing, 33(4): 1644–1660, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/pdf/
10.1287/ijoc.2020.1055.

Wang:2021:CCM

[1606] Shanshan Wang, Jinlin Li, and Sanjay Mehrotra. Chance-constrained multiple bin packing problem with an application to operating room planning. INFORMS Journal on Computing, 33(4):1661–1677, Fall 2021. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/pdf/10.1287/ijoc.2020.1010.

Anonymous:2021:AR

[1607] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 33(4):1678–1684, Fall 2021. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
pdf/10.1287/ijoc.2021.1131.

Anonymous:2021:EBd

[1608] Anonymous. Editorial Board. INFORMS Journal on Computing, 33(4): ??, Fall 2021. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/pdf/10.1287/
ijoc.2021.eb.v3304.

Smith:2022:NEa

[1609] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 34(1):1, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2021.1146.

Brooks:2022:SII

[1610] J. Paul Brooks, Ted Ralphs, and Nicola Secomandi. Special issue of INFORMS Journal on Computing — scalable reinforcement learning algorithms. INFORMS Journal on Computing, 34(1):2–3, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/abs/10.1287/
ijoc.2021.1153.

Custodio:2022:SDS

[1611] Janiele E. S. C. Custodio and Miguel A. Lejeune. Spatiotemporal data set for out-of-hospital cardiac arrests. INFORMS Journal on Computing, 34 (1):4–10, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2020.1022.

Papp:2022:AMP

[1612] Dávid Papp and Sercan Yıldız. Alfonso: Matlab package for nonsymmetric conic optimization. INFORMS Journal on Computing, 34(1): 11–19, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1058.

Selinka:2022:PAT

[1613] Gregor Selinka, Raik Stolletz, and Thomas I. Maindl. Performance approximation for time-dependent queues with generally distributed abandonments. INFORMS Journal on Computing, 34(1):20–38, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1090.

Goldsztajn:2022:SLT

[1614] Diego Goldsztajn, Sem C. Borst, Johan S. H. van Leeuwaarden, Debankur Mukherjee, and Philip A. Whiting. Self-learning threshold-based load balancing. INFORMS Journal on Computing, 34(1):39–54, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2021.1100.

Bentert:2022:PAP

[1615] Matthias Bentert, René van Bevern, André Nichterlein, Rolf Niedermeier, and Pavel V. Smirnov. Parameterized algorithms for powerefficiently connecting wireless sensor networks: Theory and experiments. INFORMS Journal on Computing, 34(1):55–75, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2020.1045.

Kergosien:2022:ELC

[1616] Yannick Kergosien, Antoine Giret, Emmanuel Néron, and Gaël Sauvanet. An efficient label-correcting algorithm for the multiobjective shortest path problem. INFORMS Journal on Computing, 34(1):76–92, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1081.

Camur:2022:SDC

[1617] Mustafa C. Camur, Thomas Sharkey, and Chrysafis Vogiatzis. The star degree centrality problem: a decomposition approach. INFORMS Journal on Computing, 34(1):93–112, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1074.

Araz:2022:AFE

[1618] Ozgur M. Araz, Mayteé Cruz-Aponte, Fernando A. Wilson, Brock W. Hanisch, and Ruth S. Margalit. An analytic framework for effective public health program design using correctional facilities. INFORMS Journal on Computing, 34(1):113–128, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2020.1056.

Zhou:2022:IMC

[1619] Liping Zhou, Na Geng, Zhibin Jiang, and Shan Jiang. Integrated multiresource capacity planning and multitype patient scheduling. INFORMS Journal on Computing, 34(1):129–149, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2020.1048.

Aprahamian:2022:OSP

[1620] Hrayer Aprahamian and Hadi El-Amine. Optimal screening of populations with heterogeneous risk profiles under the availability of multiple tests. INFORMS Journal on Computing, 34(1):150–164, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2020.1051.

Zhang:2022:MMO

[1621] Qiong Zhang, Amin Khademi, and Yongjia Song. Min-max optimal design of two-armed trials with side information. INFORMS Journal on Computing, 34(1):165–182, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1068.

Wang:2022:HFM

[1622] Kanix Wang, Walid Hussain, John R. Birge, Michael D. Schreiber, and Daniel Adelman. A high-fidelity model to predict length of stay in the neonatal intensive care unit. INFORMS Journal on Computing, 34(1): 183–195, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1062.

Zhen:2022:ROM

[1623] Jianzhe Zhen, Frans J. C. T. de Ruiter, Ernst Roos, and Dick den Hertog. Robust optimization for models with uncertain second-order cone and semidefinite programming constraints. INFORMS Journal on Computing, 34(1):196–210, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2020.1025.

Marandi:2022:ESR

[1624] Ahmadreza Marandi, Aharon Ben-Tal, Dick den Hertog, and Bertrand Melenberg. Extending the scope of robust quadratic optimization. INFORMS Journal on Computing, 34(1):211–226, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1059.

Larsen:2022:PTS

[1625] Eric Larsen, Sébastien Lachapelle, Yoshua Bengio, Emma Frejinger, Simon Lacoste-Julien, and Andrea Lodi. Predicting tactical solutions to operational planning problems under imperfect information. INFORMS Journal on Computing, 34(1):227–242, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1091.

Liers:2022:RRF

[1626] Frauke Liers, Lars Schewe, and Johannes Thürauf. Radius of robust feasibility for mixed-integer problems. INFORMS Journal on Computing, 34(1):243–261, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2020.1030.

Yu:2022:TOT

[1627] Qinxiao Yu, Yossiri Adulyasak, Louis-Martin Rousseau, Ning Zhu, and Shoufeng Ma. Team orienteering with time-varying profit. INFORMS Journal on Computing, 34(1):262–280, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2020.1026.

Hermans:2022:EAA

[1628] Ben Hermans, Roel Leus, and Jannik Matuschke. Exact and approximation algorithms for the expanding search problem. INFORMS Journal on Computing, 34(1):281–296, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2020.1047.

MacNeil:2022:IPC

[1629] Moira MacNeil and Merve Bodur. Integer programming, constraint programming, and hybrid decomposition approaches to discretizable distance geometry problems. INFORMS Journal on Computing, 34(1):297– 314, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2020.1039.

Zohali:2022:STA

[1630] Hassan Zohali, Bahman Naderi, and Vahid Roshanaei. Solving the type-2 assembly line balancing with setups using logic-based Benders decomposition. INFORMS Journal on Computing, 34(1):315–332, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2020.1015.

Delage:2022:VRS

[1631] Erick Delage and Ahmed Saif. The value of randomized solutions in mixed-integer distributionally robust optimization problems. INFORMS Journal on Computing, 34(1):333–353, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2020.1042.

Doulabi:2022:SVM

[1632] Hossein Hashemi Doulabi, Shabbir Ahmed, and George Nemhauser. State-variable modeling for a class of two-stage stochastic optimization problems. INFORMS Journal on Computing, 34(1):354–369, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2020.1044.

Tang:2022:CHI

[1633] Ziye Tang, Yang Jiao, and R. Ravi. Combinatorial heuristics for inventory routing problems. INFORMS Journal on Computing, 34(1):370– 384, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1064.

Wu:2022:PSC

[1634] Tao Wu. Predictive search for capacitated multi-item lot sizing problems. INFORMS Journal on Computing, 34(1):385–406, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1073.

Li:2022:GME

[1635] Yongzhen Li, Xueping Li, Jia Shu, Miao Song, and Kaike Zhang. A general model and efficient algorithms for reliable facility location problem under uncertain disruptions. INFORMS Journal on Computing, 34(1): 407–426, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1063.

Serra:2022:TBM

[1636] Thiago Serra, Teng Huang, Arvind U. Raghunathan, and David Bergman. Template-based minor embedding for adiabatic quantum optimization. INFORMS Journal on Computing, 34(1):427–439, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1065.

Loffler:2022:PRA

[1637] Maximilian Löffler, Nils Boysen, and Michael Schneider. Picker routing in AGV-assisted order picking systems. INFORMS Journal on Computing, 34(1):440–462, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2021.1060.

Misra:2022:LCO

[1638] Sidhant Misra, Line Roald, and Yeesian Ng. Learning for constrained optimization: Identifying optimal active constraint sets. INFORMS Journal on Computing, 34(1):463–480, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2020.1037.

Chen:2022:CAS

[1639] Shutong Chen and Weijun Xie. On cluster-aware supervised learning: Frameworks, convergent algorithms, and applications. INFORMS Journal on Computing, 34(1):481–502, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2020.1053.

Adomavicius:2022:IRE

[1640] Gediminas Adomavicius and Yaqiong Wang. Improving reliability estimation for individual numeric predictions: a machine learning approach. INFORMS Journal on Computing, 34(1):503–521, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2020.1019.

Yang:2022:AFR

[1641] Yi Yang, Kunpeng Zhang, and Yangyang Fan. Analyzing firm reports for volatility prediction: a knowledge-driven text-embedding approach. INFORMS Journal on Computing, 34(1):522–540, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2020.1046.

Zhang:2022:DPA

[1642] Zhu Zhang, Xuan Wei, Xiaolong Zheng, Qiudan Li, and Daniel Dajun Zeng. Detecting product adoption intentions via multiview deep learning. INFORMS Journal on Computing, 34(1):541–556, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1083.

Shi:2022:DSA

[1643] Zhongshun Shi, Yijie Peng, Leyuan Shi, Chun-Hung Chen, and Michael C. Fu. Dynamic sampling allocation under finite simulation budget for feasibility determination. INFORMS Journal on Computing, 34(1):557–568, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2020.1057.

Wang:2022:MSO

[1644] Songhao Wang, Szu Hui Ng, and William Benjamin Haskell. A multilevel simulation optimization approach for quantile functions. INFORMS Journal on Computing, 34(1):569–585, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2020.1049.

Zhong:2022:SPP

[1645] Ying Zhong, Shaoxuan Liu, Jun Luo, and L. Jeff Hong. Speeding up Paulson’s procedure for large-scale problems using parallel computing. INFORMS Journal on Computing, 34(1):586–606, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2020.1054.

Kleijnen:2022:STC

[1646] Jack P. C. Kleijnen and Wim C. M. van Beers. Statistical tests for crossvalidation of kriging models. INFORMS Journal on Computing, 34(1): 607–621, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1072.

Meng:2022:CGL

[1647] Qun Meng, Songhao Wang, and Szu Hui Ng. Combined global and local search for optimization with Gaussian process models. INFORMS Journal on Computing, 34(1):622–637, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1078.

Peng:2022:NLR

[1648] Yijie Peng, Li Xiao, Bernd Heidergott, L. Jeff Hong, and Henry Lam. A new likelihood ratio method for training artificial neural networks. INFORMS Journal on Computing, 34(1):638–655, January/ February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1088.

Kuryatnikova:2022:MCS

[1649] Olga Kuryatnikova, Renata Sotirov, and Juan C. Vera. The maximum kcolorable subgraph problem and related problems. INFORMS Journal on Computing, 34(1):656–669, January/February 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1086.

Anonymous:2022:EBa

[1650] Anonymous. Editorial Board. INFORMS Journal on Computing, 34 (1):??, January/February 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2022.eb.v3401.

Smith:2022:NEb

[1651] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 34(2):671, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2022.1173.

Legat:2022:MDS

[1652] Benoît Legat, Oscar Dowson, Joaquim Dias Garcia, and Miles Lubin. MathOptInterface: a data structure for mathematical optimization problems. INFORMS Journal on Computing, 34(2):672–689, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1067.

Yang:2022:OPF

[1653] Haoxiang Yang and Harsha Nagarajan. Optimal power flow in distribution networks under N − 1 disruptions: a multistage stochastic programming approach. INFORMS Journal on Computing, 34(2):690– 709, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2021.1080.

Raghavan:2022:IML

[1654] S. Raghavan and Rui Zhang. Influence maximization with latency requirements on social networks. INFORMS Journal on Computing, 34 (2):710–728, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1095.

Noyan:2022:DRO

[1655] Nilay Noyan, Gábor Rudolf, and Miguel Lejeune. Distributionally robust optimization under a decision-dependent ambiguity set with applications to machine scheduling and humanitarian logistics. INFORMS Journal on Computing, 34(2):729–751, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1096.

Shin:2022:PNS

[1656] Dongwook Shin, Mark Broadie, and Assaf Zeevi. Practical nonparametric sampling strategies for quantile-based ordinal optimization. IN-FORMS Journal on Computing, 34(2):752–768, March/April 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1071.

Liu:2022:IPO

[1657] Junming Liu, Weiwei Chen, Jingyuan Yang, Hui Xiong, and Can Chen. Iterative prediction-and-optimization for e-logistics distribution network design. INFORMS Journal on Computing, 34(2):769–789, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1107.

Hu:2022:SEE

[1658] Yuheng Hu and Yili Hong. SHEDR: an end-to-end deep neural event detection and recommendation framework for hyperlocal news using social media. INFORMS Journal on Computing, 34(2):790–806, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1112.

Bergman:2022:JIP

[1659] David Bergman, Teng Huang, Philip Brooks, Andrea Lodi, and Arvind U. Raghunathan. JANOS: an integrated predictive and prescriptive modeling framework. INFORMS Journal on Computing, 34 (2):807–816, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2020.1023.

Wu:2022:FVW

[1660] Xinyun Wu, Zhipeng Lü, and Fred Glover. A fast vertex weightingbased local search for finding minimum connected dominating sets. IN-FORMS Journal on Computing, 34(2):817–833, March/April 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1106.

Chen:2022:SVH

[1661] Liting Chen, Sebastian Wandelt, Weibin Dai, and Xiaoqian Sun. Scalable vertiport hub location selection for air taxi operations in a metropolitan region. INFORMS Journal on Computing, 34(2):834–856, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1109.

Arslan:2022:DBA

[1662] Ayşe N. Arslan and Boris Detienne. Decomposition-based approaches for a class of two-stage robust binary optimization problems. INFORMS Journal on Computing, 34(2):857–871, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1061.

Rehfeldt:2022:ESP

[1663] Daniel Rehfeldt and Thorsten Koch. On the exact solution of prizecollecting Steiner tree problems. INFORMS Journal on Computing, 34 (2):872–889, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1087.

Contardo:2022:PAA

[1664] Claudio Contardo and Jorge A. Sefair. A progressive approximation approach for the exact solution of sparse large-scale binary interdiction games. INFORMS Journal on Computing, 34(2):890–908, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1085.

Adelgren:2022:BBB

[1665] Nathan Adelgren and Akshay Gupte. Branch-and-bound for biobjective mixed-integer linear programming. INFORMS Journal on Computing, 34(2):909–933, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1092.

Hendel:2022:ESB

[1666] Gregor Hendel, Daniel Anderson, Pierre Le Bodic, and Marc E. Pfetsch. Estimating the size of branch-and-bound trees. INFORMS Journal on Computing, 34(2):934–952, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1103.

Yu:2022:ICG

[1667] Miao Yu, Viswanath Nagarajan, and Siqian Shen. Improving column generation for vehicle routing problems via random coloring and parallelization. INFORMS Journal on Computing, 34(2):953–973, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1105.

Pavlik:2022:IDD

[1668] John A. Pavlik, Edward C. Sewell, and Sheldon H. Jacobson. Iterative deepening dynamically improved bounds bidirectional search. IN-FORMS Journal on Computing, 34(2):974–989, March/April 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1116.

Bergman:2022:NMM

[1669] David Bergman, Merve Bodur, Carlos Cardonha, and Andre A. Cire. Network models for multiobjective discrete optimization. INFORMS Journal on Computing, 34(2):990–1005, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1066.

Coniglio:2022:OCR

[1670] Stefano Coniglio and Stefano Gualandi. Optimizing over the closure of rank inequalities with a small right-hand side for the maximum stable set problem via bilevel programming. INFORMS Journal on Computing, 34 (2):1006–1023, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1115.

Quezada:2022:CPA

[1671] Franco Quezada, Céline Gicquel, and Safia Kedad-Sidhoum. Combining polyhedral approaches and stochastic dual dynamic integer programming for solving the uncapacitated lot-sizing problem under uncertainty. INFORMS Journal on Computing, 34(2):1024–1041, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1118.

Warwicker:2022:CTM

[1672] John Alasdair Warwicker and Steffen Rebennack. A comparison of two mixed-integer linear programs for piecewise linear function fitting. INFORMS Journal on Computing, 34(2):1042–1047, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1114.

Li:2022:NFL

[1673] Yantong Li, Jean-François Côté, Leandro Callegari-Coelho, and Peng Wu. Novel formulations and logic-based Benders decomposition for the integrated parallel machine scheduling and location problem. INFORMS Journal on Computing, 34(2):1048–1069, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1113.

Mehrani:2022:MAP

[1674] Saharnaz Mehrani, Carlos Cardonha, and David Bergman. Models and algorithms for the bin-packing problem with minimum color fragmentation. INFORMS Journal on Computing, 34(2):1070–1085, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1120.

He:2022:DDD

[1675] Edward Yuhang He, Natashia Boland, George Nemhauser, and Martin Savelsbergh. Dynamic discretization discovery algorithms for timedependent shortest path problems. INFORMS Journal on Computing, 34 (2):1086–1114, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1084.

Badenbroek:2022:ACC

[1676] Riley Badenbroek and Etienne de Klerk. An analytic center cutting plane method to determine complete positivity of a matrix. INFORMS Journal on Computing, 34(2):1115–1125, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1108.

Zhang:2022:FOO

[1677] Dewei Zhang, Yin Liu, and Sam Davanloo Tajbakhsh. A first-order optimization algorithm for statistical learning with hierarchical sparsity structure. INFORMS Journal on Computing, 34(2):1126–1140, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1069.

Costa:2022:SCG

[1678] Luciano Costa, Claudio Contardo, Guy Desaulniers, and Julian Yarkony. Stabilized column generation via the dynamic separation of aggregated rows. INFORMS Journal on Computing, 34(2):1141–1156, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1094.

Jungwirth:2022:EBP

[1679] Alexander Jungwirth, Guy Desaulniers, Markus Frey, and Rainer Kolisch. Exact branch-price-and-cut for a hospital therapist scheduling problem with flexible service locations and time-dependent location capacity. INFORMS Journal on Computing, 34(2):1157–1175, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1119.

Nasrollahzadeh:2022:DPR

[1680] Amir Ali Nasrollahzadeh and Amin Khademi. Dynamic programming for response-adaptive dose-finding clinical trials. INFORMS Journal on Computing, 34(2):1176–1190, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1082.

Smeulders:2022:RKE

[1681] Bart Smeulders, Valentin Bartier, Yves Crama, and Frits C. R. Spieksma. Recourse in kidney exchange programs. INFORMS Journal on Computing, 34(2):1191–1206, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2021.1099.

Bai:2022:SSC

[1682] Miao Bai, Robert H. Storer, and Gregory L. Tonkay. Surgery sequencing coordination with recovery resource constraints. INFORMS Journal on Computing, 34(2):1207–1223, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1089.

Geng:2022:MAA

[1683] Na Geng and Xiaolan Xie. Managing advance admission requests for obstetric care. INFORMS Journal on Computing, 34(2):1224–1239, March/ April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1093.

Adibi:2022:SOM

[1684] Ali Adibi and Ehsan Salari. Scalable optimization methods for incorporating spatiotemporal fractionation into intensity-modulated radiotherapy planning. INFORMS Journal on Computing, 34(2):1240– 1256, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2021.1070.

Tang:2022:TRM

[1685] Shaojie Tang, Siyuan Liu, Xu Han, and Yu Qiao. Toward robust monitoring of malicious outbreaks. INFORMS Journal on Computing, 34(2): 1257–1271, March/April 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1077.

Zhang:2022:ESM

[1686] Kaike Zhang, Xueping Li, and Mingzhou Jin. Efficient solution methods for a general r-interdiction median problem with fortification. INFORMS Journal on Computing, 34(2):1272–1290, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1111.

Helmberg:2022:PES

[1687] Christoph Helmberg, Tobias Hofmann, and David Wenzel. Periodic event scheduling for automated production systems. INFORMS Journal on Computing, 34(2):1291–1304, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1101.

Anonymous:2022:EBb

[1688] Anonymous. Editorial Board. INFORMS Journal on Computing, 34(2): ??, March/April 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2022.eb.v3402.

Smith:2022:NEc

[1689] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 34(3):1305, May/June 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2022.1186.

Glover:2022:UCT

[1690] Fred W. Glover. Unforeseen consequences of “Tabu” choices — a retrospective. INFORMS Journal on Computing, 34(3):1306–1308, May/ June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1187.

Salemi:2022:SDB

[1691] Hosseinali Salemi and Austin Buchanan. Solving the distance-based critical node problem. INFORMS Journal on Computing, 34(3):1309–1326, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1136.

Chopra:2022:EGF

[1692] Sunil Chopra, Hyunwoo Park, and Sangho Shim. Extended graph formulation for the inequity aversion pricing problem on social networks. IN-FORMS Journal on Computing, 34(3):1327–1344, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1148.

Raghavan:2022:RIM

[1693] S. Raghavan and Rui Zhang. Rapid influence maximization on social networks: The positive influence dominating set problem. INFORMS Journal on Computing, 34(3):1345–1365, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1144.

Maass:2022:NOA

[1694] Kelsey Maass, Minsun Kim, and Aleksandr Aravkin. A nonconvex optimization approach to IMRT planning with dose–volume constraints. IN-FORMS Journal on Computing, 34(3):1366–1386, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1129.

Uiterkamp:2022:RCR

[1695] Martijn H. H. Schoot Uiterkamp, Marco E. T. Gerards, and Johann L. Hurink. On a reduction for a class of resource allocation problems. IN-FORMS Journal on Computing, 34(3):1387–1402, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1104.

Dang:2022:IPD

[1696] Chuangyin Dang, P. Jean-Jacques Herings, and Peixuan Li. An interiorpoint differentiable path-following method to compute stationary equilibria in stochastic games. INFORMS Journal on Computing, 34(3): 1403–1418, May/June 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1139.

Zhang:2022:LBB

[1697] Xiangyi Zhang, Lu Chen, Michel Gendreau, and André Langevin. Learning-based branch-and-price algorithms for the vehicle routing problem with time windows and two-dimensional loading constraints. IN-FORMS Journal on Computing, 34(3):1419–1436, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1110.

Happach:2022:GFA

[1698] Felix Happach, Lisa Hellerstein, and Thomas Lidbetter. A general framework for approximating min sum ordering problems. INFORMS Journal on Computing, 34(3):1437–1452, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1124.

Mahmoodian:2022:CSB

[1699] Vahid Mahmoodian, Iman Dayarian, Payman Ghasemi Saghand, Yu Zhang, and Hadi Charkhgard. A criterion space branch-and-cut algorithm for mixed integer bilinear maximum multiplicative programs. IN-FORMS Journal on Computing, 34(3):1453–1470, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1097.

Bodur:2022:IMI

[1700] Merve Bodur, Timothy C. Y. Chan, and Ian Yihang Zhu. Inverse mixed integer optimization: Polyhedral insights and trust region methods. IN-FORMS Journal on Computing, 34(3):1471–1488, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1138.

Bertsimas:2022:SAS

[1701] Dimitris Bertsimas and Ryan Cory-Wright. A scalable algorithm for sparse portfolio selection. INFORMS Journal on Computing, 34(3):1489– 1511, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2021.1127.

Pessoa:2022:EAS

[1702] Artur Alves Pessoa, Teobaldo Bulhões, Vitor Nesello, and Anand Subramanian. Exact approaches for single machine total weighted tardiness batch scheduling. INFORMS Journal on Computing, 34(3):1512–1530, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1133.

Zhang:2022:BLC

[1703] Yiling Zhang and Jin Dong. Building load control using distributionally robust chance-constrained programs with right-hand side uncertainty and the risk-adjustable variants. INFORMS Journal on Computing, 34 (3):1531–1547, May/June 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1152.

Becker:2022:GBC

[1704] Tristan Becker, Maximilian Schiffer, and Grit Walther. A general branchand-cut framework for rotating workforce scheduling. INFORMS Journal on Computing, 34(3):1548–1564, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1149.

Liu:2022:ULH

[1705] Siyuan Liu, Shaojie Tang, Jiangchuan Zheng, and Lionel M. Ni. Unsupervised learning for human mobility behaviors. INFORMS Journal on Computing, 34(3):1565–1586, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1098.

Xu:2022:EAM

[1706] Haifeng Xu, Rasha F. Kashef, Hans De Sterck, and Geoffrey Sanders. Efficient algebraic multigrid methods for multilevel overlapping coclustering of user-item relationships. INFORMS Journal on Computing, 34 (3):1587–1605, May/June 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1137.

Jena:2022:EDC

[1707] Sanjay Dominik Jena, Andrea Lodi, and Claudio Sole. On the estimation of discrete choice models to capture irrational customer behaviors. IN-FORMS Journal on Computing, 34(3):1606–1625, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1154.

Wang:2022:CRS

[1708] Tong Wang and Cynthia Rudin. Causal rule sets for identifying subgroups with enhanced treatment effects. INFORMS Journal on Computing, 34(3):1626–1643, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2021.1143.

Bi:2022:ISF

[1709] Xuan Bi, Gediminas Adomavicius, William Li, and Annie Qu. Improving sales forecasting accuracy: a tensor factorization approach with demand awareness. INFORMS Journal on Computing, 34(3):1644–1660, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1147.

Song:2022:UCR

[1710] Yicheng Song, Nachiketa Sahoo, Shuba Srinivasan, and Chrysanthos Dellarocas. Uncovering characteristic response paths of a population. IN-FORMS Journal on Computing, 34(3):1661–1680, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1121.

vanEekelen:2022:MDM

[1711] Wouter van Eekelen, Dick den Hertog, and Johan S. H. van Leeuwaarden. MAD dispersion measure makes extremal queue analysis simple. IN-FORMS Journal on Computing, 34(3):1681–1692, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1130.

Rastpour:2022:AQS

[1712] Amir Rastpour, Armann Ingolfsson, and Burhaneddin Sandıkçı. Algorithms for queueing systems with reneging and priorities modeled as quasi-birth–death processes. INFORMS Journal on Computing, 34(3): 1693–1710, May/June 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1141.

Eckman:2022:PBS

[1713] David J. Eckman and Shane G. Henderson. Posterior-based stopping rules for Bayesian ranking-and-selection procedures. INFORMS Journal on Computing, 34(3):1711–1728, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1132.

LEcuyer:2022:MCQ

[1714] Pierre L’Ecuyer, Florian Puchhammer, and Amal Ben Abdellah. Monte Carlo and quasi–Monte Carlo density estimation via conditioning. IN-FORMS Journal on Computing, 34(3):1729–1748, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1135.

Byeon:2022:BSD

[1715] Geunyeong Byeon and Pascal Van Hentenryck. Benders subproblem decomposition for bilevel problems with convex follower. INFORMS Journal on Computing, 34(3):1749–1767, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1128.

Cheramin:2022:CEA

[1716] Meysam Cheramin, Jianqiang Cheng, Ruiwei Jiang, and Kai Pan. Computationally efficient approximations for distributionally robust optimization under moment and Wasserstein ambiguity. INFORMS Journal on Computing, 34(3):1768–1794, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1123.

Lima:2022:RAS

[1717] Ricardo M. Lima, Antonio J. Conejo, Loïc Giraldi, Olivier Le Maître, Ibrahim Hoteit, and Omar M. Knio. Risk-averse stochastic programming vs. adaptive robust optimization: a virtual power plant application. IN-FORMS Journal on Computing, 34(3):1795–1818, May/June 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1157.

Pan:2022:ISO

[1718] Kai Pan and Yongpei Guan. Integrated stochastic optimal self-scheduling for two-settlement electricity markets. INFORMS Journal on Computing, 34(3):1819–1840, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2021.1150.

Anonymous:2022:EBc

[1719] Anonymous. Editorial Board. INFORMS Journal on Computing, 34 (3):??, May/June 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2022.eb.v3403.

Smith:2022:NEd

[1720] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 34(4):1841, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2022.1213.

Peng:2022:VVR

[1721] Lan Peng and Chase Murray. VeRoViz: a vehicle routing visualization toolkit. INFORMS Journal on Computing, 34(4):1842–1848, July/ August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1159.

Steever:2022:IBA

[1722] Zachary Steever, Chase Murray, Junsong Yuan, Mark Karwan, and Marco Lübbecke. An image-based approach to detecting structural similarity among mixed integer programs. INFORMS Journal on Computing, 34(4):1849–1870, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2021.1117.

Torrico:2022:DRO

[1723] Alfredo Torrico and Alejandro Toriello. Dynamic relaxations for online bipartite matching. INFORMS Journal on Computing, 34(4):1871– 1884, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2022.1168.

Biel:2022:ESP

[1724] Martin Biel and Mikael Johansson. Efficient stochastic programming in Julia. INFORMS Journal on Computing, 34(4):1885–1902, July/ August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1158.

Zhang:2022:NSR

[1725] Weiguo Zhang and Xiaolei He. A new scenario reduction method based on higher-order moments. INFORMS Journal on Computing, 34(4): 1903–1918, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2021.1155.

Huang:2022:NCD

[1726] Zhouchun Huang, Qipeng P. Zheng, and Andrew L. Liu. A nested cross decomposition algorithm for power system capacity expansion with multiscale uncertainties. INFORMS Journal on Computing, 34(4):1919– 1939, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2022.1177.

Tian:2022:IRL

[1727] Hu Tian, Xiaolong Zheng, Kang Zhao, Maggie Wenjing Liu, and Daniel Dajun Zeng. Inductive representation learning on dynamic stock co-movement graphs for stock predictions. INFORMS Journal on Computing, 34(4):1940–1957, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1172.

Chen:2022:DHP

[1728] Xi Chen, Yan Liu, and Cheng Zhang. Distinguishing homophily from peer influence through network representation learning. INFORMS Journal on Computing, 34(4):1958–1969, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1171.

Pereira:2022:ESA

[1729] Dilson Lucas Pereira, Abilio Lucena, Alexandre Salles da Cunha, and Luidi Simonetti. Exact solution algorithms for the chordless cycle problem. INFORMS Journal on Computing, 34(4):1970–1986, July/ August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1164.

Casazza:2022:EAM

[1730] Marco Casazza and Alberto Ceselli. Exact algorithms for maximum lifetime data-gathering tree in wireless sensor networks. INFORMS Journal on Computing, 34(4):1987–2002, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1175.

Mendes:2022:CED

[1731] Luis Henrique Pauleti Mendes, Fábio Luiz Usberti, and Celso Cavellucci. The capacitated and economic districting problem. INFORMS Journal on Computing, 34(4):2003–2016, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1180.

Ripsman:2022:RDA

[1732] Danielle A. Ripsman, Thomas G. Purdie, Timothy C. Y. Chan, and Houra Mahmoudzadeh. Robust direct aperture optimization for radiation therapy treatment planning. INFORMS Journal on Computing, 34 (4):2017–2038, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2022.1167.

Liu:2022:MLE

[1733] Zeyu Liu, Anahita Khojandi, Xueping Li, Akram Mohammed, Robert L. Davis, and Rishikesan Kamaleswaran. A machine learning–enabled partially observable Markov decision process framework for early sepsis prediction. INFORMS Journal on Computing, 34(4):2039–2057, July/ August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1176.

Chow:2022:TOD

[1734] Vincent Tsz Fai Chow, Zheng Cui, and Daniel Zhuoyu Long. Targetoriented distributionally robust optimization and its applications to surgery allocation. INFORMS Journal on Computing, 34(4):2058–2072, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2021.1145.

Chen:2022:VBT

[1735] Yifu Chen and Christos T. Maravelias. Variable bound tightening and valid constraints for multiperiod blending. INFORMS Journal on Computing, 34(4):2073–2090, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1140.

Selvi:2022:CMA

[1736] Aras Selvi, Aharon Ben-Tal, Ruud Brekelmans, and Dick den Hertog. Convex maximization via adjustable robust optimization. INFORMS Journal on Computing, 34(4):2091–2105, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1134.

Kuchlbauer:2022:ABM

[1737] Martina Kuchlbauer, Frauke Liers, and Michael Stingl. Adaptive bundle methods for nonlinear robust optimization. INFORMS Journal on Computing, 34(4):2106–2124, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1122.

Graham:2022:RDP

[1738] Naomi Graham, Hao Hu, Jiyoung Im, Xinxin Li, and Henry Wolkowicz. A restricted dual Peaceman–Rachford splitting method for a strengthened DNN relaxation for QAP. INFORMS Journal on Computing, 34(4): 2125–2143, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2022.1161.

Piccialli:2022:SSE

[1739] Veronica Piccialli, Antonio M. Sudoso, and Angelika Wiegele. SOS-SDP: an exact solver for minimum sum-of-squares clustering. INFORMS Journal on Computing, 34(4):2144–2162, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1166.

Daues:2022:COP

[1740] Endric Daues and Ulf Friedrich. Computing optimized path integrals for knapsack feasibility. INFORMS Journal on Computing, 34(4):2163– 2176, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2021.1142.

Martinez:2022:LBB

[1741] Karim Pérez Martínez, Yossiri Adulyasak, and Raf Jans. Logic-based Benders decomposition for integrated process configuration and production planning problems. INFORMS Journal on Computing, 34(4):2177– 2191, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2021.1079.

Wolck:2022:BPA

[1742] Martin Wölck and Stephan Meisel. Branch-and-price approaches for real-time vehicle routing with picking, loading, and soft time Windows. INFORMS Journal on Computing, 34(4):2192–2211, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2021.1151.

Arslan:2022:MSM

[1743] Ayşe N. Arslan, Michael Poss, and Marco Silva. Min-sup-min robust combinatorial optimization with few recourse solutions. INFORMS Journal on Computing, 34(4):2212–2228, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2021.1156.

Bertsimas:2022:OMI

[1744] Dimitris Bertsimas and Bartolomeo Stellato. Online mixed-integer optimization in milliseconds. INFORMS Journal on Computing, 34(4): 2229–2248, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2022.1181.

Letelier:2022:PPT

[1745] Orlando Rivera Letelier, François Clautiaux, and Ruslan Sadykov. Bin packing problem with time lags. INFORMS Journal on Computing, 34 (4):2249–2270, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2022.1165.

Castro:2022:DDD

[1746] Margarita P. Castro, Andre A. Cire, and J. Christopher Beck. Decision diagrams for discrete optimization: a survey of recent advances. INFORMS Journal on Computing, 34(4):2271–2295, July/ August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1170.

Johnson:2022:SLB

[1747] Emma S. Johnson and Santanu Subhas Dey. A scalable lower bound for the worst-case relay attack problem on the transmission grid. INFORMS Journal on Computing, 34(4):2296–2312, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1178.

Tahir:2022:ICG

[1748] Adil Tahir, Guy Desaulniers, and Issmail El Hallaoui. Integral column generation for set partitioning problems with side constraints. INFORMS Journal on Computing, 34(4):2313–2331, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1174.

Chen:2022:GLC

[1749] Rui Chen and James Luedtke. On generating Lagrangian cuts for twostage stochastic integer programs. INFORMS Journal on Computing, 34 (4):2332–2349, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2022.1185.

Hu:2022:RSL

[1750] Zhaolin Hu and L. Jeff Hong. Robust simulation with likelihood-ratio constrained input uncertainty. INFORMS Journal on Computing, 34(4): 2350–2367, July/August 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2022.1169.

Morgan:2022:RCI

[1751] Lucy E. Morgan, Luke Rhodes-Leader, and Russell R. Barton. Reducing and calibrating for input model bias in computer simulation. INFORMS Journal on Computing, 34(4):2368–2382, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1183.

Anonymous:2022:EBd

[1752] Anonymous. Editorial Board. INFORMS Journal on Computing, 34(4): ??, July/August 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2022.eb.v3404.

Smith:2022:NEe

[1753] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 34(5):2383, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/abs/10.1287/ijoc.2022.1230.

Scheinberg:2022:FDG

[1754] Katya Scheinberg. Finite difference gradient approximation: To randomize or not? INFORMS Journal on Computing, 34(5):2384–2388, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2022.1218.

Baxter:2022:HMR

[1755] Arden Baxter, Pinar Keskinocak, and Mohit Singh. Heterogeneous multiresource allocation with subset demand requests. INFORMS Journal on Computing, 34(5):2389–2399, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1204.

Bertsimas:2022:SCP

[1756] Dimitris Bertsimas and Michael Lingzhi Li. Stochastic cutting planes for data-driven optimization. INFORMS Journal on Computing, 34(5): 2400–2409, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2022.1205.

Zhen:2022:DBO

[1757] Jianzhe Zhen, Ahmadreza Marandi, Danique de Moor, Dick den Hertog, and Lieven Vandenberghe. Disjoint bilinear optimization: a two-stage robust optimization perspective. INFORMS Journal on Computing, 34 (5):2410–2427, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2022.1163.

Elci:2022:SPS

[1758] Özgün Elçi and John Hooker. Stochastic planning and scheduling with logic-based Benders decomposition. INFORMS Journal on Computing, 34(5):2428–2442, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1184.

Meng:2022:FGJ

[1759] Qingxin Meng, Keli Xiao, Dazhong Shen, Hengshu Zhu, and Hui Xiong. Fine-grained job salary benchmarking with a nonparametric Dirichlet process–based latent factor model. INFORMS Journal on Computing, 34 (5):2443–2463, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2022.1182.

Liu:2022:ENF

[1760] Menghan Liu, Erik Poppleton, Giulia Pedrielli, Petr Šulc, and Dimitri P. Bertsekas. ExpertRNA: a new framework for RNA secondary structure prediction. INFORMS Journal on Computing, 34(5):2464–2484, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2022.1188.

Keshanian:2022:FSN

[1761] Kimia Keshanian, Daniel Zantedeschi, and Kaushik Dutta. Features selection as a Nash-bargaining solution: Applications in online advertising and information systems. INFORMS Journal on Computing, 34(5):2485– 2501, September/October 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
fpi/10.1287/ijoc.2022.1190.

Gmys:2022:ESH

[1762] Jan Gmys. Exactly solving hard permutation flowshop scheduling problems on peta-scale GPU-accelerated supercomputers. INFORMS Journal on Computing, 34(5):2502–2522, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1193.

Wu:2022:IDS

[1763] Wei Wu, Manuel Iori, Silvano Martello, and Mutsunori Yagiura. An iterated dual substitution approach for binary integer programming problems under the min–max regret criterion. INFORMS Journal on Computing, 34(5):2523–2539, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1189.

Haugh:2022:PLP

[1764] Martin B. Haugh and Chun Wang. Play like the pros? Solving the game of darts as a dynamic zero-sum game. INFORMS Journal on Computing, 34(5):2540–2551, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1197.

Cappart:2022:IVO

[1765] Quentin Cappart, David Bergman, Louis-Martin Rousseau, Isabeau Prémont-Schwarz, and Augustin Parjadis. Improving variable orderings of approximate decision diagrams using reinforcement learning. IN-FORMS Journal on Computing, 34(5):2552–2570, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1194.

Gudapati:2022:NDS

[1766] Naga V. C. Gudapati, Enrico Malaguti, and Michele Monaci. Network design with service requirements: Scaling-up the size of solvable problems. INFORMS Journal on Computing, 34(5):2571–2582, September/ October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1200.

Moghaddass:2022:OFD

[1767] Ramin Moghaddass and Yongtao Guan. Optimal frameworks for detecting anomalies in sensor-intensive heterogeneous networks. INFORMS Journal on Computing, 34(5):2583–2610, September/October 2022. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1192.

Besancon:2022:FJH

[1768] Mathieu Besançon, Alejandro Carderera, and Sebastian Pokutta. FrankWolfe.jl: a high-performance and flexible toolbox for Frank– Wolfe algorithms and conditional gradients. INFORMS Journal on Computing, 34(5):2611–2620, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1191.

Ahmadi-Javid:2022:CQF

[1769] Amir Ahmadi-Javid and Pooya Hoseinpour. Convexification of queueing formulas by mixed-integer second-order cone programming: an application to a discrete location problem with congestion. INFORMS Journal on Computing, 34(5):2621–2633, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2021.1125.

Taninmis:2022:BCA

[1770] Kübra Tanınmış and Markus Sinnl. A branch-and-cut algorithm for submodular interdiction games. INFORMS Journal on Computing, 34 (5):2634–2657, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2022.1196.

Bui:2022:CFN

[1771] Quang Minh Bui, Bernard Gendron, and Margarida Carvalho. A catalog of formulations for the network pricing problem. INFORMS Journal on Computing, 34(5):2658–2674, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/fpi/10.1287/ijoc.2022.1198.

Zhang:2022:CAL

[1772] Zhao Zhang, Wei Liang, Hongmin W. Du, and Siwen Liu. Constant approximation for the lifetime scheduling problem of p-percent coverage. INFORMS Journal on Computing, 34(5):2675–2685, September/ October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1201.

Coey:2022:SNC

[1773] Chris Coey, Lea Kapelevich, and Juan Pablo Vielma. Solving natural conic formulations with Hypatia.jl. INFORMS Journal on Computing, 34(5):2686–2699, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1202.

Daryalal:2022:SRL

[1774] Maryam Daryalal and Merve Bodur. Stochastic RWA and lightpath rerouting in WDM networks. INFORMS Journal on Computing, 34(5): 2700–2719, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2022.1179.

Gupta:2022:DAS

[1775] Rishabh Gupta and Qi Zhang. Decomposition and adaptive sampling for data-driven inverse linear optimization. INFORMS Journal on Computing, 34(5):2720–2735, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1162.

Porumbel:2022:PCP

[1776] Daniel Porumbel. Projective cutting-planes for robust linear programming and cutting stock problems. INFORMS Journal on Computing, 34 (5):2736–2753, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2022.1160.

Samorani:2022:SPD

[1777] Michele Samorani, Ram Bala, Rohit Jacob, and Shuhan He. A software package and data set for the personal protective equipment matching problem during COVID-19. INFORMS Journal on Computing, 34(5): 2754–2761, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/fpi/10.1287/ijoc.2022.1203.

Zheng:2022:RLH

[1778] Zemin Zheng, Jie Zhang, and Yang Li. L_{0}-regularized learning for highdimensional additive hazards regression. INFORMS Journal on Computing, 34(5):2762–2775, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1208.

Liu:2022:RSF

[1779] Tianqi Liu, Francisco Saldanha da Gama, Shuming Wang, and Yuchen Mao. Robust stochastic facility location: Sensitivity analysis and exact solution. INFORMS Journal on Computing, 34(5):2776–2803, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2022.1206.

Seo:2022:CBC

[1780] Kiho Seo, Seulgi Joung, Chungmok Lee, and Sungsoo Park. A closest Benders cut selection scheme for accelerating the Benders decomposition algorithm. INFORMS Journal on Computing, 34(5):2804–2827, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/fpi/
10.1287/ijoc.2022.1207.

Khorramfar:2022:MPT

[1781] Rahman Khorramfar, Osman Y. Özaltın, Karl G. Kempf, and Reha Uzsoy. Managing product transitions: a bilevel programming approach. INFORMS Journal on Computing, 34(5):2828–2844, September/ October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/fpi/10.1287/
ijoc.2022.1210.

Homem-de-Mello:2022:SOA

[1782] Tito Homem de Mello, Qingxia Kong, and Rodrigo Godoy-Barba. A simulation optimization approach for the appointment scheduling problem with decision-dependent uncertainties. INFORMS Journal on Computing, 34(5):2845–2865, September/October 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/fpi/10.1287/ijoc.2022.1212.

Anonymous:2022:EBe

[1783] Anonymous. Editorial Board. INFORMS Journal on Computing, 34(5): ??, September/October 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2022.eb.v3405.

Smith:2022:NEf

[1784] Alice Smith. Note from the Editor. INFORMS Journal on Computing, 34(6):2867, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.1247.

Hicks:2022:CTC

[1785] Illya V. Hicks, Boris Brimkov, Louis Deaett, Ruth Haas, Derek Mikesell, David Roberson, and Logan Smith. Computational and theoretical challenges for computing the minimum rank of a graph. INFORMS Journal on Computing, 34(6):2868–2872, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1219.

Vayanos:2022:RRO

[1786] Phebe Vayanos, Qing Jin, and George Elissaios. ROC++: Robust optimization in C++. INFORMS Journal on Computing, 34(6):2873–2888, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.1209.

Hu:2022:SAM

[1787] Jiaqiao Hu, Yijie Peng, Gongbo Zhang, and Qi Zhang. A stochastic approximation method for simulation-based quantile optimization. IN-FORMS Journal on Computing, 34(6):2889–2907, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1214.

Liu:2022:BDT

[1788] Yanchao Liu. bsnsing: A decision tree induction method based on recursive optimal Boolean rule composition. INFORMS Journal on Computing, 34(6):2908–2929, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1225.

Hong:2022:SLS

[1789] L. Jeff Hong, Guangxin Jiang, and Ying Zhong. Solving large-scale fixedbudget ranking and selection problems. INFORMS Journal on Computing, 34(6):2930–2949, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1221.

Rhuggenaath:2022:SRP

[1790] Jason Rhuggenaath, Alp Akcay, Yingqian Zhang, and Uzay Kaymak. Setting reserve prices in second-price auctions with unobserved bids. INFORMS Journal on Computing, 34(6):2950–2967, November/ December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1199.

Costa:2022:AMC

[1791] Carina Moreira Costa, Dennis Kreber, and Martin Schmidt. An alternating method for cardinality-constrained optimization: a computational study for the best subset selection and sparse portfolio problems. INFORMS Journal on Computing, 34(6):2968–2988, November/ December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1211.

Zhang:2022:SSO

[1792] Liwei Zhang, Yule Zhang, Jia Wu, and Xiantao Xiao. Solving stochastic optimization with expectation constraints efficiently by a stochastic augmented Lagrangian-type algorithm. INFORMS Journal on Computing, 34(6):2989–3006, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1228.

Andres-Thio:2022:BSM

[1793] Nicolau Andrés-Thió, Mario Andrés Muñoz, and Kate Smith-Miles. Bifidelity surrogate modelling: Showcasing the need for new test instances. INFORMS Journal on Computing, 34(6):3007–3022, November/ December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1217.

Islam:2022:CFS

[1794] Md Saiful Islam, Md Sarowar Morshed, and Md. Noor-E-Alam. A computational framework for solving nonlinear binary optimization problems in robust causal inference. INFORMS Journal on Computing, 34(6):3023– 3041, November/December 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.1226.

Hill:2022:OSR

[1795] Alessandro Hill, Andrea J. Brickey, Italo Cipriano, Marcos Goycoolea, and Alexandra Newman. Optimization strategies for resourceconstrained project scheduling problems in underground mining. IN-FORMS Journal on Computing, 34(6):3042–3058, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1222.

Song:2022:PMS

[1796] Guopeng Song and Roel Leus. Parallel machine scheduling under uncertainty: Models and exact algorithms. INFORMS Journal on Computing, 34(6):3059–3079, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1229.

Zhou:2022:FFS

[1797] Yuwei Zhou, Sigrún Andradóttir, Seong-Hee Kim, and Chuljin Park. Finding feasible systems for subjective constraints using recycled observations. INFORMS Journal on Computing, 34(6):3080–3095, November/ December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1227.

Lin:2022:TEE

[1798] Shao-Bo Lin, Shaojie Tang, Yao Wang, and Di Wang. Toward efficient ensemble learning with structure constraints: Convergent algorithms and applications. INFORMS Journal on Computing, 34(6):3096–3116, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.1224.

Santis:2022:PBB

[1799] Marianna De Santis, Sven de Vries, Martin Schmidt, and Lukas Winkel. A penalty branch-and-bound method for mixed binary linear complementarity problems. INFORMS Journal on Computing, 34(6):3117–3133, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.1216.

Lalonde:2022:BPA

[1800] Olivier Lalonde, Jean-François Côté, and Bernard Gendron. A branchand-price algorithm for the multiple knapsack problem. INFORMS Journal on Computing, 34(6):3134–3150, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1223.

Thevenin:2022:SDD

[1801] Simon Thevenin, Yossiri Adulyasak, and Jean-François Cordeau. Stochastic dual dynamic programming for multiechelon lot sizing with component substitution. INFORMS Journal on Computing, 34(6):3151– 3169, November/December 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.1215.

Hazen:2022:AMC

[1802] Gordon B. Hazen. Augmenting Markov cohort analysis to compute (co)Variances: Implications for strength of cost-effectiveness. IN-FORMS Journal on Computing, 34(6):3170–3180, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1234.

Lu:2022:FTL

[1803] Yajun Lu, Hosseinali Salemi, Balabhaskar Balasundaram, and Austin Buchanan. On fault-tolerant low-diameter clusters in graphs. IN-FORMS Journal on Computing, 34(6):3181–3199, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1231.

Ben-Tal:2022:AMC

[1804] Aharon Ben-Tal and Ernst Roos. An algorithm for maximizing a convex function based on its minimum. INFORMS Journal on Computing, 34(6): 3200–3214, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.1238.

Yu:2022:RTO

[1805] Qinxiao Yu, Chun Cheng, and Ning Zhu. Robust team orienteering problem with decreasing profits. INFORMS Journal on Computing, 34(6): 3215–3233, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.1240.

Yin:2022:IBR

[1806] Jiateng Yin, Lixing Yang, Andrea D’Ariano, Tao Tang, and Ziyou Gao. Integrated backup rolling stock allocation and timetable rescheduling with uncertain time-variant passenger demand under disruptive events. INFORMS Journal on Computing, 34(6):3234–3258, November/ December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1233.

Ragodos:2022:DRL

[1807] Ronilo Ragodos and Tong Wang. Disjunctive rule lists. INFORMS Journal on Computing, 34(6):3259–3276, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1242.

Tang:2022:RAS

[1808] Shaojie Tang. Robust adaptive submodular maximization. IN-FORMS Journal on Computing, 34(6):3277–3291, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1239.

Lera-Romero:2022:DPT

[1809] Gonzalo Lera-Romero, Juan José Miranda Bront, and Francisco J. Soulignac. Dynamic programming for the time-dependent traveling salesman problem with time windows. INFORMS Journal on Computing, 34(6):3292–3308, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1236.

Pan:2022:PSF

[1810] Kai Pan, Ming Zhao, Chung-Lun Li, and Feng Qiu. A polyhedral study on fuel-constrained unit commitment. INFORMS Journal on Computing, 34(6):3309–3324, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1235.

Wiratchotisatian:2022:SRM

[1811] Pitchaya Wiratchotisatian, Hoda Atef Yekta, and Andrew C. Trapp. Stability representations of many-to-one matching problems: an integer optimization approach. INFORMS Journal on Computing, 34(6):3325– 3343, November/December 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.1237.

Anonymous:2022:AR

[1812] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 34(6):3344–3350, November/December 2022. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1244.

Anonymous:2022:EB

[1813] Anonymous. Editorial Board. INFORMS Journal on Computing, 34(6): ??, November/December 2022. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2022.eb.v3406.

Smith:2023:NEa

[1814] Alice Smith. Note from the Editor. INFORMS Journal on Computing, 35(1):1, January/February 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.1262.

Contardo:2023:CPB

[1815] Claudio Contardo, Andrea Lodi, and Andrea Tramontani. Cutting planes from the branch-and-bound tree: Challenges and opportunities. IN-FORMS Journal on Computing, 35(1):2–4, January/February 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1248.

Lalor:2023:PIS

[1816] John Patrick Lalor and Pedro Rodriguez. py-irt: a scalable item response theory library for Python. INFORMS Journal on Computing, 35 (1):5–13, January/February 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.1250.

Liu:2023:BCP

[1817] Yiming Liu, Yang Yu, Yu Zhang, Roberto Baldacci, Jiafu Tang, Xinggang Luo, and Wei Sun. Branch-cut-and-price for the time-dependent green vehicle routing problem with time windows. INFORMS Journal on Computing, 35(1):14–30, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1195.

Li:2023:NEA

[1818] Jiliu Li, Zhixing Luo, Roberto Baldacci, Hu Qin, and Zhou Xu. A new exact algorithm for single-commodity vehicle routing with split pickups and deliveries. INFORMS Journal on Computing, 35(1):31–49, January/ February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1249.

Hessler:2023:PDB

[1819] Katrin Heßler and Stefan Irnich. Partial dominance in branch-price-and-cut for the basic multicompartment vehicle-routing problem. INFORMS Journal on Computing, 35(1):50–65, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1255.

Zheng:2023:POD

[1820] Hua Zheng, Wei Xie, Ilya O. Ryzhov, and Dongming Xie. Policy optimization in dynamic Bayesian network hybrid models of biomanufacturing processes. INFORMS Journal on Computing, 35(1):66–82, January/ February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1232.

Ansari:2023:RMC

[1821] Mehdi Ansari, Juan S. Borrero, and Leonardo Lozano. Robust minimumcost flow problems under multiple ripple effect disruptions. INFORMS Journal on Computing, 35(1):83–103, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1243.

Zhang:2023:SAB

[1822] Yanhang Zhang, Junxian Zhu, Jin Zhu, and Xueqin Wang. A splicing approach to best subset of groups selection. INFORMS Journal on Computing, 35(1):104–119, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1241.

Moskal:2023:UAV

[1823] Michael D. Moskal II, Erdi Dasdemir, and Rajan Batta. Unmanned aerial vehicle information collection missions with uncertain characteristics. INFORMS Journal on Computing, 35(1):120–137, January/ February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1245.

Han:2023:CES

[1824] Xiao Han, Leye Wang, and Weiguo Fan. Cost-effective social media influencer marketing. INFORMS Journal on Computing, 35(1):138–157, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.1246.

Chen:2023:MLM

[1825] Zhen-Yu Chen, Zhi-Ping Fan, and Minghe Sun. Machine learning methods for data-driven demand estimation and assortment planning considering cross-selling and substitutions. INFORMS Journal on Computing, 35(1):158–177, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.1251.

Zhao:2023:RSU

[1826] Ming Zhao, Nickolas Freeman, and Kai Pan. Robust sourcing under multilevel supply risks: Analysis of random yield and capacity. INFORMS Journal on Computing, 35(1):178–195, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1254.

Wang:2023:LSI

[1827] Tan Wang and L. Jeff Hong. Large-scale inventory optimization: a recurrent neural networks-inspired simulation approach. INFORMS Journal on Computing, 35(1):196–215, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1253.

Sadana:2023:VRS

[1828] Utsav Sadana and Erick Delage. The value of randomized strategies in distributionally robust risk-averse network interdiction problems. INFORMS Journal on Computing, 35(1):216–232, January/ February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1257.

Chopra:2023:PPS

[1829] Sunil Chopra, Feng Qiu, and Sangho Shim. Parallel power system restoration. INFORMS Journal on Computing, 35(1):233–247, January/ February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1258.

Seker:2023:MAS

[1830] Oylum Şeker, Mucahit Cevik, Merve Bodur, Young Lee, and Mark Ruschin. A multiobjective approach for sector duration optimization in stereotactic radiosurgery treatment planning. INFORMS Journal on Computing, 35(1):248–264, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1252.

Anonymous:2023:EBa

[1831] Anonymous. Editorial Board. INFORMS Journal on Computing, 35 (1):??, January/February 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2023.eb.v3501.

Smith:2023:NEb

[1832] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 35(2):265, March/April 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.1278.

Parekh:2023:SBO

[1833] Ojas Parekh. Synergies between operations research and quantum information science. INFORMS Journal on Computing, 35(2):266–273, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2023.1268.

Pelegrin:2023:ACR

[1834] Mercedes Pelegrín and Martina Cerulli. Aircraft conflict resolution: a benchmark generator. INFORMS Journal on Computing, 35(2):274– 285, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.1265.

Eshghi:2023:NYS

[1835] Ashkan Eshghi, Ram D. Gopal, Hooman Hidaji, and Raymond A. Patterson. Now you see it, now you don’t: Obfuscation of online thirdparty information sharing. INFORMS Journal on Computing, 35(2): 286–303, March/April 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.1266.

Bergman:2023:OEM

[1836] David Bergman, Carlos Cardonha, Jason Imbrogno, and Leonardo Lozano. Optimizing the expected maximum of two linear functions defined on a multivariate Gaussian distribution. INFORMS Journal on Computing, 35(2):304–317, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.1259.

Wang:2023:RVA

[1837] Wenyu Wang and Christine A. Shoemaker. Reference vector assisted candidate search with aggregated surrogate for computationally expensive many objective optimization problems. INFORMS Journal on Computing, 35(2):318–334, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.1260.

Henrion:2023:RSP

[1838] Didier Henrion, Felix Kirschner, Etienne De Klerk, Milan Korda, Jean-Bernard Lasserre, and Victor Magron. Revisiting semidefinite programming approaches to options pricing: Complexity and computational perspectives. INFORMS Journal on Computing, 35(2):335–349, March/ April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1220.

Eckman:2023:DTE

[1839] David J. Eckman, Shane G. Henderson, and Sara Shashaani. Diagnostic tools for evaluating and comparing simulation-optimization algorithms. INFORMS Journal on Computing, 35(2):350–367, March/ April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1261.

Chen:2023:CSC

[1840] Zhongzhu Chen, Marcia Fampa, and Jon Lee. On computing with some convex relaxations for the maximum-entropy sampling problem. IN-FORMS Journal on Computing, 35(2):368–385, March/April 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1264.

Li:2023:CAS

[1841] Cheng Li, Siyang Gao, and Jianzhong Du. Convergence analysis of stochastic kriging-assisted simulation with random covariates. IN-FORMS Journal on Computing, 35(2):386–402, March/April 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.1263.

Valeva:2023:ABA

[1842] Silviya Valeva, Guodong Pang, Andrew J. Schaefer, and Gilles Clermont. Acuity-based allocation of ICU-downstream beds with flexible staffing. INFORMS Journal on Computing, 35(2):403–422, March/ April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1267.

Hoogendoorn:2023:IIS

[1843] Y. N. Hoogendoorn and R. Spliet. An improved integer L-shaped method for the vehicle routing problem with stochastic demands. INFORMS Journal on Computing, 35(2):423–439, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.1271.

Ray:2023:LMM

[1844] Arindam Ray, Wolfgang Jank, Kaushik Dutta, and Matthew Mullarkey. An LSTM^{+} model for managing epidemics: Using population mobility and vulnerability for forecasting COVID-19 hospital admissions. IN-FORMS Journal on Computing, 35(2):440–457, March/April 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.1269.

Aigner:2023:SAO

[1845] Kevin-Martin Aigner, Robert Burlacu, Frauke Liers, and Alexander Martin. Solving AC optimal power flow with discrete decisions to global optimality. INFORMS Journal on Computing, 35(2):458–474, March/ April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.1270.

Barkel:2023:AFC

[1846] Mathijs Barkel and Maxence Delorme. Arcflow formulations and constraint generation frameworks for the two bar charts packing problem. INFORMS Journal on Computing, 35(2):475–494, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.1256.

Eckman:2023:STS

[1847] David J. Eckman, Shane G. Henderson, and Sara Shashaani. SimOpt: a testbed for simulation-optimization experiments. INFORMS Journal on Computing, 35(2):495–508, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.1273.

Rodriguez:2023:SPN

[1848] Jose S. Rodriguez, Robert B. Parker, Carl D. Laird, Bethany L. Nicholson, John D. Siirola, and Michael L. Bynum. Scalable parallel nonlinear optimization with PyNumero and Parapint. INFORMS Journal on Computing, 35(2):509–517, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.1272.

Anonymous:2023:EBb

[1849] Anonymous. Editorial Board. INFORMS Journal on Computing, 35(2): ??, March/April 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2023.eb.v3502.

Smith:2023:NEc

[1850] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 35(3):519–520, May/June 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.1299000.

Coffrin:2023:SII

[1851] Carleton Coffrin, Elisabeth Lobe, Giacomo Nannicini, and Ojas Parekh. Special issue of INFORMS Journal on Computing — quantum computing and operations research. INFORMS Journal on Computing, 35(3):521– 522, May/June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2023.cfp.v35.n3.

Bichler:2023:LEA

[1852] Martin Bichler, Nils Kohring, and Stefan Heidekrüger. Learning equilibria in asymmetric auction games. INFORMS Journal on Computing, 35(3):523–542, May/June 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.1281.

Santini:2023:DSV

[1853] Alberto Santini, Michael Schneider, Thibaut Vidal, and Daniele Vigo. Decomposition strategies for vehicle routing heuristics. INFORMS Journal on Computing, 35(3):543–559, May/June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.1288.

Jalota:2023:ORP

[1854] Devansh Jalota, Dario Paccagnan, Maximilian Schiffer, and Marco Pavone. Online routing over parallel networks: Deterministic limits and data-driven enhancements. INFORMS Journal on Computing, 35 (3):560–577, May/June 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.1275.

Bougeret:2023:OPG

[1855] Marin Bougeret, Jérémy Omer, and Michael Poss. Optimization problems in graphs with locational uncertainty. INFORMS Journal on Computing, 35(3):578–592, May/June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.1276.

Daryalal:2023:NMP

[1856] Maryam Daryalal, Hamed Pouya, and Marc Antoine DeSantis. Network migration problem: a hybrid logic-based Benders decomposition approach. INFORMS Journal on Computing, 35(3):593–613, May/ June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.1280.

Ran:2023:IUC

[1857] Yaxuan Ran, Jiani Liu, and Yishi Zhang. Integrating users’ contextual engagements with their general preferences: An interpretable followee recommendation method. INFORMS Journal on Computing, 35(3):614– 632, May/June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2023.1284.

Xu:2023:GBS

[1858] Jingxu Xu and Zeyu Zheng. Gradient-based simulation optimization algorithms via multi-resolution system approximations. INFORMS Journal on Computing, 35(3):633–651, May/June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.1279.

Wang:2023:OET

[1859] Keliang Wang, Leonardo Lozano, Carlos Cardonha, and David Bergman. Optimizing over an ensemble of trained neural networks. INFORMS Journal on Computing, 35(3):652–674, May/June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.1285.

Wang:2023:BSA

[1860] Guanshen Wang, Yichen Cheng, Yusen Xia, Qiang Ling, and Xinlei Wang. A Bayesian semisupervised approach to keyword extraction with only positive and unlabeled data. INFORMS Journal on Computing, 35 (3):675–691, May/June 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.1283.

Kraul:2023:MLS

[1861] Sebastian Kraul, Markus Seizinger, and Jens O. Brunner. Machine learning-supported prediction of dual variables for the cutting stock problem with an application in stabilized column generation. IN-FORMS Journal on Computing, 35(3):692–709, May/June 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.1277.

Anonymous:2023:EBc

[1862] Anonymous. Editorial Board. INFORMS Journal on Computing, 35 (3):??, May/June 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2023.eb.v3503.

Smith:2023:NEd

[1863] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 35(4):711–712, July/August 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.ed.v35.n4.

Anonymous:2023:CPI

[1864] Anonymous. Call for papers — INFORMS Journal on Computing: Special issue on responsible AI and data science for social good. INFORMS Journal on Computing, 35(4):713–716, July/August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.cfp.v35.n4.

Chen:2023:RPO

[1865] Zhi Chen and Peng Xiong. RSOME in Python: An open-source package for robust stochastic optimization made easy. INFORMS Journal on Computing, 35(4):717–724, July/August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.1291.

Lai:2023:PBT

[1866] Xiangjing Lai, Jin-Kao Hao, Renbin Xiao, and Fred Glover. Perturbationbased thresholding search for packing equal circles and spheres. IN-FORMS Journal on Computing, 35(4):725–746, July/August 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.1290.

Zhao:2023:ELA

[1867] Hongke Zhao, Chuang Zhao, Xi Zhang, Nanlin Liu, Hengshu Zhu, Qi Liu, and Hui Xiong. An ensemble learning approach with gradient resampling for class-imbalance problems. INFORMS Journal on Computing, 35(4): 747–763, July/August 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.1274.

Liu:2023:MCD

[1868] Jiapeng Liu, Miłosz Kadziński, and Xiuwu Liao. Modeling contingent decision behavior: A Bayesian nonparametric preference-learning approach. INFORMS Journal on Computing, 35(4):764–785, July/ August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.1292.

Wang:2023:TDP

[1869] Xiao-Jun Wang, Tao Liu, and Weiguo Fan. TGVx: Dynamic personalized POI deep recommendation model. INFORMS Journal on Computing, 35(4):786–796, July/August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.1286.

Wu:2023:CNR

[1870] Can Wu, Ying Cui, Donghui Li, and Defeng Sun. Convex and nonconvex risk-based linear regression at scale. INFORMS Journal on Computing, 35(4):797–816, July/August 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.1282.

Naderi:2023:MIP

[1871] Bahman Naderi, Rubén Ruiz, and Vahid Roshanaei. Mixed-integer programming vs. constraint programming for shop scheduling problems:

New results and outlook. INFORMS Journal on Computing, 35(4):817– 843, July/August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2023.1287.

Pham:2023:PBA

[1872] Tu San Pham, Antoine Legrain, Patrick De Causmaecker, and Louis-Martin Rousseau. A prediction-based approach for online dynamic appointment scheduling: A case study in radiotherapy treatment. IN-FORMS Journal on Computing, 35(4):844–868, July/August 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.1289.

Severitt:2023:ESD

[1873] Marvin Severitt and Paul Manns. Efficient solution of discrete subproblems arising in integer optimal control with total variation regularization. INFORMS Journal on Computing, 35(4):869–885, July/ August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.1294.

Zhang:2023:OSR

[1874] Wei Zhang, Kai Wang, Alexandre Jacquillat, and Shuaian Wang. Optimized scenario reduction: Solving large-scale stochastic programs with quality guarantees. INFORMS Journal on Computing, 35(4):886–908, July/August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2023.1295.

Anonymous:2023:EBd

[1875] Anonymous. Editorial Board. INFORMS Journal on Computing, 35(4): ??, July/August 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2023.eb.v3504.

Smith:2023:NEe

[1876] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 35(5):909, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.ed.v35.n5.

Bochkarev:2023:ANO

[1877] Alexey A. Bochkarev and J. Cole Smith. On aligning non-order-associated binary decision diagrams. INFORMS Journal on Computing, 35(5):910–928, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.1293.

Baxter:2023:HMR

[1878] Arden Baxter, Pinar Keskinocak, and Mohit Singh. Heterogeneous multiresource planning and allocation under stochastic demand. INFORMS Journal on Computing, 35(5):929–951, September/October 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.1298.

Bertsimas:2023:IMC

[1879] Dimitris Bertsimas and Michael Lingzhi Li. Interpretable matrix completion: a discrete optimization approach. INFORMS Journal on Computing, 35(5):952–965, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0022.

Blanco:2023:HLP

[1880] Víctor Blanco, Elena Fernández, and Yolanda Hinojosa. Hub location with protection under interhub link failures. INFORMS Journal on Computing, 35(5):966–985, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.1296.

Xiao:2023:BBA

[1881] Jin Xiao, Yuhang Tian, Yanlin Jia, Xiaoyi Jiang, Lean Yu, and Shouyang Wang. Black-box attack-based security evaluation framework for credit card fraud detection models. INFORMS Journal on Computing, 35(5): 986–1001, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.1297.

Ruan:2023:ADR

[1882] Haolin Ruan, Zhi Chen, and Chin Pang Ho. Adjustable distributionally robust optimization with infinitely constrained ambiguity sets. IN-FORMS Journal on Computing, 35(5):1002–1023, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0181.

Ghaddar:2023:LSB

[1883] Bissan Ghaddar, Ignacio Gómez-Casares, Julio González-Díaz, Brais González-Rodríguez, Beatriz Pateiro-López, and Sofía Rodríguez-Ballesteros. Learning for spatial branching: an algorithm selection approach. INFORMS Journal on Computing, 35(5):1024–1043, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2022.0090.

Wen:2023:SDR

[1884] Canhong Wen, Zhenduo Li, Ruipeng Dong, Yijin Ni, and Wenliang Pan. Simultaneous dimension reduction and variable selection for multinomial logistic regression. INFORMS Journal on Computing, 35(5):1044–1060, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0132.

Zanetti:2023:IPI

[1885] Filippo Zanetti and Jacek Gondzio. An interior point–inspired algorithm for linear programs arising in discrete optimal transport. INFORMS Journal on Computing, 35(5):1061–1078, September/October 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0184.

Hertrich:2023:PGS

[1886] Christoph Hertrich and Martin Skutella. Provably good solutions to the knapsack problem via neural networks of bounded size. INFORMS Journal on Computing, 35(5):1079–1097, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2021.0225.

Liu:2023:CES

[1887] Junming Liu, Mingfei Teng, Weiwei Chen, and Hui Xiong. A costeffective sequential route recommender system for taxi drivers. IN-FORMS Journal on Computing, 35(5):1098–1119, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0112.

Liu:2023:GDR

[1888] Feng Liu, Zhi Chen, and Shuming Wang. Globalized distributionally robust counterpart. INFORMS Journal on Computing, 35(5):1120–1142, September/October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0274.

Dragotto:2023:ZRA

[1889] Gabriele Dragotto and Rosario Scatamacchia. The zero regrets algorithm: Optimizing over pure Nash equilibria via integer programming. INFORMS Journal on Computing, 35(5):1143–1160, September/ October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0282.

Bhattacharya:2023:NRS

[1890] Arnab Bhattacharya, Jeffrey P. Kharoufeh, and Bo Zeng. A nonconvex regularization scheme for the stochastic dual dynamic programming algorithm. INFORMS Journal on Computing, 35(5):1161–1178, September/ October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0255.

Lai:2023:AAK

[1891] Xiaofan Lai, Liang Xu, Zhou Xu, and Yang Du. An approximation algorithm for k-depot split delivery vehicle routing problem. INFORMS Journal on Computing, 35(5):1179–1194, September/October 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2021.0193.

Lamontagne:2023:OEV

[1892] Steven Lamontagne, Margarida Carvalho, Emma Frejinger, Bernard Gendron, Miguel F. Anjos, and Ribal Atallah. Optimising electric vehicle charging station placement using advanced discrete choice models. INFORMS Journal on Computing, 35(5):1195–1213, September/ October 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0185.

Anonymous:2023:EBe

[1893] Anonymous. Editorial Board. INFORMS Journal on Computing, 35(5): ??, September/October 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2023.eb.v3505.

Smith:2023:NEf

[1894] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 35(6):1215–1217, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.ed.v35.n6.

Chen:2023:SDB

[1895] Xiaotie Chen and David L. Woodruff. Software for data-based stochastic programming using bootstrap estimation. INFORMS Journal on Computing, 35(6):1218–1224, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0253.

Bertsimas:2023:PML

[1896] Dimitris Bertsimas and Cheol Woo Kim. A prescriptive machine learning approach to mixed-integer convex optimization. INFORMS Journal on Computing, 35(6):1225–1241, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0188.

Bhoopalam:2023:POB

[1897] Anirudh Kishore Bhoopalam, Niels Agatz, and Rob Zuidwijk. Platoon optimization based on truck pairs. INFORMS Journal on Computing, 35(6):1242–1260, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2020.0302.

Zhang:2023:AOS

[1898] Gongbo Zhang, Yijie Peng, Jianghua Zhang, and Enlu Zhou. Asymptotically optimal sampling policy for selecting top-m alternatives. IN-FORMS Journal on Computing, 35(6):1261–1285, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0333.

Luo:2023:HSC

[1899] Yuchen Luo, Bruce Golden, and Rui Zhang. The hot spot coverage patrol problem: Formulations and solution approaches. INFORMS Journal on Computing, 35(6):1286–1307, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0192.

Bottcher:2023:CDS

[1900] Lucas Böttcher, Thomas Asikis, and Ioannis Fragkos. Control of dual-sourcing inventory systems using recurrent neural networks. IN-FORMS Journal on Computing, 35(6):1308–1328, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0136.

Gleixner:2023:PPP

[1901] Ambros Gleixner, Leona Gottwald, and Alexander Hoen. PaPILO: A parallel presolving library for integer and linear optimization with multiprecision support. INFORMS Journal on Computing, 35(6):1329–1341, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0171.

Goyal:2023:DRA

[1902] Akshit Goyal, Yiling Zhang, and Chuan He. Decision rule approaches for pessimistic bilevel linear programs under moment ambiguity with facility location applications. INFORMS Journal on Computing, 35(6): 1342–1360, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0168.

Zhao:2023:DRC

[1903] Yue Zhao, Zhi Chen, and Zhenzhen Zhang. Distributionally robust chance-constrained p-hub center problem. INFORMS Journal on Computing, 35(6):1361–1382, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0113.

Kim:2023:LSE

[1904] Jongeun Kim, Sven Leyffer, and Prasanna Balaprakash. Learning symbolic expressions: Mixed-integer formulations, cuts, and heuristics. INFORMS Journal on Computing, 35(6):1383–1403, November/ December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0050.

Walter:2023:RSP

[1905] Matthias Walter. Recognizing series-parallel matrices in linear time. IN-FORMS Journal on Computing, 35(6):1404–1418, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0233.

Fahimi:2023:OCE

[1906] Hamed Fahimi and Claude-Guy Quimper. Overload-checking and edgefinding for robust cumulative scheduling. INFORMS Journal on Computing, 35(6):1419–1438, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2021.0138.

vandenHeuvel:2023:DAS

[1907] Wilco van den Heuvel, Semra Ağralı, and Z. Caner Taşkın. A decomposition algorithm for single and multiobjective integrated market selection and production planning. INFORMS Journal on Computing, 35(6):1439– 1453, November/December 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0053.

Porras:2023:TCS

[1908] Álvaro Porras, Concepción Domínguez, Juan Miguel Morales, and Salvador Pineda. Tight and compact sample average approximation for joint chance-constrained problems with applications to optimal power flow. INFORMS Journal on Computing, 35(6):1454–1469, November/ December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0302.

Hwang:2023:MLS

[1909] Hark-Chin Hwang, Wilco van den Heuvel, and Albert P. M. Wagelmans. Multilevel lot-sizing with inventory bounds. INFORMS Journal on Computing, 35(6):1470–1490, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0216.

Wen:2023:TF

[1910] Canhong Wen, Xueqin Wang, and Aijun Zhang. ℓ_{0} trend filtering. INFORMS Journal on Computing, 35(6):1491–1510, November/ December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0313.

Eibelshauser:2023:LST

[1911] Steffen Eibelshäuser, Victor Klockmann, David Poensgen, and Alicia von Schenk. The logarithmic stochastic tracing procedure: a homotopy method to compute stationary equilibria of stochastic games. IN-FORMS Journal on Computing, 35(6):1511–1526, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0360.

Anonymous:2023:AR

[1912] Anonymous. Appreciation to reviewers, 2023. INFORMS Journal on Computing, 35(6):1527–1532, November/December 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.reviewthx.
v35.n6.

Anonymous:2023:EBf

[1913] Anonymous. Editorial Board. INFORMS Journal on Computing, 35(6): ??, November/December 2023. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2023.eb.v3506.

Smith:2024:NEa

[1914] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 36(1):1–2, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2024.ed.v36.n1.

Bock:2024:RRW

[1915] Stefan Bock and Nils Boysen. Routing replenishment workers: The prize collecting traveling salesman problem in scattered storage warehouses. INFORMS Journal on Computing, 36(1):3–20, January/ February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0173.

Michini:2024:SCA

[1916] Carla Michini, Peter Ohmann, Ben Liblit, and Jeff Linderoth. A setcovering approach to customized coverage instrumentation. INFORMS Journal on Computing, 36(1):21–38, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2021.0349.

Zhou:2024:DCN

[1917] Yangming Zhou, Jiaqi Li, Jin-Kao Hao, and Fred Glover. Detecting critical nodes in sparse graphs via “Reduce–Solve–Combine” memetic search. INFORMS Journal on Computing, 36(1):39–60, January/ February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0130.

Ahanor:2024:DNM

[1918] Izuwa Ahanor, Hugh Medal, and Andrew C. Trapp. DiversiTree: A new method to efficiently compute diverse sets of near-optimal solutions to mixed-integer optimization problems. INFORMS Journal on Computing, 36(1):61–77, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0164.

Lefebvre:2024:ARO

[1919] Henri Lefebvre, Enrico Malaguti, and Michele Monaci. Adjustable robust optimization with discrete uncertainty. INFORMS Journal on Computing, 36(1):78–96, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0086.

Li:2024:DOD

[1920] Yongchun Li, Marcia Fampa, Jon Lee, Feng Qiu, Weijun Xie, and Rui Yao. D-optimal data fusion: Exact and approximation algorithms. INFORMS Journal on Computing, 36(1):97–120, January/ February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0235.

Huang:2024:SOR

[1921] Yeu-Shiang Huang, Chih-Chiang Fang, Chun-Hsuan Chou, and Tzu-Liang (Bill) Tseng. A study on optimal release schedule for multiversion software. INFORMS Journal on Computing, 36(1):121–140, January/ February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0141.

Baldacci:2024:NEA

[1922] Roberto Baldacci, Stefano Coniglio, Jean-François Cordeau, and Fabio Furini. A numerically exact algorithm for the bin-packing problem.

INFORMS Journal on Computing, 36(1):141–162, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0257.

Li:2024:ASD

[1923] Tianyi Li and Munther A. Dahleh. Automation of strategic data prioritization in system model calibration: Sensor placement. INFORMS Journal on Computing, 36(1):163–184, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0128.

Yan:2024:FOU

[1924] Yikai Yan, Chaoyue Niu, Yucheng Ding, Zhenzhe Zheng, Shaojie Tang, Qinya Li, Fan Wu, Chengfei Lyu, Yanghe Feng, and Guihai Chen. Federated optimization under intermittent client availability. INFORMS Journal on Computing, 36(1):185–202, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0057.

Larsen:2024:FCI

[1925] Eric Larsen, Emma Frejinger, Bernard Gendron, and Andrea Lodi. Fast continuous and integer L-shaped heuristics through supervised learning. INFORMS Journal on Computing, 36(1):203–223, January/ February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0175.

Bianchessi:2024:RWR

[1926] Nicola Bianchessi, Timo Gschwind, and Stefan Irnich. Resource-window reduction by reduced costs in path-based formulations for routing and scheduling problems. INFORMS Journal on Computing, 36(1):224–244, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0214.

Escobedo:2024:EMF

[1927] Adolfo R. Escobedo. Exact matrix factorization updates for nonlinear programming. INFORMS Journal on Computing, 36(1):245–265, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2021.0331.

Li:2024:MCM

[1928] Yajuan Li, Zachary T. Kaplan, and Marvin K. Nakayama. Monte Carlo methods for economic capital. INFORMS Journal on Computing, 36(1): 266–284, January/February 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2021.0261.

Forget:2024:EBB

[1929] Nicolas Forget and Sophie N. Parragh. Enhancing branch-and-bound for multiobjective 0–1 programming. INFORMS Journal on Computing, 36 (1):285–304, January/February 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0299.

Anonymous:2024:EBa

[1930] Anonymous. Editorial Board. INFORMS Journal on Computing, 36 (1):??, January/February 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2024.eb.v3601.

Smith:2024:NEb

[1931] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 36(2):305–307, March/April 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2024.ed.v36.n2.

Zhou:2024:HSR

[1932] Yangming Zhou, Jin-Kao Hao, and Zhen Li. Heuristic search for rank aggregation with application to label ranking. INFORMS Journal on Computing, 36(2):308–326, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0019.

Garcia:2024:BJM

[1933] Joaquim Dias Garcia, Guilherme Bodin, and Alexandre Street. BilevelJuMP.jl: Modeling and solving bilevel optimization problems in Julia. INFORMS Journal on Computing, 36(2):327–335, March/ April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0135.

Raghunathan:2024:SMT

[1934] Arvind U. Raghunathan, David Bergman, John N. Hooker, Thiago Serra, and Shingo Kobori. Seamless multimodal transportation scheduling. IN-FORMS Journal on Computing, 36(2):336–358, March/April 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2019.0163.

Tang:2024:GEA

[1935] Shaojie Tang and Jing Yuan. Group equality in adaptive submodular maximization. INFORMS Journal on Computing, 36(2):359–376, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2022.0384.

Cao:2024:DPF

[1936] Yiyin Cao, Yin Chen, and Chuangyin Dang. A differentiable pathfollowing method with a compact formulation to compute proper equilibria. INFORMS Journal on Computing, 36(2):377–396, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0148.

Krug:2024:CBA

[1937] Richard Krug, Günter Leugering, Alexander Martin, Martin Schmidt, and Dieter Weninger. A consensus-based alternating direction method for mixed-integer and PDE-constrained gas transport problems. IN-FORMS Journal on Computing, 36(2):397–416, March/April 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0319.

Sinjorgo:2024:SMS

[1938] Lennart Sinjorgo and Renata Sotirov. On solving MAX-SAT using sum of squares. INFORMS Journal on Computing, 36(2):417–433, March/ April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0036.

Ajayi:2024:CCO

[1939] Temitayo Ajayi, Seyedmohammadhossein Hosseinian, Andrew J. Schaefer, and Clifton D. Fuller. Combination chemotherapy optimization with discrete dosing. INFORMS Journal on Computing, 36(2):434–455, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2022.0207.

Besancon:2024:FDO

[1940] Mathieu Besançon, Joaquim Dias Garcia, Benoît Legat, and Akshay Sharma. Flexible differentiable optimization via model transformations. INFORMS Journal on Computing, 36(2):456–478, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0283.

Leitner:2024:EMC

[1941] Markus Leitner, Andrea Lodi, Roberto Roberti, and Claudio Sole. An exact method for (constrained) assortment optimization problems with product costs. INFORMS Journal on Computing, 36(2):479–494, March/ April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0262.

Tan:2024:BNM

[1942] Yi Tan, Prakash P. Shenoy, Ben Sherwood, Catherine Shenoy, Melinda Gaddy, and Mary E. Oehlert. Bayesian network models for PTSD screening in veterans. INFORMS Journal on Computing, 36(2):495– 509, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2021.0174.

Zhang:2024:EAG

[1943] Yaoyao Zhang, Chaojie Zhu, Shaojie Tang, Yingli Ran, Ding-Zhu Du, and Zhao Zhang. Evolutionary algorithm on general cover with theoretically guaranteed approximation ratio. INFORMS Journal on Computing, 36(2):510–525, March/April 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0327.

Fan:2024:DRA

[1944] Xiangyi Fan and Grani A. Hanasusanto. A decision rule approach for two-stage data-driven distributionally robust optimization problems with random recourse. INFORMS Journal on Computing, 36(2):526–542, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2021.0306.

Sun:2024:DMG

[1945] Defeng Sun, Lixin Tang, Roberto Baldacci, and Zihan Chen. A decomposition method for the group-based quay crane scheduling problem. IN-FORMS Journal on Computing, 36(2):543–570, March/April 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0298.

Lin:2024:DGM

[1946] Hao Lin, Guannan Liu, Junjie Wu, and J. Leon Zhao. Deterring the gray market: Product diversion detection via learning disentangled representations of multivariate time series. INFORMS Journal on Computing, 36 (2):571–586, March/April 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0155.

Xu:2024:EVR

[1947] Jiajun Xu and Suvrajeet Sen. Ensemble variance reduction methods for stochastic mixed-integer programming and their application to the stochastic facility location problem. INFORMS Journal on Computing, 36(2):587–599, March/April 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2021.0324.

Bent:2024:ICM

[1948] Russell Bent, Byron Tasseff, and Carleton Coffrin. InfrastructureModels: Composable multi-infrastructure optimization in Julia. INFORMS Journal on Computing, 36(2):600–615, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0118.

Aghasi:2024:DLI

[1949] Alireza Aghasi, Arun Rai, and Yusen Xia. A deep learning and image processing pipeline for object characterization in firm operations. IN-FORMS Journal on Computing, 36(2):616–634, March/April 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0260.

Liu:2024:SGC

[1950] Lindong Liu, Xiangtong Qi, and Zhou Xu. Stabilizing grand cooperation via cost adjustment: an inverse optimization approach. IN-FORMS Journal on Computing, 36(2):635–656, March/April 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0268.

DAlessandro:2024:DAA

[1951] Pietro D’Alessandro, Manlio Gaudioso, Giovanni Giallombardo, and Giovanna Miglionico. The descent–ascent algorithm for DC programming. INFORMS Journal on Computing, 36(2):657–671, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0142.

Ales:2024:CCP

[1952] Zacharie Ales, Céline Engelbeen, and Rosa Figueiredo. Correlation clustering problem under mediation. INFORMS Journal on Computing, 36 (2):672–689, March/April 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0129.

Xu:2024:EGO

[1953] Wei Xu, Jie Tang, Ka Fai Cedric Yiu, and Jian Wen Peng. An efficient global optimal method for cardinality constrained portfolio optimization. INFORMS Journal on Computing, 36(2):690–704, March/ April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0344.

Anonymous:2024:EBb

[1954] Anonymous. Editorial Board. INFORMS Journal on Computing, 36(2): ??, March/April 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2024.eb.v3602.

Smith:2024:NEc

[1955] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 36(3):705–707, May/June 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2024.ed.v36.n3.

Rajabalizadeh:2024:SCC

[1956] Atefeh Rajabalizadeh and Danial Davarnia. Solving a class of cutgenerating linear programs via machine learning. INFORMS Journal on Computing, 36(3):708–722, May/June 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0241.

Balvert:2024:IRE

[1957] Marleen Balvert. Iterative rule extension for logic analysis of data: an MILP-based heuristic to derive interpretable binary classifiers from large data sets. INFORMS Journal on Computing, 36(3):723–741, May/ June 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2021.0284.

Qin:2024:PRI

[1958] Jindong Qin, Pan Zheng, and Xiaojun Wang. Product redesign and innovation based on online reviews: a multistage combined search method. INFORMS Journal on Computing, 36(3):742–765, May/June 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0333.

Baron:2024:SMS

[1959] Opher Baron, Dmitry Krass, Arik Senderovich, and Eliran Sherzer. Supervised ML for solving the GI/GI/1 queue. INFORMS Journal on Computing, 36(3):766–786, May/June 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0263.

Behroozi:2024:LVI

[1960] Mehdi Behroozi. Largest volume inscribed rectangles in convex sets defined by finite number of inequalities. INFORMS Journal on Computing, 36(3):787–819, May/June 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0239.

Dai:2024:SED

[1961] Yongzheng Dai and Chen Chen. Sparsity-exploiting distributed projections onto a simplex. INFORMS Journal on Computing, 36(3):820–835, May/June 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0328.

Hermelin:2024:MWN

[1962] Danny Hermelin, Hendrik Molter, and Dvir Shabtay. Minimizing the weighted number of tardy jobs via (max,+)-convolutions. IN-FORMS Journal on Computing, 36(3):836–848, May/June 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0307.

Wang:2024:CGS

[1963] Shanshan Wang and Erick Delage. A column generation scheme for distributionally robust multi-item newsvendor problems. INFORMS Journal on Computing, 36(3):849–867, May/June 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0010.

vanDoornmalen:2024:EPT

[1964] Jasper van Doornmalen and Christopher Hojny. Efficient propagation techniques for handling cyclic symmetries in binary programs. IN-FORMS Journal on Computing, 36(3):868–883, May/June 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0060.

Gu:2024:SSS

[1965] Xiaoyi Gu, Santanu S. Dey, and Jean-Philippe P. Richard. Solving sparse separable bilinear programs using lifted bilinear cover inequalities. IN-FORMS Journal on Computing, 36(3):884–899, May/June 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0230.

Loger:2024:AKL

[1966] Benoît Loger, Alexandre Dolgui, Fabien Lehuédé, and Guillaume Massonnet. Approximate kernel learning uncertainty set for robust combinatorial optimization. INFORMS Journal on Computing, 36(3):900–917, May/June 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0330.

Bansal:2024:SFR

[1967] Ankit Bansal, Jean-Philippe Richard, Bjorn P. Berg, and Yu-Li Huang. A sequential follower refinement algorithm for robust surgery scheduling. INFORMS Journal on Computing, 36(3):918–937, May/June 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0191.

Anonymous:2024:EBc

[1968] Anonymous. Editorial Board. INFORMS Journal on Computing, 36 (3):??, May/June 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2024.eb.v3603.

Smith:2024:NEd

[1969] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 36(4):939–940, July/August 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2024.ed.v36.n4.

Kullman:2024:ISS

[1970] Nicholas D. Kullman, Jorge E. Mendoza, and Ted K. Ralphs. Introduction to the special section on software tools for vehicle routing. INFORMS Journal on Computing, 36(4):941–942, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.ed.v36.n4.2.

Wouda:2024:PHP

[1971] Niels A. Wouda, Leon Lan, and Wouter Kool. PyVRP: a highperformance VRP solver package. INFORMS Journal on Computing, 36 (4):943–955, July/August 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0055.

Errami:2024:VPL

[1972] Najib Errami, Eduardo Queiroga, Ruslan Sadykov, and Eduardo Uchoa. VRPSolverEasy: a Python library for the exact solution of a rich vehicle routing problem. INFORMS Journal on Computing, 36(4):956–965, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2023.0103.

Klein:2024:ROS

[1973] Patrick S. Klein and Maximilian Schiffer. RoutingBlocks: an open-source Python package for vehicle routing problems with intermediate stops. IN-FORMS Journal on Computing, 36(4):966–973, July/August 2024. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0104.

Maximo:2024:AIA

[1974] Vinícius R. Máximo, Jean-François Cordeau, and Mariá C. V. Nascimento. AILS-II: an adaptive iterated local search heuristic for the largescale capacitated vehicle routing problem. INFORMS Journal on Computing, 36(4):974–986, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.0106.

Kim:2024:NSA

[1975] Hyeonah Kim, Jinkyoo Park, and Changhyun Kwon. A neural separation algorithm for the rounded capacity inequalities. INFORMS Journal on Computing, 36(4):987–1005, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0310.

You:2024:ADP

[1976] Fan You and Thomas Vossen. An approximate dynamic programming approach to dynamic stochastic matching. INFORMS Journal on Computing, 36(4):1006–1022, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2021.0203.

Song:2024:SAI

[1977] Eunhye Song, Henry Lam, and Russell R. Barton. A shrinkage approach to improve direct bootstrap resampling under input uncertainty. INFORMS Journal on Computing, 36(4):1023–1039, July/ August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0044.

Tasseff:2024:PRO

[1978] Byron Tasseff, Russell Bent, Carleton Coffrin, Clayton Barrows, Devon Sigler, Jonathan Stickel, Ahmed S. Zamzam, Yang Liu, and Pascal Van Hentenryck. Polyhedral relaxations for optimal pump scheduling of potable water distribution networks. INFORMS Journal on Computing, 36(4):1040–1063, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0233.

Blufstein:2024:DSS

[1979] Marcos Blufstein, Gonzalo Lera-Romero, and Francisco J. Soulignac. Decremental state-space relaxations for the basic traveling salesman problem with a drone. INFORMS Journal on Computing, 36(4):1064– 1083, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0390.

Cattaruzza:2024:EHS

[1980] Diego Cattaruzza, Martine Labbé, Matteo Petris, Marius Roland, and Martin Schmidt. Exact and heuristic solution techniques for mixedinteger quantile minimization problems. INFORMS Journal on Computing, 36(4):1084–1107, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0105.

Legrain:2024:DPA

[1981] Antoine Legrain and Jérémy Omer. A dedicated pricing algorithm to solve a large family of nurse scheduling problems with branchand-price. INFORMS Journal on Computing, 36(4):1108–1128, July/ August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0019.

Sugishita:2024:UML

[1982] Nagisa Sugishita, Andreas Grothey, and Ken McKinnon. Use of machine learning models to warmstart column generation for unit commitment. INFORMS Journal on Computing, 36(4):1129–1146, July/ August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0140.

Anonymous:2024:EBd

[1983] Anonymous. Editorial Board. INFORMS Journal on Computing, 36(4): ??, July/August 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2024.eb.v3604.

Smith:2024:NEe

[1984] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 36(5):1147–1149, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2024.ed.v36.n5.

Schmitt:2024:QOM

[1985] Christian Schmitt and Bismark Singh. Quadratic optimization models for balancing preferential access and fairness: Formulations and optimality conditions. INFORMS Journal on Computing, 36(5):1150–1167, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0308.

Jiang:2024:RTD

[1986] Guangxin Jiang, L. Jeff Hong, and Haihui Shen. Real-time derivative pricing and hedging with consistent metamodels. INFORMS Journal on Computing, 36(5):1168–1189, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0292.

Du:2024:AIM

[1987] Hongmin W. Du, Yingfan L. Du, and Zhao Zhang. Adaptive influence maximization: Adaptability via nonadaptability. INFORMS Journal on Computing, 36(5):1190–1200, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0267.

Wang:2024:ASD

[1988] Shuming Wang, Jun Li, Marcus Ang, and Tsan Sheng Ng. Appointment scheduling with delay tolerance heterogeneity. INFORMS Journal on Computing, 36(5):1201–1224, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0025.

Li:2024:SBP

[1989] Yu-Wei Li, Gui-Hua Lin, and Xide Zhu. Solving bilevel programs based on lower-level Mond–Weir duality. INFORMS Journal on Computing, 36 (5):1225–1241, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.0108.

Liu:2024:CEA

[1990] Xiaoping Liu and Xiao-Bai Li. Cost-effective acquisition of first-party data for business analytics. INFORMS Journal on Computing, 36(5): 1242–1260, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0037.

Dehghanian:2024:ISO

[1991] Amin Dehghanian, Yujia Xie, and Nicoleta Serban. Identifying socially optimal equilibria using combinatorial properties of Nash equilibria in bimatrix games. INFORMS Journal on Computing, 36(5):1261–1286, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0072.

Wu:2024:EMP

[1992] Tao Wu. Exact method for production hub location. INFORMS Journal on Computing, 36(5):1287–1315, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0339.

Maragno:2024:FRC

[1993] Donato Maragno, Jannis Kurtz, Tabea E. Röber, Rob Goedhart, Ş. İlker Birbil, and Dick den Hertog. Finding regions of counterfactual explanations via robust optimization. INFORMS Journal on Computing, 36 (5):1316–1334, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.0153.

Okumusoglu:2024:IPM

[1994] Bahar Cennet Okumuşoğlu, Beste Basciftci, and Burak Kocuk. An integrated predictive maintenance and operations scheduling framework for power systems under failure uncertainty. INFORMS Journal on Computing, 36(5):1335–1358, September/October 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0154.

Anonymous:2024:EBe

[1995] Anonymous. Editorial Board. INFORMS Journal on Computing, 36(5): ??, September/October 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2024.eb.v3605.

Anonymous:2024:NE

[1996] Anonymous. Note from the Editor. INFORMS Journal on Computing, 36(6):1359–1361, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2024.ed.v36.n6.

Kahr:2024:IPS

[1997] Michael Kahr, Markus Leitner, and Ivana Ljubić. The impact of passive social media viewers in influence maximization. INFORMS Journal on Computing, 36(6):1362–1381, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0047.

Shi:2024:CLD

[1998] Chenbo Shi, Mohsen Emadikhiav, Leonardo Lozano, and David Bergman. Constraint learning to define trust regions in optimization over pre-trained predictive models. INFORMS Journal on Computing, 36(6): 1382–1399, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0312.

Zhang:2024:LLB

[1999] Zhu (Drew) Zhang, Jie Yuan, and Amulya Gupta. Let the laser beam connect the dots: Forecasting and narrating stock market volatility. INFORMS Journal on Computing, 36(6):1400–1416, November/ December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0055.

Hessler:2024:ESS

[2000] Katrin Heßler and Stefan Irnich. Exact solution of the single-picker routing problem with scattered storage. INFORMS Journal on Computing, 36(6):1417–1435, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0075.

Lin:2024:UFC

[2001] Yun Hui Lin, Qingyun Tian, and Yanlu Zhao. Unified framework for choice-based facility location problem. INFORMS Journal on Computing, 36(6):1436–1458, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0366.

Guan:2024:PBF

[2002] Hongzhao Guan, Beste Basciftci, and Pascal Van Hentenryck. Pathbased formulations for the design of on-demand multimodal transit systems with adoption awareness. INFORMS Journal on Computing, 36(6): 1459–1480, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.0014.

Liang:2024:FMN

[2003] Guo Liang, Kun Zhang, and Jun Luo. A FAST method for nested estimation. INFORMS Journal on Computing, 36(6):1481–1500, November/ December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0118.

Bayani:2024:DBF

[2004] Mahdis Bayani, Borzou Rostami, Yossiri Adulyasak, and Louis-Martin Rousseau. A dual bounding framework through cost splitting for binary quadratic optimization. INFORMS Journal on Computing, 36(6):1501– 1521, November/December 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2021.0186.

Coppe:2024:DDB

[2005] Vianney Coppé, Xavier Gillard, and Pierre Schaus. Decision diagrambased branch-and-bound with caching for dominance and suboptimality detection. INFORMS Journal on Computing, 36(6):1522–1542, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528

(electronic). URL https://pubsonline.informs.org/doi/full/10.
1287/ijoc.2022.0340.

Kocyigit:2024:RMS

[2006] Çağıl Koçyiğit, Daniel Kuhn, and Napat Rujeerapaiboon. Regret minimization and separation in multi-bidder, multi-item auctions. IN-FORMS Journal on Computing, 36(6):1543–1561, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0275.

Xavier:2024:DFT

[2007] Álinson Santos Xavier, Santanu Subhas Dey, and Feng Qiu. Decomposable formulation of transmission constraints for decentralized power systems optimization. INFORMS Journal on Computing, 36(6):1562– 1578, November/December 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0326.

Halbig:2024:COC

[2008] Katrin Halbig, Lukas Hümbs, Florian Rösel, Lars Schewe, and Dieter Weninger. Computing optimality certificates for convex mixed-integer nonlinear problems. INFORMS Journal on Computing, 36(6):1579–1610, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0099.

Wang:2024:CRZ

[2009] Tianyu Wang and Yasong Feng. Convergence rates of zeroth order gradient descent for Łojasiewicz functions. INFORMS Journal on Computing, 36(6):1611–1633, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0247.

Zhang:2024:CBM

[2010] Jiachen Zhang, Youcef Magnouche, Pierre Bauguion, Sebastien Martin, and J. Christopher Beck. Computing bipath multicommodity flows with constraint programming–based branch-and-price-and-cut. IN-FORMS Journal on Computing, 36(6):1634–1653, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0128.

Lendl:2024:RNO

[2011] Stefan Lendl, Ulrich Pferschy, and Elena Rener. Rescheduling with new orders under bounded disruption. INFORMS Journal on Computing, 36(6):1654–1675, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0038.

Artigues:2024:CTR

[2012] Christian Artigues, Emmanuel Hébrard, Alain Quilliot, and Hélène Toussaint. The continuous time-resource trade-off scheduling problem with time Windows. INFORMS Journal on Computing, 36(6):1676–1695, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0142.

Kowalczyk:2024:FBF

[2013] Daniel Kowalczyk, Roel Leus, Christopher Hojny, and Stefan Røpke. A flow-based formulation for parallel machine scheduling using decision diagrams. INFORMS Journal on Computing, 36(6):1696–1714, November/ December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0301.

Steever:2024:GBA

[2014] Zachary Steever, Kyle Hunt, Mark Karwan, Junsong Yuan, and Chase C. Murray. A graph-based approach for relating integer programs. IN-FORMS Journal on Computing, 36(6):1715–1736, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0255.

Fullner:2024:FVU

[2015] Christian Füllner, Peter Kirst, Hendrik Otto, and Steffen Rebennack. Feasibility verification and upper bound computation in global minimization using approximate active index sets. INFORMS Journal on Computing, 36(6):1737–1756, November/December 2024. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0162.

Anonymous:2024:EBf

[2016] Anonymous. Editorial Board. INFORMS Journal on Computing, 36(6): ??, November/December 2024. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2024.eb.v3606.

Smith:2025:NEa

[2017] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 27(1):1, January/February 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2025.ed.v37.n1.2.

Coffrin:2025:ISI

[2018] Carleton Coffrin, Elisabeth Lobe, Giacomo Nannicini, and Ojas Parekh. Introduction to the special issue on quantum computing and operations research. INFORMS Journal on Computing, 27(1):2, January/ February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2025.ed.v37.n1.

Karimi:2025:EII

[2019] Mehdi Karimi and Levent Tunçel. Efficient implementation of interiorpoint methods for quantum relative entropy. INFORMS Journal on Computing, 27(1):3–21, January/February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2024.0570.

Wagner:2025:ONS

[2020] Friedrich Wagner, Daniel J. Egger, and Frauke Liers. Optimized noise suppression for quantum circuits. INFORMS Journal on Computing, 27(1):22–41, January/February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2024.0551.

Jung:2025:NMN

[2021] Jihye Jung and In-Chan Choi. A new multicommodity network flow model and branch and cut for optimal quantum Boolean circuit synthesis. INFORMS Journal on Computing, 27(1):42–62, January/February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.0562.

Larson:2025:NNA

[2022] Jeffrey Larson, Matt Menickelly, and Jiahao Shi. A novel noise-aware classical optimizer for variational quantum algorithms. INFORMS Journal on Computing, 27(1):63–85, January/February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2024.0578.

Fei:2025:BQC

[2023] Xinyu Fei, Lucas T. Brady, Jeffrey Larson, Sven Leyffer, and Siqian Shen. Binary quantum control optimization with uncertain Hamiltonians. INFORMS Journal on Computing, 27(1):86–106, January/ February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.0560.

Kushnir:2025:QSN

[2024] Samuel Kushnir, Jiaqi Leng, Yuxiang Peng, Lei Fan, and Xiaodi Wu. QH-DOPT: A software for nonlinear optimization with quantum Hamiltonian descent. INFORMS Journal on Computing, 27(1):107–124, January/ February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.0587.

Ha:2025:TSE

[2025] Yunsoo Ha, Sara Shashaani, and Matt Menickelly. Two-stage estimation and variance modeling for latency-constrained variational quantum algorithms. INFORMS Journal on Computing, 27(1):125–145, January/ February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.0575.

Katial:2025:IDP

[2026] Vivek Katial, Kate Smith-Miles, Charles Hill, and Lloyd Hollenberg. On the instance dependence of parameter initialization for the quantum approximate optimization algorithm: Insights via instance space analysis. INFORMS Journal on Computing, 27(1):146–171, January/ February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.0564.

Karahalios:2025:QIB

[2027] Anthony Karahalios, Sridhar Tayur, Ananth Tenneti, Amirreza Pashapour, F. Sibel Salman, and Barış Yıldız. A quantum-inspired bilevel optimization algorithm for the first responder network design problem. INFORMS Journal on Computing, 27(1):172–188, January/ February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.0574.

Anonymous:2025:AR

[2028] Anonymous. Appreciation to reviewers. INFORMS Journal on Computing, 27(1):189–196, January/February 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2025.thx.v37.n1.

Anonymous:2025:EBa

[2029] Anonymous. Editorial Board. INFORMS Journal on Computing, 27 (1):??, January/February 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2025.eb.v3701.

Smith:2025:NEb

[2030] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 37(2):197–198, March/April 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2025.ed.v37.n2.

Du:2025:SMS

[2031] Yu Du, Fred Glover, Gary Kochenberger, Rick Hennig, Haibo Wang, and Amit Hulandageri. Solving the minimum sum coloring problem: Alternative models, exact solvers, and metaheuristics. INFORMS Journal on Computing, 37(2):199–211, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0334.

Im:2025:SFO

[2032] Hyungki Im and Paul Grigas. Stochastic first-order algorithms for constrained distributionally robust optimization. INFORMS Journal on Computing, 37(2):212–229, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0167.

Cherkesly:2025:RDD

[2033] Marilène Cherkesly, Claudio Contardo, and Matthieu Gruson. Ranking decomposition for the discrete ordered median problem. INFORMS Journal on Computing, 37(2):230–248, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0059.

Caskurlu:2025:MTC

[2034] Bugra Caskurlu, K. Subramani, Utku Umur Acikalin, Alvaro Velasquez, and Piotr Wojciechowski. Models for test cost minimization in database migration. INFORMS Journal on Computing, 37(2):249–269, March/ April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0021.

Austgen:2025:CTS

[2035] Brent Austgen, Erhan Kutanoglu, John J. Hasenbein, and Surya Santoso. Comparisons of two-stage models for flood mitigation of electrical substations. INFORMS Journal on Computing, 37(2):270–292, March/ April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0125.

Guillet:2025:CCR

[2036] Marianne Guillet and Maximilian Schiffer. Coordinating charging request allocation between self-interested navigation service platforms. IN-FORMS Journal on Computing, 37(2):293–314, March/April 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0269.

Du:2025:TRL

[2037] Mu Du, Hongtao Yu, and Nan Kong. Transfer reinforcement learning for mixed observability Markov decision processes with time-varying interval-valued parameters and its application in pandemic control. IN-FORMS Journal on Computing, 37(2):315–337, March/April 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0236.

Deng:2025:EAD

[2038] Qi Deng, Qing Feng, Wenzhi Gao, Dongdong Ge, Bo Jiang, Yuntian Jiang, Jingsong Liu, Tianhao Liu, Chenyu Xue, Yinyu Ye, and Chuwen Zhang. An enhanced alternating direction method of multipliers-based interior point method for linear and conic optimization. INFORMS Journal on Computing, 37(2):338–359, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0017.

Chen:2025:OIP

[2039] Zhi-Long Chen. Online integrated production and distribution scheduling: Review and extensions. INFORMS Journal on Computing, 37(2): 360–380, March/April 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0305.

Jiang:2025:TII

[2040] Nan Jiang and Weijun Xie. The terminator: an integration of inner and outer approximations for solving Wasserstein distributionally robust chance constrained programs via variable fixing. INFORMS Journal on Computing, 37(2):381–412, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0299.

Rong:2025:SAD

[2041] Dongling Rong, Xinyu Sun, Meilin Zhang, and Shuangchi He. Satisficing approach to on-demand ride matching. INFORMS Journal on Computing, 37(2):413–427, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2021.0210.

Li:2025:RPP

[2042] Wenhui Li, Xin Zhou, Ruipeng Dong, and Zemin Zheng. Robust parallel pursuit for large-scale association network learning. INFORMS Journal on Computing, 37(2):428–445, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0181.

Wang:2025:EFL

[2043] Yaqiong Wang, Junjie Wu, Zhiang Wu, and Gediminas Adomavicius. Efficient and flexible long-tail recommendation using cosine patterns. IN-FORMS Journal on Computing, 37(2):446–464, March/April 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0194.

Li:2025:FPA

[2044] Qiudan Li, David Jingjun Xu, Haoda Qian, Linzi Wang, Minjie Yuan, and Daniel Dajun Zeng. A fusion pretrained approach for identifying the cause of sarcasm remarks. INFORMS Journal on Computing, 37 (2):465–479, March/April 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0285.

Yang:2025:IEE

[2045] Jin Yang, Guangxin Jiang, Yinan Wang, and Ying Chen. An intelligent end-to-end neural architecture search framework for electricity forecasting model development. INFORMS Journal on Computing, 37(2): 480–501, March/April 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0034.

Anonymous:2025:RMR

[2046] Anonymous. Recognition of meritorious reviewers, 2024. INFORMS Journal on Computing, 37(2):502, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2025.merits.
v37.n2.

Anonymous:2025:EBb

[2047] Anonymous. Editorial board. INFORMS Journal on Computing, 37(2): ??, March/April 2025. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2025.eb.v3702.

Smith:2025:NEc

[2048] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 37(3):??, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2025.ed.v37.n3.

Reich:2025:MPO

[2049] Daniel Reich and Leah Frye. Microgrid Planner: an open-source software platform. INFORMS Journal on Computing, 37(3):503–515, May/ June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0336.

Ioannou:2025:PJE

[2050] Ekaterini Ioannou, Konstantinos Nikoletos, and George Papadakis. py-JedAI: a library with resolution-related structures and procedures for products. INFORMS Journal on Computing, 37(3):516–530, May/ June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0410.

Ma:2025:LHM

[2051] Simin Ma, Amin Dehghanian, Gian-Gabriel Garcia, and Nicoleta Serban. Learning hidden Markov models with structured transition dynamics. INFORMS Journal on Computing, 37(3):531–556, May/June 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0342.

Postek:2025:FOA

[2052] Krzysztof Postek and Shimrit Shtern. First-order algorithms for robust optimization problems via convex–concave saddle-point Lagrangian reformulation. INFORMS Journal on Computing, 37(3):557–581, May/ June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0200.

Li:2025:EAA

[2053] Yongchun Li and Weijun Xie. Exact and approximation algorithms for sparse principal component analysis. INFORMS Journal on Computing, 37(3):582–602, May/June 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2022.0372.

Legault:2025:MFA

[2054] Robin Legault and Emma Frejinger. A model-free approach for solving choice-based competitive facility location problems using simulation and submodularity. INFORMS Journal on Computing, 37(3):603–622, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0280.

Bernardelli:2025:MLE

[2055] Ambrogio Maria Bernardelli, Stefano Gualandi, Simone Milanesi, Hoong Chuin Lau, and Neil Yorke-Smith. Multiobjective linear ensembles for robust and sparse training of few-bit neural networks. IN-FORMS Journal on Computing, 37(3):623–643, May/June 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0281.

Julien:2025:MLK

[2056] Esther Julien, Krzysztof Postek, and Ş. İlker Birbil. Machine learning for K-adaptability in two-stage robust optimization. INFORMS Journal on Computing, 37(3):644–665, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0314.

Clarkson:2025:COS

[2057] Jake Clarkson and Kyle Y. Lin. Computing optimal strategies for a search game in discrete locations. INFORMS Journal on Computing, 37 (3):666–683, May/June 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0155.

Liu:2025:CHS

[2058] Xuanqi Liu and Ke-Wei Huang. Controlling homophily in social network regression analysis by machine learning. INFORMS Journal on Computing, 37(3):684–702, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2022.0287.

Feng:2025:RMA

[2059] Ziheng Feng, Baihua He, Tianfa Xie, Xinyu Zhang, and Xianpeng Zong. Ranking model averaging: Ranking based on model averaging. IN-FORMS Journal on Computing, 37(3):703–722, May/June 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0257.

Feng:2025:ENS

[2060] Ben Mingbin Feng and Eunhye Song. Efficient nested simulation experiment design via the likelihood ratio method. INFORMS Journal on Computing, 37(3):723–742, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0392.

He:2025:ESR

[2061] Xiaolei He and Weiguo Zhang. An efficient scenario reduction method for problems with higher moment coherent risk measures. INFORMS Journal on Computing, 37(3):743–760, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2022.0375.

Hill:2025:EPS

[2062] Alessandro Hill and Thomas W. M. Vossen. Efficient project scheduling with autonomous learning opportunities. INFORMS Journal on Computing, 37(3):761–783, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.0107.

Anonymous:2025:EBc

[2063] Anonymous. Editorial board. INFORMS Journal on Computing, 37 (3):??, May/June 2025. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2025.eb.v3703.

Smith:2025:NEd

[2064] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 37(4):??, July/August 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2025.ed.v37.n4.

Liu:2025:ENS

[2065] Xiaotian Liu, Yijie Peng, Gongbo Zhang, and Ruihan Zhou. An efficient node selection policy for Monte Carlo tree search with neural networks. INFORMS Journal on Computing, 37(4):785–807, July/ August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0307.

Yamin:2025:EVR

[2066] Daniel Yamín, Guy Desaulniers, and Jorge E. Mendoza. The electric vehicle routing and overnight charging scheduling problem on a multigraph. INFORMS Journal on Computing, 37(4):808–830, July/ August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0404.

Hewitt:2025:PPU

[2067] Mike Hewitt and Giovanni Pantuso. Production planning under demand and endogenous supply uncertainty. INFORMS Journal on Computing, 37(4):831–855, July/August 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0067.

Cui:2025:DMU

[2068] Xiangyu Cui, Rujun Jiang, Yun Shi, Rufeng Xiao, and Yifan Yan. Decision making under cumulative prospect theory: an alternating direction method of multipliers. INFORMS Journal on Computing, 37(4): 856–873, July/August 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0243.

Zhang:2025:CTE

[2069] Haoting Zhang, Donglin Zhan, James Anderson, Rhonda Righter, and Zeyu Zheng. Clustering then estimation of spatio-temporal self-exciting processes. INFORMS Journal on Computing, 37(4):874–893, July/ August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0351.

Chen:2025:DSK

[2070] Zhen-Yu Chen and Minghe Sun. Deep stacking kernel machines for the data-driven multi-item, one-warehouse, multiretailer problems with backlog and lost sales. INFORMS Journal on Computing, 37(4):894– 916, July/August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/full/
10.1287/ijoc.2022.0365.

Chen:2025:PSW

[2071] Caihua Chen, Junhao Tao, and Yang Zhan. Pairwise stability in weighted network formation games: Selection and computation. IN-FORMS Journal on Computing, 37(4):917–932, July/August 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2024.0546.

Eifler:2025:CPB

[2072] Leon Eifler, Jules Nicolas-Thouvenin, and Ambros Gleixner. Combining precision boosting with LP iterative refinement for exact linear optimization. INFORMS Journal on Computing, 37(4):933–944, July/ August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0409.

Chen:2025:FUT

[2073] Xinyu Chen, Xi-Le Zhao, and Chun Cheng. Forecasting urban traffic states with sparse data using Hankel temporal matrix factorization. IN-FORMS Journal on Computing, 37(4):945–961, July/August 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

//pubsonline.informs.org/doi/full/10.1287/ijoc.2022.0197.

Burkowski:2025:PRS

[2074] Forbes Burkowski, Haesol Im, and Henry Wolkowicz. A Peaceman– Rachford splitting method for the protein side-chain positioning problem. INFORMS Journal on Computing, 37(4):962–976, July/August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0094.

Zhang:2025:DID

[2075] Jiachen Zhang and J. Christopher Beck. Domain-independent dynamic programming and constraint programming approaches for assembly line balancing problems with setups. INFORMS Journal on Computing, 37 (4):977–997, July/August 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2024.0603.

Qiu:2025:FDI

[2076] Yongqin Qiu, Yuanxing Chen, Kan Fang, Lean Yu, and Kuangnan Fang. Fraud detection by integrating multisource heterogeneous presenceonly data. INFORMS Journal on Computing, 37(4):998–1017, July/ August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0366.

Lai:2025:EOM

[2077] Xiangjing Lai, Zhenheng Lin, Jin-Kao Hao, and Qinghua Wu. An efficient optimization model and tabu search-based global optimization approach for the continuous p-dispersion problem. INFORMS Journal on Computing, 37(4):1018–1043, July/August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0089.

Baumann:2025:ACC

[2078] Philipp Baumann and Dorit S. Hochbaum. An algorithm for clustering with confidence-based must-link and cannot-link constraints. INFORMS Journal on Computing, 37(4):1044–1068, July/August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0419.

Zhong:2025:IDC

[2079] Haonan Zhong, Foad Mahdavi Pajouh, Sergiy Butenko, and Oleg A. Prokopyev. On interdicting dense clusters in a network. INFORMS Journal on Computing, 37(4):1069–1086, July/August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0027.

Lai:2025:SCI

[2080] Qidong Lai, Guangwu Liu, Bingfeng Zhang, and Kun Zhang. Simulating confidence intervals for conditional value-at-risk via least-squares metamodels. INFORMS Journal on Computing, 37(4):1087–1105, July/ August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0394.

Zhou:2025:NAA

[2081] Jiao Zhou, Yingli Ran, Panos M. Pardalos, Zhao Zhang, Shaojie Tang, and Ding-Zhu Du. A new approximation algorithm for minimum-weight (1,m)–connected dominating set. INFORMS Journal on Computing, 37 (4):1106–1120, July/August 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0306.

Manns:2025:DSI

[2082] Paul Manns and Marvin Severitt. On discrete subproblems in integer optimal control with total variation regularization in two dimensions. INFORMS Journal on Computing, 37(4):1121–1141, July/ August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2024.0680.

Anonymous:2025:EBd

[2083] Anonymous. Editorial board. INFORMS Journal on Computing, 37(4): ??, July/August 2025. CODEN ???? ISSN 1091-9856 (print), 1526-

5528 (electronic). URL https://pubsonline.informs.org/doi/abs/
10.1287/ijoc.2025.eb.v3704.

Smith:2025:NEe

[2084] Alice E. Smith. Note from the Editor. INFORMS Journal on Computing, 37(5):??, September/October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2025.ed.v37.n5.

Nannicini:2025:IQC

[2085] Giacomo Nannicini. Introduction to quantum computing area. IN-FORMS Journal on Computing, 37(5):??, September/October 2025. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:/

/pubsonline.informs.org/doi/full/10.1287/ijoc.2025.qcintro.
v37.n5.

Yu:2025:VRA

[2086] Xian Yu and Siqian Shen. On the value of risk-averse multistage stochastic programming in capacity planning. INFORMS Journal on Computing, 37(5):1143–1162, September/October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0396.

Bertsimas:2025:SBD

[2087] Dimitris Bertsimas, Ryan Cory-Wright, Jean Pauphilet, and Periklis Petridis. A stochastic Benders decomposition scheme for large-scale stochastic network design. INFORMS Journal on Computing, 37(5): 1163–1181, September/October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.0074.

Qu:2025:ESQ

[2088] Yan Qu, Angelos Dassios, Anxin Liu, and Hongbiao Zhao. Exact simulation of quadratic intensity models. INFORMS Journal on Computing, 37 (5):1182–1201, September/October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.informs.

org/doi/full/10.1287/ijoc.2023.0323.

Lancia:2025:ACS

[2089] Giuseppe Lancia and Paolo Vidoni. Average case subquadratic exact and heuristic procedures for the traveling salesman 2-OPT neighborhood. INFORMS Journal on Computing, 37(5):1202–1222, September/ October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0169.

Helfrich:2025:ECC

[2090] Stephan Helfrich, Stefan Ruzika, and Clemens Thielen. Efficiently constructing convex approximation sets in multiobjective optimization problems. INFORMS Journal on Computing, 37(5):1223–1241, September/ October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0220.

Ocampo-Giraldo:2025:IEA

[2091] Daniel M. Ocampo-Giraldo, Ana M. Anaya-Arenas, and Claudio Contardo. An iterative exact algorithm over a time-expanded network for the transportation of biomedical samples. INFORMS Journal on Computing, 37(5):1242–1266, September/October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://pubsonline.

informs.org/doi/full/10.1287/ijoc.2023.0061.

Huang:2025:ICB

[2092] Dian Huang, Zhaofang Mao, Kan Fang, Enyuan Fu, and Michael L. Pinedo. An improved combinatorial Benders decomposition algorithm for the human-robot collaborative assembly line balancing problem. IN-FORMS Journal on Computing, 37(5):1267–1283, September/October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0279.

Khandan:2025:IPE

[2093] Melika Khandan and Pooya Hoseinpour. Incorporating promotional effects in sales planning of the retail industry using geometric programming. INFORMS Journal on Computing, 37(5):1284–1305, September/ October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0275.

Halman:2025:FPT

[2094] Nir Halman and Giacomo Nannicini. Fully polynomial time approximation schemes for robust multistage decision making. INFORMS Journal on Computing, 37(5):1306–1327, September/October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/doi/full/10.1287/ijoc.2023.0126.

Zhang:2025:BPC

[2095] Rui Zhang. Branch-and-price for the capacitated autonomous vehicle assisted delivery problem. INFORMS Journal on Computing, 37(5):1328– 1349, September/October 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0177.

Zhou:2025:RFS

[2096] Xin Zhou, Yang Li, Zemin Zheng, Jie Wu, and Jiarui Zhang. Reproducible feature selection for high-dimensional measurement error models. INFORMS Journal on Computing, 37(5):1350–1368, September/ October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2023.0282.

Liu:2025:DDO

[2097] Junming Liu, Weiwei Chen, and Leilei Sun. A data-driven optimization framework for static rebalancing operations in bike sharing systems. INFORMS Journal on Computing, 37(5):1369–1390, September/ October 2025. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (elec-

tronic). URL https://pubsonline.informs.org/doi/full/10.1287/
ijoc.2022.0182.

Li:2025:FSG

[2098] Tie Li, Gang Kou, Yi Peng, and Philip S. Yu. Feature selection and grouping effect analysis for credit evaluation via regularized diagonal distance metric learning. INFORMS Journal on Computing, 37(5):1391– 1412, September/October 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0322.

Yang:2025:CFT

[2099] Tianyu Yang, Jiongbai Liu, Tasnim Ibn Faiz, Chrysafis Vogiatzis, and Md. Noor-E-Alam. Computational framework for target tracking information fusion problems. INFORMS Journal on Computing, 37(5):1413– 1432, September/October 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
full/10.1287/ijoc.2023.0016.

Anonymous:2025:EBe

[2100] Anonymous. Editorial board. INFORMS Journal on Computing, 37(5): ??, September/October 2025. CODEN ???? ISSN 1091-9856 (print),

1526-5528 (electronic). URL https://pubsonline.informs.org/doi/
abs/10.1287/ijoc.2025.eb.v3705.